Maxima Manual

1 2 Maxima Manual

1 Lists

1.1 Introduction to Lists

Lists are the basic building block for maxima and lisp. All data types other than arrays, hash tables, numbers are represented as lisp lists, These lisp lists have the form

((mplus) $A 2)

to indicate an expression A+2. At maxima level one would see the infix notation A+2. Maxima also has lists which are printed as

[1, 2, 7, x+y]

for a list with 4 elements. Internally this corresponds to a lisp list of the form

((mlist) 1  2  7  ((mplus)  $X $Y ))

The flag which denotes the type field of the maxima expression is a list itself, since after it has been through the simplifier the list would become

((mlist simp) 1 2 7 ((mplus simp) $X $Y))

1.2 Definitions for Lists

Function APPEND (list1, list2, ...)

returns a single list of the elements of list1 followed by the elements of list2,... AP-PEND also works on general expressions, e.g. APPEND(F(A,B), F(C,D,E)); -> F(A,B,C,D,E). Do EXAMPLE(APPEND); for an example.

Function ATOM (exp)

is TRUE if exp is atomic (i.e. a number or name) else FALSE. Thus ATOM(5) is TRUE while ATOM(A[1]) and ATOM(SIN(X)) are FALSE. (Assuming A[1] and X are unbound.)

Function CONS (exp, list)

returns a new list constructed of the element exp as its first element, followed by the elements of list. CONS also works on other expressions, e.g. CONS(X, F(A,B,C)); -> F(X,A,B,C).

Function COPYLIST (L)

creates a copy of the list L.

Function DELETE (exp1, exp2)

removes all occurrences of exp1 from exp2. Exp1 may be a term of exp2 (if it is a sum) or a factor of exp2 (if it is a product).

     (C1)  DELETE(SIN(X),X+SIN(X)+Y);
     (D1)               Y + X

DELETE(exp1, exp2, integer) removes the first integer occurrences of exp1 from exp2. Of course, if there are fewer than integer occurrences of exp1 in exp2 then all occurrences will be deleted.

Function ENDCONS (exp, list)

returns a new list consisting of the elements of list followed by exp. ENDCONS also works on general expressions, e.g. ENDCONS(X, F(A,B,C)); -> F(A,B,C,X).

Function FIRST (exp)

yields the first part of exp which may result in the first element of a list, the first row of a matrix, the first term of a sum, etc. Note that FIRST and its related functions, REST and LAST, work on the form of exp which is displayed not the form which is typed on input. If the variable INFLAG [FALSE] is set to TRUE however, these functions will look at the internal form of exp. Note that the simplifier re-orders expressions. Thus FIRST(X+Y) will be X if INFLAG is TRUE and Y if INFLAG is FALSE. (FIRST(Y+X) gives the same results).

Function GET (a, i)

retrieves the user property indicated by i associated with atom a or returns FALSE if a doesn’t have property i.

     (C1) PUT(%E,'TRANSCENDENTAL,'TYPE);
     (D1)          TRANSCENDENTAL
     (C2) PUT(%PI,'TRANSCENDENTAL,'TYPE)$
     (C3) PUT(%I,'ALGEBRAIC,'TYPE)$
     (C4) TYPEOF(EXP) := BLOCK([Q],
                             IF NUMBERP(EXP)
                             THEN RETURN('ALGEBRAIC),
                             IF NOT ATOM(EXP)
                             THEN RETURN(MAPLIST('TYPEOF, EXP)),
                             Q : GET(EXP, 'TYPE),
                             IF Q=FALSE
     THEN ERRCATCH(ERROR(EXP,"is not numeric.")) ELSE Q)$
     (C5) TYPEOF(2*%E+X*%PI);
     X is not numeric.
     (D5)       [[TRANSCENDENTAL, []], [ALGEBRAIC, TRANSCENDENTAL]]
     (C6) TYPEOF(2*%E+%PI);
     (D6)          [TRANSCENDENTAL, [ALGEBRAIC, TRANSCENDENTAL]]

Function LAST (exp)

yields the last part (term, row, element, etc.) of the exp.

Function LENGTH (exp)

gives (by default) the number of parts in the external (displayed) form of exp. For lists this is the number of elements, for matrices it is the number of rows, and for sums it is the number of terms. (See DISPFORM). The LENGTH command is affected by the INFLAG switch [default FALSE]. So, e.g. LENGTH(A/(B*C)); gives 2 if INFLAG is FALSE (Assuming EXPTDISPFLAG is TRUE), but 3 if INFLAG is TRUE (the internal representation is essentially A*B^-1*C^-1).

Variable LISTARITH

default: [TRUE] - if FALSE causes any arithmetic operations with lists to be suppressed; when TRUE, list-matrix operations are contagious causing lists to be converted to matrices yielding a result which is always a matrix. However, list-list operations should return lists.

Function LISTP (exp)

is TRUE if exp is a list else FALSE.

Function MAKELIST (exp,var,lo,hi)

returns a list as value. MAKELIST may be called as MAKELIST(exp,var,lo,hi) ["lo" and "hi" must be integers], or as MAKELIST(exp,var,list). In the first case MAKELIST is analogous to SUM, whereas in the second case MAKELIST is similar to MAP. Examples:

        MAKELIST(CONCAT(X,I),I,1,6) yields [X1,X2,X3,X4,X5,X6]
        MAKELIST(X=Y,Y,[A,B,C]) yields [X=A,X=B,X=C]

Function MEMBER (exp, list)

returns TRUE if exp occurs as a member of list (not within a member). Otherwise FALSE is returned. Member also works on non-list expressions, e.g. MEMBER(B, F(A,B,C)); -> TRUE.

Function REST (exp, n)

yields exp with its first n elements removed if n is positive and its last -n elements removed if n is negative. If n is 1 it may be omitted. Exp may be a list, matrix, or other expression.

Function REVERSE (list)

reverses the order of the members of the list (not the members themselves). RE-VERSE also works on general expressions, e.g. REVERSE(A=B); gives B=A. RE-VERSE default: [FALSE] - in the Plotting functions, if TRUE cause a left-handed coordinate system to be assumed.

6 Maxima Manual

2 Differentiation

2.1 Definitions for Differentiation

Function ANTID (G,X,U(X))

A routine for evaluating integrals of expressions involving an arbitrary unspecified function and its derivatives. It may be used by LOAD(ANTID); , after which, the function ANTIDIFF may be used. E.g. ANTIDIFF(G,X,U(X)); where G is the expression involving U(X) (U(X) arbitrary) and its derivatives, whose integral with respect to X is desired. The functions NONZEROANDFREEOF and LINEAR are also defined, as well as ANTID. ANTID is the same as ANTIDIFF except that it returns a list of two parts, the first part is the integrated part of the expression and the second part of the list is the non-integrable remainder.

Function ANTIDIFF -

See ANTID.

property ATOMGRAD

- the atomic gradient property of an expression. May be set by GRADEF.

Function ATVALUE (form, list, value)

enables the user to assign the boundary value value to form at the points specified by list.

     (C1) ATVALUE(F(X,Y),[X=0,Y=1],A**2)$
         The form must be a function, f(v1,v2,...), or a derivative,

DIFF(f(v1,v2,...),vi,ni,vj,nj,...) in which the functional arguments explicitly appear (ni is the order of differentiation with respect vi). The list of equations determine the "boundary" at which the value is given; list may be a list of equations, as above, or a single equation, vi = expr. The symbols @1, @2,... will be used to represent the functional variables v1,v2,... when atvalues are displayed. PRINTPROPS([f1, f2,...], ATVALUE) will display the atvalues of the functions f1,f2,... as specified in previously given uses of the ATVALUE function. If the list contains just one element then the element can be given without being in a list. If a first argument of ALL is given then atvalues for all functions which have them will be displayed. Do EXAMPLE(ATVALUE); for an example.

Function CARTAN -

The exterior calculus of differential forms is a basic tool of differential geometry developed by Elie Cartan and has important applications in the theory of partial differential equations. The present implementation is due to F.B. Estabrook and H.D. Wahlquist. The program is self-explanatory and can be accessed by doing batch("cartan"); which will give a description with examples.

Function DELTA (t)

This is the Dirac Delta function. Currently only LAPLACE knows about the DELTA function:

(C1) LAPLACE(DELTA(T-A)*SIN(B*T),T,S);
Is A positive, negative or zero?
POS;
                                          - A S
(D1)                           SIN(A B) %E

Variable DEPENDENCIES

default: [] - the list of atoms which have functional dependencies (set up by the DE-PENDS or GRADEF functions). The command DEPENDENCIES has been replaced by the DEPENDS command. Do DESCRIBE(DEPENDS);

Function DEPENDS (funlist1,varlist1,funlist2,varlist2,...)

declares functional dependencies for variables to be used by DIFF.

DEPENDS([F,G],[X,Y],[R,S],[U,V,W],U,T)

informs DIFF that F and G depend on X and Y, that R and S depend on U,V, and W, and that U depends on T. The arguments to DEPENDS are evaluated. The variables in each funlist are declared to depend on all the variables in the next varlist. A funlist can contain the name of an atomic variable or array. In the latter case, it is assumed that all the elements of the array depend on all the variables in the succeeding varlist. Initially, DIFF(F,X) is 0; executing DEPENDS(F,X) causes future differentiations of F with respect to X to give dF/dX or Y (if DERIVABBREV:TRUE).

(C1) DEPENDS([F,G],[X,Y],[R,S],[U,V,W],U,T);
(D1)           [F(X, Y), G(X, Y), R(U, V, W), S(U, V, W), U(T)]
(C2) DEPENDENCIES;
(D2)           [F(X, Y), G(X, Y), R(U, V, W), S(U, V, W), U(T)]
(C3) DIFF(R.S,U);
                               dR           dS
(D3)                           -- . S + R . --
                               dU           dU

Since MACSYMA knows the chain rule for symbolic derivatives, it takes advantage of the given dependencies as follows:

(C4) DIFF(R.S,T);
                           dR dU             dS dU
(D4)                      (-- --) . S + R . (-- --)
                           dU dT             dU dT
If we set
(C5) DERIVABBREV:TRUE;
(D5)                                 TRUE
then re-executing the command C4, we obtain
(C6) ''C4;
(D6)                      (R  U ) . S + R . (S  U )
                            U  T              U  T

To eliminate a previously declared dependency, the REMOVE command can be used. For example, to say that R no longer depends on U as declared in C1, the user can type

     REMOVE(R,DEPENDENCY)

This will eliminate all dependencies that may have been declared for R.

     (C7) REMOVE(R,DEPENDENCY);
     (D7)                                 DONE
     (C8) ''C4;
     (D8)                             R . (S  U )
                                       U    T

CAVEAT: DIFF is the only MACSYMA command which uses DEPENDENCIES information. The arguments to INTEGRATE, LAPLACE, etc. must be given their dependencies explicitly in the command, e.g., INTEGRATE(F(X),X).

Variable DERIVABBREV

default: [FALSE] if TRUE will cause derivatives to display as subscripts.

Function DERIVDEGREE (exp, dv, iv)

finds the highest degree of the derivative of the dependent variable dv with respect to the independent variable iv occuring in exp.

     (C1) 'DIFF(Y,X,2)+'DIFF(Y,Z,3)*2+'DIFF(Y,X)*X**2$
     (C2) DERIVDEGREE(%,Y,X);
     (D2)                           2

Function DERIVLIST (var1,...,vark)

causes only differentiations with respect to the indicated variables, within the EV command.

Variable DERIVSUBST

default: [FALSE] - controls non-syntactic substitutions such as

     SUBST(X,'DIFF(Y,T),'DIFF(Y,T,2));

If DERIVSUBST is set to true, this gives ’DIFF(X,T).

special symbol DIFF

[flag] for ev causes all differentiations indicated in exp to be performed.

Function DIFF (exp, v1, n1, v2, n2, ...)

DIFF differentiates exp with respect to each vi, ni times. If just the first derivative with respect to one variable is desired then the form DIFF(exp,v) may be used. If the noun form of the function is required (as, for example, when writing a differential equation), ’DIFF should be used and this will display in a two dimensional format. DERIVABBREV[FALSE] if TRUE will cause derivatives to display as subscripts.

DIFF(exp) gives the "total differential", that is, the sum of the derivatives of exp with respect to each of its variables times the function DEL of the variable. No further simplification of DEL is offered.

     (C1) DIFF(EXP(F(X)),X,2);
                                  2
                           F(X)  d             F(X)  d         2
     (D1)                %E     (--- F(X)) + %E     (-- (F(X)))
                                   2                 dX
                                 dX
     (C2) DERIVABBREV:TRUE$
     (C3) 'INTEGRATE(F(X,Y),Y,G(X),H(X));
                                     H(X)
                                    /
                                    [
     (D3)                           I     F(X, Y) dY
                                    ]
                                    /
                                     G(X)
     (C4) DIFF(%,X);
                 H(X)
                /
                [
     (D4)       I     F(X, Y)  dY + F(X, H(X)) H(X)  - F(X, G(X)) G(X)
                ]            X                     X                  X
                /
                 G(X)

For the tensor package, the following modifications have been incorporated: 1) the derivatives of any indexed objects in exp will have the variables vi appended as additional arguments. Then all the derivative indices will be sorted. 2) the vi may be integers from 1 up to the value of the variable DIMENSION[default value: 4]. This will cause the differentiation to be carried out wrt the vith member of the list COOR-DINATES which should be set to a list of the names of the coordinates, e.g., [x,y,z,t]. If COORDINATES is bound to an atomic variable, then that variable subscripted by vi will be used for the variable of differentiation. This permits an array of coordinate names or subscripted names like X[1], X[2],... to be used. If COORDINATES has not been assigned a value, then the variables will be treated as in 1) above.

Function DSCALAR (function)

applies the scalar d’Alembertian to the scalar function.

     (C41) DEPENDENCIES(FIELD(R));
     (D41)                           [FIELD(R)]
     (C42) DSCALAR(FIELD);
     (D43)
         -M
       %E  ((FIELD  N - FIELD  M + 2 FIELD   ) R + 4 FIELD )
                  R  R       R  R         R R             R

     - -----------------------------------------------------
                                  2 R

Function EXPRESS (expression)

The result uses the noun form of any derivatives arising from expansion of the vector differential operators. To force evaluation of these derivatives, the built-in EV function can be used together with the DIFF evflag, after using the built-in DEPENDS function to establish any new implicit dependencies.

Function GENDIFF

Sometimes DIFF(E,X,N) can be reduced even though N is symbolic.

     batch("gendif.mc")$

and you can try, for example,

     DIFF(%E^(A*X),X,Q)

by using GENDIFF rather than DIFF. Unevaluable items come out quoted. Some items are in terms of "GENFACT", which see.

Function GRADEF (f(x1, ..., xn), g1, ..., gn)

defines the derivatives of the function f with respect to its n arguments. That is, df/dxi = gi, etc. If fewer than n gradients, say i, are given, then they refer to the first i arguments of f. The xi are merely dummy variables as in function definition headers and are used to indicate the ith argument of f. All arguments to GRADEF except the first are evaluated so that if g is a defined function then it is invoked and the result is used. Gradients are needed when, for example, a function is not known explicitly but its first derivatives are and it is desired to obtain higher order derivatives. GRADEF may also be used to redefine the derivatives of MACSYMA’s predefined functions (e.g. GRADEF(SIN(X),SQRT(1-SIN(X)**2)) ). It is not permissible to use GRADEF on subscripted functions. GRADEFS is a list of the functions which have been given gradients by use of the GRADEF command (i.e. GRADEF(f(x1, ..., xn), g1, ..., gn)). PRINTPROPS([f1,f2,...],GRADEF) may be used to display the gradefs of the functions f1,f2,.. GRADEF(a,v,exp) may be used to state that the derivative of the atomic variable a with respect to v is exp. This automatically does a DEPENDS(a,v). PRINTPROPS([a1,a2,...],ATOMGRAD) may be used to display the atomic gradient properties of a1,a2,...

Variable GRADEFS

default: [] - a list of the functions which have been given gradients by use of the GRADEF command (i.e. GRADEF(f(x1, ..., xn), g1, ..., gn).)

Function LAPLACE (exp, ovar, lvar)

takes the Laplace transform of exp with respect to the variable ovar and transform parameter lvar. Exp may only involve the functions EXP, LOG, SIN, COS, SINH, COSH, and ERF. It may also be a linear, constant coefficient differential equation in which case ATVALUE of the dependent variable will be used. These may be supplied either before or after the transform is taken. Since the initial conditions must be specified at zero, if one has boundary conditions imposed elsewhere he can impose these on the general solution and eliminate the constants by solving the general solution for them and substituting their values back. Exp may also involve convolution integrals. Functional relationships must be explicitly represented in order for LAPLACE to work properly. That is, if F depends on X and Y it must be written as F(X,Y) wherever F occurs as in LAPLACE(’DIFF(F(X,Y),X),X,S). LAPLACE is not affected by DEPENDENCIES set up with the DEPENDS command.

     (C1) LAPLACE(%E**(2*T+A)*SIN(T)*T,T,S);
                          A
                      2 %E  (S - 2)
     (D1)            ---------------
                             2     2
                     ((S - 2)  + 1)

Function UNDIFF (exp)

returns an expression equivalent to exp but with all derivatives of indexed objects replaced by the noun form of the DIFF function with arguments which would yield that indexed object if the differentiation were carried out. This is useful when it is desired to replace a differentiated indexed object with some function definition and then carry out the differentiation by saying EV(...,DIFF).

3 Integration

3.1 Introduction to Integration

MACSYMA has several routines for handling integration. The INTEGRATE command makes use of most of them. There is also the ANTID package, which handles an unspecified function (and its derivatives, of course). For numerical uses, there is the ROMBERG function, and the IMSL version of Romberg, DCADRE. There is also an adaptave integrator which uses the Newton-Cotes 8 panel quadrature rule, called QUANC8. Hypergeometric Functions are being worked on, do DESCRIBE(SPECINT); for details. Generally speaking, MACSYMA only handles integrals which are integrable in terms of the "elementary functions" (rational functions, trigonometrics, logs, exponentials, radicals, etc.) and a few extensions (error function, dilogarithm). It does not handle integrals in terms of unknown functions such as g(x) and h(x).

3.2 Definitions for Integration

Function CHANGEVAR (exp,f(x,y),y,x)

makes the change of variable given by f(x,y) = 0 in all integrals occurring in exp with integration with respect to x; y is the new variable.

              (C1) 'INTEGRATE(%E**SQRT(A*Y),Y,0,4);
                                  4
                                 /
                                 [    SQRT(A) SQRT(Y)
              (D1)               I (%E               ) dY
                                 ]
                                 /
                                  0
              (C2) CHANGEVAR(D1,Y-Z^2/A,Z,Y);
                                             2 SQRT(A)
                                            /
                                            [              Z
                                          2 I          Z %E  dZ
                                            ]
                                            /
                                             0
              (D4)                        ---------------------
                                                    A

CHANGEVAR may also be used to changes in the indices of a sum or product. However, it must be realized that when a change is made in a sum or product, this change must be a shift, i.e. I=J+ ..., not a higher degree function. E.g.

          (C3) SUM(A[I]*X^(I-2),I,0,INF);
                                         INF
                                         ====
                                         \         I - 2

     (D3)                            >    A  X
                                    /      I
                                    ====
                                    I = 0
     (C4) CHANGEVAR(%,I-2-N,N,I);
                                   INF
                                   ====
                                   \               N
     (D4)                           >      A      X
                                   /        N + 2
                                   ====
                                   N = - 2

Function DBLINT (’F,’R,’S,a,b)

a double-integral routine which was written in top-level macsyma and then translated and compiled to machine code. Use LOAD(DBLINT); to access this package. It uses the Simpson’s Rule method in both the x and y directions to calculate /B /S(X) | | | | F(X,Y) DY DX . | | /A /R(X) The function F(X,Y) must be a translated or compiled function of two variables, and R(X) and S(X) must each be a translated or compiled function of one variable, while a and b must be floating point numbers. The routine has two global variables which determine the number of divisions of the x and y intervals: DBLINT X and DBLINT Y, both of which are initially 10, and can be changed independently to other integer values (there are 2*DBLINT X+1 points computed in the x direction, and 2*DBLINT Y+1 in the y direction). The routine subdivides the X axis and then for each value of X it first computes R(X) and S(X); then the Y axis between R(X) and S(X) is subdivided and the integral along the Y axis is performed using Simpson’s Rule; then the integral along the X axis is done using Simpson’s Rule with the function values being the Y-integrals. This procedure may be numerically unstable for a great variety of reasons, but is reasonably fast: avoid using it on highly oscillatory functions and functions with singularities (poles or branch points in the region). The Y integrals depend on how far apart R(X) and S(X) are, so if the distance S(X)-R(X) varies rapidly with X, there may be substantial errors arising from truncation with different step-sizes in the various Y integrals. One can increase DBLINT X and DBLINT Y in an effort to improve the coverage of the region, at the expense of computation time. The function values are not saved, so if the function is very time-consuming, you will have to wait for recomputation if you change anything (sorry). It is required that the functions F, R, and S be either translated or compiled prior to calling DBLINT. This will result in orders of magnitude speed improvement over interpreted code in many cases! Ask LPH (or GJC) about using these numerical aids. The file SHARE1;DBLINT DEMO can be run in batch or demo mode to illustrate the usage on a sample problem; the file SHARE1;DBLNT DEMO1 is an extension of the DEMO which also makes use of other numerical aids, FLOATDEFUNK and QUANC8. Please send all bug notes and questions to LPH Function DEFINT (exp, var, low, high)

DEFinite INTegration, the same as INTEGRATE(exp,var,low,high).

Function ERF (X)

the error function, whose derivative is: 2*EXP(-X^2)/SQRT(%PI).

Variable ERFFLAG

default: [TRUE] if FALSE prevents RISCH from introducing the ERF function in the answer if there were none in the integrand to begin with.

Variable ERRINTSCE

default: [TRUE] - If a call to the INTSCE routine is not of the form

     EXP(A*X+B)*COS(C*X)^N*SIN(C*X)

then the regular integration program will be invoked if the switch ERRINTSCE[TRUE] is TRUE. If it is FALSE then INTSCE will err out.

Function ILT (exp, lvar, ovar)

takes the inverse Laplace transform of exp with respect to lvar and parameter ovar. exp must be a ratio of polynomials whose denominator has only linear and quadratic factors. By using the functions LAPLACE and ILT together with the SOLVE or LINSOLVE functions the user can solve a single differential or convolution integral equation or a set of them.

     (C1) 'INTEGRATE(SINH(A*X)*F(T-X),X,0,T)+B*F(T)=T**2;
                   T
                  /
                  [                                     2
     (D1)         I (SINH(A X) F(T - X)) dX + B F(T) = T
                  ]
                  /
                  0
     (C2) LAPLACE(%,T,S);
                 A LAPLACE(F(T), T, S)
     (D2)        ---------------------
                         2    2
                        S  - A
                                               2
                     + B LAPLACE(F(T), T, S) = --
                                                3
                                               S
     (C3) LINSOLVE([%],['LAPLACE(F(T),T,S)]);
     SOLUTION
                                             2      2
                                          2 S  - 2 A
     (E3)       LAPLACE(F(T), T, S) = --------------------
                                         5         2     3
                                      B S  + (A - A  B) S
     (D3)                         [E3]

 (C4) ILT(E3,S,T);
 IS  A B (A B - 1)  POSITIVE, NEGATIVE, OR ZERO?
 POS;
                                        2
                        SQRT(A) SQRT(A B  - B) T
                 2 COSH(------------------------)
                                   B
 (D4)  F(T) =  - --------------------------------
                                A
               2
            A T             2
         + ------- + ------------------
           A B - 1    3  2      2
                     A  B  - 2 A  B + A

Function INTEGRATE (exp, var)

integrates exp with respect to var or returns an integral expression (the noun form) if it cannot perform the integration (see note 1 below). Roughly speaking three stages are used:

• (1) INTEGRATE sees if the integrand is of the form F(G(X))*DIFF(G(X),X) by testing whether the derivative of some subexpression (i.e. G(X) in the above case) divides the integrand. If so it looks up F in a table of integrals and substitutes G(X) for X in the integral of F. This may make use of gradients in taking the derivative. (If an unknown function appears in the integrand it must be eliminated in this stage or else INTEGRATE will return the noun form of the integrand.)

• (2) INTEGRATE tries to match the integrand to a form for which a specific method can be used, e.g. trigonometric substitutions.

• (3) If the first two stages fail it uses the Risch algorithm. Functional relationships must be explicitly represented in order for INTEGRATE to work properly. INTEGRATE is not affected by DEPENDENCIES set up with the DEPENDS command. INTEGRATE(exp, var, low, high) finds the definite integral of exp with respect to var from low to high or returns the noun form if it cannot perform the integration. The limits should not contain var. Several methods are used, including direct substitution in the indefinite integral and contour integration. Improper integrals may use the names INF for positive infinity and MINF for negative infinity. If an integral "form" is desired for manipulation (for example, an integral which cannot be computed until some numbers are substituted for some parameters), the noun form ’INTEGRATE may be used and this will display with an integral sign. (See Note 1 below.) The function LDEFINT uses LIMIT to evaluate the integral at the lower and upper limits. Sometimes during integration the user may be asked what the sign of an expression is. Suitable responses are POS;, ZERO;, or NEG;.

 (C1) INTEGRATE(SIN(X)**3,X);

                 COS (X)
(D1)             ------- - COS(X)
                    3
(C2) INTEGRATE(X**A/(X+1)**(5/2),X,0,INF);
IS  A + 1  POSITIVE, NEGATIVE, OR ZERO?
POS;
IS  2 A - 3  POSITIVE, NEGATIVE, OR ZERO?
NEG;
                            3
(D2)            BETA(A + 1, - - A)
                            2
(C3) GRADEF(Q(X),SIN(X**2));
(D3)                              Q(X)
(C4) DIFF(LOG(Q(R(X))),X);
                           d             2
                          (-- R(X)) SIN(R (X))
                           dX
(D4)                      --------------------
                                Q(R(X))
(C5) INTEGRATE(%,X);
(D5)                          LOG(Q(R(X)))

(Note 1) The fact that MACSYMA does not perform certain integrals does not always imply that the integral does not exist in closed form. In the example below the integration call returns the noun form but the integral can be found fairly easily. For example, one can compute the roots of X^3+X+1 = 0 to rewrite the integrand in the form

1/((X-A)*(X-B)*(X-C))

where A, B and C are the roots. MACSYMA will integrate this equivalent form although the integral is quite complicated.

(C6) INTEGRATE(1/(X^3+X+1),X);
                          /
                          [     1
(D6)                      I ---------- dX
                          ]  3
                          / X  + X + 1

Variable INTEGRATION CONSTANT COUNTER

- a counter which is updated each time a constant of integration (called by MAC-SYMA, e.g., "INTEGRATIONCONSTANT1") is introduced into an expression by indefinite integration of an equation.

Function INTSCE (expr,var)

INTSCE LISP contains a routine, written by Richard Bogen, for integrating products of sines,cosines and exponentials of the form

  EXP(A*X+B)*COS(C*X)^N*SIN(C*X)^M

The call is INTSCE(expr,var) expr may be any expression, but if it is not in the above form then the regular integration program will be invoked if the switch ERRINTSCE[TRUE] is TRUE. If it is FALSE then INTSCE will err out.

Function LDEFINT (exp,var,ll,ul)

yields the definite integral of exp by using LIMIT to evaluate the indefinite integral of exp with respect to var at the upper limit ul and at the lower limit ll.

Function POTENTIAL (givengradient)

The calculation makes use of the global variable

     POTENTIALZEROLOC[0]

which must be NONLIST or of the form

     [indeterminatej=expressionj, indeterminatek=expressionk, ...]

the former being equivalent to the nonlist expression for all right-hand sides in the latter. The indicated right-hand sides are used as the lower limit of integration. The success of the integrations may depend upon their values and order. POTENTIALZE-ROLOC is initially set to 0.

Function QQ

- The file SHARE1;QQ FASL (which may be loaded with LOAD("QQ");) contains a function QUANC8 which can take either 3 or 4 arguments. The 3 arg version computes the integral of the function specified as the first argument over the interval from lo to hi as in QUANC8(’function name,lo,hi); . The function name should be quoted. The 4 arg version will compute the integral of the function or expression (first arg) with respect to the variable (second arg) over the interval from lo to hi as in QUANC8(<f(x) or expression in x>,x,lo,hi). The method used is the Newton-Cotes 8th order polynomial quadrature, and the routine is adaptive. It will thus spend time dividing the interval only when necessary to achieve the error conditions specified by the global variables QUANC8 RELERR (default value=1.0e-4) and QUANC8 ABSERR (default value=1.0e-8) which give the relative error test: |integral(function)-computed value|< quanc8 relerr*|integral(function)| and the absolute error test: |integral(function)-computed value|<quanc8 abserr. Do PRINT-FILE(QQ,USAGE,SHARE1) for details.

Function QUANC8 (’function name,lo,hi)

An adaptive integrator, available in SHARE1;QQ FASL. DEMO and USAGE files are provided. The method is to use Newton-Cotes 8-panel quadrature rule, hence the function name QUANC8, available in 3 or 4 arg versions. Absolute and relative error checks are used. To use it do LOAD("QQ"); For more details do DESCRIBE(QQ); .

Function RESIDUE (exp, var, val)

computes the residue in the complex plane of the expression exp when the variable var assumes the value val. The residue is the coefficient of (var-val)**(-1) in the Laurent series for exp.

(C1) RESIDUE(S/(S**2+A**2),S,A*%I);
                        1
(D1)                    -
                        2
(C2) RESIDUE(SIN(A*X)/X**4,X,0);
                        3
                       A
(D2)                 - --
                       6

Function RISCH (exp, var)

integrates exp with respect to var using the transcendental case of the Risch algorithm. (The algebraic case of the Risch algorithm has not been implemented.) This currently handles the cases of nested exponentials and logarithms which the main part of INTEGRATE can’t do. INTEGRATE will automatically apply RISCH if given these cases. ERFFLAG[TRUE] - if FALSE prevents RISCH from introducing the ERF function in the answer if there were none in the integrand to begin with.

(C1) RISCH(X^2*ERF(X),X);
            2     2
         - X     X             3           2
       %E     (%E   SQRT(%PI) X  ERF(X) + X  + 1)
(D1)   ------------------------------------------
                      3 SQRT(%PI)
(C2) DIFF(%,X),RATSIMP;
                                 2
(D2)                            X  ERF(X)

Function ROMBERG (exp,var,ll,ul)

or ROMBERG(exp,ll,ul) - Romberg Integration. You need not load in any file to use ROMBERG, it is autoloading. There are two ways to use this function. The first is an inefficient way like the definite integral version of INTEGRATE: ROMBERG(<integrand>,<variable of integration>,<lower limit>, <upper limit>);

Examples:
        ROMBERG(SIN(Y),Y,1,%PI);
                TIME= 39 MSEC.          1.5403023
        F(X):=1/(X^5+X+1);
        ROMBERG(F(X),X,1.5,0);
                TIME= 162 MSEC.         - 0.75293843

The second is an efficient way that is used as follows:

ROMBERG(<function name>,<lower limit>,<upper limit>);
Example:
F(X):=(MODE_DECLARE([FUNCTION(F),X],FLOAT),1/(X^5+X+1));
TRANSLATE(F);

     ROMBERG(F,1.5,0);
             TIME= 13 MSEC.          - 0.75293843

The first argument must be a TRANSLATEd or compiled function. (If it is compiled it must be declared to return a FLONUM.) If the first argument is not already TRANSLATEd, ROMBERG will not attempt to TRANSLATE it but will give an error. The accuracy of the integration is governed by the global variables ROMBERGTOL (default value 1.E-4) and ROMBERGIT (default value 11). ROMBERG will return a result if the relative difference in successive approximations is less than ROMBERGTOL. It will try halving the stepsize ROMBERGIT times before it gives up. The number of iterations and function evaluations which ROMBERG will do is governed by ROMBERGABS and ROMBERGMIN, do DE-SCRIBE(ROMBERGABS,ROMBERGMIN); for details. ROMBERG may be called recursively and thus can do double and triple integrals.

     Example:
     INTEGRATE(INTEGRATE(X*Y/(X+Y),Y,0,X/2),X,1,3);
                             13/3 (2 LOG(2/3) + 1)
     %,NUMER;
                             0.81930233
     DEFINE_VARIABLE(X,0.0,FLOAT,"Global variable in function F")$
     F(Y):=(MODE_DECLARE(Y,FLOAT), X*Y/(X+Y) )$
     G(X):=ROMBERG('F,0,X/2)$
     ROMBERG(G,1,3);
                              0.8193023

The advantage with this way is that the function F can be used for other purposes, like plotting. The disadvantage is that you have to think up a name for both the function F and its free variable X. Or, without the global:

             G1(X):=(MODE_DECLARE(X,FLOAT), ROMBERG(X*Y/(X+Y),Y,0,X/2))$
             ROMBERG(G1,1,3);
                                     0.8193023

The advantage here is shortness.

             Q(A,B):=ROMBERG(ROMBERG(X*Y/(X+Y),Y,0,X/2),X,A,B)$
             Q(1,3);
                                     0.8193023

It is even shorter this way, and the variables do not need to be declared because they are in the context of ROMBERG. Use of ROMBERG for multiple integrals can have great disadvantages, though. The amount of extra calculation needed because of the geometric information thrown away by expressing multiple integrals this way can be incredible. The user should be sure to understand and use the ROMBERGTOL and ROMBERGIT switches. (The IMSL version of Romberg integration is now available in Macsyma. Do DESCRIBE(DCADRE); for more information.)

Variable ROMBERGABS

default: [0.0] (0.0B0) Assuming that successive estimates produced by ROMBERG are Y[0], Y[1], Y[2] etc., then ROMBERG will return after N iterations if (roughly speaking) (ABS(Y[N]-Y[N-1]) <= ROMBERGABS OR ABS(Y[N]-Y[N-1])/(IF Y[N]=0.0 THEN 1.0 ELSE Y[N]) <= ROMBERGTOL) is TRUE. (The condition on the number of iterations given by ROMBERGMIN must also be satisfied.) Thus if ROMBER-GABS is 0.0 (the default) you just get the relative error test. The usefulness of the additional variable comes when you want to perform an integral, where the dominant contribution comes from a small region. Then you can do the integral over the small dominant region first, using the relative accuracy check, followed by the integral over the rest of the region using the absolute accuracy check. Example: Suppose you want to compute

        Integral(exp(-x),x,0,50)

(numerically) with a relative accuracy of 1 part in 10000000. Define the function. N is a counter, so we can see how many function evaluations were needed.

     F(X):=(MODE_DECLARE(N,INTEGER,X,FLOAT),N:N+1,EXP(-X))$
     TRANSLATE(F)$
       /* First of all try doing the whole integral at once */
     BLOCK([ROMBERGTOL:1.E-6,ROMBERABS:0.],N:0,ROMBERG(F,0,50));
                   ==> 1.00000003
     N; ==> 257  /* Number of function evaluations*/

Now do the integral intelligently, by first doing Integral(exp(-x),x,0,10) and then setting ROMBERGABS to 1.E-6*(this partial integral).

     BLOCK([ROMBERGTOL:1.E-6,ROMBERGABS:0.,SUM:0.],
       N:0,SUM:ROMBERG(F,0,10),ROMBERGABS:SUM*ROMBERGTOL,ROMBERGTOL:0.,
           SUM+ROMBERG(F,10,50));  ==> 1.00000001  /* Same as before */
     N;  ==> 130

So if F(X) were a function that took a long time to compute, the second method would be about 2 times quicker.

Variable ROMBERGIT

default: [11] - The accuracy of the ROMBERG integration command is governed by the global variables ROMBERGTOL[1.E-4] and ROMBERGIT[11]. ROMBERG will return a result if the relative difference in successive approximations is less than ROMBERGTOL. It will try halving the stepsize ROMBERGIT times before it gives up.

Variable ROMBERGMIN

default: [0] - governs the minimum number of function evaluations that ROMBERG will make. ROMBERG will evaluate its first arg. at least 2^(ROMBERGMIN+2)+1 times. This is useful for integrating oscillatory functions, when the normal converge test might sometimes wrongly pass.

Variable ROMBERGTOL

default: [1.E-4] - The accuracy of the ROMBERG integration command is governed by the global variables ROMBERGTOL[1.E-4] and ROMBERGIT[11]. ROMBERG will return a result if the relative difference in successive approximations is less than ROMBERGTOL. It will try halving the stepsize ROMBERGIT times before it gives up.

Function TLDEFINT (exp,var,ll,ul)

is just LDEFINT with TLIMSWITCH set to TRUE.

22 Maxima Manual

4 Polynomials

4.1 Introduction to Polynomials

Polynomials are stored in maxima either in General Form or as Cannonical Rational Expressions (CRE) form. The latter is a standard form, and is used internally by operations such as factor, ratsimp, and so on.

Canonical Rational Expressions constitute a kind of representation which is especially suitable for expanded polynomials and rational functions (as well as for partially factored polynomials and rational functions when RATFAC[FALSE] is set to TRUE). In this CRE form an ordering of variables (from most to least main) is assumed for each expression. Polynomials are represented recursively by a list consisting of the main variable followed by a series of pairs of expressions, one for each term of the polynomial. The first member of each pair is the exponent of the main variable in that term and the second member is the coefficient of that term which could be a number or a polynomial in another variable again represented in this form. Thus the principal part of the CRE form of 3*X^2-1 is (X 2 3 0 -1) and that of 2*X*Y+X-3 is (Y 1 (X 1 2) 0 (X 1 1 0 -3)) assuming Y is the main variable, and is (X 1 (Y 1 2 0 1) 0 -3) assuming X is the main variable. "Main"-ness is usually determined by reverse alphabetical order. The "variables" of a CRE expression needn’t be atomic. In fact any subexpression whose main operator is not + - * / or ^with integer power will be considered a "variable" of the expression (in CRE form) in which it occurs. For example the CRE variables of the expression X+SIN(X+1)+2*SQRT(X)+1 are X, SQRT(X), and SIN(X+1). If the user does not specify an ordering of variables by using the RATVARS function MACSYMA will choose an alphabetic one. In general, CRE’s represent rational expressions, that is, ratios of polynomials, where the numerator and denominator have no common factors, and the denominator is positive. The internal form is essentially a pair of polynomials (the numerator and denominator) preceded by the variable ordering list. If an expression to be displayed is in CRE form or if it contains any subexpressions in CRE form, the symbol /R/ will follow the line label. See the RAT function for converting an expression to CRE form. An extended CRE form is used for the representation of Taylor series. The notion of a rational expression is extended so that the exponents of the variables can be positive or negative rational numbers rather than just positive integers and the coefficients can themselves be rational expressions as described above rather than just polynomials. These are represented internally by a recursive polynomial form which is similar to and is a generalization of CRE form, but carries additional information such as the degree of truncation. As with CRE form, the symbol /T/ follows the line label of such expressions.

4.2 Definitions for Polynomials

Variable ALGEBRAIC

default: [FALSE] must be set to TRUE in order for the simplification of algebraic integers to take effect.

Variable BERLEFACT

default: [TRUE] if FALSE then the Kronecker factoring algorithm will be used otherwise the Berlekamp algorithm, which is the default, will be used.

Function BEZOUT (p1, p2, var)

an alternative to the RESULTANT command. It returns a matrix. DETERMINANT of this matrix is the desired resultant.

Function BOTHCOEF (exp, var)

returns a list whose first member is the coefficient of var in exp (as found by RAT-COEF if exp is in CRE form otherwise by COEFF) and whose second member is the remaining part of exp. That is, [A,B] where exp=A*var+B.

     (C1) ISLINEAR(EXP,VAR):=BLOCK([C],
             C:BOTHCOEF(RAT(EXP,VAR),VAR),
             IS(FREEOF(VAR,C) AND C[1]#0))$
     (C2) ISLINEAR((R**2-(X-R)**2)/X,X);
     (D2)                              TRUE

Function COEFF (exp, v, n)

obtains the coefficient of v**n in exp. n may be omitted if it is 1. v may be an atom, or complete subexpression of exp e.g., X, SIN(X), A[I+1], X+Y, etc. (In the last case the expression (X+Y) should occur in exp). Sometimes it may be necessary to expand or factor exp in order to make v^n explicit. This is not done automatically by COEFF.

     (C1) COEFF(2*A*TAN(X)+TAN(X)+B=5*TAN(X)+3,TAN(X));
     (D1)                         2 A + 1 = 5
     (C2) COEFF(Y+X*%E**X+1,X,0);
     (D2)                            Y + 1

Function COMBINE (exp)

simplifies the sum exp by combining terms with the same denominator into a single term.

Function CONTENT (p1, var1, ..., varn)

returns a list whose first element is the greatest common divisor of the coefficients of the terms of the polynomial p1 in the variable varn (this is the content) and whose second element is the polynomial p1 divided by the content.

     (C1) CONTENT(2*X*Y+4*X**2*Y**2,Y);
     (D1)            [2*X, 2*X*Y**2+Y].

Function DENOM (exp)

returns the denominator of the rational expression exp.

Function DIVIDE (p1, p2, var1, ..., varn)

computes the quotient and remainder of the polynomial p1 divided by the polynomial p2, in a main polynomial variable, varn. The other variables are as in the RATVARS function. The result is a list whose first element is the quotient and whose second element is the remainder.

     (C1) DIVIDE(X+Y,X-Y,X);
     (D1)                        [1, 2 Y]
     (C2) DIVIDE(X+Y,X-Y);
     (D2)                      [ - 1, 2 X]

(Note that Y is the main variable in C2)

Function ELIMINATE ([eq1,eq2,...,eqn],[v1,v2,...,vk])

eliminates variables from equations (or expressions assumed equal to zero) by taking successive resultants. This returns a list of n-k expressions with the k variables v1,...,vk eliminated. First v1 is eliminated yielding n-1 expressions, then v2 is, etc. If k=n then a single expression in a list is returned free of the variables v1,...,vk. In this case SOLVE is called to solve the last resultant for the last variable. Example:

     (C1) EXP1:2*X^2+Y*X+Z;
                                         2
     (D1)                   Z + X Y + 2 X
     (C2) EXP2:3*X+5*Y-Z-1;
     (D2)                - Z + 5 Y + 3 X - 1
     (C3) EXP3:Z^2+X-Y^2+5;
                            2    2
     (D3)                  Z  - Y  + X + 5
     (C4) ELIMINATE([EXP3,EXP2,EXP1],[Y,Z]);
                 8         7         6          5          4
     (D3) [7425 X  - 1170 X  + 1299 X  + 12076 X  + 22887 X
                                    3         2
                            - 5154 X  - 1291 X  + 7688 X + 15376]

Function EZGCD (p1, p2, ...)

gives a list whose first element is the g.c.d of the polynomials p1,p2,... and whose remaining elements are the polynomials divided by the g.c.d. This always uses the EZGCD algorithm.

Variable FACEXPAND

default: [TRUE] controls whether the irreducible factors returned by FACTOR are in expanded (the default) or recursive (normal CRE) form.

Function FACTCOMB (exp)

tries to combine the coefficients of factorials in exp with the factorials themselves by converting, for example, (N+1)*N! into (N+1)!. SUMSPLITFACT[TRUE] if set to FALSE will cause MINFACTORIAL to be applied after a FACTCOMB.

     (C1) (N+1)^B*N!^B;
                                           B   B
     (D1)                           (N + 1)  N!
     (C2) FACTCOMB(%);

Function FACTOR (exp)

factors the expression exp, containing any number of variables or functions, into factors irreducible over the integers. FACTOR(exp, p) factors exp over the field of integers with an element adjoined whose minimum polynomial is p. FACTOR-FLAG[FALSE] if FALSE suppresses the factoring of integer factors of rational expressions. DONTFACTOR may be set to a list of variables with respect to which factoring is not to occur. (It is initially empty). Factoring also will not take place with respect to any variables which are less important (using the variable ordering assumed for CRE form) than those on the DONTFACTOR list. SAVEFACTORS[FALSE] if TRUE causes the factors of an expression which is a product of factors to be saved by certain functions in order to speed up later factorizations of expressions containing some of the same factors. BERLEFACT[TRUE] if FALSE then the Kronecker factoring algorithm will be used otherwise the Berlekamp algorithm, which is the default, will be used. INTFACLIM[1000] is the largest divisor which will be tried when factoring a bignum integer. If set to FALSE (this is the case when the user calls FACTOR explicitly), or if the integer is a fixnum (i.e. fits in one machine word), complete factorization of the integer will be attempted. The user’s setting of INTFACLIM is used for internal calls to FACTOR. Thus, INTFACLIM may be reset to prevent MACSYMA from taking an inordinately long time factoring large integers. NEWFAC[FALSE] may be set to true to use the new factoring routines. Do EXAMPLE(FACTOR); for examples.

Variable FACTORFLAG

default: [FALSE] if FALSE suppresses the factoring of integer factors of rational expressions.

Function FACTOROUT (exp,var1,var2,...)

rearranges the sum exp into a sum of terms of the form f(var1,var2,...)*g where g is a product of expressions not containing the vari’s and f is factored.

Function FACTORSUM (exp)

tries to group terms in factors of exp which are sums into groups of terms such that their sum is factorable. It can recover the result of EXPAND((X+Y)^2+(Z+W)^2) but it can’t recover EXPAND((X+1)^2+(X+Y)^2) because the terms have variables in common.

     (C1) (X+1)*((U+V)^2+A*(W+Z)^2),EXPAND;
           2      2                            2      2
     (D1) A X Z  + A Z  + 2 A W X Z + 2 A W Z + A W  X + V  X
                          2        2    2            2

             + 2 U V X + U  X + A W  + V  + 2 U V + U
     (C2) FACTORSUM(%);
                                        2          2
     (D2)                 (X + 1) (A (Z + W)  + (V + U) )

Function FASTTIMES (p1, p2)

multiplies the polynomials p1 and p2 by using a special algorithm for multiplication of polynomials. They should be multivariate, dense, and nearly the same size. Classical multiplication is of order N*M where N and M are the degrees. FASTTIMES is of order MAX(N,M)**1.585.

Function FULLRATSIMP (exp)

When non-rational expressions are involved, one call to RATSIMP followed as is usual by non-rational ("general") simplification may not be sufficient to return a simplified result. Sometimes, more than one such call may be necessary. The command FULL-RATSIMP makes this process convenient. FULLRATSIMP repeatedly applies RAT-SIMP followed by non-rational simplification to an expression until no further change occurs. For example, consider For the expression EXP: (X^(A/2)+1)^2*(X^(A/2)- 1)^2/(X^A-1) . RATSIMP(EXP); gives (X^(2*A)-2*X^A+1)/(X^A-1) . FULLRAT-SIMP(EXP); gives X^A-1 . The problem may be seen by looking at RAT(EXP); which gives ((X^(A/2))^4-2*(X^(A/2))^2+1)/(X^A-1) . FULLRATSIMP(exp,var1,...,varn) takes one or more arguments similar to RATSIMP and RAT.

Function FULLRATSUBST (a,b,c)

is the same as RATSUBST except that it calls itself recursively on its result until that result stops changing. This function is useful when the replacement expression and the replaced expression have one or more variables in common. FULLRATSUBST will also accept its arguments in the format of LRATSUBST. That is, the first argument may be a single substitution equation or a list of such equations, while the second argument is the expression being processed. There is a demo available by DEMO("lrats.dem"); .

Function GCD (p1, p2, var1, ...)

computes the greatest common divisor of p1 and p2. The flag GCD[SPMOD] determines which algorithm is employed. Setting GCD to EZ, EEZ, SUBRES, RED, or SPMOD selects the EZGCD, New EEZ GCD, subresultant PRS, reduced, or modular algorithm, respectively. If GCD:FALSE then GCD(p1,p2,var) will always return 1 for all var. Many functions (e.g. RATSIMP, FACTOR, etc.) cause gcd’s to be taken implicitly. For homogeneous polynomials it is recommended that GCD:SUBRES be used. To take the gcd when an algebraic is present, e.g. GCD(X^2-2*SQRT(2)*X+2,X-SQRT(2)); , ALGEBRAIC must be TRUE and GCD must not be EZ. SUBRES is a new algorithm, and people who have been using the RED setting should probably change it to SUBRES. The GCD flag, default: [SPMOD], if FALSE will also prevent the greatest common divisor from being taken when expressions are converted to CRE form. This will sometimes speed the calculation if gcds are not required.

Function GCFACTOR (n)

factors the gaussian integer n over the gaussians, i.e. numbers of the form a + b i where a and b are rational integers (i.e. ordinary integers). Factors are normalized by making a and b non-negative.

Function GFACTOR (exp)

factors the polynomial exp over the Gaussian integers (i. e. with SQRT(-1) = %I adjoined). This is like FACTOR(exp,A**2+1) where A is %I.

     (C1)  GFACTOR(X**4-1);
     (D1)        (X - 1) (X + 1) (X + %I) (X - %I)

Function GFACTORSUM (exp)

is similar to FACTORSUM but applies GFACTOR instead of FACTOR.

Function HIPOW (exp, v)

the highest explicit exponent of v in exp. Sometimes it may be necessary to expand exp since this is not done automatically by HIPOW. Thus HIPOW(Y**3*X**2+X*Y**4,X) is 2.

Variable INTFACLIM

default: [1000] is the largest divisor which will be tried when factoring a bignum integer. If set to FALSE (this is the case when the user calls FACTOR explicitly), or if the integer is a fixnum (i.e. fits in one machine word), complete factorization of the integer will be attempted. The user’s setting of INTFACLIM is used for internal calls to FACTOR. Thus, INTFACLIM may be reset to prevent MACSYMA from taking an inordinately long time factoring large integers.

Variable KEEPFLOAT

default: [FALSE] - if set to TRUE will prevent floating point numbers from being rationalized when expressions which contain them are converted to CRE form.

Function LRATSUBST (list,exp)

is analogous to SUBST(list of equations,exp) except that it uses RATSUBST instead of SUBST. The first argument of LRATSUBST must be an equation or a list of equations identical in format to that accepted by SUBST (see DESCRIBE(SUBST);). The substitutions are made in the order given by the list of equations, that is, from left to right. A demo is available by doing DEMO("lrats.dem"); .

Variable MODULUS

default: [FALSE] - if set to a positive prime p, then all arithmetic in the rational function routines will be done modulo p. That is all integers will be reduced to less than p/2 in absolute value (if p=2 then all integers are reduced to 1 or 0). This is the so called "balanced" modulus system, e.g. N MOD 5 = -2, -1, 0, 1, or 2. Warning: If EXP is already in CRE form when you reset MODULUS, then you may need to re-rat EXP, e.g. EXP:RAT(RATDISREP(EXP)), in order to get correct results. (If MODULUS is set to a positive non-prime integer, this setting will be accepted, but a warning will be given.)

Variable NEWFAC

default: [FALSE], if TRUE then FACTOR will use the new factoring routines.

Function NUM (exp)

obtains the numerator, exp1, of the rational expression exp = exp1/exp2.

Function QUOTIENT (p1, p2, var1, ...)

computes the quotient of the polynomial p1 divided by the polynomial p2.

Function RAT (exp, v1, ..., vn)

converts exp to CRE form by expanding and combining all terms over a common denominator and cancelling out the greatest common divisor of the numerator and denominator as well as converting floating point numbers to rational numbers within a tolerance of RATEPSILON[2.0E-8]. The variables are ordered according to the v1,...,vn as in RATVARS, if these are specified. RAT does not generally simplify functions other than + , - , * , / , and exponentiation to an integer power whereas RATSIMP does handle these cases. Note that atoms (numbers and names) in CRE form are not the same as they are in the general form. Thus RAT(X)- X results in RAT(0) which has a different internal representation than 0. RATFAC[FALSE] when TRUE invokes a partially factored form for CRE rational expressions. During rational operations the expression is maintained as fully factored as possible without an actual call to the factor package. This should always save space and may save some time in some computations. The numerator and denominator are still made relatively prime (e.g. RAT((X^2 -1)^4/(X+1)^2); yields (X-1)^4*(X+1)^2), but the factors within each part may not be relatively prime. RATPRINT[TRUE] if FALSE suppresses the printout of the message informing the user of the conversion of floating point numbers to rational numbers. KEEPFLOAT[FALSE] if TRUE prevents floating point numbers from being converted to rational numbers. (Also see the RATEXPAND and RATSIMP functions.)

     (C1) ((X-2*Y)**4/(X**2-4*Y**2)**2+1)*(Y+A)*(2*Y+X)
             /(4*Y**2+X**2);
                                                4
                                       (X - 2 Y)
                   (Y + A) (2 Y + X) (------------ + 1)
                                        2      2 2
                                      (X  - 4 Y )
     (D1)          ------------------------------------
                                   2    2
                                4 Y  + X
     (C2) RAT(%,Y,A,X);
                                 2 A + 2 Y
     (D2)/R/                     ---------
                                  X + 2 Y

Variable RATALGDENOM

default: [TRUE] - if TRUE allows rationalization of denominators wrt. radicals to take effect. To do this one must use CRE form in algebraic mode.

Function RATCOEF (exp, v, n)

returns the coefficient, C, of the expression v**n in the expression exp. n may be omitted if it is 1. C will be free (except possibly in a non-rational sense) of the variables in v. If no coefficient of this type exists, zero will be returned. RATCOEF expands and rationally simplifies its first argument and thus it may produce answers different from those of COEFF which is purely syntactic. Thus RATCOEF((X+1)/Y+X,X) returns (Y+1)/Y whereas COEFF returns 1. RATCOEF(exp,v,0), viewing exp as a sum, gives a sum of those terms which do not contain v. Therefore if v occurs to any negative powers, RATCOEF should not be used. Since exp is rationally simplified before it is examined, coefficients may not appear quite the way they were envisioned.

     (C1) S:A*X+B*X+5$
     (C2) RATCOEF(S,A+B);
     (D2)               X

Function RATDENOM (exp)

obtains the denominator of the rational expression exp. If exp is in general form then the DENOM function should be used instead, unless one wishes to get a CRE result.

Variable RATDENOMDIVIDE

default: [TRUE] - if FALSE will stop the splitting up of the terms of the numerator of RATEXPANDed expressions from occurring.

Function RATDIFF (exp, var)

differentiates the rational expression exp (which must be a ratio of polynomials or a polynomial in the variable var) with respect to var. For rational expressions this is much faster than DIFF. The result is left in CRE form. However, RATDIFF should not be used on factored CRE forms; use DIFF instead for such expressions.

     (C1) (4*X**3+10*X-11)/(X**5+5);
                                              3
                                           4 X  + 10 X - 11
     (D1)                                  ----------------
                                                 5
                                                X
     (C2) MODULUS:3$
     (C3) MOD(D1);
                                     2
                                    X  + X - 1
     (D3)                      --------------------

                                4    3    2
                               X  + X  + X  + X + 1
     (C4) RATDIFF(D1,X);
                               5    4    3
                              X  - X  - X  + X - 1
     (D4)                ------------------------------
                          8    7    5    4    3
                         X  - X  + X  - X  + X  - X + 1

Function RATDISREP (exp)

changes its argument from CRE form to general form. This is sometimes convenient if one wishes to stop the "contagion", or use rational functions in non-rational contexts. Most CRE functions will work on either CRE or non-CRE expressions, but the answers may take different forms. If RATDISREP is given a non-CRE for an argument, it returns its argument unchanged. See also TOTALDISREP.

Variable RATEPSILON

default: [2.0E-8] - the tolerance used in the conversion of floating point numbers to rational numbers.

Function RATEXPAND (exp)

expands exp by multiplying out products of sums and exponentiated sums, combining fractions over a common denominator, cancelling the greatest common divisor of the numerator and denominator, then splitting the numerator (if a sum) into its respective terms divided by the denominator. This is accomplished by converting exp to CRE form and then back to general form. The switch RATEXPAND, default: [FALSE], if TRUE will cause CRE expressions to be fully expanded when they are converted back to general form or displayed, while if it is FALSE then they will be put into a recursive form. (see RATSIMP) RATDENOMDIVIDE[TRUE] - if FALSE will stop the splitting up of the terms of the numerator of RATEXPANDed expressions from occurring. KEEPFLOAT[FALSE] if set to TRUE will prevent floating point numbers from being rationalized when expressions which contain them are converted to CRE form.

     (C1) RATEXPAND((2*X-3*Y)**3);
                     3         2       2        3
     (D1)      - 27 Y  + 54 X Y  - 36 X  Y + 8 X
     (C2) (X-1)/(X+1)**2+1/(X-1);
                                   X - 1       1
     (D2)                         -------- + -----
                                         2   X - 1
                                  (X + 1)
     (C3) EXPAND(D2);
                              X              1           1
     (D3)                ------------ - ------------ + -----
                          2              2             X - 1

                         X  + 2 X + 1   X  + 2 X + 1
     (C4) RATEXPAND(D2);
                                  2
                               2 X                 2
     (D4)                 --------------- + ---------------
                           3    2            3    2
                          X  + X  - X - 1   X  + X  - X - 1

Variable RATFAC

default: [FALSE] - when TRUE invokes a partially factored form for CRE rational expressions. During rational operations the expression is maintained as fully factored as possible without an actual call to the factor package. This should always save space and may save some time in some computations. The numerator and denominator are still made relatively prime, for example RAT((X^2 -1)^4/(X+1)^2); yields (X-1)^4*(X+1)^2), but the factors within each part may not be relatively prime. In the CTENSR (Component Tensor Manipulation) Package, if RATFAC is TRUE, it causes the Ricci, Einstein, Riemann, and Weyl tensors and the Scalar Curvature to be factored automatically. ** This should only be set for cases where the tensorial components are known to consist of few terms **. Note: The RATFAC and RATWEIGHT schemes are incompatible and may not both be used at the same time.

Function RATNUMER (exp)

obtains the numerator of the rational expression exp. If exp is in general form then the NUM function should be used instead, unless one wishes to get a CRE result.

Function RATNUMP (exp)

is TRUE if exp is a rational number (includes integers) else FALSE.

Function RATP (exp)

is TRUE if exp is in CRE or extended CRE form else FALSE.

Variable RATPRINT

default: [TRUE] - if FALSE suppresses the printout of the message informing the user of the conversion of floating point numbers to rational numbers.

Function RATSIMP (exp)

rationally" simplifies (similar to RATEXPAND) the expression exp and all of its subexpressions including the arguments to non- rational functions. The result is returned as the quotient of two polynomials in a recursive form, i.e. the coefficients of the main variable are polynomials in the other variables. Variables may, as in RATEXPAND, include non-rational functions (e.g. SIN(X**2+1) ) but with RAT-SIMP, the arguments to non-rational functions are rationally simplified. Note that RATSIMP is affected by some of the variables which affect RATEXPAND. RAT-SIMP(exp,v1,v2,...,vn) - enables rational simplification with the specification of variable ordering as in RATVARS. RATSIMPEXPONS[FALSE] - if TRUE will cause exponents of expressions to be RATSIMPed automatically during simplification.

     (C1) SIN(X/(X^2+X))=%E^((LOG(X)+1)**2-LOG(X)**2);
                                                2          2
                         X          (LOG(X) + 1)  - LOG (X)
     (D1)          SIN(------) = %E
                        2
                       X  + X
     (C2) RATSIMP(%);
                                       1          2
     (D2)                        SIN(-----) = %E X
                                     X + 1
     (C3) ((X-1)**(3/2)-(X+1)*SQRT(X-1))/SQRT((X-1)*(X+1));
                            3/2
                     (X - 1)    - SQRT(X - 1) (X + 1)
     (D3)            --------------------------------
                         SQRT(X - 1) SQRT(X + 1)
     (C4) RATSIMP(%);
                                      2
     (D4)                      - -----------
                                 SQRT(X + 1)
     (C5)  X**(A+1/A),RATSIMPEXPONS:TRUE;
                         2
                        A  + 1
                        ------
                          A
     (D5)              X

Variable RATSIMPEXPONS

default: [FALSE] - if TRUE will cause exponents of expressions to be RATSIMPed automatically during simplification.

Function RATSUBST (a, b, c)

substitutes a for b in c. b may be a sum, product, power, etc. RATSUBST knows something of the meaning of expressions whereas SUBST does a purely syntactic substitution. Thus SUBST(A,X+Y,X+Y+Z) returns X+Y+Z whereas RATSUBST would return Z+A. RADSUBSTFLAG[FALSE] if TRUE permits RATSUBST to make substitutions like U for SQRT(X) in X. Do EXAMPLE(RATSUBST); for examples.

Function RATVARS (var1, var2, ..., varn)

forms its n arguments into a list in which the rightmost variable varn will be the main variable of future rational expressions in which it occurs, and the other variables will follow in sequence. If a variable is missing from the RATVARS list, it will be given lower priority than the leftmost variable var1. The arguments to RATVARS can be either variables or non-rational functions (e.g. SIN(X)). The variable RATVARS is a list of the arguments which have been given to this function.

Function RATWEIGHT (v1, w1, ..., vn, wn)

assigns a weight of wi to the variable vi. This causes a term to be replaced by 0 if its weight exceeds the value of the variable RATWTLVL [default is FALSE which means no truncation]. The weight of a term is the sum of the products of the weight of a variable in the term times its power. Thus the weight of 3*v1**2*v2 is 2*w1+w2. This truncation occurs only when multiplying or exponentiating CRE forms of expressions.

     (C5) RATWEIGHT(A,1,B,1);
     (D5)                         [[B, 1], [A, 1]]
     (C6) EXP1:RAT(A+B+1)$
     (C7) %**2;
                           2                  2
     (D7)/R/              B  + (2 A + 2) B + A  + 2 A + 1
     (C8) RATWTLVL:1$
     (C9) EXP1**2;
     (D9)/R/                       2 B + 2 A + 1

Note: The RATFAC and RATWEIGHT schemes are incompatible and may not both be used at the same time.

Variable RATWEIGHTS

- a list of weight assignments (set up by the RATWEIGHT function), RATWEIGHTS; or RATWEIGHT(); will show you the list.

     KILL(...,RATWEIGHTS)

and

     SAVE(...,RATWEIGHTS);

both work.

Variable RATWEYL

default: [] - one of the switches controlling the simplification of components of the Weyl conformal tensor; if TRUE, then the components will be rationally simplified; if FACRAT is TRUE then the results will be factored as well.

Variable RATWTLVL

default: [FALSE] - used in combination with the RATWEIGHT function to control the truncation of rational (CRE form) expressions (for the default value of FALSE, no truncation occurs).

Function REMAINDER (p1, p2, var1, ...)

computes the remainder of the polynomial p1 divided by the polynomial p2.

Function RESULTANT (p1, p2, var)

computes the resultant of the two polynomials p1 and p2, eliminating the variable var. The resultant is a determinant of the coefficients of var in p1 and p2 which equals zero if and only if p1 and p2 have a non-constant factor in common. If p1 or p2 can be factored, it may be desirable to call FACTOR before calling RESULTANT. RESUL-TANT[SUBRES] - controls which algorithm will be used to compute the resultant.

SUBRES for subresultant prs [the default], MOD for modular resultant algorithm, and RED for reduced prs. On most problems SUBRES should be best. On some large degree univariate or bivariate problems MOD may be better. Another alternative is the BEZOUT command which takes the same arguments as RESULTANT and returns a matrix. DETERMINANT of this matrix is the desired resultant.

Variable SAVEFACTORS

default: [FALSE] - if TRUE causes the factors of an expression which is a product of factors to be saved by certain functions in order to speed up later factorizations of expressions containing some of the same factors.

Function SQFR (exp)

is similar to FACTOR except that the polynomial factors are "square-free." That is, they have factors only of degree one. This algorithm, which is also used by the first stage of FACTOR, utilizes the fact that a polynomial has in common with its nth derivative all its factors of degree > n. Thus by taking gcds with the polynomial of the derivatives with respect to each variable in the polynomial, all factors of degree > 1 can be found.

     (C1) SQFR(4*X**4+4*X**3-3*X**2-4*X-1);
                                  2               2
     (D1)                       (X  - 1) (2 X + 1)

Function TELLRAT (poly)

adds to the ring of algebraic integers known to MACSYMA, the element which is the solution of the polynomial with integer coefficients. MACSYMA initially knows about %I and all roots of integers. TELLRAT(X); means substitute 0 for X in rational functions. There is a command UNTELLRAT which takes kernels and removes TELLRAT properties. When TELLRATing a multivariate polynomial, e.g. TELLRAT(X^2-Y^2);, there would be an ambiguity as to whether to substitute Y^2 for X^2 or vice versa. The system will pick a particular ordering, but if the user wants to specify which, e.g. TELLRAT(Y^2=X^2); provides a syntax which says replace Y^2 by X^2. TELLRAT and UNTELLRAT both can take any number of arguments, and TELLRAT(); returns a list of the current substitutions. Note: When you TELL-RAT reducible polynomials, you want to be careful not to attempt to rationalize a denominator with a zero divisor. E.g. TELLRAT(W^3-1)$ ALGEBRAIC:TRUE$ RAT(1/(W^2-W)); will give "quotient by zero". This error can be avoided by setting RATALGDENOM:FALSE$. ALGEBRAIC[FALSE] must be set to TRUE in order for the simplification of algebraic integers to take effect. Do EXAMPLE(TELLRAT); for examples.

Function TOTALDISREP (exp)

converts every subexpression of exp from CRE to general form. If exp is itself in CRE form then this is identical to RATDISREP but if not then RATDISREP would return exp unchanged while TOTALDISREP would "totally disrep" it. This is useful for ratdisrepping expressions e.g., equations, lists, matrices, etc. which have some subexpressions in CRE form.

36 Maxima Manual

Function UNTELLRAT (x)

takes kernels and removes TELLRAT properties.

5 Simplification

5.1 Definitions for Simplification

Function APPLY NOUNS (exp)

causes the application of noun forms in an expression. E.g. EXP:’DIFF(X^2/2,X); APPLY NOUNS(EXP); gives X. This gives the same result as EV(EXP,NOUNS); except that it can be faster and use less storage. It also can be used in translated code, where EV may cause problems. Note that it is called APPLY NOUNS, not EV NOUNS, because what it does is to APPLY the rules corresponding to the nounform operators, which is not evaluation.

Variable ASKEXP

default: [] contains the expression upon which ASKSIGN is called. A user may enter a MACSYMA break with ^A and inspect this expression in order to answer questions asked by ASKSIGN.

Function ASKINTEGER (exp,<optional-arg>)

exp is any valid macsyma expression and optional-arg is EVEN,ODD,INTEGER and defaults to INTEGER if not given. This function attempts to determine from the data-base whether exp is EVEN, ODD or just an INTEGER. It will ask the user if it cannot tell otherwise and attempt to install the information in the data-base if possible.

Function ASKSIGN (exp)

first attempts to determine whether the specified expression is positive, negative, or zero. If it cannot, it asks the user the necessary questions to complete its deduction. The user’s answer is recorded in the data base for the duration of the current computation (one "C-line"). The value of ASKSIGN is one of POS, NEG, or ZERO.

Variable DEMOIVRE

default: [FALSE] if TRUE will cause

     %E^(A+B*%I) ==> %E^A*(COS(B)+%I*SIN(B))

if B is free of %I. A and B are not expanded. DEMOIVRE:TRUE; is the way to reverse the effect of EXPONENTIALIZE:TRUE; DEMOIVRE(exp) will cause the conversion without setting the switch or having to re-evaluate the expression with EV.

Variable DOMAIN

default: [REAL] - if set to COMPLEX, SQRT(X^2) will remain SQRT(X^2) instead of returning ABS(X). The notion of a "domain" of simplification is still in its infancy, and controls little more than this at the moment.

Function EXPAND (exp)

will cause products of sums and exponentiated sums to be multiplied out, numerators of rational expressions which are sums to be split into their respective terms, and multiplication (commutative and non-commutative) to be distributed over addition at all levels of exp. For polynomials one should usually use RATEXPAND which uses a more efficient algorithm (see DESCRIBE(RATEXPAND);). MAXNEGEX[1000] and MAXPOSEX[1000] control the maximum negative and positive exponents, respectively, which will expand. EXPAND(exp,p,n) expands exp, using p for MAXPOSEX and n for MAXNEGEX. This is useful in order to expand part but not all of an expression. EXPON[0] - the exponent of the largest negative power which is automatically expanded (independent of calls to EXPAND). For example if EXPON is 4 then (X+1)**(-5) will not be automatically expanded. EXPOP[0] - the highest positive exponent which is automatically expanded. Thus (X+1)**3, when typed, will be automatically expanded only if EXPOP is greater than or equal to 3. If it is desired to have (X+1)**N expanded where N is greater than EXPOP then executing EX-PAND((X+1)**N) will work only if MAXPOSEX is not less than N. The EXPAND flag used with EV (see EV) causes expansion. The file SHARE1;FACEXP FASL contains several related functions (FACSUM and COLLECTTERMS are two) that provide the user with the ability to structure expressions by controlled expansion. Brief function descriptions are available in SHARE1;FACEXP USAGE. A demo is available by doing BATCH("facexp.mc")$ .

Function EXPANDWRT (exp,var1,var2,...)

expands exp with respect to the vari. All products involving the vari appear explicitly. The form returned will be free of products of sums of expressions that are not free of the vari. The vari may be variables, operators, or expressions. By default, denominators are not expanded, but this can be controlled by means of the switch EXPANDWRT DENOM. Do LOAD(STOPEX); to use this function.

Variable EXPANDWRT DENOM

default:[FALSE] controls the treatment of rational expressions by EXPANDWRT. If TRUE, then both the numerator and denominator of the expression will be expanded according to the arguments of EXPANDWRT, but if EXPANDWRT DENOM is FALSE, then only the numerator will be expanded in that way. Do LOAD(STOPEX) to use.

Function EXPANDWRT FACTORED (exp, var1, var2, ..., varN)

is similar to EXPANDWRT, but treats expressions that are products somewhat differently. EXPANDWRT FACTORED will perform the required expansion only on those factors of exp that contain the variables in its argument list argument list. Do LOAD(STOPEX) to use this function.

Variable EXPON

default: [0] - the exponent of the largest negative power which is automatically expanded (independent of calls to EXPAND). For example if EXPON is 4 then (X+1)**(-5) will not be automatically expanded.

Variable EXPONENTIALIZE

default: [FALSE] if TRUE will cause all circular and hyperbolic functions to be converted to exponential form. (Setting DEMOIVRE:TRUE; will reverse the effect.) EXPONENTIALIZE(exp) will cause the conversion to exponential form of an expression without setting the switch or having to re-evaluate the expression with EV.

Variable EXPOP

default: [0] - the highest positive exponent which is automatically expanded. Thus (X+1)**3, when typed, will be automatically expanded only if EXPOP is greater than or equal to 3. If it is desired to have (X+1)**n expanded where n is greater than EXPOP then executing EXPAND((X+1)**n) will work only if MAXPOSEX is not less than n.

Variable FACTLIM

default: [-1] gives the highest factorial which is automatically expanded. If it is -1 then all integers are expanded.

Function INTOSUM (expr)

will take all things that a summation is multiplied by, and put them inside the summation. If the index is used in the outside expression, then the function tries to find a reasonable index, the same as it does for SUMCONTRACT. This is essentially the reverse idea of the OUTATIVE property of summations, but note that it does not remove this property, it only bypasses it. In some cases, a SCANMAP(MULTTHRU,expr) may be necessary before the INTOSUM.

declaration LASSOCIATIVE

- If DECLARE(G,LASSOCIATIVE); is done, this tells the simplifier that G is leftassociative. E.g. G(G(A,B),G(C,D)) will simplify to G(G(G(A,B),C),D).

declaration LINEAR

- One of MACSYMA’s OPPROPERTIES. For univariate f so declared, "expansion" F(X+Y) -> F(X)+F(Y), F(A*X) -> A*F(X) takes place where A is a "constant". For functions F of >=2 args, "linearity" is defined to be as in the case of ’SUM or ’INTEGRATE, i.e. F(A*X+B,X) -> A*F(X,X)+B*F(1,X) for A,B FREEOF X. (LINEAR is just ADDITIVE + OUTATIVE.)

declaration MAINVAR

- You may DECLARE variables to be MAINVAR. The ordering scale for atoms is essentially: numbers < constants (e.g. %E,%PI) < scalars < other variables < mainvars. E.g. compare EXPAND((X+Y)^4); with (DECLARE(X,MAINVAR), EX-PAND((X+Y)^4)); . (Note: Care should be taken if you elect to use the above feature. E.g. if you subtract an expression in which X is a MAINVAR from one in which X isn’t a MAINVAR, resimplification e.g. with EV(expression,SIMP) may be necessary if cancellation is to occur. Also, if you SAVE an expression in which X is a MAINVAR, you probably should also SAVE X.)

Variable MAXAPPLYDEPTH

default: [10000] - the maximum depth to which APPLY1 and APPLY2 will delve.

Variable MAXAPPLYHEIGHT

default: [10000] - the maximum height to which APPLYB1 will reach before giving up.

Variable MAXNEGEX

default: [1000] - the largest negative exponent which will be expanded by the EX-PAND command (see also MAXPOSEX).

Variable MAXPOSEX

default: [1000] - the largest exponent which will be expanded with the EXPAND command (see also MAXNEGEX).

declaration MULTIPLICATIVE

- If DECLARE(F,MULTIPLICATIVE) has been executed, then: (1) If F is univariate, whenever the simplifier encounters F applied to a product, F will be distributed over that product. I.e. F(X*Y); will simplify to F(X)*F(Y). (2) If F is a function of 2 or more arguments, multiplicativity is defined as multiplicativity in the first argument to F, i.e. F(G(X)*H(X),X); will simplify to F(G(X),X)*F(H(X),X). This simplification does not occur when F is applied to expressions of the form PRODUCT(X[I],I,lowerlimit,upper-limit).

Variable NEGDISTRIB

default: [TRUE] - when TRUE allows -1 to be distributed over an expression. E.g. -(X+Y) becomes -Y-X. Setting it to FALSE will allow -(X+Y) to be displayed like that. This is sometimes useful but be very careful: like the SIMP flag, this is one flag you do not want to set to FALSE as a matter of course or necessarily for other than local use in your MACSYMA.

Variable NEGSUMDISPFLAG

default: [TRUE] - when TRUE, X-Y displays as X-Y instead of as -Y+X. Setting it to FALSE causes the special check in display for the difference of two expressions to not be done. One application is that thus A+%I*B and A-%I*B may both be displayed the same way.

special symbol NOEVAL

- suppresses the evaluation phase of EV. This is useful in conjunction with other switches and in causing expressions to be resimplified without being reevaluated.

declaration NOUN

- One of the options of the DECLARE command. It makes a function so DECLAREd a "noun", meaning that it won’t be evaluated automatically.

Variable NOUNDISP

default: [FALSE] - if TRUE will cause NOUNs to display with a single quote. This switch is always TRUE when displaying function definitions.

special symbol NOUNS

(EVFLAG) when used as an option to the EV command, converts all "noun" forms occurring in the expression being EV’d to "verbs", i.e. evaluates them. See also NOUN, NOUNIFY, VERB, and VERBIFY.

special symbol NUMER

causes some mathematical functions (including exponentiation) with numerical arguments to be evaluated in floating point. It causes variables in exp which have been given numervals to be replaced by their values. It also sets the FLOAT switch on.

Function NUMERVAL (var1, exp1, var2, exp2, ...)

declares vari to have a numerval of expi which is evaluated and substituted for the variable in any expressions in which the variable occurs if the NUMER flag is TRUE. (see the EV function).

Variable OPPROPERTIES

- the list of the special operator-properties handled by the MACSYMA simplifier: LINEAR, ADDITIVE, MULTIPLICATIVE, OUTATIVE, EVENFUN, ODDFUN, COMMUTATIVE, SYMMETRIC, ANTISYMMETRIC, NARY, LASSOCIATIVE, and RASSOCIATIVE.

Variable OPSUBST

default:[TRUE] - if FALSE, SUBST will not attempt to substitute into the operator of an expression. E.g. (OPSUBST:FALSE, SUBST(X^2,R,R+R[0])); will work.

declaration OUTATIVE

- If DECLARE(F,OUTATIVE) has been executed, then: (1) If F is univariate, whenever the simplifier encounters F applied to a product, that product will be partitioned into factors that are constant and factors that are not and the constant factors will be pulled out. I.e. F(A*X); will simplify to A*F(X) where A is a constant. Non-atomic constant factors will not be pulled out. (2) If F is a function of 2 or more arguments, outativity is defined as in the case of ’SUM or ’INTEGRATE, i.e. F(A*G(X),X); will simplify to A*F(G(X),X) for A free-of X. Initially, ’SUM, ’INTEGRATE, and ’LIMIT are declared to be OUTATIVE.

declaration POSFUN

- POSitive FUNction, e.g. DECLARE(F,POSFUN); IS(F(X)>0); -> TRUE.

Variable PRODHACK

default: [FALSE] - if set to TRUE then PRODUCT(F(I),I,3,1); will yield 1/F(2), by the identity PRODUCT(F(I),I,A,B) = 1/PRODUCT(F(I),I,B+1,A-1) when A>B.

Function RADCAN (exp)

simplifies exp, which can contain logs, exponentials, and radicals, by converting it into a form which is canonical over a large class of expressions and a given ordering of variables; that is, all functionally equivalent forms are mapped into a unique form. For a somewhat larger class of expressions, RADCAN produces a regular form. Two equivalent expressions in this class will not necessarily have the same appearance, but their difference will be simplified by RADCAN to zero. For some expressions RADCAN can be quite time consuming. This is the cost of exploring certain relationships among the components of the expression for simplifications based on factoring and partialfraction expansions of exponents. %E TO NUMLOG (default: [FALSE]) - when set to TRUE, for "r" some rational number, and "x" some expression, %E^(r*LOG(x)) will be simplified into x^r . RADEXPAND[TRUE] when set to FALSE will inhibit certain transformations: RADCAN(SQRT(1-X)) will remain SQRT(1-X) and will not become %I SQRT(X-1). RADCAN(SQRT(X^2-2*X+1)) will remain SQRT(X^2-2*X + 1) and will not be transformed to X- 1. Do EXAMPLE(RADCAN); for examples.

Variable RADEXPAND

default: [TRUE] - if set to ALL will cause nth roots of factors of a product which are powers of n to be pulled outside of the radical. E.g. if RADEXPAND is ALL, SQRT(16*X^2) will become 4*X . More particularly, consider SQRT(X^2). (a) If RADEXPAND is ALL or ASSUME(X>0) has been done, SQRT(X^2) will become X.

(b) If RADEXPAND is TRUE and DOMAIN is REAL (its default), SQRT(X^2) will become ABS(X). (c) If RADEXPAND is FALSE, or RADEXPAND is TRUE and DOMAIN is COMPLEX, SQRT(X^2) will be returned. (The notion of DOMAIN with settings of REAL or COMPLEX is still in its infancy. Note that its setting here only matters when RADEXPAND is TRUE.)

Variable RADPRODEXPAND

- this switch has been renamed RADEXPAND.

Variable RADSUBSTFLAG

default: [FALSE] - if TRUE permits RATSUBST to make substitutions such as U for SQRT(X) in X.

declaration RASSOCIATIVE

- If DECLARE(G,RASSOCIATIVE); is done, this tells the simplifier that G is rightassociative. E.g. G(G(A,B),G(C,D)) will simplify to G(A,G(B,G(C,D))).

Function SCSIMP (exp,rule1, rule2,...,rulen)

Sequential Comparative Simplification [Stoute]) takes an expression (its first argument) and a set of identities, or rules (its other arguments) and tries simplifying. If a smaller expression is obtained, the process repeats. Otherwise after all simplifications are tried, it returns the original answer. For examples, try EXAMPLE(SCSIMP); .

Function SIMP

causes exp to be simplified regardless of the setting of the switch SIMP which inhibits simplification if FALSE.

Variable SIMPSUM

default: [FALSE] - if TRUE, the result of a SUM is simplified. This simplification may sometimes be able to produce a closed form. If SIMPSUM is FALSE or if ’SUM is used, the value is a sum noun form which is a representation of the sigma notation used in mathematics.

Function SUMCONTRACT (expr)

will combine all sums of an addition that have upper and lower bounds that differ by constants. The result will be an expression containing one summation for each set of such summations added to all appropriate extra terms that had to be extracted to form this sum. SUMCONTRACT will combine all compatible sums and use one of the indices from one of the sums if it can, and then try to form a reasonable index if it cannot use any supplied. It may be necessary to do an INTOSUM(expr) before the SUMCONTRACT.

Variable SUMEXPAND

default: [FALSE] if TRUE, products of sums and exponentiated sums are converted into nested sums. For example:

             SUMEXPAND:TRUE$
             SUM(F(I),I,0,M)*SUM(G(J),J,0,N); ->
                             'SUM('SUM(F(I1)*G(I2),I2,0,N),I1,0,M)
             SUM(F(I),I,0,M)^2; -> 'SUM('SUM(F(I3)*F(I4),I4,0,M),I3,0,M)

If FALSE, they are left alone. See also CAUCHYSUM.

Variable SUMHACK

default: [FALSE] - if set to TRUE then SUM(F(I),I,3,1); will yield -F(2), by the identity SUM(F(I),I,A,B) = - SUM(F(I),I,B+1,A-1) when A>B.

Variable SUMSPLITFACT

default: [TRUE] - if set to FALSE will cause MINFACTORIAL to be applied after a FACTCOMB.

declaration SYMMETRIC

- If DECLARE(H,SYMMETRIC); is done, this tells the simplifier that H is a symmetric function. E.g. H(X,Z,Y) will simplify to H(X, Y, Z). This is the same as COMMUTATIVE.

Function UNKNOWN (exp)

returns TRUE iff exp contains an operator or function not known to the built-in simplifier.

44 Maxima Manual

6 Expressions

6.1 Introduction to Expressions

6.2 ASSIGNMENT

- There are two assignment operators in MACSYMA, : and :: . E.g. A:3 sets the variable A to 3. :: assigns the value of the expression on its right to the value of the quantity on its left, which must evaluate to an atomic variable or subscripted variable.

6.3 COMPLEX

- A complex expression is specified in MACSYMA by adding the real part of the expression to %I times the imaginary part. Thus the roots of the equation X^2-4*X+13=0 are 2+3*%I and 2-3*%I. Note that simplification of products of complex expressions can be effected by expanding the product. Simplification of quotients, roots, and other functions of complex expressions can usually be accomplished by using the REALPART, IMAGPART, RECTFORM, POLARFORM, ABS, CARG functions.

6.4 INEQUALITY

- MACSYMA has the usual inequality operators: less than: < greater than: > greater than or equal to: >= less than or equal to: <=

6.5 SYNTAX

- It is possible to add new operators to MACSYMA (infix, prefix, postfix, unary, or matchfix with given precedences), to remove existing operators, or to redefine the precedence of existing operators. While MACSYMA’s syntax should be adequate for most ordinary applications, it is possible to define new operators or eliminate predefined ones that get in the user’s way. The extension mechanism is rather straightforward and should be evident from the examples below.

     (C1) PREFIX("DDX")$
     (C2) DDX Y$             means                   "DDX"(Y)
     (C3) INFIX("<-")$
     (C4) A<-DDX Y$          means               "<-"(A,"DDX"(Y))

For each of the types of operator except SPECIAL, there is a corresponding creation function that will give the lexeme specified the corresponding parsing properties. Thus "PREFIX("DDX")" will make "DDX" a prefix operator just like "-" or "NOT". Of course, certain extension functions require additional information such as the matching keyword for a matchfix operator. In addition, binding powers and parts of speech must be specified for all keywords defined. This is done by passing additional arguments to the extension functions. If a user does not specify these additional parameters, MAC-SYMA will assign default values. The six extension functions with binding powers and parts of speech defaults (enclosed in brackets) are summarized below. PREFIX(operator, rbp[180], rpos[ANY], pos[ANY]) POSTFIX(operator, lbp[180], lpos[ANY], pos[ANY]) INFIX(operator, lbp[180], rbp[180], lpos[ANY], rpos[ANY],pos[ANY]) NARY(operator, bp[180], argpos[ANY], pos[ANY]) NOFIX(operator, pos[ANY]) MATCHFIX(operator, match, argpos[ANY], pos[ANY]) The defaults have been provided so that a user who does not wish to concern himself with parts of speech or binding powers may simply omit those arguments to the extension functions. Thus the following are all equivalent. PREFIX("DDX",180,ANY,ANY)$ PREFIX("DDX",180)$ PREFIX("DDX")$ It is also possible to remove the syntax properties of an operator by using the functions REMOVE or KILL. Specifically, "REMOVE("DDX",OP)" or "KILL("DDX")" will return "DDX" to operand status; but in the second case all the other properties of "DDX" will also be removed.

6.6 Definitions for Expressions

Function AT (exp, list)

will evaluate exp (which may be any expression) with the variables assuming the values as specified for them in the list of equations or the single equation similar to that given to the ATVALUE function. If a subexpression depends on any of the variables in list but it hasn’t had an atvalue specified and it can’t be evaluated then a noun form of the AT will be returned which will display in a two-dimensional form. Do EXAMPLE(AT); for an example.

Function BOX (expr)

returns expr enclosed in a box. The box is actually part of the expression.

     BOX(expr,label)

encloses expr in a labelled box. label is a name which will be truncated in display if it is too long. BOXCHAR["] - is the character used to draw the box in this and in the DPART and LPART functions.

Variable BOXCHAR

default: ["] is the character used to draw the box in the BOX and in the DPART and LPART functions.

special operator CONSTANT

- makes ai a constant as is %PI.

Function CONSTANTP (exp)

is TRUE if exp is a constant (i.e. composed of numbers and %PI, %E, %I or any variables bound to a constant or DECLAREd constant) else FALSE. Any function whose arguments are constant is also considered to be a constant.

Function CONTRACT (exp)

carries out all possible contractions in exp, which may be any well-formed combination of sums and products. This function uses the information given to the DEFCON function. Since all tensors are considered to be symmetric in all indices, the indices are sorted into alphabetical order. Also all dummy indices are renamed using the symbols !1,!2,... to permit the expression to be simplified as much as possible by reducing equivalent terms to a canonical form. For best results exp should be fully expanded. RATEXPAND is the fastest way to expand products and powers of sums if there are no variables in the denominators of the terms. The GCD switch should be FALSE if gcd cancellations are unnecessary.

Function DECLARE (a1, f1, a2, f2, ...)

gives the atom ai the flag fi. The ai’s and fi’s may also be lists of atoms and flags respectively in which case each of the atoms gets all of the properties. The possible flags and their meanings are: CONSTANT - makes ai a constant as is %PI. MAINVAR - makes ai a MAINVAR. The ordering scale for atoms: numbers < constants (e.g. %E,%PI) < scalars < other variables < mainvars. SCALAR - makes ai a scalar. NONSCALAR - makes ai behave as does a list or matrix with respect to the dot operator. NOUN - makes the function ai a noun so that it won’t be evaluated automatically. EVFUN - makes ai known to the EV function so that it will get applied if its name is mentioned. Initial evfuns are

     FACTOR, TRIGEXPAND,
     TRIGREDUCE, BFLOAT, RATSIMP, RATEXPAND, and RADCAN

EVFLAG - makes ai known to the EV function so that it will be bound to TRUE during the execution of EV if it is mentioned. Initial evflags are

     FLOAT, PRED, SIMP, NUMER, DETOUT, EXPONENTIALIZE, DEMOIVRE,
     KEEPFLOAT, LISTARITH, TRIGEXPAND, SIMPSUM, ALGEBRAIC,
     RATALGDENOM, FACTORFLAG, %EMODE, LOGARC, LOGNUMER,
     RADEXPAND, RATSIMPEXPONS, RATMX, RATFAC, INFEVAL, %ENUMER,
     PROGRAMMODE, LOGNEGINT, LOGABS, LETRAT, HALFANGLES,
     EXPTISOLATE, ISOLATE_WRT_TIMES, SUMEXPAND, CAUCHYSUM,
     NUMER_PBRANCH, M1PBRANCH, DOTSCRULES, and LOGEXPAND

BINDTEST - causes ai to signal an error if it ever is used in a computation unbound. DECLARE([var1, var2, ...], BINDTEST) causes MACSYMA to give an error message whenever any of the vari occur unbound in a computation. MACSYMA currently recognizes and uses the following features of objects:

     EVEN, ODD, INTEGER, RATIONAL, IRRATIONAL, REAL, IMAGINARY,
     and COMPLEX

he useful features of functions include:

      INCREASING,
     DECREASING, ODDFUN (odd function), EVENFUN (even function),
     COMMUTATIVE (or SYMMETRIC), ANTISYMMETRIC, LASSOCIATIVE and
     RASSOCIATIVE

DECLARE(F,INCREASING) is in all respects equivalent to

     ASSUME(KIND(F,INCREASING))

The ai and fi may also be lists of objects or features. The command

     FEATUREP(object,feature)

may be used to determine if an object has been DECLAREd to have "feature". See DESCRIBE(FEATURES); .

Function DISOLATE (exp, var1, var2, ..., varN)

is similar to ISOLATE(exp, var) (Do DESCRIBE(ISOLATE);) except that it enables the user to isolate more than one variable simultaneously. This might be useful, for example, if one were attempting to change variables in a multiple integration, and that variable change involved two or more of the integration variables. To access this function, do LOAD(DISOL)$ . A demo is available by DEMO("disol"); .

Function DISPFORM (exp)

returns the external representation of exp (wrt its main operator). This should be useful in conjunction with PART which also deals with the external representation. Suppose EXP is -A . Then the internal representation of EXP is "*"(-1,A), while the external representation is "-"(A). DISPFORM(exp,ALL) converts the entire expression (not just the top-level) to external format. For example, if EXP:SIN(SQRT(X)), then FREEOF(SQRT,EXP) and FREEOF(SQRT,DISPFORM(EXP)) give TRUE, while FREEOF(SQRT,DISPFORM(EXP,ALL)) gives FALSE.

Function DISTRIB (exp)

distributes sums over products. It differs from EXPAND in that it works at only the top level of an expression, i.e. it doesn’t recurse and it is faster than EXPAND. It differs from MULTTHRU in that it expands all sums at that level. For example, DISTRIB((A+B)*(C+D)) -> A C + A D + B C + B D MULTTHRU ((A+B)*(C+D)) -> (A + B) C + (A + B) D DISTRIB (1/((A+B)*(C+D))) -> 1/ ((A+B) *(C+D)) EXPAND(1/((A+B)*(C+D)),1,0) -> 1/(A C + A D + B C + B D)

Function DPART (exp, n1, ..., nk)

selects the same subexpression as PART, but instead of just returning that subexpression as its value, it returns the whole expression with the selected subexpression displayed inside a box. The box is actually part of the expression.

     (C1) DPART(X+Y/Z**2,1,2,1);
                            Y
     (D1)                  ---- + X
                              2
                          *****
                          * Z *
                          *****

Function EXP (X)

the exponential function. It is represented internally as %E^X. DEMOIVRE[FALSE] - if TRUE will cause %E^(A+B*%I) to become %E^A*(COS(B)+%I*SIN(B)) if B is free of %I. A and B are not expanded. %EMODE[TRUE] - when TRUE %E^(%PI*%I*X) will be simplified as follows: it will become COS(%PI*X)+%I*SIN(%PI*X) if X is an integer or a multiple of 1/2, 1/3, 1/4, or 1/6 and thus will simplify further. For other numerical X it will become %E^(%PI*%I*Y) where Y is X-2*k for some integer k such that ABS(Y)<1. If %EMODE is FALSE no simplification of %E^(%PI*%I*X) will take place. %ENUMER[FALSE] - when TRUE will cause %E to be converted into 2.718... whenever NUMER is TRUE. The default is that this conversion will take place only if the exponent in %E^X evaluates to a number.

Variable EXPTISOLATE

default: [FALSE] if TRUE will cause ISOLATE(expr,var); to examine exponents of atoms (like %E) which contain var.

Variable EXPTSUBST

default: [FALSE] if TRUE permits substitutions such as Y for %E**X in %E**(A*X) to take place.

Function FREEOF (x1, x2, ..., exp)

yields TRUE if the xi do not occur in exp and FALSE otherwise. The xi are atoms or they may be subscripted names, functions (e.g. SIN(X) ), or operators enclosed in "s. If ’var’ is a "dummy variable" of ’exp’, then FREEOF(var,exp); will return TRUE. "Dummy variables" are mathematical things like the index of a sum or product, the limit variable, and the definite integration variable. Example: FREEOF(I,’SUM(F(I),I,0,N)); returns TRUE. Do EXAMPLE(FREEOF); for more examples.

Function GENFACT (X, Y, Z)

is the generalized factorial of X which is: X*(X-Z)*(X-2*Z)*...*(X-(Y-1)*Z). Thus, for integral X, GENFACT(X,X,1)=X! and GENFACT(X,X/2,2)=X!!

Function IMAGPART (exp)

returns the imaginary part of the expression exp.

Function INDICES (exp)

returns a list of two elements. The first is a list of the free indices in exp (those that occur only once); the second is the list of dummy indices in exp (those that occur exactly twice).

Function INFIX (op)

- INFIX operators are used to denote functions of two arguments, one given before the operator and one after, e.g. A^2 . The INFIX("x") function is a syntax extention function to declare x to be an INFIX operator. Do DESCRIBE(SYNTAX); for more details.

Variable INFLAG

default: [FALSE] if set to TRUE, the functions for part extraction will look at the internal form of exp. Note that the simplifier re-orders expressions. Thus FIRST(X+Y) will be X if INFLAG is TRUE and Y if INFLAG is FALSE. (FIRST(Y+X) gives the same results). Also, setting INFLAG to TRUE and calling PART/SUBSTPART is the same as calling INPART/SUBSTINPART. Functions affected by the setting of INFLAG are: PART, SUBSTPART, FIRST, REST, LAST, LENGTH, the FOR ... IN construct, MAP, FULLMAP, MAPLIST, REVEAL and PICKAPART.

Function INPART (exp, n1, ..., nk)

is similar to PART but works on the internal representation of the expression rather than the displayed form and thus may be faster since no formatting is done. Care should be taken with respect to the order of subexpressions in sums and products (since the order of variables in the internal form is often different from that in the displayed form) and in dealing with unary minus, subtraction, and division (since these operators are removed from the expression). PART(X+Y,0) or INPART(X+Y,0) yield +, though in order to refer to the operator it must be enclosed in "s. For example ...IF INPART(D9,0)="+" THEN ...

     (C1)  X+Y+W*Z;
     (D1)                  W Z + Y + X
     (C2)  INPART(D1,3,2);
     (D2)                  Z
     (C3)  PART(D1,1,2);
     (D3)                  Z
     (C4) 'LIMIT(F(X)**G(X+1),X,0,MINUS);
                                           G(X + 1)
     (D4)                      LIMIT   F(X)
                               X ->0-
     (C5) INPART(%,1,2);
     (D5)                            G(X + 1)

Function ISOLATE (exp, var)

returns exp with subexpressions which are sums and which do not contain var replaced by intermediate expression labels (these being atomic symbols like E1, E2, ...). This is often useful to avoid unnecessary expansion of subexpressions which don’t contain the variable of interest. Since the intermediate labels are bound to the subexpressions they can all be substituted back by evaluating the expression in which they occur. EXPTISOLATE[FALSE] if TRUE will cause ISOLATE to examine exponents of atoms (like %E) which contain var. ISOLATE WRT TIMES[FALSE] if TRUE, then ISOLATE will also isolate wrt products. E.g. compare both settings of the switch on ISOLATE(EXPAND((A+B+C)^2),C); . Do EXAMPLE(ISOLATE); for examples.

Variable ISOLATE WRT TIMES

default: [FALSE] - if set to TRUE, then ISOLATE will also isolate wrt products. E.g. compare both settings of the switch on ISOLATE(EXPAND((A+B+C)^2),C); .

Variable LISTCONSTVARS

default: [FALSE] - if TRUE will cause LISTOFVARS to include %E, %PI, %I, and any variables declared constant in the list it returns if they appear in the expression LISTOFVARS is called on. The default is to omit these.

Variable LISTDUMMYVARS

default: [TRUE] - if FALSE, "dummy variables" in the expression will not be included in the list returned by LISTOFVARS. (The meaning of "dummy variables" is as given in DESCRIBE(FREEOF): "Dummy variables" are mathematical things like the index of a sum or product, the limit variable, and the definite integration variable.) Example: LISTOFVARS(’SUM(F(I),I,0,N)); gives [I,N] if LISTDUMMYVARS is TRUE, and [N] if LISTDUMMYVARS is FALSE.

Function LISTOFVARS (exp)

yields a list of the variables in exp. LISTCONSTVARS[FALSE] if TRUE will cause LISTOFVARS to include %E, %PI, %I, and any variables declared constant in the list it returns if they appear in exp. The default is to omit these.

     (C1) LISTOFVARS(F(X[1]+Y)/G**(2+A));
     (D1)                            [X[1], Y, A, G]

Function LOPOW (exp, v)

the lowest exponent of v which explicitly appears in exp. Thus

     LOPOW((X+Y)**2+(X+Y)**A,X+Y) ==> MIN(A,2)

.

Function LPART (label, expr, n1, ..., nk)

is similar to DPART but uses a labelled box. A labelled box is similar to the one produced by DPART but it has a name in the top line.

Function MULTTHRU (exp)

multiplies a factor (which should be a sum) of exp by the other factors of exp. That is exp is f1*f2*...*fn where at least one factor, say fi, is a sum of terms. Each term in that sum is multiplied by the other factors in the product. (Namely all the factors except fi). MULTTHRU does not expand exponentiated sums. This function is the fastest way to distribute products (commutative or noncommutative) over sums. Since quotients are represented as products MULTTHRU can be used to divide sums by products as well. MULTTHRU(exp1, exp2) multiplies each term in exp2 (which should be a sum or an equation) by exp1. If exp1 is not itself a sum then this form is equivalent to MULTTHRU(exp1*exp2).

     (C1) X/(X-Y)**2-1/(X-Y)-F(X)/(X-Y)**3;
                    1        X         F(X)
     (D1)       - ----- + -------- - --------
                  X - Y          2          3
                          (X - Y)    (X - Y)

     (C2) MULTTHRU((X-Y)**3,%);
                         2
     (D2)       - (X - Y)  + X (X - Y) - F(X)
     (C3) RATEXPAND(D2);
                                2
     (D3)                    - Y  + X Y - F(X)
     (C4) ((A+B)**10*S**2+2*A*B*S+(A*B)**2)/(A*B*S**2);
                              10  2              2  2
                     (B  + A )   S  + 2 A B S + A  B
     (D4)            --------------------------------
                                        2
                                   A B S
     (C5) MULTTHRU(%);
                                               10
                             2   A B   (B  + A)
     (D5)                    - + --- + -------
                             S    2      A B
                                 S
     (notice that (B+A)**10 is not expanded)
     (C6) MULTTHRU(A.(B+C.(D+E)+F));
     (D6)                A . F + A . (C . (E + D)) + A . B
     (compare with similar example under EXPAND)

Function NOUNIFY (f)

returns the noun form of the function name f. This is needed if one wishes to refer to the name of a verb function as if it were a noun. Note that some verb functions will return their noun forms if they can’t be evaluated for certain arguments. This is also the form returned if a function call is preceded by a quote.

Function NTERMS (exp)

gives the number of terms that exp would have if it were fully expanded out and no cancellations or combination of terms occurred. Note that expressions like SIN(E), SQRT(E), EXP(E), etc. count as just one term regardless of how many terms E has (if it is a sum).

Function OPTIMIZE (exp)

returns an expression that produces the same value and side effects as exp but does so more efficiently by avoiding the recomputation of common subexpressions. OP-TIMIZE also has the side effect of "collapsing" its argument so that all common subexpressions are shared. Do EXAMPLE(OPTIMIZE); for examples.

Variable OPTIMPREFIX

default: [%] - The prefix used for generated symbols by the OPTIMIZE command.

Function ORDERGREAT (V1, ..., Vn)

sets up aliases for the variables V1, ..., Vn such that V1 > V2 > ... > Vn > any other variable not mentioned as an argument. See also ORDERLESS. Caveat: do EXAMPLE(ORDERGREAT); for some specifics.

Function ORDERGREATP (exp1,exp2)

returns TRUE if exp2 precedes exp1 in the ordering set up with the ORDERGREAT function (see DESCRIBE(ORDERGREAT);).

Function ORDERLESS (V1, ..., Vn)

sets up aliases for the variables V1, ..., Vn such that V1 < V2 < ... < Vn < any other variable not mentioned as an argument. Thus the complete ordering scale is: numerical constants < declared constants < declared scalars < first argument to ORDERLESS < ... < last argument to ORDERLESS < variables which begin with A < ... < variables which begin with Z < last argument to ORDERGREAT < ... < first argument to ORDERGREAT < declared MAINVARs. Caveat: do EXAMPLE(ORDERLESS); for some specifics. For another ordering scheme, see DESCRIBE(MAINVAR);.

Function ORDERLESSP (exp1,exp2)

returns TRUE if exp1 precedes exp2 in the ordering set up by the ORDERLESS command (see DESCRIBE(ORDERLESS);).

Function PART (exp, n1, ..., nk)

deals with the displayed form of exp. It obtains the part of exp as specified by the indices n1,...,nk. First part n1 of exp is obtained, then part n2 of that, etc. The result is part nk of ... part n2 of part n1 of exp. Thus PART(Z+2*Y,2,1) yields 2. PART can be used to obtain an element of a list, a row of a matrix, etc. If the last argument to a Part function is a list of indices then several subexpressions are picked out, each one corresponding to an index of the list. Thus PART(X+Y+Z,[1,3]) is Z+X. PIECE holds the last expression selected when using the Part functions. It is set during the execution of the function and thus may be referred to in the function itself as shown below. If PARTSWITCH[FALSE] is set to TRUE then END is returned when a selected part of an expression doesn’t exist, otherwise an error message is given. For examples, do EXAMPLE(PART);

Function PARTITION (exp, var)

returns a list of two expressions. They are (1) the factors of exp (if it is a product), the terms of exp (if it is a sum), or the list (if it is a list) which don’t contain var and,

(2) the factors, terms, or list which do.

     (C1) PARTITION(2*A*X*F(X),X);
     (D1)                 [ 2 A , X F(X) ]
     (C2) PARTITION(A+B,X);
     (D2)                 [ A + B , 0 ]
     (C3) PARTITION([A,B,F(A),C],A);
     (D3)                [[B,C],[A,F(A)]]

Variable PARTSWITCH

default: [FALSE] - if set to TRUE then END is returned when a selected part of an expression doesn’t exist, otherwise an error message is given.

Function PICKAPART (exp,depth)

will assign E labels to all subexpressions of exp down to the specified integer depth. This is useful for dealing with large expressions and for automatically assigning parts of an expression to a variable without having to use the part functions.

     (C1) EXP:(A+B)/2+SIN(X^2)/3-LOG(1+SQRT(X+1));
                                                      2
                                                 SIN(X )   B + A
     (D1)               - LOG(SQRT(X + 1) + 1) + ------- + -----
                                                    3        2
     (C2) PICKAPART(%,1);
     (E2)                    - LOG(SQRT(X + 1) + 1)
                                         2
                                    SIN(X )
     (E3)                           -------
                                       3
                                     B + A
     (E4)                            -----
                                       2
     (D4)                         E4 + E3 + E2

Variable PIECE

- holds the last expression selected when using the Part functions. It is set during the execution of the function and thus may be referred to in the function itself.

Function POWERS (expr, var)

gives the powers of var occuring in expr. To use it, do LOAD(POWERS);. For details on usage, do PRINTFILE("powers.usg");.

Function PRODUCT (exp, ind, lo, hi)

gives the product of the values of exp as the index ind varies from lo to hi. The evaluation is similar to that of SUM. No simplification of products is available at this time. If hi is one less than lo, we have an "empty product" and PRODUCT returns 1 rather than erring out. Also see DESCRIBE(PRODHACK).

     (C1)  PRODUCT(X+I*(I+1)/2,I,1,4);
     (D1)             (X + 1) (X + 3) (X + 6) (X + 10)

Function REALPART (exp)

gives the real part of exp. REALPART and IMAGPART will work on expressions involving trigonometic and hyperbolic functions, as well as SQRT, LOG, and exponentiation.

Function RECTFORM (exp)

returns an expression of the form A + B*%I, where A and B are purely real.

Function REMBOX (expr, arg)

removes boxes from expr according to arg. If arg is UNLABELED then all unlabelled boxes are removed. If arg is the name of some label then only boxes with that label are removed. If arg is omitted then all boxes labelled and unlabelled are removed.

Function SUM (exp, ind, lo, hi)

performs a summation of the values of exp as the index ind varies from lo to hi. If the upper and lower limits differ by an integer then each term in the sum is evaluated and added together. Otherwise, if the SIMPSUM [FALSE] is TRUE the result is simplified. This simplification may sometimes be able to produce a closed form. If SIMPSUM is FALSE or if ’SUM is used, the value is a sum noun form which is a representation of the sigma notation used in mathematics. If hi is one less than lo, we have an "empty sum" and SUM returns 0 rather than erring out. Sums may be differentiated, added, subtracted, or multiplied with some automatic simplification being performed. Also see DESCRIBE(SUMHACK). CAUCHYSUM[FALSE] when TRUE causes the Cauchy product to be used when multiplying sums together rather than the usual product. In the Cauchy product the index of the inner summation is a function of the index of the outer one rather than varying independently. GENINDEX[I] is the alphabetic prefix used to generate the next variable of summation. GENSUM-NUM[0] is the numeric suffix used to generate the next variable of summation. If it is set to FALSE then the index will consist only of GENINDEX with no numeric suffix. Do EXAMPLE(SUM); for examples. See also SUMCONTRACT, INTOSUM, BASHINDICES, and NICEINDICES.

special symbol VERB

- the opposite of "noun", i.e. a function form which "does something" ("action" - for most functions the usual case). E.g. INTEGRATE integrates a function, unless it is DECLAREd to be a "noun", in which case it represents the INTEGRAL of the function. See NOUN, NOUNIFY, and VERBIFY.

Function VERBIFY (f)

returns the function name f in its verb form (See also VERB, NOUN, and NOUNIFY).

56 Maxima Manual

7 Function Definition

7.1 Introduction to Function Definition

7.2 FUNCTION

- To define a function in MACSYMA you use the := operator. E.g.

     F(X):=SIN(X)

defines a function F. Anonmyous functions may also be created using LAMBDA. For example

     lambda([i,j], ... )

can be used instead of F where

     F(I,J):=BLOCK([], ... );
     MAP(LAMBDA([I],I+1),L)

would return a list with 1 added to each term.

See also DEFINE.

7.3 MACROS

The Macsyma Macro facility is now available. BUILDQ is a MACSYMA function for constructing pieces of code to be executed. It will be very useful in conjunction with macros, but may have other applications as well. More documentation will be forthcoming as we receive feedback from users saying what questions are not answered by the existing documentation.

7.4 OPTIMIZATION

When using TRANSLATE and generating code with MACSYMA, there are a number of techniques which can save time and be helpful. Do DEMO("optimu.dem") for a demonstration. In particular, the function FLOATDEFUNK from TRANSL;OPTIMU FASL, creates a function definition from a math-like expression, but it optimizes it (with OPTIMIZE) and puts in the MODE DECLAREations needed to COMPILE correctly. (This can be done by hand, of course). The demo will only run in a fresh macsyma.

7.5 Definitions for Function Definition

Function APPLY (function, list)

gives the result of applying the function to the list of its arguments. This is useful when it is desired to compute the arguments to a function before applying that function. For example, if L is the list [1, 5, -10.2, 4, 3], then APPLY(MIN,L) gives -10.2. APPLY is also useful when calling functions which do not have their arguments evaluated if it is desired to cause evaluation of them. For example, if FILESPEC is a variable bound to the list [TEST, CASE] then APPLY(CLOSEFILE,FILESPEC) is equivalent to CLOSEFILE(TEST,CASE). In general the first argument to APPLY should be preceded by a ’ to make it evaluate to itself. Since some atomic variables have the same name as certain functions the values of the variable would be used rather than the function because APPLY has its first argument evaluated as well as its second.

Function BINDTEST (ai)

causes ai to signal an error if it ever is used in a computation unbound.

Function BLOCK ([v1,...,vk], statement1,...,statementj)

Blocks in MACSYMA are somewhat analogous to subroutines in FORTRAN or procedures in ALGOL or PL/I. Blocks are like compound statements but also enable the user to label statements within the block and to assign "dummy" variables to values which are local to the block. The vi are variables which are local to the BLOCK and the stmti are any MACSYMA expressions. If no variables are to be made local then the list may be omitted. A block uses these local variables to avoid conflict with variables having the same names used outside of the block (i.e. global to the block). In this case, upon entry to the block, the global values are saved onto a stack and are inaccessible while the block is being executed. The local variables then are unbound so that they evaluate to themselves. They may be bound to arbitrary values within the block but when the block is exited the saved values are restored to these variables. The values created in the block for these local variables are lost. Where a variable is used within a block and is not in the list of local variables for that block it will be the same as the variable used outside of the block. If it is desired to save and restore other local properties besides VALUE, for example ARRAY (except for complete arrays), FUNCTION, DEPENDENCIES, ATVALUE, MATCHDECLARE, ATOM-GRAD, CONSTANT, and NONSCALAR then the function LOCAL should be used inside of the block with arguments being the names of the variables. The value of the block is the value of the last statement or the value of the argument to the function RETURN which may be used to exit explicitly from the block. The function GO may be used to transfer control to the statement of the block that is tagged with the argument to GO. To tag a statement, precede it by an atomic argument as another statement in the BLOCK. For example: BLOCK([X],X:1,LOOP,X:X+1,...,GO(LOOP),...). The argument to GO must be the name of a tag appearing within the BLOCK. One cannot use GO to transfer to a tag in a BLOCK other than the one containing the GO. Blocks typically appear on the right side of a function definition but can be used in other places as well.

Function BREAK (arg1, ...)

will evaluate and print its arguments and will then cause a (MACSYMA-BREAK) at which point the user can examine and change his environment. Upon typing EXIT; the computation resumes. Control-A (^A) will enter a MACSYMA-BREAK from any point interactively. EXIT; will continue the computation. Control-X may be used inside the MACSYMA-BREAK to quit locally, without quitting the main computation.

Macro BUILDQ

- See DESCRIBE(MACROS); .

Function CATCH (exp1,...,expn)

evaluates its arguments one by one; if the structure of the expi leads to the evaluation of an expression of the form THROW(arg), then the value of the CATCH is the value of THROW(arg). This "non-local return" thus goes through any depth of nesting to the nearest enclosing CATCH. There must be a CATCH corresponding to a THROW, else an error is generated. If the evaluation of the expi does not lead to the evaluation of any THROW then the value of the CATCH is the value of expn.

     (C1) G(L):=CATCH(MAP(LAMBDA([X],
            IF X<0 THEN THROW(X) ELSE F(X)),L))$
     (C2) G([1,2,3,7]);
     (D2)                     [F(1), F(2), F(3), F(7)]
     (C3) G([1,2,-3,7]);
     (D3)                                - 3

The function G returns a list of F of each element of L if L consists only of non-negative numbers; otherwise, G "catches" the first negative element of L and "throws" it up.

Function COMPFILE ([filespec], f1, f2, ..., fn)

Compiles functions fi into the file "filespec". For convenience, see the COMPILE function.

Variable COMPGRIND

default: [FALSE] when TRUE function definitions output by COMPFILE are prettyprinted.

Function COMPILE (f)

The COMPILE command is a convenience feature in macsyma. It handles the calling of the function COMPFILE, which translates macsyma functions into lisp, the calling of the lisp compiler on the file produced by COMPFILE, and the loading of the output of the compiler, know as a FASL file, into the macsyma. It also checks the compiler comment listing output file for certain common errors. Do PRINT-FILE(MCOMPI,DOC,MAXDOC); for more details. COMPILE(); causes macsyma to prompt for arguments. COMPILE(function1,function2,...); compiles the functions, it uses the name of function1 as the first name of the file to put the lisp output. COMPILE(ALL); or COMPILE(FUNCTIONS); will compile all functions. COMPILE([file-name],function1,function2,...); N.B. all arguments are evaluated, just like a normal function (it is a normal function!). Therefore, if you have variables with the same name as part of the file you can not ignore that fact.

Function COMPILE LISP FILE ("input filename")

which takes an optional second argument of "output filename," can be used in conjunction with

     TRANSLATE_FILE("filename").

For convenience you might define

     Compile_and_load(FILENAME):=
         LOAD(COMPILE_LISP_FILE(TRANSLATE_FILE(FILENAME)[2]))[2]);

These file-oriented commands are to be preferred over the use of COMPILE, COMP-FILE, and the TRANSLATE SAVE combination.

Function DEFINE (f(x1, ...), body)

is equivalent to f(x1,...):=”body but when used inside functions it happens at execution time rather than at the time of definition of the function which contains it.

Function DEFINE VARIABLE

(name,default-binding,mode,optional-documentation) introduces a global variable into the MACSYMA environment. This is for user-written packages, which are often translated or compiled. Thus

     DEFINE_VARIABLE(FOO,TRUE,BOOLEAN);

does the following:

(1) MODE DECLARE(FOO,BOOLEAN); sets it up for the translator.

(2) If the variable is unbound, it sets it: FOO:TRUE.

(3) DECLARE(FOO,SPECIAL); declares it special.

(4) Sets up an assign property for it to make sure that it never gets set to a value of the wrong mode. E.g. FOO:44 would be an error once FOO is defined BOOLEAN. See DESCRIBE(MODE DECLARE); for a list of the possible "modes". The optional 4th argument is a documentation string. When TRANSLATE FILE is used on a package which includes documentation strings, a second file is output in addition to the LISP file which will contain the documentation strings, formatted suitably for use in manuals, usage files, or (for instance) DESCRIBE. With any variable which has been DEFINE VARIABLE’d with mode other than ANY, you can give a VALUE CHECK property, which is a function of one argument called on the value the user is trying to set the variable to.

     PUT('G5,LAMBDA([U],IF U#'G5 THEN ERROR("Don't set G5")),
            'VALUE_CHECK);

Use DEFINE VARIABLE(G5,’G5,ANY CHECK, "this ain’t supposed to be set by anyone but me.") ANY CHECK is a mode which means the same as ANY, but which keeps DEFINE VARIABLE from optimizing away the assign property.

Function DISPFUN (f1, f2, ...)

displays the definition of the user defined functions f1, f2, ... which may also be the names of array associated functions, subscripted functions, or functions with constant subscripts which are the same as those used when the functions were defined. DISPFUN(ALL) will display all user defined functions as given on the FUNCTIONS and ARRAYS lists except subscripted functions with constant subscripts. E.g. if the user has defined a function F(x), DISPFUN(F); will display the definition.

Variable FUNCTIONS

default: [] - all user defined functions (set up by f(x):=...).

Function FUNDEF (functionname)

returns the function definition associated with "functionname". FUNDEF(fnname); is similar to DISPFUN(fnname); except that FUNDEF does not invoke display.

Function FUNMAKE (name,[arg1,...,argn])

returns name(arg1,...,argn) without calling the function name.

Function LOCAL (v1, v2, ...)

causes the variables v1,v2,... to be local with respect to all the properties in the statement in which this function is used. LOCAL may only be used in BLOCKs, in the body of function definitions or LAMBDA expressions, or in the EV function and only one occurrence is permitted in each. LOCAL is independent of CONTEXT.

Variable MACROEXPANSION

default:[FALSE] - Controls advanced features which affect the efficiency of macros. Possible settings: FALSE – Macros expand normally each time they are called. EX-PAND – The first time a particular call is evaluated, the expansion is "remembered" internally, so that it doesn’t have to be recomputed on subsequent calls making subsequent calls faster. The macro call still GRINDs and DISPLAYs normally, however extra memory is required to remember all of the expansions. DISPLACE – The first time a particular call is evaluated, the expansion is substituted for the call. This requires slightly less storage than when MACROEXPANSION is set to EXPAND and is just as fast, but has the disadvantage that the original macro call is no longer remembered and hence the expansion will be seen if DISPLAY or GRIND is called. See documentation for TRANSLATE and MACROS for more details.

Variable MODE CHECKP

default: [TRUE] - If TRUE, MODE DECLARE checks the modes of bound variables.

Variable MODE CHECK ERRORP

default: [FALSE] - If TRUE, MODE DECLARE calls error.

Variable MODE CHECK WARNP

default: [TRUE] - If TRUE, mode errors are described.

Function MODE DECLARE (y1, mode1, y2, mode2, ...)

MODEDECLARE is a synonym for this. MODE DECLARE is used to declare the modes of variables and functions for subsequent translation or compilation of functions. Its arguments are pairs consisting of a variable yi, and a mode which is one of BOOLEAN, FIXNUM, NUMBER, RATIONAL, or FLOAT. Each yi may also be a list of variables all of which are declared to have modei. If yi is an array, and if every element of the array which is referenced has a value then ARRAY(yi, COMPLETE, dim1, dim2, ...) rather than

     ARRAY(yi, dim1, dim2, ...)

should be used when first declaring the bounds of the array. If all the elements of the array are of mode FIXNUM (FLOAT), use FIXNUM (FLOAT) instead of COMPLETE. Also if every element of the array is of the same mode, say m, then

     MODE_DECLARE(COMPLETEARRAY(yi),m))

should be used for efficient translation. Also numeric code using arrays can be made to run faster by declaring the expected size of the array, as in:

     MODE_DECLARE(COMPLETEARRAY(A[10,10]),FLOAT)

for a floating point number array which is 10 x 10. Additionally one may declare the mode of the result of a function by using FUNCTION(F1,F2,...) as an argument; here F1,F2,... are the names of functions. For example the expression,

     MODE_DECLARE([FUNCTION(F1,F2,...),X],FIXNUM,Q,
                                     COMPLETEARRAY(Q),FLOAT)

declares that X and the values returned by F1,F2,... are single-word integers and that Q is an array of floating point numbers. MODE DECLARE is used either immediately inside of a function definition or at top-level for global variables. Do PRINTFILE(MCOMPI,DOC,MAXDOC); for some examples of the use of MODE DECLARE in translation and compilation.

Function MODE IDENTITY (arg1,arg2)

A special form used with MODE DECLARE and MACROS to delcare, e.g., a list of lists of flonums, or other compound data object. The first argument to MODE IDENTITY is a primitive value mode name as given to MODE DECLARE (i.e. [FLOAT,FIXNUM,NUMBER, LIST,ANY]), and the second argument is an expression which is evaluated and returned as the value of MODE IDENTITY. However, if the return value is not allowed by the mode declared in the first argument, an error or warning is signalled. The important thing is that the MODE of the expression as determined by the MACSYMA to Lisp translator, will be that given as the first argument, independent of anything that goes on in the second argument. E.g. X:3.3; MODE IDENTITY(FIXNUM,X); is an error. MODE IDENTITY(FLONUM,X) returns 3.3 . This has a number of uses, e.g., if you knew that FIRST(L) returned a number then you might write MODE IDENTITY(NUMBER,FIRST(L)). However, a more efficient way to do it would be to define a new primitive,

     FIRSTNUMB(X)::=BUILDQ([X],MODE_IDENTITY(NUMBER,X))$

and use FIRSTNUMB every time you take the first of a list of numbers.

Variable TRANSBIND

default: [FALSE] - if TRUE removes global declarations in the local context. This applies to variables which are formal parameters to functions which one is TRANSLATEing from MACSYMA code to LISP.

Variable TRANSCOMPILE

default:[FALSE] - if true, TRANSLATE will generate the declarations necessary for possible compilation. The COMPFILE command uses TRANSCOMPILE:TRUE;.

Function TRANSLATE (f1, f2, ...)

translates the user defined functions f1,f2,... from the MACSYMA language to LISP (i.e. it makes them EXPRs). This results in a gain in speed when they are called. There is now a version of macsyma with the macsyma to lisp translator pre-loaded into it. It is available by typing :TM (for TranslateMacsyma) at DDT level. When given a file name, E.g. :TM GJC;TMTEST > , it gives that file to the function TRANSLATE FILE, and proceeds without further user interaction. If no file name is given, :TM gives a regular macsyma "(C1)" line. P.s. A user init file with second name "TM" will be loaded if it exists. You may just want to link this to your macsyma init file. Functions to be translated should include a call to MODE DECLARE at the beginning when possible in order to produce more efficient code. For example:

     F(X1,X2,...):=BLOCK([v1,v2,...],
             MODE_DECLARE(v1,mode1,v2,mode2,...),...)

where the X1,X2,... are the parameters to the function and the v1,v2,... are the local variables. The names of translated functions are removed from the FUNCTIONS list if SAVEDEF is FALSE (see below) and are added to the PROPS lists. Functions should not be translated unless they are fully debugged. Also, expressions are assumed simplified; if they are not, correct but non- optimal code gets generated. Thus, the user should not set the SIMP switch to FALSE which inhibits simplification of the expressions to be translated. The switch TRANSLATE, default: [FALSE], If TRUE, causes automatic translation of a user’s function to LISP. Note that translated functions may not run identically to the way they did before translation as certain incompatabilities may exist between the LISP and MACSYMA versions. Principally, the RAT function with more than one argument and the RATVARS function should not be used if any variables are MODE DECLAREd CRE. Also the PREDERROR:FALSE setting will not translate. SAVEDEF[TRUE] - if TRUE will cause the MACSYMA version of a user function to remain when the function is TRANSLATEd. This permits the definition to be displayed by DISPFUN and allows the function to be edited. TRANSRUN[TRUE] - if FALSE will cause the interpreted version of all functions to be run (provided they are still around) rather than the translated version. One can translate functions stored in a file by giving TRANSLATE an argument which is a file specification. This is a list of the form [fn1,fn2,DSK,dir] where fn1 fn2 is the name of the file of MACSYMA functions, and dir is the name of a file directory. The result returned by TRANSLATE is a list of the names of the functions TRANSLATEd. In the case of a file translation the corresponding element of the list is a list of the first and second new file names containing the LISP code resulting from the translation. This will be fn1 LISP on the disk directory dir. The file of LISP code may be read into MACSYMA by using the LOADFILE function.

Function TRANSLATE FILE (file)

translates a file of MACSYMA code into a file of LISP code. It takes one or two arguments. The first argument is the name of the MACSYMA file, and the optional second argument is the name of the LISP file to produce. The second argument defaults to the first argument with second file name the value of TR OUTPUT FILE DEFAULT which defaults to TRLISP. For example: TRANSLATE FILE("test.mc")); will translate "test.mc" to "test.LISP". Also produced is a file of translator warning messages of various degrees of severity. The second file name is always UNLISP. This file contains valuable (albeit obsure for some) information for tracking down bugs in translated code. Do APROPOS(TR ) to get a list of TR (for TRANSLATE) switches. In summary, TRANSLATE FILE("foo.mc"), LOADFILE("foo.LISP") is "=" to BATCH("foo.mc") modulo certain restrictions (the use of ” and % for example).

Variable TRANSRUN

default: [TRUE] - if FALSE will cause the interpreted version of all functions to be run (provided they are still around) rather than the translated version.

Variable TR ARRAY AS REF

default: [TRUE] - If TRUE runtime code uses the value of the variable as the array.

Variable TR BOUND FUNCTION APPLYP

default: [TRUE] - Gives a warning if a bound variable is found being used as a function.

Variable TR FILE TTY MESSAGESP

default: [FALSE] - Determines whether messages generated by TRANSLATE FILE during translation of a file will be sent to the TTY. If FALSE (the default), messages about translation of the file are only inserted into the UNLISP file. If TRUE, the messages are sent to the TTY and are also inserted into the UNLISP file.

Variable TR FLOAT CAN BRANCH COMPLEX

default: [TRUE] - States whether the arc functions might return complex results. The arc functions are SQRT, LOG, ACOS, etc. e.g. When it is TRUE then ACOS(X) will be of mode ANY even if X is of mode FLOAT. When FALSE then ACOS(X) will be of mode FLOAT if and only if X is of mode FLOAT.

Variable TR FUNCTION CALL DEFAULT

default: [GENERAL] - FALSE means give up and call MEVAL, EXPR means assume Lisp fixed arg function. GENERAL, the default gives code good for MEXPRS and MLEXPRS but not MACROS. GENERAL assures variable bindings are correct in compiled code. In GENERAL mode, when translating F(X), if F is a bound variable, then it assumes that APPLY(F,[X]) is meant, and translates a such, with apropriate warning. There is no need to turn this off. With the default settings, no warning messages implies full compatibility of translated and compiled code with the macsyma interpreter.

Variable TR GEN TAGS

default: [FALSE] - If TRUE, TRANSLATE FILE generates a TAGS file for use by the text editor.

Variable TR NUMER

default: [FALSE] - If TRUE numer properties are used for atoms which have them, e.g. %PI.

Variable TR OPTIMIZE MAX LOOP

default: [100] - The maximum number of times the macro-expansion and optimization pass of the translator will loop in considering a form. This is to catch MACRO expansion errors, and non-terminating optimization properties.

Variable TR OUTPUT FILE DEFAULT

default: [TRLISP] - This is the second file name to be used for translated lisp output.

Variable TR PREDICATE BRAIN DAMAGE

default: [FALSE] - If TRUE, output possible multiple evaluations in an attempt to interface to the COMPARE package.

Variable TR SEMICOMPILE

default: [FALSE] - If TRUE TRANSLATE FILE and COMPFILE output forms which will be macroexpanded but not compiled into machine code by the lisp compiler.

Variable TR STATE VARS

default:

     [TRANSCOMPILE, TR_SEMICOMPILE,
     TR_WARN_UNDECLARED, TR_WARN_MEVAL, TR_WARN_FEXPR, TR_WARN_MODE,
     TR_WARN_UNDEFINED_VARIABLE, TR_FUNCTION_CALL_DEFAULT,
      TR_ARRAY_AS_REF,TR_NUMER]

The list of the switches that affect the form of the translated output. This information is useful to system people when trying to debug the translator. By comparing the translated product to what should have been produced for a given state, it is possible to track down bugs.

Variable TR TRUE NAME OF FILE BEING TRANSLATED

default: [FALSE] is bound to the quoted string form of the true name of the file most recently translated by TRANSLATE FILE.

Variable TR VERSION

- The version number of the translator.

Function TR WARNINGS GET ()

Prints a list of warnings which have been given by the translator during the current translation.

Variable TR WARN BAD FUNCTION CALLS

default: [TRUE] - Gives a warning when when function calls are being made which may not be correct due to improper declarations that were made at translate time.

Variable TR WARN FEXPR

default: [COMPFILE] - Gives a warning if any FEXPRs are encountered. FEXPRs should not normally be output in translated code, all legitimate special program forms are translated.

Variable TR WARN MEVAL

default: [COMPFILE] - Gives a warning if the function MEVAL gets called. If MEVAL is called that indicates problems in the translation.

Variable TR WARN MODE

default: [ALL] - Gives a warning when variables are assigned values inappropriate for their mode.

Variable TR WARN UNDECLARED

default: [COMPILE] - Determines when to send warnings about undeclared variables to the TTY.

Variable TR WARN UNDEFINED VARIABLE

default: [ALL] - Gives a warning when undefined global variables are seen.

Variable TR WINDY

default: [TRUE] - Generate "helpfull" comments and programming hints.

Variable UNDECLAREDWARN

default: [COMPFILE] - A switch in the Translator. There are four relevant settings: SETTING | ACTION ———————————————————— FALSE | never print warning messages. COMPFILE | warn when in COMPFILE TRANS-LATE | warn when in TRANSLATE and when TRANSLATE:TRUE ALL | warn in COMPFILE and TRANSLATE ———————————————————— Do MODE DECLARE(<variable>,ANY) to declare a variable to be a general macsyma variable (i.e. not limited to being FLOAT or FIXNUM). The extra work in declaring all your variables in code to be compiled should pay off.

Function COMPILE FILE (filename,&optional-outfile)

It takes filename which contains macsyma code, and translates this to lisp and then compiles the result. It returns a list of four files (the original file,translation, notes on translation and the compiled code).

Function DECLARE TRANSLATED (FN1,FN2..)

When translating a file of macsyma code to lisp, it is important for the translator to know which functions it sees in the file are to be called as translated or compiled functions, and which ones are just macsyma functions or undefined. Putting this declaration at the top of the file, lets it know that although a symbol does which does not yet have a lisp function value, will have one at call time. (MFUNCTION-CALL fn arg1 arg2.. ) is generated when the translator does not know fn is going to be a lisp function.

8 Input and Output

8.1 Introduction to Input and Output

8.2 FILES

- A file is simply an area on a particular storage device which contains data or text. The only storage devices which are used on the MC machine are disks and tapes. Files on the disks are figuratively grouped into "directories". A directory is just a list of all the files stored under a given user name. Do DESCRIBE(FILEOP); for details of how you may inspect your files using MACSYMA. Other commands which deal with files are: SAVE, FASSAVE, STORE, LOAD, LOADFILE, RESTORE, UNSTORE, STRINGOUT, BATCH, BATCON, DEMO, WRITEFILE, CLOSEFILE, DELFILE, REMFILE, and AP-PENDFILE.

8.3 PLAYBACK

It is possible to play back the input lines in a temporary scroll down window, and so not lose ones current work. This can be done by typing Function E. A numeric argument tells it the line number to start at, otherwise it will go back about 40 lines.

8.4 Definitions for Input and Output

Variable %

The last D-line computed by MACSYMA (whether or not it was printed out). (See also %%.)

Variable %%

The value of the last computation performed while in a (MACSYMA-BREAK). Also may be used in compound statements in the nth statement to refer to the value of the (n-1)th statement. E.g. F(N):=(INTEGRATE(X^N,X),SUBST(3,X,%%)- SUBST(2,X,%%)); is in essence equivalent to F(N):=BLOCK([%%], %%:INTE-GRATE(X^N,X), SUBST(3,X,%%)-SUBST(2,X,%%)); This will also work for communicating between the (n-1)th and nth (non-atomic) BLOCK statements.

Variable %EDISPFLAG

default: [FALSE] - if TRUE, MACSYMA displays %E to a negative exponent as a quotient, i.e. %E^-X as 1/%E^X.

Function %TH (i)

is the ith previous computation. That is, if the next expression to be computed is D(j) this is D(j-i). This is useful in BATCH files or for referring to a group of D expressions. For example, if SUM is initialized to 0 then FOR I:1 THRU 10 DO SUM:SUM+%TH(I) will set SUM to the sum of the last ten D expressions.

special symbol "?"

- As prefix to a function or variable name, signifies that the function or variable is a LISP token, not a MACSYMA token. Two question marks typed together, ??, will flush the current MACSYMA command line.

Variable ABSBOXCHAR

default: [!] is the character used to draw absolute value signs around expressions which are more than a single line high.

Function APPENDFILE (filename1, filename2, DSK, directory)

is like WRITEFILE(DSK,directory) but appends to the file whose name is specified by the first two arguments. A subsequent CLOSEFILE will delete the original file and rename the appended file.

Function BACKUP ()

To "back up" and see what you did, see PLAYBACK.

Function BATCH (file-specification)

reads in and evaluates MACSYMA command lines from a file - A facility for executing command lines stored on a disk file rather than in the usual on-line mode. This facility has several uses, namely to provide a reservoir for working command lines, for giving error-free demonstrations, or helping in organizing one’s thinking in complex problem-solving situations where modifications may be done via a text editor. A batch file consists of a set of MACSYMA command lines, each with its terminating ; or $, which may be further separated by spaces, carriage- returns, form-feeds, and the like. The BATCH function calls for reading in the command lines from the file one at a time, echoing them on the user console, and executing them in turn. Control is returned to the user console only when serious errors occur or when the end of the file is met. Of course, the user may quit out of the file-processing by typing control-G at any point. BATCH files may be created using a text editor or by use of the STRINGOUT command. Do DESCRIBE(STRINGOUT) for details DESCRIBE(FILE); and DESCRIBE(FILES); have additional information on how the file argument is interpreted, and files in general.

Variable BATCHKILL

default: [FALSE] if TRUE then the effect of all previous BATCH files is nullified because a KILL(ALL) and a RESET() will be done automatically when the next one is read in. If BATCHKILL is bound to any other atom then a KILL of the value of BATCHKILL will be done.

Function BATCHLOAD (file-specification)

Batches in the file silently without terminal output or labels.

Function BATCON (argument)

continues BATCHing in a file which was interrupted.

Variable BATCOUNT

default: [0] may be set to the number of the last expression BATCHed in from a file. Thus BATCON(BATCOUNT-1) will resume BATCHing from the expression before the last BATCHed in from before.

Variable BOTHCASES

default: [FALSE] if TRUE will cause MACSYMA to retain lower case text as well as upper case. Note, however, that the names of any MACSYMA special variables or functions must be typed in upper case. (For historical reasons this is also a function, but it should be used as a switch. Please use it as described here, not as a function.)

Variable CHANGE FILEDEFAULTS

default: [TRUE] on PDP10 systems, and FALSE elsewhere. Controls whether the user doing a LOADFILE or BATCH has his file defaults changed to the file LOADFILEd or BATCHed. The TRUE setting is for people who like DDT-style file defaulting. The FALSE setting is for people who like the conventions of other operating systems, who like LISP-style file defaulting, or who write packages which do LOADFILEs or BATCHes which should not interfere with their user’s file defaults.

Function CLOSEFILE (filename1, filename2)

closes a file opened by WRITEFILE and gives it the name filename1 filename2. (On a Lisp Machine one need only say CLOSEFILE();.) Thus to save a file consisting of the display of all input and output during some part of a session with MACSYMA the user issues a WRITEFILE, transacts with MACSYMA, then issues a CLOSEFILE. The user can also issue the PLAYBACK function after a WRITEFILE to save the display of previous transactions. (Note that what is saved this way is a copy of the display of expressions not the expressions themselves). To save the actual expression in internal form the SAVE function may be used. The expression can then be brought back into MACSYMA via the LOADFILE function. To save the expression in a linear form which may then be BATCHed in later, the STRINGOUT function is used.

Function COLLAPSE (expression)

collapses" its argument by causing all of its common (i.e. equal) subexpressions to share (i.e. use the same cells), thereby saving space. (COLLAPSE is a subroutine used by the OPTIMIZE command.) Thus, calling COLLAPSE may be useful before using FASSAVE or after loading in a SAVE file. You can collapse several expressions together by using COLLAPSE([expr1,...,exprN])$. Similarly, you can collapse the elements of the array A by doing COLLAPSE(LISTARRAY(’A))$.

Function CONCAT (arg1, arg2, ...)

evaluates its arguments and returns the concatenation of their values resulting in a name or a quoted string the type being given by that of the first argument. Thus if X is bound to 1 and D is unbound then CONCAT(X,2)="12" and CONCAT(D,X+1)=D2.

Function SCONCAT (arg1, arg2, ...)

evaluates its arguments and concatenates them into a string. Unlike CONCAT, the arguments do NOT need to be atoms. The result is a Common Lisp String.

     (C5) sconcat("xx[",3,"]:",expand((x+y)^3));
     (D5)   xx[3]:Y^3+3*X*Y^2+3*X^2*Y+X^3

The resulting string could be used in conjunction with print.

Variable CURSORDISP

default: [TRUE] If TRUE, causes expressions to be drawn by the displayer in logical sequence. This only works with a console which can do cursor movement. If FALSE, expressions are simply printed line by line. CURSORDISP is FALSE when a WRITEFILE is in effect.

Variable DIREC

- The value of this variable is the default file directory for SAVE, STORE, FASSAVE, and STRINGOUT. It is initialized to the user’s login name, if he has a disk directory, and to one of the USERSi directories otherwise. DIREC determines to what directory disk files will be written.

Function DISP (expr1,expr2, ...)

is like DISPLAY but only the value of the arguments are displayed rather than equations. This is useful for complicated arguments which don’t have names or where only the value of the argument is of interest and not the name.

Function DISPCON (tensor1,tensor2,...)

displays the contraction properties of the tensori as were given to DEFCON. DISP-CON(ALL) displays all the contraction properties which were defined.

Function DISPLAY (expr1, expr2, ...)

displays equations whose left side is expri unevaluated, and whose right side is the value of the expression centered on the line. This function is useful in blocks and FOR statements in order to have intermediate results displayed. The arguments to DISPLAY are usually atoms, subscripted variables, or function calls. (see the DISP function)

     (C1) DISPLAY(B[1,2]);
                                           2
                              B     = X - X
                               1, 2
     (D1)                            DONE

Variable DISPLAY2D

default: [TRUE] - if set to FALSE will cause the standard display to be a string (1dimensional) form rather than a display (2-dimensional) form. This may be of benefit for users on printing consoles who would like to conserve paper.

Variable DISPLAY FORMAT INTERNAL

default: [FALSE] - if set to TRUE will cause expressions to be displayed without being transformed in ways that hide the internal mathematical representation. The display then corresponds to what the INPART command returns rather than the PART command. Examples:

     User     PART       INPART
     a-b;      A - B     A + (- 1) B
                A            - 1
     a/b;       -         A B
                B
                            1/2
     sqrt(x);   SQRT(X)    X
               4 X        4
     X*4/3;    ---        - X
                3         3

Function DISPTERMS (expr)

displays its argument in parts one below the other. That is, first the operator of ’expr’ is displayed, then each term in a sum, or factor in a product, or part of a more general expression is displayed separately. This is useful if expr is too large to be otherwise displayed. For example if P1, P2, ... are very large expressions then the display program may run out of storage space in trying to display P1+P2+... all at once. However, DISPTERMS(P1+P2+...) will display P1, then below it P2, etc. When not using DISPTERMS, if an exponential expression is too wide to be displayed as A**B it will appear as EXPT(A,B) (or as NCEXPT(A,B) in the case of A^^B).

Variable DSKALL

default: [] If TRUE will cause values, functions, arrays, and rules to be written periodically onto the disk in addition to labelled expressions. TRUE is the default value whereas if DSKALL is FALSE then only labelled expresions will be written.

Variable ERROR SIZE

default: [20 for a display terminal, 10 for others]. controls the size of error messages. For example, let U:(C^D^E+B+A)/(COS(X-1)+1); . U has an error size of 24. So if ERROR SIZE has value < 24 then

     (C1) ERROR("The function", FOO,"doesn't like", U,"as input.");
     prints as:
     The function FOO doesn't like ERREXP1 as input.
     If ERROR_SIZE>24 then as:
                                      E
                                     D
                                    C   + B + A
     The function FOO doesn't like -------------- as input.
                                   COS(X - 1) + 1

Expressions larger than ERROR SIZE are replaced by symbols, and the symbols are set to the expressions. The symbols are taken from the user-settable list

     ERROR_SYMS:[ERREXP1,ERREXP2,ERREXP3]

. The default value of this switch might change depending on user experience. If you find the defaults either too big or two small for your tastes, send mail to MACSYMA.

Variable ERROR SYMS

default: [ERREXP1,ERREXP2,ERREXP3] - In error messages, expressions larger than ERROR SIZE are replaced by symbols, and the symbols are set to the expressions. The symbols are taken from the list ERROR SYMS, and are initially ERREXP1, ERREXP2, ERREXP3, etc. After an error message is printed, e.g. "The function FOO doesn’t like ERREXP1 as input.", the user can type ERREXP1; to see the expression. ERROR SYMS may be set by the user to a different set of symbols, if desired.

Function EXPT (A,B)

if an exponential expression is too wide to be displayed as A^B it will appear as EXPT(A,B) (or as NCEXPT(A,B) in the case of A^^B).

Variable EXPTDISPFLAG

default: [TRUE] - if TRUE, MACSYMA displays expressions with negative exponents using quotients e.g., X**(-1) as 1/X.

Function FASSAVE (args)

is similar to SAVE but produces a FASL file in which the sharing of subexpressions which are shared in core is preserved in the file created. Hence, expressions which have common subexpressions will consume less space when loaded back from a file created by FASSAVE rather than by SAVE. Files created with FASSAVE are reloaded using LOADFILE, just as files created with SAVE. FASSAVE returns a list of the form [<name of file>,<size of file in blocks>,...] where ... are the things saved. Warnings are printed out in the case of large files. FASSAVE may be used while a WRITEFILE is in progress.

Function FILEDEFAULTS ()

returns the current default filename, in whatever format is used by the specific Macsyma implementation. (See DESCRIBE(FILE) for what that format is.) This is the file specification used by LOADFILE, BATCH, and a number of other file-accessing commands. FILEDEFAULTS(’file) - sets the current filedefaults to "file". The argument to FILEDEFAULTS is evaulated as it is anticipated that the command will be used mainly in programs. The "file" need not be a real file, so one can use this function e.g. if one’s real purpose is to set only the "device" field to something, where one does not care about the settings of the other fields.

Variable FILENAME

default: [] - The value of this variable is the first name of the files which are generated by the automatic disk storage scheme. The default value is the first three characters of the user’s login name concatenated with the lowest unused integer, e.g. ECR1.

Function FILENAME MERGE ("filename1","filename2",...)

; merges together filenames. What this means is that it returns "filename1" except that missing components come from the corresponding components of "filename2", and if they are missing there, then from "filename3".

Variable FILENUM

default: [0] - The default second file name for files generated by SAVE, STRINGOUT, or FASSAVE if no file names are specified by the user. It is an integer, and is incremented by one each time a new file is written.

Variable FILE SEARCH

- this is a list of files naming directories to search by LOAD and a number of other functions. The default value of this is a list of the various SHARE directories used by Macsyma. FILE SEARCH("filename"); searches on those directories and devices specified by the FILE SEARCH variable, and returns the name of the first file it finds. This function is invoked by the LOAD function, which is why LOAD("FFT") finds and loads DSK:SHARE;FFT FASL. You may do FILE SEARCH:CONS("devdir",FILE SEARCH); to add other directories to the search rules, where "devdir" is a file naming only a device and a directory. In NIL Macsyma, for instance, if one had a directory for a personal library of Macsyma routines, "devdir" would be something like "USRD$:[myname.MAXLIB]".

Variable FILE STRING PRINT

default: [FALSE] on MC, [TRUE] elsewhere. If TRUE, filenames are output as strings; if FALSE, as lists. For example, the message when an out of core file is loaded into MACSYMA (e.g. the LIMIT package), appears on MC in list format as LIMIT FASL DSK MACSYM being loaded and in string format as: DSK:MACSYM;LIMIT FASL being loaded The string format is like the top level (DDT) file specifications.

Function FILE TYPE ("filename")

; returns FASL, LISP, or MACSYMA, depending on what kind of file it is. FASL means a compiled Lisp file, which normally has an extension of .VAS in NIL.

Function GRIND (arg)

prints out arg in a more readable format than the STRING command. It returns a Dline as value. The GRIND switch, default: [FALSE], if TRUE will cause the STRING, STRINGOUT, and PLAYBACK commands to use "grind" mode instead of "string" mode. For PLAYBACK, "grind" mode can also be turned on (for processing input lines) by specifying GRIND as an option.

Variable IBASE

default: [10] - the base for inputting numbers.

Variable INCHAR

default: [C] - the alphabetic prefix of the names of expressions typed by the user.

Function LDISP (expr1,expr2,...)

is like DISP but also generates intermediate labels.

Function LDISPLAY (expr1,expr2,...)

is like DISPLAY but also generates intermediate labels.

Variable LINECHAR

default: [E] - the alphabetic prefix of the names of intermediate displayed expressions.

Variable LINEDISP

default: [TRUE] - Allows the use of line graphics in the drawing of equations on those systems which support them (e.g. the Lisp Machine). This can be disabled by setting LINEDISP to FALSE. It is automatically disabled during WRITEFILE.

Variable LINEL

default: [] - the number of characters which are printed on a line. It is initially set by MACSYMA to the line length of the type of terminal being used (as far as is known) but may be reset at any time by the user. The user may have to reset it in DDT with :TCTYP as well.

Function LOAD ("filename")

; takes one argument, a filename represented as a "string" (i.e. inside quotation marks), or as list (e.g. inside square brackets), and locates and loads in the indicated file. If no directory is specified, it then searches the SHAREi directories and any other directories listed in the FILE SEARCH variable and loads the indicated file. LOAD("EIGEN") will load the eigen package without the need for the user to be aware of the details of whether the package was compiled, translated, saved, or fassaved, i.e. LOAD will work on both LOADILEable and BATCHable files. Note: LOAD will use BATCHLOAD if it finds the file is BATCHable (which means that it will BATCH the file in "silently" without terminal output or labels). Other MAC-SYMA commands to load in files are: LOADFILE, RESTORE, BATCH, and DEMO. Do DESCRIBE(command); for details. LOADFILE and RESTORE work for files written with SAVE; BATCH and DEMO for those files written with STRINGOUT or created as lists of commands with a text editor. If load can’t find the file, check the value FILE SEARCH to make sure that it contains an appropriate template.

     (C4) load("eigen");
     MACSYMA BUG: Unknown file type NIL
     Error: macsyma error
     Error signalled by MEVAL1.
     Broken at $LOAD.  Type :H for Help.
     MAXIMA>>:q
     By examining the file system we find the file is actually in
     /public/maxima/share/eigen.mc.   So we add that to the file_search
     path.   This can be done at start up (see sys-init.lsp in the
     system directory or, else it can be done and then the system resaved
     once it has been customized for local directories and pathnames.
     At lisp level we would do
     (in-package "MAXIMA")
     (setq $file_search ($append (list '(mlist)
             "/tmp/foo.mac" "/tmp/foo.mc") $file_search))
     and at maxima level:

     (C5) file_search:append(["/public/maxima/share/foo.mc"],
             file_search)$
     (C6) load("eigen");
     batching /usr/public/maxima/share/eigen.mc
     (D6)          #/public/maxima/share/eigen.mc
     (C7) eigenvalues(matrix([a,b],[c,d]));
            2     2
            - SQRT(D  - 2 A D + 4 B C + A ) + D + A
     (D7) [[---------------------------------------,
       2
           2    2
     SQRT(D  - 2 A D + 4 B C + A ) + D + A
     -------------------------------------], [1, 1]]
       2
     (C8)

Function LOADFILE (filename)

loads a file as designated by its arguments. This function may be used to bring back quantities that were stored from a prior MACSYMA session by use of the SAVE or STORE functions. Specify the pathname as on your operating system. For unix this would be "/home/wfs/foo.mc" for example.

Variable LOADPRINT

default: [TRUE] - governs the printing of messages accompanying loading of files. The following options are available: TRUE means always print the message; ’LOADFILE means print only when the LOADFILE command is used; ’AUTOLOAD means print only when a file is automatically loaded in (e.g. the integration file SIN FASL); FALSE means never print the loading message.

Function NOSTRING (arg)

displays all input lines when playing back rather than STRINGing them. If arg is GRIND then the display will be in a more readable format. One may include any number of options as in PLAYBACK([5,10],20,TIME,SLOW).

Variable OBASE

default: [10] the base for display of numbers.

Variable OUTCHAR

default: [D] - the alphabetic prefix of the names of outputted expressions.

Variable PACKAGEFILE

default:[FALSE] - Package designers who use SAVE, FASSAVE, or TRANSLATE to create packages (files) for others to use may want to set PACKAGEFILE:TRUE$ to prevent information from being added to MACSYMA’s information-lists (e.g. VAL-UES, FUNCTIONS) except where necessary when the file is loaded in. In this way, the contents of the package will not get in the user’s way when he adds his own data. Note that this will not solve the problem of possible name conflicts. Also note that the flag simply affects what is output to the package file. Setting the flag to TRUE is also useful for creating MACSYMA init files.

Variable PARSEWINDOW

default:[10] - the maximum number of "lexical tokens" that are printed out on each side of the error-point when a syntax (parsing) error occurs. This option is especially useful on slow terminals. Setting it to -1 causes the entire input string to be printed out when an error occurs.

Variable PFEFORMAT

default: [FALSE] - if TRUE will cause rational numbers to display in a linear form and denominators which are integers to display as rational number multipliers.

Function PRINT (exp1, exp2, ...)

evaluates and displays its arguments one after the other "on a line" starting at the leftmost position. If expi is unbound or is preceded by a single quote or is enclosed in "s then it is printed literally. For example, PRINT("THE VALUE OF X IS ",X). The value returned by PRINT is the value of its last argument. No intermediate lines are generated. (For "printing" files, see the PRINTFILE function.)

Function READ (string1, ...)

prints its arguments, then reads in and evaluates one expression. For example: A:READ("ENTER THE NUMBER OF VALUES").

Function READONLY (string1,...)

prints its arguments, then reads in an expression (which in contrast to READ is not evaluated).

Function REVEAL (exp,depth)

will display exp to the specified integer depth with the length of each part indicated. Sums will be displayed as Sum(n) and products as Product(n) where n is the number of subparts of the sum or product. Exponentials will be displayed as Expt.

     (C1) INTEGRATE(1/(X^3+2),X)$
     (C2) REVEAL(%,2);
     (D2)                    Negterm + Quotient + Quotient
     (C3) REVEAL(D1,3);
                                          ATAN         LOG
     (D3)                 - Quotient + ---------- + ----------
                                       Product(2)   Product(2)

Variable RMXCHAR

default: []] - The character used to display the (right) delimiter of a matrix (see also LMXCHAR).

Function SAVE ([optional file spec],arg1, arg2,...,argi)

saves quantities described by its arguments on disk and keeps them in core also. The optional file spec is in the form NAME1,NAME2,DSK,DIRECTORY and is enclosed in square brackets, e.g. [JAMU,FUNCS,DSK,USERS2]. Do DESCRIBE(FILES) for more details on file specifications. If you omit the file spec, MACSYMA will select a file name for you consisting of the first three letters of your login name followed by a digit (the lowest digit which will give a unique file name). The second file name will be a number. The arg’s are the expressions to be SAVEd. ALL is the simplest, but note that saving ALL will save the entire contents of your MACSYMA, which in the case of a large computation may result in a file which is too large to be reloaded. VALUES, FUNCTIONS, or any other items on the INFOLISTS (do DESCRIBE(INFOLISTS); for the list) may be SAVEd, as may functions and variables by name. C and D lines may also be saved, but it is better to give them explicit names, which may be done in the command line, e.g. SAVE(RES1=D15); Files saved with SAVE should be reloaded with LOADFILE. SAVE returns a list of the form [<name of file>,<size of file in blocks>,...] where ... are the things saved. Warnings are printed out in the case of large files, or if an empty file is accidently generated. It is possible to do a SAVE when a WRITEFILE is in progress.

Variable SAVEDEF

default: [TRUE] - if TRUE will cause the MACSYMA version of a user function to remain when the function is TRANSLATEd. This permits the definition to be displayed by DISPFUN and allows the function to be edited. If SAVEDEF is FALSE, the names of translated functions are removed from the FUNCTIONS list.

Function SHOW (exp)

will display exp with the indexed objects in it shown having covariant indices as subscripts,contravariant indices as superscripts. The derivative indices will be displayed as subscripts, separated from the covariant indices by a comma.

Function SHOWRATVARS (exp)

returns a list of the RATVARS (CRE variables) of exp.

Variable STARDISP

default: [FALSE] - if TRUE will cause multiplication to be displayed explicitly with an * between operands.

Function STRING (expr)

converts expr to MACSYMA’s linear notation (similar to FORTRAN’s) just as if it had been typed in and puts expr into the buffer for possible editing (in which case expr is usually Ci) The STRING’ed expression should not be used in a computation.

Function STRINGOUT (args)

will output an expression to a file in a linear format. Such files are then used by the BATCH or DEMO commands. STRINGOUT(file-specification, A1, A2, ...) outputs to a file given by file-specification ([filename1,filename2,DSK, directory]) the values given by A1,A2,.. in a MACSYMA readable format. The file-specification may be omitted, in which case the default values will be used. The Ai are usually C labels or may be INPUT meaning the value of all C labels. Another option is to make ai FUNCTIONS which will cause all of the user’s function definitions to be strungout (i.e. all those retrieved by DISPFUN(ALL)). Likewise the ai may be VALUES, and all the variables to which the user has assigned values will be strungout. ai may also be a list [m,n] which means to stringout all labels in the range m through n inclusive. This function may be used to create a file of FORTRAN statements by doing some simple editing on the strungout expressions. If the GRIND switch is set to TRUE, then STRINGOUT will use GRIND format instead of STRING format. Note: a STRINGOUT may be done while a WRITEFILE is in progress.

Function SYSTEM(command)

Execute COMMAND as a subprocess. The command will be passed to the default shell for execution. System is not supported by all operating systems, but generally exists in the unix environment. if hist is a list of frequencies which you wish to plot as a bar graph using xgraph.

           with_stdout("_hist.out",
                for i:1 thru length(hist) do (
                  print(i,hist[i]))),
           system("xgraph -bar -brw .7 -nl < _hist.out")

In order to make the plot be done in the background (returning control to maxima) and remove the temporary file after it is done do:

     system("(xgraph -bar -brw .7 -nl < _hist.out;  rm -f _hist.out)&")

Variable TTYOFF

default: [FALSE] - if TRUE stops printing output to the console.

macro WITH STDOUT(file,stmt1,stmt2,...)

Opens file and then evaluates stmt1, stmt2, .... Any printing to standard output goes to the file instead of the terminal. It always returns FALSE.

     mygnuplot(f,var,range,number_ticks):=
      block([numer:true],
      with_stdout("/tmp/gnu",
        for x:range[1] thru range[2] step
                     (range[2]-range[1])/number_ticks
           do (print(x,at(f,var=x)))),
      system("echo \"set data style lines; set title '",
             f,"' ;plot '/tmp/gnu'
     ;pause 10 \" | gnuplot"));
     (C8) with_stdout("/home/wfs/joe",

      n:10,
      for i:8 thru n
        do(print("factorial(",i,") gives ",i!)));
(D8)       FALSE
(C9) system("cat /home/wfs/joe");
factorial( 8 ) gives  40320
factorial( 9 ) gives  362880
factorial( 10 ) gives  3628800
(D9)         0

Function WRITEFILE (DSK, directory)

opens up a file for writing. On a Lisp Machine one uses WRITEFILE("filename"). All interaction between the user and MACSYMA is then recorded in this file, just as it is on the console. Such a file is a transcript of the session, and is not reloadable or batchable into MACSYMA again. (See also CLOSEFILE.)

80 Maxima Manual

9 Program Flow

9.1 Introduction to Program Flow

MACSYMA provides a DO loop for iteration, as well as more primitive constructs such as GO.

9.2 Definitions for Program Flow

Variable BACKTRACE

default: [] (when DEBUGMODE:ALL has been done) has as value a list of all functions currently entered.

special operator DO

- The DO statement is used for performing iteration. Due to its great generality the DO statement will be described in two parts. First the usual form will be given which is analogous to that used in several other programming languages (FORTRAN, ALGOL, PL/I, etc.); then the other features will be mentioned. 1. There are three variants of this form that differ only in their terminating conditions. They are:

• (a) FOR variable : initial-value STEP increment THRU limit DO body

• (b) FOR variable : initial-value STEP increment WHILE condition DO body • (c) FOR variable : initial-value STEP increment UNLESS condition DO body

(Alternatively, the STEP may be given after the termination condition or limit.) The initial-value, increment, limit, and body can be any expressions. If the increment is 1 then "STEP 1" may be omitted. The execution of the DO statement proceeds by first assigning the initial-value to the variable (henceforth called the control-variable). Then: (1) If the control-variable has exceeded the limit of a THRU specification, or if the condition of the UNLESS is TRUE, or if the condition of the WHILE is FALSE then the DO terminates. (2) The body is evaluated. (3) The increment is added to the control-variable. The process from (1) to (3) is performed repeatedly until the termination condition is satisfied. One may also give several termination conditions in which case the DO terminates when any of them is satisfied. In general the THRU test is satisfied when the control-variable is greater than the limit if the increment was non-negative, or when the control-variable is less than the limit if the increment was negative. The increment and limit may be non-numeric expressions as long as this inequality can be determined. However, unless the increment is syntactically negative (e.g. is a negative number) at the time the DO statement is input, MACSYMA assumes it will be positive when the DO is executed. If it is not positive, then the DO may not terminate properly. Note that the limit, increment, and termination condition are evaluated each time through the loop. Thus if any of these involve much computation, and yield a result that does not change during all the executions of the body, then it is more efficient to set a variable to their value prior to the DO and use this variable in the DO form. The value normally returned by a DO statement is the atom DONE, as every statement in MACSYMA returns a value. However, the function RETURN may be used inside the body to exit the DO prematurely and give it any desired value. Note however that a RETURN within a DO that occurs in a BLOCK will exit only the DO and not the BLOCK. Note also that the GO function may not be used to exit from a DO into a surrounding BLOCK. The control-variable is always local to the DO and thus any variable may be used without affecting the value of a variable with the same name outside of the DO. The control-variable is unbound after the DO terminates.

(C1)   FOR A:-3 THRU 26 STEP 7 DO LDISPLAY(A)$
(E1)          A = -3
(E2)          A =  4
(E3)          A = 11
(E4)          A = 18
(E5)          A = 25

The function LDISPLAY generates intermediate labels; DISPLAY does not.

(C6)   S:0$
(C7)   FOR I:1 WHILE I<=10 DO S:S+I;
(D7)          DONE
(C8)   S;
(D8)          55

Note that the condition in C7 is equivalent to UNLESS I > 10 and also THRU 10

(C9)   SERIES:1$
(C10)  TERM:EXP(SIN(X))$
(C11)  FOR P:1 UNLESS P>7 DO
          (TERM:DIFF(TERM,X)/P,
          SERIES:SERIES+SUBST(X=0,TERM)*X^P)$
(C12)   SERIES;
                7    6     5    4    2
(D12)          X    X     X    X    X
               -- - --- - -- - -- + -- + X + 1
               96   240   15   8    2
which gives 8 terms of the Taylor series for e^sin(x).
(C13) POLY:0$
(C14) FOR I:1 THRU 5 DO
        FOR J:I STEP -1 THRU 1 DO
           POLY:POLY+I*X^J$
(C15) POLY;
              5      4       3       2
(D15)      5 X  + 9 X  + 12 X  + 14 X  + 15 X
(C16) GUESS:-3.0$
(C17) FOR I:1 THRU 10 DO (GUESS:SUBST(GUESS,X,.5*(X+10/X)),
         IF ABS(GUESS^2-10)<.00005 THEN RETURN(GUESS));
(D17)                  - 3.1622807

This example computes the negative square root of 10 using the Newton- Raphson iteration a maximum of 10 times. Had the convergence criterion not been met the value returned would have been "DONE". Additional Forms of the DO Statement Instead of always adding a quantity to the control-variable one may sometimes wish to change it in some other way for each iteration. In this case one may use "NEXT expression" instead of "STEP increment". This will cause the control-variable to be set to the result of evaluating expression each time through the loop.

     (C1)  FOR COUNT:2 NEXT 3*COUNT THRU 20
              DO DISPLAY(COUNT)$
                         COUNT = 2
                         COUNT = 6
                         COUNT = 18

As an alternative to FOR variable:value ...DO... the syntax FOR variable FROM value ...DO... may be used. This permits the "FROM value" to be placed after the step or next value or after the termination condition. If "FROM value" is omitted then 1 is used as the initial value. Sometimes one may be interested in performing an iteration where the control-variable is never actually used. It is thus permissible to give only the termination conditions omitting the initialization and updating information as in the following example to compute the square-root of 5 using a poor initial guess.

     (C1) X:1000
     (C2)  THRU 10 WHILE X#0.0 DO X:.5*(X+5.0/X)$
     (C3) X;
     (D3)               2.236068

If it is desired one may even omit the termination conditions entirely and just give "DO body" which will continue to evaluate the body indefinitely. In this case the function RETURN should be used to terminate execution of the DO.

     (C1) NEWTON(F,GUESS):=
       BLOCK([NUMER,Y],
             LOCAL(DF),
             NUMER:TRUE,
             DEFINE(DF(X),DIFF(F(X),X)),
             DO (Y:DF(GUESS),
                 IF Y=0.0 THEN ERROR("Derivative at:",GUESS," is zero."),
                 GUESS:GUESS-F(GUESS)/Y,
                 IF ABS(F(GUESS))<5.0E-6 THEN RETURN(GUESS)))$
     (C2) SQR(X):=X^2-5.0$
     (C3) NEWTON(SQR,1000);
     (D3)                    2.236068

(Note that RETURN, when executed, causes the current value of GUESS to be returned as the value of the DO. The BLOCK is exited and this value of the DO is returned as the value of the BLOCK because the DO is the last statement in the block.) One other form of the DO is available in MACSYMA. The syntax is:

     FOR variable IN list [end-tests] DO body

The members of the list are any expressions which will successively be assigned to the variable on each iteration of the body. The optional end-tests can be used to terminate execution of the DO; otherwise it will terminate when the list is exhausted or when a RETURN is executed in the body. (In fact, list may be any non-atomic expression, and successive parts are taken.)

     (C1)  FOR F IN [LOG, RHO, ATAN] DO LDISP(F(1))$
     (E1)                                  0
     (E2)                                RHO(1)
                                          %PI
     (E3)                                 ---
                                           4
     (C4) EV(E3,NUMER);
     (D4)                             0.78539816

Function ERRCATCH (exp1, exp2, ...)

evaluates its arguments one by one and returns a list of the value of the last one if no error occurs. If an error occurs in the evaluation of any arguments, ERRCATCH "catches" the error and immediately returns [] (the empty list). This function is useful in BATCH files where one suspects an error might occur which would otherwise have terminated the BATCH if the error weren’t caught.

Variable ERREXP

default: [ERREXP] When an error occurs in the course of a computation, MACSYMA prints out an error message and terminates the computation. ERREXP is set to the offending expression and the message "ERREXP contains the offending expression" is printed. The user can then type ERREXP; to see this and hopefully find the problem.

Function ERROR (arg1, arg2, ...)

will evaluate and print its arguments and then will cause an error return to top level MACSYMA or to the nearest enclosing ERRCATCH. This is useful for breaking out of nested functions if an error condition is detected, or wherever one can’t type control-^. The variable ERROR is set to a list describing the error, the first of it being a string of text, and the rest the objects in question. ERRORMSG(); is the preferred way to see the last error message. ERRORFUN default: [FALSE] - if set to the name of a function of no arguments will cause that function to be executed whenever an error occurs. This is useful in BATCH files where the user may want his MACSYMA killed or his terminal logged out if an error occurs. In these cases ERRORFUN would be set to QUIT or LOGOUT.

Variable ERRORFUN

default: [FALSE] - if set to the name of a function of no arguments will cause that function to be executed whenever an error occurs. This is useful in BATCH files where the user may want his MACSYMA killed or his terminal logged out if an error occurs. In these cases ERRORFUN would be set to QUIT or LOGOUT.

Function ERRORMSG ()

reprints the last error message. This is very helpful if you are using a display console and the message has gone off the screen. The variable ERROR is set to a list describing the error, the first of it being a string of text, and the rest the objects in question. TTYINTFUN:LAMBDA([],ERRORMSG(),PRINT(""))$ will set up the user-interrupt character (^U) to reprint the message.

special operator FOR

- Used in iterations, do DESCRIBE("DO"); for a description of MACSYMA’s iteration facilities.

Function GO (tag)

is used within a BLOCK to transfer control to the statement of the block which is tagged with the argument to GO. To tag a statement, precede it by an atomic argument as another statement in the BLOCK. For example:

     BLOCK([X],X:1,LOOP,X+1,...,GO(LOOP),...)

. The argument to GO must be the name of a tag appearing in the same BLOCK. One cannot use GO to transfer to tag in a BLOCK other than the one containing the GO.

special operator IF

- The IF statement is used for conditional execution. The syntax is:

         IF condition THEN expression1 ELSE expression2.

The result of an IF statement is expression1 if condition is true and expression2 if it is false. expression1 and expression2 are any MACSYMA expressions (including nested IF statements), and condition is an expression which evaluates to TRUE or FALSE and is composed of relational and logical operators which are as follows:

     Operator name       Symbol      Type
     greater than        >           relational infix
     equal to            = , EQUAL   "  "
     not equal to        #           "  "
     less than           <           "  "
     greater than        >=
       or equal to                   "  "
     less than           <=
       or equal to                   "  "
     and                 AND         logical infix
     or                  OR          "  "
     not                 NOT         logical prefix

Function LISPDEBUGMODE ()

LISPDEBUGMODE(); DEBUGPRINTMODE(); and DEBUG(); make available to the user debugging features used by systems programmers. These tools are powerful, and although some conventions are different from the usual macsyma level it is felt their use is very intuitive. [Some printout may be verbose for slow terminals, there are switches for controlling this.] These commands were designed for the user who must debug translated macsyma code, as such they are a boon. See MACDOC;TRDEBG USAGE for more information. For more help, consult GJC.

Function MAP (fn, exp1, exp2, ...)

returns an expression whose leading operator is the same as that of the expi but whose subparts are the results of applying fn to the corresponding subparts of the expi. Fn is either the name of a function of n arguments (where n is the number of expi) or is a LAMBDA form of n arguments. MAPERROR[TRUE] - if FALSE will cause all of the mapping functions to (1) stop when they finish going down the shortest expi if not all of the expi are of the same length and (2) apply fn to [exp1, exp2,...] if the expi are not all the same type of object. If MAPERROR is TRUE then an error message will be given in the above two instances. One of the uses of this function is to MAP a function (e.g. PARTFRAC) onto each term of a very large expression where it ordinarily wouldn’t be possible to use the function on the entire expression due to an exhaustion of list storage space in the course of the computation.

     (C1) MAP(F,X+A*Y+B*Z);
     (D1)                        F(B Z) + F(A Y) + F(X)
     (C2) MAP(LAMBDA([U],PARTFRAC(U,X)),X+1/(X^3+4*X^2+5*X+2));
                                1       1        1
     (D2)                     ----- - ----- + -------- + X
                              X + 2   X + 1          2
                                              (X + 1)
     (C3) MAP(RATSIMP, X/(X^2+X)+(Y^2+Y)/Y);
                                           1
     (D3)                            Y + ----- + 1
                                         X + 1
     (C4) MAP("=",[A,B],[-0.5,3]);
     (D4)                          [A = - 0.5, B = 3]

Function MAPATOM (expr)

is TRUE if and only if expr is treated by the MAPping routines as an "atom", a unit. "Mapatoms" are atoms, numbers (including rational numbers), and subscripted variables.

Variable MAPERROR

default: [TRUE] - if FALSE will cause all of the mapping functions, for example

     MAP(fn,exp1,exp2,...))

to (1) stop when they finish going down the shortest expi if not all of the expi are of the same length and (2) apply fn to [exp1, exp2,...] if the expi are not all the same type of object. If MAPERROR is TRUE then an error message will be given in the above two instances.

Function MAPLIST (fn, exp1, exp2, ...)

yields a list of the applications of fn to the parts of the expi. This differs from MAP(fn,exp1,exp2,...) which returns an expression with the same main operator as expi has (except for simplifications and the case where MAP does an APPLY). Fn is of the same form as in MAP.

Variable PREDERROR

default: [TRUE] - If TRUE, an error message is signalled whenever the predicate of an IF statement or an IS function fails to evaluate to either TRUE or FALSE.

If FALSE, UNKNOWN is returned instead in this case. The PREDERROR:FALSE mode is not supported in translated code.

Function RETURN (value)

may be used to exit explicitly from a BLOCK, bringing its argument. Do DE-SCRIBE(BLOCK); for more information.

Function SCANMAP (function,exp)

recursively applies function to exp, in a "top down" manner. This is most useful when "complete" factorization is desired, for example:

     (C1) EXP:(A^2+2*A+1)*Y + X^2$
     (C2) SCANMAP(FACTOR,EXP);
                                         2      2
     (D2)                         (A + 1)  Y + X

Note the way in which SCANMAP applies the given function FACTOR to the constituent subexpressions of exp; if another form of exp is presented to SCANMAP then the result may be different. Thus, D2 is not recovered when SCANMAP is applied to the expanded form of exp:

     (C3) SCANMAP(FACTOR,EXPAND(EXP));
                                2                  2
     (D3)                      A  Y + 2 A Y + Y + X

Here is another example of the way in which SCANMAP recursively applies a given function to all subexpressions, including exponents:

     (C4) EXPR : U*V^(A*X+B) + C$
     (C5) SCANMAP('F, EXPR);
                         F(F(F(A) F(X)) + F(B))
     (D5) F(F(F(U) F(F(V)                      )) + F(C))

SCANMAP(function,expression,BOTTOMUP) applies function to exp in a "bottomup" manner. E.g., for undefined F,

     SCANMAP(F,A*X+B) ->
        F(A*X+B) -> F(F(A*X)+F(B)) -> F(F(F(A)*F(X))+F(B))
     SCANMAP(F,A*X+B,BOTTOMUP) -> F(A)*F(X)+F(B)
         -> F(F(A)*F(X))+F(B) ->
          F(F(F(A)*F(X))+F(B))

In this case, you get the same answer both ways.

Function THROW (exp)

evaluates exp and throws the value back to the most recent CATCH. THROW is used with CATCH as a structured nonlocal exit mechanism.

88 Maxima Manual

10 Plotting

10.1 Definitions for Plotting

Function CLOSEPS ()

This should usually becalled at the end of a sequence of plotting commands. It closes the current output stream PSTREAM, and sets it to nil. It also may be called at the start of a plot, to ensure pstream is closed if it was open. All commands which write to pstream, open it if necessary. CLOSEPS is separate from the other plotting commands, since we may want to plot 2 ranges or superimpose several plots, and so must keep the stream open.

Variable PLOT OPTIONS

Members of this list indicate defaults for plotting. They may be altered using SET PLOT OPTION

     [X, - 3, 3]
     [Y, - 3, 3]

are the x range and y range respectively. [TRANSFORM XY, FALSE] if not false, should be the output of

     make_transform([x,y,z], [f1(x,y,z),f2(x,y,z),f3(x,y,z)])

which produces a transformation from 3 space to 3 space, which will be applied to the graph. A built in one is polar xy which gives the same as

     make_transform([r,th,z],[r*cos(th),r*sin(th),z])

[RUN VIEWER,TRUE] if not false, means run the viewer software dont just output a data file. [GRID,30,30] means plot3d should divide the x range into 30 intervals and similarly the y range. [COLOUR Z,false] aplies to colouring done with plot format ps. [PLOT FORMAT,ZIC] is for plot3d and currently ZIC, GNUPLOT, PS, and GE-OMVIEW are supported. There are good quality public domain viewers for these formats. They are izic, gnuplot, ghostview, and geomview. izic is available by ftp from zenon.inria.fr. Contact one of {fournier,kajler,mourrain}@sophia.inria.fr. It has beautiful colour gouraud shading, and very fast wireframe. It runs on X windows. geomview is from the geometry institute in minnesota, and is available by anonymous ftp from there. It is currently not quite as pretty as izic, but provides excellent support for multiple objects and multiple lights. gnuplot is everywhere as is ghostview. We also provide mgnuplot a tcl interface for gnuplot, which lets you rotate the plots using the mouse and a scale.

Function PLOT3D (expr,xrange,yrange,...,options,..) Function PLOT3D ([expr1,expr2,expr3],xrange,yrange,...,options,..)

     plot3d(2^(-u^2+v^2),[u,-5,5],[v,-7,7]);

would plot z = 2^(-u^2+v^2) with u and v varying in [-5,5] and [-7,7] respectively, and with u on the x axis, and v on the y axis. An example of the second pattern of arguments is

     plot3d([cos(x)*(3+y*cos(x/2)),sin(x)*(3+y*cos(x/2)),y*sin(x/2)],
        [x,-%pi,%pi],[y,-1,1],['grid,50,15])

which will plot a moebius band, parametrized by the 3 expressions given as the first argument to plot3d. An additional optional argument [grid,50,15] gives the grid number of rectangles in the x direction and y direction.

     /* REal part of z ^1/3 */
      plot3d(r^.33*cos(th/3),[r,0,1],[th,0,6*%pi],
          ['grid,12,80],['plot_format,ps],
          ['transform_xy,polar_to_xy],['view_direction,1,1,1.4],
          ['colour_z,true])

Here the View direction indicates the direction from which we take a projection. We actually do this from infinitely far away, but parallel to the line from view direction to the origin. This is currently only used in ’ps’ plot format, since the other viewers allow interactive rotating of the object. Another example is a moebius band:

     plot3d([cos(x)*(3+y*cos(x/2)),
               sin(x)*(3+y*cos(x/2)),y*sin(x/2)],
                [x,-%pi,%pi],[y,-1,1],['grid,50,15]);

or a klein bottle:

     plot3d([5*cos(x)*(cos(x/2)*cos(y)+sin(x/2)*sin(2*y)+3.0) - 10.0,
               -5*sin(x)*(cos(x/2)*cos(y)+sin(x/2)*sin(2*y)+3.0),
                5*(-sin(x/2)*cos(y)+cos(x/2)*sin(2*y))],
                [x,-%pi,%pi],[y,-%pi,%pi],['grid,40,40])

or a torus

     plot3d([cos(y)*(10.0+6*cos(x)),
                sin(y)*(10.0+6*cos(x)),
                -6*sin(x)], [x,0,2*%pi],[y,0,2*%pi],
                 ['grid,40,40])

We can output to gnplot too:

      plot3d(2^(x^2-y^2),[x,-1,1],[y,-2,2],[plot_format,gnuplot])

Function PLOT2D PS (expr,range)

writes to pstream a sequence of postscript commands which plot EXPR for RANGE. EXPR should be an expression of 1 variable. RANGE should be of the form [variable,min,max] over which to plot expr. see CLOSEPS.

Function PLOT2D (expr,range,...,options,..) Function PLOT2D ([expr1,expr2,..,exprn],xrange,...,options,..)

EXPR is an expression to be plotted on y axis as a function of 1 variable. RANGE is of the form [var,min,max] and expr is assumed to be an expression to be plotted against VAR. In the second form of the function a list of expressions may be given to plot against. Truncation in the y direction will be performed, for the default y range. It may be specified as an option or using SET PLOT OPTION.

     plot2d(sin(x),[x,-5,5]);
     plot2d(sec(x),[x,-2,2],[y,-20,20],[nticks,200]);

Function xgraph curves(list)

graphs the list of ‘point sets’ given in list by using xgraph. A point set may be of the form

     [x0,y0,x1,y1,x2,y2,...] or
     [[x0,y0],[x1,y1],....]

A point set may also contain symbols which give labels or other information.

      xgraph_curves([pt_set1,pt_set2,pt_set3]);

would graph the three point sets as three curves.

     pt_set:append(["NoLines: True","LargePixels: true"],
             [x0,y0,x1,y1,...])

would make the point set [and subsequent ones], have no lines between points, and to use large pixels. See the man page on xgraph for more options to specify.

     pt_set:append([concat("\"","x^2+y")],[x0,y0,x1,y1,...])

would make there be a "label" of "x^2+y" for this particular point set. The " at the beginning is what tells xgraph this is a label.

     pt_set:append([concat("TitleText: Sample Data")],[x0,...])

would make the main title of the plot be "Sample Data" instead of "Maxima PLot". A temporary file ‘xgraph-out’ is used.

Function SET PLOT OPTION (option)

option is of the format of one of the elements of the PLOT OPTIONS list. Thus

     SET_PLOT_OPTION([grid,30,40])

would change the default grid used by plot3d. Note that if the symbol grid has a value, then you should quote it here:

     SET_PLOT_OPTION(['grid,30,40])

so that the value will not be substituted.

Function PSDRAW CURVE (ptlist)

Draws a curve connecting the points in PTLIST. The latter may be of the form [x0,y0,x1,y1,...] or [[x0,y0],[x1,y1],...] The function JOIN is handy for taking a list of x’s and a list of y’s and splicing them together. PSDRAW CURVE simply invokes the more primitive function PSCURVE. Here is the definition:

     (defun $psdraw_curve (lis)
       (p "newpath")
       ($pscurve lis)
       (p "stroke"))

?DRAW2D may also be used to produce a list

      points1:?draw2d(1/x,[.05,10],.03)

92 Maxima Manual

Function PSCOM (com)

COM will be inserted in the poscript file eg

            pscom("4.5 72 mul 5.5 72 mul translate  14 14 scale");

11 Floating Point

11.1 Definitions for Floating Point

Function BFFAC (exp,n)

BFLOAT version of the Factorial (shifted Gamma) function. The 2nd argument is how many digits to retain and return, it’s a good idea to request a couple of extra. This function is available by doing LOAD(BFFAC); .

Variable ALGEPSILON

The default value is 10^-8. The value of ALGEPSILON is used by ALGSYS.

Function BFLOAT (X)

converts all numbers and functions of numbers to bigfloat numbers. Setting FP-PREC[16] to N, sets the bigfloat precision to N digits. If FLOAT2BF[FALSE] is FALSE a warning message is printed when a floating point number is converted into a bigfloat number (since this may lead to loss of precision).

Function BFLOATP (exp)

is TRUE if exp is a bigfloat number else FALSE.

Function BFPSI (n,z,fpprec)

gives polygammas of real arg and integer order. For digamma, BFPSI0(z,fpprec) is more direct. Note -BFPSI0(1,fpprec) provides BFLOATed %GAMMA. To use this do LOAD(BFFAC);

Variable BFTORAT

default: [FALSE] controls the conversion of bfloats to rational numbers. If

     BFTORAT:FALSE

RATEPSILON will be used to control the conversion (this results in relatively small rational numbers). If

     BFTORAT:TRUE

, the rational number generated will accurately represent the bfloat.

Variable BFTRUNC

default: [TRUE] causes trailing zeroes in non-zero bigfloat numbers not to be displayed. Thus, if BFTRUNC:FALSE, BFLOAT(1); displays as 1.000000000000000B0. Otherwise, this is displayed as 1.0B0.

Function CBFAC (z,fpprec)

a factorial for complex bfloats. It may be used by doing LOAD(BFAC); For more details see share2/bfac.usg.

Function FLOAT (exp)

converts integers, rational numbers and bigfloats in exp to floating point numbers. It is also an EVFLAG, FLOAT causes non-integral rational numbers and bigfloat numbers to be converted to floating point.

Variable FLOAT2BF

default: [FALSE] if FALSE, a warning message is printed when a floating point number is converted into a bigfloat number (since this may lead to loss of precision).

Function FLOATDEFUNK

- is a utility for making floating point functions from mathematical expression. It will take the input expression and FLOAT it, then OPTIMIZE it, and then insert MODE DECLAREations for all the variables. This is THE way to use ROMBERG, PLOT2, INTERPOLATE, etc. e.g. EXP:some-hairy-macsyma-expression;

     FLOATDEFUNK('F,['X],EXP);

will define the function F(X) for you. (Do PRINTFILE(MCOMPI,DOC,MAXDOC); for more details.)

Function FLOATNUMP (exp)

is TRUE if exp is a floating point number else FALSE.

Variable FPPREC

default: [16] - Floating Point PRECision. Can be set to an integer representing the desired precision.

Variable FPPRINTPREC

default: [0] - The number of digits to print when printing a bigfloat number, making it possible to compute with a large number of digits of precision, but have the answer printed out with a smaller number of digits. If FPPRINTPREC is 0 (the default), or >= FPPREC, then the value of FPPREC controls the number of digits used for printing. However, if FPPRINTPREC has a value between 2 and FPPREC-1, then it controls the number of digits used. (The minimal number of digits used is 2, one to the left of the point and one to the right. The value 1 for FPPRINTPREC is illegal.)

Function ?ROUND (x,&optional-divisor)

round the floating point X to the nearest integer. The argument must be a regular system float, not a bigfloat. The ? beginning the name indicates this is normal common lisp function.

          (C3) ?round(-2.8);
          (D3)        - 3

Function ?TRUNCATE (x,&optional-divisor)

truncate the floating point X towards 0, to become an integer. The argument must be a regular system float, not a bigfloat. The ? beginning the name indicates this is normal common lisp function.

(C4) ?truncate(-2.8);
(D4)        - 2
(C5) ?truncate(2.4);
(D5)         2
(C6) ?truncate(2.8);
(D6)         2

Variable ZUNDERFLOW

default: [TRUE] - if FALSE, an error will be signaled if floating point underflow occurs. Currently in NIL Macsyma, all floating-point underflow, floating-point overflow, and division-by-zero errors signal errors, and this switch is ignored.

96 Maxima Manual

12 Numerical

12.1 Introduction to Numerical

12.2 DCADRE

The following is obsolete. To make an interface to fortran libraries in the current MAXIMA look at the examples in "maxima/src/fortdef.lsp" - The IMSL version of Romberg integration is now available in Macsyma. For documentation, Do PRINT-FILE(DCADRE,USAGE,IMSL1); . For a demo, do batch("dcadre.mc"); This is a numerical integration package using cautious, adaptive Romberg extrapolation. The DCADRE package is written to call the IMSL fortran library routine DCADRE. This is documentation for that program. Send bugs/comments to KMP To load this package, do

       LOADFILE("imsl")$

For a demo of this package, do

       batch("dcadre.mc");

The worker function takes the following syntax: IMSL ROMBERG(fn,low,hi) where fn is a function of 1 argument; low and hi should be the lower and upper bounds of integration. fn must return floating point values. IMSL ROMBERG(exp,var,low,hi) where exp should be integrated over the range var=low to hi. The result of evaluating exp must always be a floating point number. FAST IMSL ROMBERG(fn,low,hi) This function does no error checking but may achieve a speed gain over the IMSL ROMBERG function. It expects that fn is a Lisp function (or translated Macsyma function) which accepts a floating point argument and that it always returns a floating point value.

Returns either [SUCCESS, answer, error] where answer is the result of the integration and error is the estimated bound on the absolute error of the output, DCADRE, as described in PURPOSE below. or [WARNING, n, answer, error] where n is a warning code, answer is the answer, and error is the estimated bound on the absolute error of the output, DCADRE, as described in PURPOSE below. The following warnings may occur: 65 = One or more singularities were successfully handled. 66 = In some subinterval(s), the estimate of the integral was accepted merely because the estimated error was small, even though no regular behavior was recognized. or [ERROR, errorcode] where error code is the IMSL-generated error code. The following error codes may occur: 131 = Failure due to insufficient internal working storage. 132 = Failure. This may be due to too much noise in function (relative to the given error requirements) or due to an ill-behaved integrand. 133 = RERR is greater than 0.1 or less than 0.0 or is too small for the precision of the machine.

The following flags have an influence upon the operation of IMSL ROMBERG – ROMBERG AERR [Default 1.0E-5] – Desired absolute error in answer. ROMBERG RERR [Default 0.0] – Desired relative error in the answer. Note: If IMSL signals an error, a message will be printed on the user’s console stating

the nature of the error. (This error message may be supressed by setting IMSLVERBOSE to FALSE.)

Note: Because this uses a translated Fortran routine, it may not be recursively invoked. It does not call itself, but the user should be aware that he may not type ^A in the middle of an IMSL ROMBERG computation, begin another calculation using the same package, and expect to win – IMSL ROMBERG will complain if it was already doing one project when you invoke it. This should cause minimal problems.

Purpose (modified version of the IMSL documentation) —————————————— ———-

DCADRE attempts to solve the following problem: Given a real-valued function F of one argument, two real numbers A and B, find a number

DCADRE such that:

 |   / B               |        [                              | / B      | ]
 |   [                 |        [                              | [        | ]
 |   I F(x)dx - DCADRE | <= max [ ROMBERG_AERR, ROMBERG_RERR * | I F(x)dx | ]
 |   ]                 |        [                              | ]        | ]
 |   / A               |        [                              | / A      | ]

Algorithm (modified version of the IMSL documentation) This routine uses a scheme whereby DCADRE is computed as the sum of estimates for

the integral of F(x) over suitably chosen subintervals of the given interval of integration. Starting with the interval of integration itself as the first such subinterval, cautious Romberg extrapolation is used to find an acceptable estimate on a given subinterval. If this attempt fails, the subinterval is divided into two subintervals of equal length, each of which is considered separately. Programming Notes (modified version of the IMSL documentation)

• 1. DCADRE (the translated-Fortran base for IMSL ROMBERG) can, in many cases, handle jump discontinuities and certain algebraic discontinuities. See reference for full details.

• 2. The relative error parameter ROMBERG RERR must be in the interval [0.0,0.1]. For example, ROMBERG RERR=0.1 indicates that the estimate of the intergral is to be correct to one digit, where as ROMBERG RERR=1.0E-4 calls for four digits of accuracy. If DCADRE determines that the relative accuracy requirement cannot be satisfied, IER is set to 133 (ROMBERG RERR should be large enough that, when added to 100.0, the result is a number greater than 100.0 (this will not be true of very tiny floating point numbers due to the nature of machine arithmetic)).

• 3. The absolute error parameter, ROMBERG AERR, should be nonnegative. In order to give a reasonable value for ROMBERG AERR, the user must know the approximate magnitude of the integral being computed. In many cases, it is satisfactory to use AERR=0.0. In this case, only the relative error requirement is satisfied in the compuatation.

• 4. We quote from the reference, “A very cautious man would accept DCADRE only if IER [the warning or error code] is 0 or 65. The merely reasonable man would keep the faith even if IER is 66. The adventurous man is quite often right in accepting DCADRE even if the IER is 131 or 132.” Even when IER is not 0, DCADRE returns the best estimate that has been computed.

For references on this technique, see de Boor, Calr, “CADRE: An Algorithm for Numerical Quadrature,” Mathematical Software (John R. Rice, Ed.), New York, Academic Press, 1971, Chapter 7.

12.3 ELLIPT

- A package on the SHARE directory for Numerical routines for Elliptic Functions and Complete Elliptic Integrals. (Notation of Abramowitz and Stegun, Chs 16 and 17) Do LOAD(ELLIPT); to use this package. At present all arguments MUST be floating point. You’ll get nonsense otherwise. Be warned. The functions available are: Jacobian elliptic functions

AM(U,M) - amplitude with modulus M
AM1(U,M1) - ampiltude with complementary modulus M1
AM(U,M):=AM1(U,1-M); so use AM1 if M ~ 1
SN(U,M):=SIN(AM(U,M));
CN(U,M):=COS(AM(U,M));
DN(U,M):=SQRT(1-M*SN(U,M)^2);
(These functions come defined like this.  Others CD, NS etc.  may be
similarly defined.)
Complete Elliptic Integrals
ELLIPTK(M) - Complete elliptic integral of first kind
ELLIPTK1(M1) - Same but with complementary modulus.
ELLIPTK(M):=ELLIPTK1(1-M); so use if M ~ 1
ELLIPTE(M) - Complete elliptic integral of second kind
ELLIPTE1(M1) - Same but with complementary modulus.
ELLIPTE(M):=ELLIPTE1(1-M); so use if M ~ 1

12.4 FOURIER

- There is a Fast Fourier Transform package, do DESCRIBE(FFT) for details. There is also a Fourier Series package. It may be loaded with LOAD(FOURIE). It will also calculate Fourier integral coefficients and has various other functions to do such things as replace all occurrences of F(ARG) by ARG in expression (like changing ABS(a*x+b) to a*x+b). Do PRINTFILE(FOURIE,USAGE,DSK,SHARE1); for a list of the functions included.

12.5 NDIFFQ

a package residing on the SHARE directory for numerical solutions of differential equations. LOAD("NDIFFQ"); will load it in for use. An example of its use would be:

Define_Variable(N,0.3,FLOAT);
Define_Variable(H,0.175,FLOAT);
F(X,E):=(Mode_Declare([X,E],FLOAT),N*EXP(X)/(E+X^(2*H)*EXP(H*X)));
Compile(F);
Array([X,E],FLOAT,35);
Init_Float_Array(X,1.0E-3,6.85); /* Fills X with the interval */
E[0]:5.0;                        /* Initial condition */
Runge_Kutta(F,X,E);              /* Solve it */
Graph2(X,E);                     /* Graph the solution */

p.s. Runge Kutta(F,X,E,E Prime) would be the call for a second-order equation.

12.6 Definitions for Numerical

Function FFT (real-array, imag-array)

Fast Fourier Transform. This package may be loaded by doing LOAD(FFT); There is also an IFT command, for Inverse Fourier Transform. These functions perform a (complex) fast fourier transform on either 1 or 2 dimensional FLOATING-POINT arrays, obtained by:

     ARRAY(<ary>,FLOAT,<dim1>); or
     ARRAY(<ary>,FLOAT,<dim1>,<dim2>);

For 1D arrays

     <dim1> = 2^n-1

and for 2D arrays

     <dim1>=<dim2>=2^n-1

(i.e. the array is square). (Recall that MACSYMA arrays are indexed from a 0 origin so that there will be 2^n and (2^n)^2 arrays elements in the above two cases.) This package also contains two other functions, POLARTORECT and RECTTOPOLAR. Do DESCRIBE(cmd) for details. For details on the implementation, do PRINT-FILE(FFT,USAGE,SHARE); .

Variable FORTINDENT

default: [0] - controls the left margin indentation of expressions printed out by the FORTRAN command. 0 gives normal printout (i.e. 6 spaces), and positive values will causes the expressions to be printed farther to the right.

Function FORTMX (name,matrix)

converts a MACSYMA matrix into a sequence of FORTRAN assignment statements of the form name(i,j)=<corresponding matrix element>. This command is now obsolete. FORTMX(name,matrix); may now be done as FORTRAN(name=matrix);. (If "name" is bound, FORTRAN(’name=matrix); may be necessary.) Please convert code that uses the FORTMX command as it may be flushed some day.

Function FORTRAN (exp)

converts exp into a FORTRAN linear expression in legal FORTRAN with 6 spaces inserted at the beginning of each line, continuation lines, and ** rather than ^for exponentiation. When the option FORTSPACES[FALSE] is TRUE, the FORTRAN command fills out to 80 columns using spaces. If FORTRAN is called on a bound symbolic atom, e.g. FORTRAN(X); where X:A*B$ has been done, then X={value of X}, e.g. X=A*B will be generated. In particular, if e.g. M:MATRIX(...); has been done, then FORTRAN(M); will generate the appropriate assignment statements of the form name(i,j)=<corresponding matrix element>. FORTINDENT[0] controls the left margin of expressions printed out, 0 is the normal margin (i.e. indented 6 spaces), increasing it will cause the expression to be printed further to the right.

Variable FORTSPACES

default: [FALSE] - if TRUE, the FORTRAN command fills out to 80 columns using spaces.

Function HORNER (exp, var)

will convert exp into a rearranged representation as in Horner’s rule, using var as the main variable if it is specified. Var may also be omitted in which case the main variable of the CRE form of exp is used. HORNER sometimes improves stability if expr is to be numerically evaluated. It is also useful if MACSYMA is used to generate programs to be run in FORTRAN (see DESCRIBE(STRINGOUT);)

     (C1) 1.0E-20*X^2-5.5*X+5.2E20;
                                     2
     (D1)                   1.0E-20 X  - 5.5 X + 5.2E+20
     (C2) HORNER(%,X),KEEPFLOAT:TRUE;
     (D2)                  X (1.0E-20 X - 5.5) + 5.2E+20
     (C3) D1,X=1.0E20;
     ARITHMETIC OVERFLOW
     (C4) D2,X=1.0E20;
     (D4)                          6.9999999E+19

Function IFT (real-array, imag-array)

Inverse Fourier Transform. Do LOAD(FFT); to load in this package. These functions (FFT and IFT) perform a (complex) fast fourier transform on either 1 or 2 dimensional FLOATING-POINT arrays, obtained by: ARRAY(<ary>,FLOAT,<dim1>); or ARRAY(<ary>,FLOAT,<dim1>,<dim2>); For 1D arrays <dim1> must equal 2^n-1, and for 2D arrays <dim1>=<dim2>=2^n-1 (i.e. the array is square). (Recall that MACSYMA arrays are indexed from a 0 origin so that there will be 2^n and (2^n)^2 arrays elements in the above two cases.) For details on the implementation, do PRINTFILE(FFT,USAGE,SHARE); .

Function INTERPOLATE (func,x,a,b)

finds the zero of func as x varies. The last two args give the range to look in. The function must have a different sign at each endpoint. If this condition is not met, the action of the of the function is governed by INTPOLERROR[TRUE]). If INTPOL-ERROR is TRUE then an error occurs, otherwise the value of INTPOLERROR is returned (thus for plotting INTPOLERROR might be set to 0.0). Otherwise (given that MACSYMA can evaluate the first argument in the specified range, and that it is continuous) INTERPOLATE is guaranteed to come up with the zero (or one of them if there is more than one zero). The accuracy of INTERPOLATE is governed by INTPOLABS[0.0] and INTPOLREL[0.0] which must be non-negative floating point numbers. INTERPOLATE will stop when the first arg evaluates to something less than or equal to INTPOLABS or if successive approximants to the root differ by no more than INTPOLREL * <one of the approximants>. The default values of INT-POLABS and INTPOLREL are 0.0 so INTERPOLATE gets as good an answer as is possible with the single precision arithmetic we have. The first arg may be an equation. The order of the last two args is irrelevant. Thus

     INTERPOLATE(SIN(X)=X/2,X,%PI,.1);
        is equivalent to
     INTERPOLATE(SIN(X)=X/2,X,.1,%PI);

The method used is a binary search in the range specified by the last two args. When it thinks the function is close enough to being linear, it starts using linear interpolation. An alternative syntax has been added to interpolate, this replaces the first two arguments by a function name. The function MUST be TRANSLATEd or compiled function of one argument. No checking of the result is done, so make sure the function returns a floating point number.

     F(X):=(MODE_DECLARE(X,FLOAT),SIN(X)-X/2.0);
     INTERPOLATE(SIN(X)-X/2,X,0.1,%PI)       time= 60 msec
     INTERPOLATE(F(X),X,0.1,%PI);            time= 68 msec
     TRANSLATE(F);
     INTERPOLATE(F(X),X,0.1,%PI);            time= 26 msec
     INTERPOLATE(F,0.1,%PI);                 time=  5 msec

There is also a Newton method interpolation routine, do DESCRIBE(NEWTON); .

Variable INTPOLABS

default: [0.0] - The accuracy of the INTERPOLATE command is governed by INT-POLABS[0.0] and INTPOLREL[0.0] which must be non-negative floating point numbers. INTERPOLATE will stop when the first arg evaluates to something less than or equal to INTPOLABS or if successive approximants to the root differ by no more than INTPOLREL * <one of the approximants>. The default values of INTPOLABS and INTPOLREL are 0.0 so INTERPOLATE gets as good an answer as is possible with the single precision arithmetic we have.

Variable INTPOLERROR

default: [TRUE] - Governs the behavior of INTERPOLATE. When INTERPOLATE is called, it determines whether or not the function to be interpolated satisfies the condition that the values of the function at the endpoints of the interpolation interval are opposite in sign. If they are of opposite sign, the interpolation proceeds. If they are of like sign, and INTPOLERROR is TRUE, then an error is signaled. If they are of like sign and INTPOLERROR is not TRUE, the value of INTPOLERROR is returned. Thus for plotting, INTPOLERROR might be set to 0.0.

Variable INTPOLREL

default: [0.0] - The accuracy of the INTERPOLATE command is governed by INT-POLABS[0.0] and INTPOLREL[0.0] which must be non-negative floating point numbers. INTERPOLATE will stop when the first arg evaluates to something less than or equal to INTPOLABS or if successive approximants to the root differ by no more than INTPOLREL * <one of the approximants>. The default values of INTPOLABS and INTPOLREL are 0.0 so INTERPOLATE gets as good an answer as is possible with the single precision arithmetic we have.

Function NEWTON (exp,var,X0,eps)

The file NEWTON 1 on the SHARE directory contains a function which will do interpolation using Newton’s method. It may be accessed by LOAD(NEWTON); . The Newton method can do things that INTERPOLATE will refuse to handle, since IN-TERPOLATE requires that everything evaluate to a flonum. Thus NEWTON(x^2a^2,x,a/2,a^2/100); will say that it can’t tell if flonum*a^2<a^2/100. Doing AS-SUME(a>0); and then doing NEWTON again works. You get x=a+<small flonum>*a which is symbolic all the way. INTERPOLATE(x^2-a^2,x,a/2,2*a); complains that .5*a is not flonum... An adaptive integrator which uses the Newton-Cotes 8 panel quadrature rule is available in SHARE1;QQ FASL. Do DESCRIBE(QQ) for details.

Function POLARTORECT (magnitude-array, phase-array)

converts from magnitude and phase form into real and imaginary form putting the real part in the magnitude array and the imaginary part into the phase array

     <real>=<magnitude>*COS(<phase>) ==>
       <imaginary>=<magnitude>*SIN(<phase>

This function is part of the FFT package. Do LOAD(FFT); to use it. Like FFT and IFT this function accepts 1 or 2 dimensional arrays. However, the array dimensions need not be a power of 2, nor need the 2D arrays be square.

Function RECTTOPOLAR (real-array, imag-array)

undoes POLARTORECT. The phase is given in the range from -%PI to %PI. This function is part of the FFT package. Do LOAD(FFT); to use it. Like FFT and IFT this function accepts 1 or 2 dimensional arrays. However, the array dimensions need not be a power of 2, nor need the 2D arrays be square.

104 Maxima Manual

13 Trigonometric

13.1 Introduction to Trigonometric

- MACSYMA has many Trig functions defined. Not all Trig identities are programmed, but it is possible for the user to add many of them using the pattern matching capabilities of the system. The Trig functions defined in MACSYMA are: ACOS, ACOSH, ACOT, ACOTH, ACSC, ACSCH, ASEC, ASECH, ASIN, ASINH, ATAN, ATANH, COS, COSH, COT, COTH, CSC, CSCH, SEC, SECH, SIN, SINH, TAN, and TANH. There are a number of commands especially for handling Trig functions, see TRIGEXPAND, TRIGREDUCE, and the switch TRIGSIGN. Two SHARE packages extend the simplification rules built into MACSYMA, NTRIG and ATRIG1. Do DESCRIBE(cmd) for details.

13.2 Definitions for Trigonometric

Function ACOS

- Arc Cosine

Function ACOSH

- Hyperbolic Arc Cosine

Function ACOT

- Arc Cotangent

Function ACOTH

- Hyperbolic Arc Cotangent

Function ACSC

- Arc Cosecant

Function ACSCH

- Hyperbolic Arc Cosecant

Function ASEC

- Arc Secant

Function ASECH

- Hyperbolic Arc Secant

Function ASIN

- Arc Sine

Function ASINH

- Hyperbolic Arc Sine

Function ATAN

- Arc Tangent

Function ATAN2 (Y,X)

yields the value of ATAN(Y/X) in the interval -%PI to %PI.

Function ATANH

- Hyperbolic Arc Tangent

Function ATRIG1

- SHARE1;ATRIG1 FASL contains several additional simplification rules for inverse trig functions. Together with rules already known to Macsyma, the following angles are fully implemented: 0, %PI/6, %PI/4, %PI/3, and %PI/2. Corresponding angles in the other three quadrants are also available. Do LOAD(ATRIG1); to use them.

Function COS

- Cosine

Function COSH

- Hyperbolic Cosine

Function COT

- Cotangent

Function COTH

- Hyperbolic Cotangent

Function CSC

- Cosecant

Function CSCH

- Hyperbolic Cosecant

Variable HALFANGLES

default: [FALSE] - if TRUE causes half-angles to be simplified away.

Function SEC

- Secant

Function SECH

- Hyperbolic Secant

Function SIN

- Sine

Function SINH

- Hyperbolic Sine

Function TAN

- Tangent

Function TANH

- Hyperbolic Tangent

Function TRIGEXPAND (exp)

expands trigonometric and hyperbolic functions of sums of angles and of multiple angles occurring in exp. For best results, exp should be expanded. To enhance user control of simplification, this function expands only one level at a time, expanding sums of angles or multiple angles. To obtain full expansion into sines and cosines immediately, set the switch TRIGEXPAND:TRUE. TRIGEXPAND default: [FALSE] - if TRUE causes expansion of all expressions containing SINs and COSs occurring subsequently. HALFANGLES[FALSE] - if TRUE causes half-angles to be simplified away. TRIGEXPANDPLUS[TRUE] - controls the "sum" rule for TRIGEXPAND, expansion of sums (e.g. SIN(X+Y)) will take place only if TRIGEXPANDPLUS is TRUE. TRIGEXPANDTIMES[TRUE] - controls the "product" rule for TRIGEXPAND, expansion of products (e.g. SIN(2*X)) will take place only if TRIGEXPANDTIMES is TRUE.

     (C1) X+SIN(3*X)/SIN(X),TRIGEXPAND=TRUE,EXPAND;
                                   2           2
     (D1)                     - SIN (X) + 3 COS (X) + X
     (C2) TRIGEXPAND(SIN(10*X+Y));
     (D2)               COS(10 X) SIN(Y) + SIN(10 X) COS(Y)

Variable TRIGEXPANDPLUS

default: [TRUE] - controls the "sum" rule for TRIGEXPAND. Thus, when the TRIG-EXPAND command is used or the TRIGEXPAND switch set to TRUE, expansion of sums (e.g. SIN(X+Y)) will take place only if TRIGEXPANDPLUS is TRUE.

Variable TRIGEXPANDTIMES

default: [TRUE] - controls the "product" rule for TRIGEXPAND. Thus, when the TRIGEXPAND command is used or the TRIGEXPAND switch set to TRUE, expansion of products (e.g. SIN(2*X)) will take place only if TRIGEXPANDTIMES is TRUE.

Variable TRIGINVERSES

default: [ALL] - controls the simplification of the composition of trig and hyperbolic functions with their inverse functions: If ALL, both e.g. ATAN(TAN(X)) and TAN(ATAN(X)) simplify to X. If TRUE, the arcfunction(function(x)) simplification is turned off. If FALSE, both the arcfun(fun(x)) and fun(arcfun(x)) simplifications are turned off.

Function TRIGREDUCE (exp, var)

combines products and powers of trigonometric and hyperbolic SINs and COSs of var into those of multiples of var. It also tries to eliminate these functions when they occur in denominators. If var is omitted then all variables in exp are used. Also see the POISSIMP function (6.6).

     (C4) TRIGREDUCE(-SIN(X)^2+3*COS(X)^2+X);
     (D4)                        2 COS(2 X) + X + 1
     The trigonometric simplification routines will use declared
     information in some simple cases.  Declarations about variables are
     used as follows, e.g.
     (C5) DECLARE(J, INTEGER, E, EVEN, O, ODD)$
     (C6) SIN(X + (E + 1/2)*%PI)$
     (D6)                      COS(X)
     (C7) SIN(X + (O + 1/2) %PI);
     (D7)                     - COS(X)

Variable TRIGSIGN

default: [TRUE] - if TRUE permits simplification of negative arguments to trigonometric functions. E.g., SIN(-X) will become -SIN(X) only if TRIGSIGN is TRUE.

Function TRIGSIMP (expr)

employs the identities sin(x)^2 + cos(x)^2 = 1 and cosh(x)^2 - sinh(x)^2 = 1 to simplify expressions containing tan, sec, etc. to sin, cos, sinh, cosh so that further simplification may be obtained by using TRIGREDUCE on the result. Some examples may be seen by doing DEMO("trgsmp.dem"); . See also the TRIGSUM function.

Function TRIGRAT (trigexp)

gives a canonical simplifyed quasilinear form of a trigonometrical expression; trigexp is a rational fraction of several sin, cos or tan, the arguments of them are linear forms in some variables (or kernels) and %pi/n (n integer) with integer coefficients. The result is a simplifyed fraction with numerator and denominator linear in sin and cos. Thus TRIGRAT linearize always when it is possible.(written by D. Lazard).

     (c1) trigrat(sin(3*a)/sin(a+%pi/3));
     (d1)          sqrt(3) sin(2 a) + cos(2 a) - 1

Here is another example (for which the function was intended); see [Davenport, Siret, Tournier, Calcul Formel, Masson (or in english, Addison-Wesley), section 1.5.5, Morley theorem). Timings are on VAX 780.

     (c4)   c:%pi/3-a-b;
        %pi
     (d4)   - b - a + ---
         3
     (c5)   bc:sin(a)*sin(3*c)/sin(a+b);

     sin(a) sin(3 b + 3 a)
(d5)       ---------------------
  sin(b + a)
(c6)   ba:bc,c=a,a=c$
(c7)   ac2:ba^2+bc^2-2*bc*ba*cos(b);
2       2
     sin (a) sin (3 b + 3 a)
(d7) -----------------------
      2
   sin (b + a)
%pi
   2 sin(a) sin(3 a) cos(b) sin(b + a - ---) sin(3 b + 3 a)
 3
 - --------------------------------------------------------
   %pi
   sin(a - ---) sin(b + a)
    3
      2         2   %pi
   sin (3 a) sin (b + a - ---)
   3
 + ---------------------------
     2    %pi
  sin (a - ---)
    3
(c9)   trigrat(ac2);
Totaltime= 65866 msec.  GCtime= 7716 msec.
(d9)
- (sqrt(3) sin(4 b + 4 a) - cos(4 b + 4 a)
- 2 sqrt(3) sin(4 b + 2 a)
+ 2 cos(4 b + 2 a) - 2 sqrt(3) sin(2 b + 4 a) + 2 cos(2 b + 4 a)
+ 4 sqrt(3) sin(2 b + 2 a) - 8 cos(2 b + 2 a) - 4 cos(2 b - 2 a)
+ sqrt(3) sin(4 b) - cos(4 b) - 2 sqrt(3) sin(2 b) + 10 cos(2 b)
+ sqrt(3) sin(4 a) - cos(4 a) - 2 sqrt(3) sin(2 a) + 10 cos(2 a)
   - 9)/4

110 Maxima Manual

14 Equations

14.1 Definitions for Equations

Variable %RNUM LIST

default: [] - When %R variables are introduced in solutions by the ALGSYS command, they are added to %RNUM LIST in the order they are created. This is convenient for doing substitutions into the solution later on. It’s recommended to use this list rather than doing CONCAT(’%R,J).

Variable ALGEXACT

default: [FALSE] affects the behavior of ALGSYS as follows: If ALGEXACT is TRUE, ALGSYS always calls SOLVE and then uses REALROOTS on SOLVE’s failures. If ALGEXACT is FALSE, SOLVE is called only if the eliminant was not univariate, or if it was a quadratic or biquadratic. Thus ALGEXACT:TRUE doesn’t guarantee only exact solutions, just that ALGSYS will first try as hard as it can to give exact solutions, and only yield approximations when all else fails.

Function ALGSYS ([exp1, exp2, ...], [var1, var2, ...])

solves the list of simultaneous polynomials or polynomial equations (which can be nonlinear) for the list of variables. The symbols %R1, %R2, etc. will be used to represent arbitrary parameters when needed for the solution (the variable %RNUM LIST holds these). In the process described below, ALGSYS is entered recursively if necessary. The method is as follows: (1) First the equations are FACTORed and split into subsystems. (2) For each subsystem Si, an equation E and a variable var are selected (the var is chosen to have lowest nonzero degree). Then the resultant of E and Ej with respect to var is computed for each of the remaining equations Ej in the subsystem Si. This yields a new subsystem S’i in one fewer variables (var has been eliminated). The process now returns to (1). (3) Eventually, a subsystem consisting of a single equation is obtained. If the equation is multivariate and no approximations in the form of floating point numbers have been introduced, then SOLVE is called to find an exact solution. (The user should realize that SOLVE may not be able to produce a solution or if it does the solution may be a very large expression.) If the equation is univariate and is either linear, quadratic, or bi-quadratic, then again SOLVE is called if no approximations have been introduced. If approximations have been introduced or the equation is not univariate and neither linear, quadratic, or bi-quadratic, then if the switch REALONLY[FALSE] is TRUE, the function REALROOTS is called to find the real-valued solutions. If REALONLY:FALSE then ALLROOTS is called which looks for real and complex-valued solutions. If ALGSYS produces a solution which has fewer significant digits than required, the user can change the value of ALGEPSILON[10^8] to a higher value. If ALGEXACT[FALSE] is set to TRUE, SOLVE will always be called. (4) Finally, the solutions obtained in step (3) are re-inserted into previous levels and the solution process returns to (1). The user should be aware of several caveats: When ALGSYS encounters a multivariate equation which contains floating point approximations (usually due to its failing to find exact solutions at an earlier stage), then it does not attempt to apply exact methods to such equations and instead prints the message: "ALGSYS cannot solve - system too complicated." Interactions with RADCAN can produce large or complicated expressions. In that case, the user may use PICKAPART or REVEAL to analyze the solution. Occasionally, RADCAN may introduce an apparent %I into a solution which is actually real-valued. Do EXAMPLE(ALGSYS); for examples.

Function ALLROOTS (poly)

finds all the real and complex roots of the real polynomial poly which must be univariate and may be an equation, e.g. poly=0. For complex polynomials an algorithm by Jenkins and Traub is used (Algorithm 419, Comm. ACM, vol. 15, (1972), p. 97). For real polynomials the algorithm used is due to Jenkins (Algorithm 493, TOMS, vol. 1, (1975), p.178). The flag POLYFACTOR[FALSE] when true causes ALLROOTS to factor the polynomial over the real numbers if the polynomial is real, or over the complex numbers, if the polynomial is complex. ALLROOTS may give inaccurate results in case of multiple roots. (If poly is real and you get inaccurate answers, you may want to try ALLROOTS(%I*poly);) Do EXAMPLE(ALLROOTS); for an example. ALLROOTS rejects non-polynomials. It requires that the numerator after RATting should be a polynomial, and it requires that the denominator be at most a complex number. As a result of this ALLROOTS will always return an equivalent (but factored) expression, if POLYFACTOR is TRUE.

Variable BACKSUBST

default: [TRUE] if set to FALSE will prevent back substitution after the equations have been triangularized. This may be necessary in very big problems where back substitution would cause the generation of extremely large expressions. (On MC this could cause storage capacity to be exceeded.)

Variable BREAKUP

default: [TRUE] if FALSE will cause SOLVE to express the solutions of cubic or quartic equations as single expressions rather than as made up of several common subexpressions which is the default. BREAKUP:TRUE only works when PROGRAM-MODE is FALSE.

Function DIMENSION (equation or list of equations)

The file "share1/dimen.mc" contains functions for automatic dimensional analysis. LOAD(DIMEN); will load it up for you. There is a demonstration available in share1/dimen.dem. Do DEMO("dimen"); to run it.

Variable DISPFLAG

default: [TRUE] if set to FALSE within a BLOCK will inhibit the display of output generated by the solve functions called from within the BLOCK. Termination of the BLOCK with a dollar sign, $, sets DISPFLAG to FALSE.

Function FUNCSOLVE (eqn,g(t))

gives [g(t) = ...] or [], depending on whether or not there exists a rational fcn g(t) satisfying eqn, which must be a first order, linear polynomial in (for this case) g(t) and g(t+1).

     (C1) FUNCSOLVE((N+1)*FOO(N)-(N+3)*FOO(N+1)/(N+1) =
     (N-1)/(N+2),FOO(N));
                                        N
     (D1)               FOO(N) = ---------------
                                 (N + 1) (N + 2)

Warning: this is a very rudimentary implementation–many safety checks and obvious generalizations are missing.

Variable GLOBALSOLVE

default: [FALSE] if set to TRUE then variables which are SOLVEd for will be set to the solution of the set of simultaneous equations.

Function IEQN (ie,unk,tech,n,guess)

Integral Equation solving routine. Do LOAD(INTEQN); to access it. CAVEAT: To free some storage, a KILL(LABELS) is included in this file. Therefore, before loading the integral equation package, the user should give names to any expressions he wants to keep. ie is the integral equation; unk is the unknown function; tech is the technique to be tried from those given above (tech = FIRST means: try the first technique which finds a solution; tech = ALL means: try all applicable techniques); n is the maximum number of terms to take for TAYLOR, NEUMANN, FIRSTKINDSERIES, or FREDSERIES (it is also the maximum depth of recursion for the differentiation method); guess is the initial guess for NEUMANN or FIRSTKINDSERIES. Default values for the 2nd thru 5th parameters are: unk: P(X), where P is the first function encountered in an integrand which is unknown to MACSYMA and X is the variable which occurs as an argument to the first occurrence of P found outside of an integral in the case of SECONDKIND equations, or is the only other variable besides the variable of integration in FIRSTKIND equations. If the attempt to search for X fails, the user will be asked to supply the independent variable; tech: FIRST; n: 1; guess: NONE, which will cause NEUMANN and FIRSTKINDSERIES to use F(X) as an initial guess.

Variable IEQNPRINT

default: [TRUE] - governs the behavior of the result returned by the IEQN command (which see). If IEQNPRINT is set to FALSE, the lists returned by the IEQN function are of the form [SOLUTION, TECHNIQUE USED, NTERMS, FLAG] where FLAG is absent if the solution is exact. Otherwise, it is the word APPROXIMATE or INCOMPLETE corresponding to an inexact or non-closed form solution, respectively. If a series method was used, NTERMS gives the number of terms taken (which could be less than the n given to IEQN if an error prevented generation of further terms).

Function LHS (eqn)

the left side of the equation eqn.

Function LINSOLVE ([exp1, exp2, ...], [var1, var2, ...])

solves the list of simultaneous linear equations for the list of variables. The expi must each be polynomials in the variables and may be equations. If GLOBAL-SOLVE[FALSE] is set to TRUE then variables which are SOLVEd for will be set to the solution of the set of simultaneous equations. BACKSUBST[TRUE] if set to FALSE will prevent back substitution after the equations have been triangularized. This may be necessary in very big problems where back substitution would cause the generation of extremely large expressions. (On MC this could cause the storage capacity to be exceeded.) LINSOLVE PARAMS[TRUE] If TRUE, LINSOLVE also generates the %Ri symbols used to represent arbitrary parameters described in the manual under ALGSYS. If FALSE, LINSOLVE behaves as before, i.e. when it meets up with an under-determined system of equations, it solves for some of the variables in terms of others.

     (C1) X+Z=Y$
     (C2) 2*A*X-Y=2*A**2$
     (C3) Y-2*Z=2$
     (C4) LINSOLVE([D1,D2,D3],[X,Y,Z]),GLOBALSOLVE:TRUE;
     SOLUTION
     (E4)                            X : A + 1
     (E5)                             Y : 2 A
     (E6)                            Z : A - 1
     (D6)                          [E4, E5, E6]

Variable LINSOLVEWARN

default: [TRUE] - if FALSE will cause the message "Dependent equations eliminated" to be suppressed.

Variable LINSOLVE PARAMS

default: [TRUE] - If TRUE, LINSOLVE also generates the %Ri symbols used to represent arbitrary parameters described in the manual under ALGSYS. If FALSE, LINSOLVE behaves as before, i.e. when it meets up with an under-determined system of equations, it solves for some of the variables in terms of others.

Variable MULTIPLICITIES

default: [NOT SET YET] - will be set to a list of the multiplicities of the individual solutions returned by SOLVE or REALROOTS.

Function NROOTS (poly, low, high)

finds the number of real roots of the real univariate polynomial poly in the half-open interval (low,high]. The endpoints of the interval may also be MINF,INF respectively for minus infinity and plus infinity. The method of Sturm sequences is used.

     (C1) POLY1:X**10-2*X**4+1/2$
     (C2) NROOTS(POLY1,-6,9.1);
     RAT REPLACED 0.5 BY 1/2 = 0.5
     (D2)                               4

Function NTHROOT (p,n)

where p is a polynomial with integer coefficients and n is a positive integer returns q, a polynomial over the integers, such that q^n=p or prints an error message indicating that p is not a perfect nth power. This routine is much faster than FACTOR or even SQFR.

Variable PROGRAMMODE

default: [TRUE] - when FALSE will cause SOLVE, REALROOTS, ALLROOTS, and LINSOLVE to print E-labels (intermediate line labels) to label answers. When TRUE, SOLVE, etc. return answers as elements in a list. (Except when BACKSUBST is set to FALSE, in which case PROGRAMMODE:FALSE is also used.)

Variable REALONLY

default: [FALSE] - if TRUE causes ALGSYS to return only those solutions which are free of %I.

Function REALROOTS (poly, bound)

finds all of the real roots of the real univariate polynomial poly within a tolerance of bound which, if less than 1, causes all integral roots to be found exactly. The parameter bound may be arbitrarily small in order to achieve any desired accuracy. The first argument may also be an equation. REALROOTS sets MULTIPLICI-TIES, useful in case of multiple roots. REALROOTS(poly) is equivalent to REAL-ROOTS(poly,ROOTSEPSILON). ROOTSEPSILON[1.0E-7] is a real number used to establish the confidence interval for the roots. Do EXAMPLE(REALROOTS); for an example.

Function RHS (eqn)

the right side of the equation eqn.

Variable ROOTSCONMODE

default: [TRUE] - Determines the behavior of the ROOTSCONTRACT command. Do DESCRIBE(ROOTSCONTRACT); for details.

Function ROOTSCONTRACT (exp)

converts products of roots into roots of products. For example,

     ROOTSCONTRACT(SQRT(X)*Y^(3/2)) ==> SQRT(X*Y^3)

When RADEXPAND is TRUE and DOMAIN is REAL (their defaults), ROOTSCON-TRACT converts ABS into SQRT, e.g.

     ROOTSCONTRACT(ABS(X)*SQRT(Y)) ==> SQRT(X^2*Y)

There is an option ROOTSCONMODE (default value TRUE), affecting ROOTSCON-TRACT as follows:

     Problem            Value of        Result of applying
                       ROOTSCONMODE        ROOTSCONTRACT

     X^(1/2)*Y^(3/2)      FALSE          (X*Y^3)^(1/2)
     X^(1/2)*Y^(1/4)      FALSE          X^(1/2)*Y^(1/4)
     X^(1/2)*Y^(1/4)      TRUE           (X*Y^(1/2))^(1/2)
     X^(1/2)*Y^(1/3)      TRUE           X^(1/2)*Y^(1/3)
     X^(1/2)*Y^(1/4)      ALL            (X^2*Y)^(1/4)
     X^(1/2)*Y^(1/3)      ALL            (X^3*Y^2)^(1/6)

The above examples and more may be tried out by typing

     EXAMPLE(ROOTSCONTRACT);

When ROOTSCONMODE is FALSE, ROOTSCONTRACT contracts only wrt rational number exponents whose denominators are the same. The key to the ROOTSCONMODE:TRUE$ examples is simply that 2 divides into 4 but not into 3. ROOTSCONMODE:ALL$ involves taking the lcm (least common multiple) of the denominators of the exponents. ROOTSCONTRACT uses RATSIMP in a manner similar to LOGCONTRACT (see the manual).

Variable ROOTSEPSILON

default: [1.0E-7] - a real number used to establish the confidence interval for the roots found by the REALROOTS function.

Function SOLVE (exp, var)

solves the algebraic equation exp for the variable var and returns a list of solution equations in var. If exp is not an equation, it is assumed to be an expression to be set equal to zero. Var may be a function (e.g. F(X)), or other non-atomic expression except a sum or product. It may be omitted if exp contains only one variable. Exp may be a rational expression, and may contain trigonometric functions, exponentials, etc. The following method is used: Let E be the expression and X be the variable. If E is linear in X then it is trivially solved for X. Otherwise if E is of the form A*X**N+B then the result is (-B/A)**(1/N) times the Nth roots of unity. If E is not linear in X then the gcd of the exponents of X in E (say N) is divided into the exponents and the multiplicity of the roots is multiplied by N. Then SOLVE is called again on the result. If E factors then SOLVE is called on each of the factors. Finally SOLVE will use the quadratic, cubic, or quartic formulas where necessary. In the case where E is a polynomial in some function of the variable to be solved for, say F(X), then it is first solved for F(X) (call the result C), then the equation F(X)=C can be solved for X provided the inverse of the function F is known. BREAKUP[TRUE] if FALSE will cause SOLVE to express the solutions of cubic or quartic equations as single expressions rather than as made up of several common subexpressions which is the default. MULTIPLICITIES[NOT SET YET] - will be set to a list of the multiplicities of the individual solutions returned by SOLVE, REALROOTS, or ALLROOTS. Try APROPOS(SOLVE) for the switches which affect SOLVE. DESCRIBE may then by used on the individual switch names if their purpose is not clear. SOLVE([eq1, ..., eqn], [v1, ..., vn]) solves a system of simultaneous (linear or non-linear) polynomial equations by calling LINSOLVE or ALGSYS and returns a list of the solution lists in the variables. In the case of LINSOLVE this list would contain a single list of solutions. It takes two lists as arguments. The first list (eqi, i=1,...,n) represents the equations to be solved; the second list is a list of the unknowns to be determined. If the total number of variables in the equations is equal to the number of equations, the second argument-list may be omitted. For linear systems if the given equations are not compatible, the message INCONSISTENT will be displayed (see the SOLVE INCONSISTENT ERROR switch); if no unique solution exists, then SINGULAR will be displayed. For examples, do EXAMPLE(SOLVE);

Variable SOLVEDECOMPOSES

default: [TRUE] - if TRUE, will induce SOLVE to use POLYDECOMP (see POLY-DECOMP) in attempting to solve polynomials.

Variable SOLVEEXPLICIT

default: [FALSE] - if TRUE, inhibits SOLVE from returning implicit solutions i.e. of the form F(x)=0.

Variable SOLVEFACTORS

default: [TRUE] - if FALSE then SOLVE will not try to factor the expression. The FALSE setting may be desired in some cases where factoring is not necessary.

Variable SOLVENULLWARN

default: [TRUE] - if TRUE the user will be warned if he calls SOLVE with either a null equation list or a null variable list. For example, SOLVE([],[]); would print two warning messages and return [].

Variable SOLVERADCAN

default: [FALSE] - if TRUE then SOLVE will use RADCAN which will make SOLVE slower but will allow certain problems containing exponentials and logs to be solved.

Variable SOLVETRIGWARN

default: [TRUE] - if set to FALSE will inhibit printing by SOLVE of the warning message saying that it is using inverse trigonometric functions to solve the equation, and thereby losing solutions.

Variable SOLVE INCONSISTENT ERROR

default: [TRUE] - If TRUE, SOLVE and LINSOLVE give an error if they meet up with a set of inconsistent linear equations, e.g. SOLVE([A+B=1,A+B=2]). If FALSE, they return [] in this case. (This is the new mode, previously gotten only by calling ALGSYS.)

Function ZRPOLY

- The IMSL ZRPOLY routine for finding the zeros of simple polynomials (single variable, real coefficients, non-negative integer exponents), using the Jenkins-Traub technique. To use it, do: LOADFILE("imsl"); The command is POLY-ROOTS(polynomial); For more info, do: PRINTFILE("zrpoly.usg"); For a demo do: DEMO("zrpoly.dem"); For general info on MACSYMA-IMSL packages, PRINT-FILE(IMSL,USAGE,SHARE2);

Function ZSOLVE

- For those who can make use of approximate numerical solutions to problems, there is a package which calls a routine which has been translated from the IMSL fortran library to solve N simultaneous non-linear equations in N unknowns. It uses black-box techniques that probably aren’t desirable if an exact solution can be obtained from one of the smarter solvers (LINSOLVE, ALGSYS, etc). But for things that the other solvers don’t attempt to handle, this can probably give some very useful results. For documentation, do PRINTFILE("zsolve.usg");. For a demo do batch("zsolve.mc")$

15 Arrays and Tables

15.1 Definitions for Arrays and Tables

Function ARRAY (name, dim1, dim2, ..., dimk)

This sets up a k-dimensional array. A maximum of five dimensions may be used. The subscripts for the ith dimension are the integers running from 0 to dimi. If the user assigns to a subscripted variable before declaring the corresponding array, an undeclared array is set up. If the user has more than one array to be set up the same way, they may all be set up at the same time, by ARRAY([list-of-names],dim1, dim2, ..., dimk). Undeclared arrays, otherwise known as hashed arrays (because hash coding is done on the subscripts), are more general than declared arrays. The user does not declare their maximum size, and they grow dynamically by hashing as more elements are assigned values. The subscripts of undeclared arrays need not even be numbers. However, unless an array is rather sparse, it is probably more efficient to declare it when possible than to leave it undeclared. The ARRAY function can be used to transform an undeclared array into a declared array.

Function ARRAYAPPLY (array,[sub1, ... ,subk])

is like APPLY except the first argument is an array.

Function ARRAYINFO (a)

returns a list of information about the array a. For hashed arrays it returns a list of "HASHED", the number of subscripts, and the subscripts of every element which has a value. For declared arrays it returns a list of "DECLARED", the number of subscripts, and the bounds that were given the the ARRAY function when it was called on a. Do EXAMPLE(ARRAYINFO); for an example.

Function ARRAYMAKE (name,[i1,i2,...])

returns name[i1,i2,...].

Variable ARRAYS

default: [] a list of all the arrays that have been allocated, both declared and undeclared. Functions which deal with arrays are: ARRAY, ARRAYAPPLY, ARRAY-INFO, ARRAYMAKE, FILLARRAY, LISTARRAY, and REARRAY.

Function BASHINDICES (expr)

- transforms the expression expr by giving each summation and product a unique index. This gives CHANGEVAR greater precision when it is working with summations or products. The form of the unique index is J<number>. The quantity <number> is determined by referring to GENSUMNUM, which can be changed by the user. For example, GENSUMNUM:0$ resets it.

Function FILLARRAY (array,list-or-array)

fills array from list-or-array. If array is a floating-point (integer) array then list-or-array should be either a list of floating-point (integer) numbers or another floatingpoint (integer) array. If the dimensions of the arrays are different array is filled in row-major order. If there are not enough elements in list-or-array the last element is used to fill out the rest of array. If there are too many the remaining ones are thrown away. FILLARRAY returns its first argument.

Function GETCHAR (a, i)

returns the ith character of the quoted string or atomic name a. This function is useful in manipulating the LABELS list.

Function LISTARRAY (array)

returns a list of the elements of a declared or hashed array. the order is row-major. Elements which you have not defined yet will be represented by #####.

Function MAKE ARRAY (type,dim1,dim2,...,dimn)

- creates an array. "type" may be ’ANY, ’FLONUM, ’FIXNUM, ’HASHED or ’FUNCTIONAL. This is similar to the ARRAY command, except that the created array is a functional array object. The advantage of this over ARRAY is that it doesn’t have a name, and once a pointer to it goes away, it will also go away. e.g. Y:MAKE ARRAY(....); Y now points to an object which takes up space, but do Y:FALSE, and Y no longer points to that object, so the object will get garbage collected. Note: the "dimi" here are different from the ARRAY command, since they go from 0 to i-1, i.e. a "dimension" of 10 means you have elements from 0 to 9. Y:MAKE ARRAY(’FUNCTIONAL,’F,’HASHED,1) - The second argument to MAKE ARRAY in this case is the function to call to calculate array elements, and the rest of the arguments are passed recursively to MAKE ARRAY to generate the "memory" for the array function object.

Function REARRAY (array,dim1, ... ,dimk)

can be used to change the size or dimensions of an array. The new array will be filled with the elements of the old one in row-major order. If the old array was too small, FALSE, 0.0 or 0 will be used to fill the remaining elements, depending on the type of the array. The type of the array cannot be changed.

Function REMARRAY (name1, name2, ...)

removes arrays and array associated functions and frees the storage occupied. If name is ALL then all arrays are removed. It may be necessary to use this function if it is desired to redefine the values in a hashed array.

Variable USE FAST ARRAYS

[TRUE on Lispm] - If TRUE then only two types of arrays are recognized.

1) The art-q array (t in common lisp) which may have several dimensions indexed by integers, and may hold any lisp or macsyma object as an entry. To construct such an array, enter A:MAKE ARRAY(ANY,3,4); then A will have as value, an array with twelve slots, and the indexing is zero based.

2) The Hash table array which is the default type of array created if one does B[X+1]:Y^2 (and B is not already an array,a list, or a matrix– if it were one of these an error would be caused since x+1 would not be a valid subscript for an art-q array,a list or a matrix ). Its indices (also known as keys) may be any object. It only takes ONE KEY at a time (B[X+1,U]:Y would ignore the u) Referencing is done by B[X+1]==> Y^2. Of course the key may be a list, eg B[[x+1,u]]:y would be valid. This is in- compatible with the old Macsyma hash arrays, but saves consing. An advantage of storing the arrays as values of the symbol is that the usual conventions about local variables of a function apply to arrays as well. The Hash table type also uses less consing and is more efficient than the old type of macsyma hashar. To obtain consistent behaviour in translated and compiled code set TRANS-LATE FAST ARRAYS [TRUE] to be TRUE.

122 Maxima Manual

16 Series

16.1 Introduction to Series

Maxima contains functions Taylor and Powerseries for finding the sseries of differentiable functions. It also has tools such as Nusum capable of finding the closed form of some series. Operations such as addition and multiplication work as usual on series. This section presents the various global various variables which control the expansion.

16.2 Definitions for Series

Variable CAUCHYSUM

default: [FALSE] - When multiplying together sums with INF as their upper limit, if SUMEXPAND is TRUE and CAUCHYSUM is set to TRUE then the Cauchy product will be used rather than the usual product. In the Cauchy product the index of the inner summation is a function of the index of the outer one rather than varying independently. That is: SUM(F(I),I,0,INF)*SUM(G(J),J,0,INF) becomes SUM(SUM(F(I)*G(J-I),I,0,J),J,0,INF)

Function DEFTAYLOR (function, exp)

allows the user to define the Taylor series (about 0) of an arbitrary function of one variable as exp which may be a polynomial in that variable or which may be given implicitly as a power series using the SUM function. In order to display the information given to DEFTAYLOR one can use POWERSERIES(F(X),X,0). (see below).

     (C1) DEFTAYLOR(F(X),X**2+SUM(X**I/(2**I*I!**2),
             I,4,INF));
     (D1)                          [F]
     (C2) TAYLOR(%E**SQRT(F(X)),X,0,4);
                          2         3          4
                         X    3073 X    12817 X
     (D2)/R/     1 + X + -- + ------- + -------- +  . . .
                         2     18432     307200

Variable MAXTAYORDER

default: [TRUE] - if TRUE, then during algebraic manipulation of (truncated) Taylor series, TAYLOR will try to retain as many terms as are certain to be correct.

Function NICEINDICES (expr)

will take the expression and change all the indices of sums and products to something easily understandable. It makes each index it can "I" , unless "I" is in the internal expression, in which case it sequentially tries J,K,L,M,N,I0,I1,I2,I3,I4,... until it finds a legal index.

Variable NICEINDICESPREF

default: [I,J,K,L,M,N] - the list which NICEINDICES uses to find indices for sums and products. This allows the user to set the order of preference of how NICEINDICES finds the "nice indices". E.g. NICEINDICESPREF:[Q,R,S,T,INDEX]$. Then if NICEINDICES finds that it cannot use any of these as indices in a particular summation, it uses the first as a base to try and tack on numbers. Here, if the list is exhausted, Q0, then Q1, etc, will be tried.

Function NUSUM (exp,var,low,high)

performs indefinite summation of exp with respect to var using a decision procedure due to R.W. Gosper. exp and the potential answer must be expressible as products of nth powers, factorials, binomials, and rational functions. The terms "definite" and "indefinite summation" are used analogously to "definite" and "indefinite integration". To sum indefinitely means to give a closed form for the sum over intervals of variable length, not just e.g. 0 to inf. Thus, since there is no formula for the general partial sum of the binomial series, NUSUM can’t do it.

Function PADE (taylor-series,num-deg-bound,denom-deg-bound)

returns a list of all rational functions which have the given taylor-series expansion where the sum of the degrees of the numerator and the denominator is less than or equal to the truncation level of the power series, i.e. are "best" approximants, and which additionally satisfy the specified degree bounds. Its first argument must be a univariate taylor-series; the second and third are positive integers specifying degree bounds on the numerator and denominator. PADE’s first argument can also be a Laurent series, and the degree bounds can be INF which causes all rational functions whose total degree is less than or equal to the length of the power series to be returned. Total degree is num-degree + denom-degree. Length of a power series is "truncation level" + 1 - minimum(0,"order of series").

     (C15) ff:taylor(1+x+x^2+x^3,x,0,3);
          2   3
     (D15)/T/      1 + X + X  + X  + . . .
     (C16) pade(ff,1,1);
     1
     (D16)     [- -----]
           X - 1
     (c1) ff:taylor(-(83787*X^10-45552*X^9-187296*X^8
                       +387072*X^7+86016*X^6-1507328*X^5
       +1966080*X^4+4194304*X^3-25165824*X^2
                       +67108864*X-134217728)
            /134217728,x,0,10);
     (C25) PADE(ff,4,4);
     (D25) []

There is no rational function of degree 4 numerator/denominator, with this power series expansion. You must in general have degree of the numerator and degree of the denominator adding up to at least the degree of the power series, in order to have enough unknown coefficients to solve.

     (C26) PADE(ff,5,5);

(D26) [-(520256329*X^5-96719020632*X^4-489651410240*X^3
           -1619100813312*X^2 -2176885157888*X-2386516803584)
/(47041365435*X^5+381702613848*X^4+1360678489152*X^3
                +2856700692480*X^2
        +3370143559680*X+2386516803584)]

Variable POWERDISP

default: [FALSE] - if TRUE will cause sums to be displayed with their terms in the reverse order. Thus polynomials would display as truncated power series, i.e., with the lowest power first.

Function POWERSERIES (exp, var, pt)

generates the general form of the power series expansion for exp in the variable var about the point pt (which may be INF for infinity). If POWERSERIES is unable to expand exp, the TAYLOR function may give the first several terms of the series. VER-BOSE[FALSE] - if TRUE will cause comments about the progress of POWERSERIES to be printed as the execution of it proceeds.

(C1) VERBOSE:TRUE$
(C2) POWERSERIES(LOG(SIN(X)/X),X,0);
Can't expand
                                 LOG(SIN(X))
So we'll try again after applying the rule:
                                        d
                                      / -- (SIN(X))
                                      [ dX
                        LOG(SIN(X)) = I ----------- dX
                                      ]   SIN(X)
                                      /
In the first simplification we have returned:
                             /
                             [
                             I COT(X) dX - LOG(X)
                             ]
                             /
                    INF
                    ====        I1  2 I1             2 I1
                    \      (- 1)   2     BERN(2 I1) X
                     >     ------------------------------
                    /                I1 (2 I1)!
                    ====
                    I1 = 1
(D2)                -------------------------------------

Variable PSEXPAND

default: [FALSE] - if TRUE will cause extended rational function expressions to display fully expanded. (RATEXPAND will also cause this.) If FALSE, multivariate expressions will be displayed just as in the rational function package. If PSEX-PAND:MULTI, then terms with the same total degree in the variables are grouped together.

Function REVERT (expression,variable)

Does reversion of Taylor Series. "Variable" is the variable the original Taylor expansion is in. Do LOAD(REVERT) to access this function. Try

     REVERT2(expression,variable,hipower)

also. REVERT only works on expansions around 0.

Function SRRAT (exp)

this command has been renamed to TAYTORAT.

Function TAYLOR (exp, var, pt, pow)

expands the expression exp in a truncated Taylor series (or Laurent series, if required) in the variable var around the point pt. The terms through (var-pt)**pow are generated. If exp is of the form f(var)/g(var) and g(var) has no terms up to degree pow then TAYLOR will try to expand g(var) up to degree 2*pow. If there are still no non-zero terms TAYLOR will keep doubling the degree of the expansion of g(var) until reaching pow*2**n where n is the value of the variable TAY-LORDEPTH[3]. If MAXTAYORDER[FALSE] is set to TRUE, then during algebraic manipulation of (truncated) Taylor series, TAYLOR will try to retain as many terms as are certain to be correct. Do EXAMPLE(TAYLOR); for examples. TAY-LOR(exp,[var1,pt1,ord1],[var2,pt2,ord2],...) returns a truncated power series in the variables vari about the points pti, truncated at ordi. PSEXPAND[FALSE] if TRUE will cause extended rational function expressions to display fully expanded. (RAT-EXPAND will also cause this.) If FALSE, multivariate expressions will be displayed just as in the rational function package. If PSEXPAND:MULTI, then terms with the same total degree in the variables are grouped together. TAYLOR(exp, [var1, var2, . . .], pt, ord) where each of pt and ord may be replaced by a list which will correspond to the list of variables. that is, the nth items on each of the lists will be associated together. TAYLOR(exp, [x,pt,ord,ASYMP]) will give an expansion of exp in negative powers of (x-pt). The highest order term will be (x-pt)^(-ord). The ASYMP is a syntactic device and not to be assigned to. See also the TAYLOR LOGEXPAND switch for controlling expansion.

Variable TAYLORDEPTH

default: [3] - If there are still no non-zero terms TAYLOR will keep doubling the degree of the expansion of g(var) until reaching pow*2**n where n is the value of the variable TAYLORDEPTH[3].

Function TAYLORINFO (exp)

returns FALSE if exp is not a Taylor series. Otherwise, a list of lists is returned describing the particulars of the Taylor expansion. For example,

     (C3) TAYLOR((1-Y^2)/(1-X),X,0,3,[Y,A,INF]);
                  2                        2
     (D3)/R/ 1 - A  - 2 A (Y - A) - (Y - A)
                         2                        2
                 + (1 - A  - 2 A (Y - A) - (Y - A) ) X
              2                        2   2
      + (1 - A  - 2 A (Y - A) - (Y - A) ) X
                         2                        2   3
                 + (1 - A  - 2 A (Y - A) - (Y - A) ) X
          + . . .
     (C4) TAYLORINFO(D3);
     (D4)                       [[Y, A, INF], [X, 0, 3]]

Function TAYLORP (exp)

a predicate function which returns TRUE if and only if the expression ’exp’ is in Taylor series representation.

Variable TAYLOR LOGEXPAND

default: [TRUE] controls expansions of logarithms in TAYLOR series. When TRUE all log’s are expanded fully so that zero-recognition problems involving logarithmic identities do not disturb the expansion process. However, this scheme is not always mathematically correct since it ignores branch information. If TAY-LOR LOGEXPAND is set to FALSE, then the only expansion of log’s that will occur is that necessary to obtain a formal power series.

Variable TAYLOR ORDER COEFFICIENTS

default: [TRUE] controls the ordering of coefficients in the expression. The default (TRUE) is that coefficients of taylor series will be ordered canonically.

Function TAYLOR SIMPLIFIER

- A function of one argument which TAYLOR uses to simplify coefficients of power series.

Variable TAYLOR TRUNCATE POLYNOMIALS

default: [TRUE] When FALSE polynomials input to TAYLOR are considered to have infinite precison; otherwise (the default) they are truncated based upon the input truncation levels.

Function TAYTORAT (exp)

converts exp from TAYLOR form to CRE form, i.e. it is like RAT(RATDISREP(exp)) although much faster.

Function TRUNC (exp)

causes exp which is in general representation to be displayed as if its sums were truncated Taylor series. E.g. compare EXP1:X^2+X+1; with EXP2:TRUNC(X^2+X+1); . Note that IS(EXP1=EXP2); gives TRUE.

Function UNSUM (fun,n)

is the first backward difference fun(n) - fun(n-1).

(C1) G(P):=P*4^N/BINOMIAL(2*N,N);
                                            N
                                         P 4
(D1)                       G(P) := ----------------
                                   BINOMIAL(2 N, N)
(C2) G(N^4);
                                     4  N
                                    N  4
(D2)                           ----------------
                               BINOMIAL(2 N, N)
(C3) NUSUM(D2,N,0,N);
                         4        3       2              N
          2 (N + 1) (63 N  + 112 N  + 18 N  - 22 N + 3) 4      2
(D3)      ------------------------------------------------ - ------
                        693 BINOMIAL(2 N, N)                 3 11 7
(C4) UNSUM(%,N);
                                     4  N
                                    N  4
(D4)                           ----------------
                               BINOMIAL(2 N, N)

Variable VERBOSE

default: [FALSE] - if TRUE will cause comments about the progress of POW-ERSERIES to be printed as the execution of it proceeds.

17 Symmetries

17.1 Definitions for Symmetries

Function COMP2PUI (n, l)

re’alise le passage des fonctions syme’triques comple‘tes, donnee’s dans la liste l, aux fonctions syme’triques e’le’mentaires de 0 a‘ n. Si la liste l contient moins de n+1 e’le’ments les valeurs formelles viennent la completer. Le premier e’le’ment de la liste l donne le cardinal de l’alphabet si il existe, sinon on le met e’gal a n.

     COMP2PUI(3,[4,g]);
                          2          3
                [4, g, - g  + 2 h2, g  - 3 h2 g + 3 h3]

Function CONT2PART (pc,lvar)

rend le polyno^me partitionne’ associe’ a‘ la forme contracte’e pc dont les variables sont dans lvar.

     pc : 2*a^3*b*x^4*y + x^5$
     CONT2PART(pc,[x,y]);
                              3
                         [[2 a  b, 4, 1], [1, 5]]

Autres fonctions de changements de repre’sentations : CONTRACT, EXPLOSE, PART2CONT, PARTPOL, TCONTRACT, TPARTPOL.

Function CONTRACT (psym,lvar)

rend une forme contracte’e (i.e. un mono^me par orbite sous l’action du groupe syme’trique) du polyno^me psym en les variables contenues dans la liste lvar. La fonction EXPLOSE re’alise l’ope’ration inverse. La fonction TCONTRACT teste en plus la syme’trie du polyno^me.

     psym : EXPLOSE(2*a^3*b*x^4*y,[x,y,z]);
             3      4      3      4      3    4
          2 a  b y z  + 2 a  b x z  + 2 a  b y  z
                     3    4        3      4      3    4
                + 2 a  b x  z + 2 a  b x y  + 2 a  b x  y
     CONTRACT(psym,[x,y,z]);
                             3    4
                          2 a  b x  y

Autres fonctions de changements de repre’sentations : CONT2PART, EXPLOSE, PART2CONT, PARTPOL, TCONTRACT, TPARTPOL.

Function DIRECT ([P1,...,Pn],y,f,[lvar1,...,lvarn])

calcul l’image directe (voir M. GIUSTI,D. LAZARD et A. VALIBOUZE, ISSAC 1988, Rome) associe’e a‘ la fonction f, en les listes de variables lvar1,...,lvarn, et aux polyno^mes P1,...,Pn d’une variable y. l’arite’ de la fonction f est importante pour le calcul. Ainsi, si l’expression de f ne depend pas d’une variable, non seulement il est inutile de donner cette variable mais cela diminue conside’rablement lees calculs si on ne le fait pas.

DIRECT([z^2  - e1* z + e2, z^2  - f1* z + f2], z, b*v + a*u,
              [[u, v], [a, b]]);
                    2                         2           2
                   z  - e1 f1 z - 4 e2 f2 + e1  f2 + e2 f1
DIRECT([z^3-e1*z^2+e2*z-e3,z^2  - f1* z + f2], z, b*v + a*u,
              [[u, v], [a, b]]);
 6            5            4       2     4          2  4
Y  - 2 E1 F1 Y  - 6 E2 F2 Y  + 2 E1  F2 Y  + 2 E2 F1  Y
    2   2  4
+ E1  F1  Y
              3                  3       3        3          3  3
+ 9 E3 F1 F2 Y  + 5 E1 E2 F1 F2 Y  - 2 E1  F1 F2 Y  - 2 E3 F1  Y
    3  3       2   2  2       2      2  2     4   2  2
- 2 E1 E2 F1  Y  + 9 E2  F2  Y  - 6 E1  E2 F2  Y  + E1  F2  Y
    2     2       2   2     2       2      2     2
- 9 E1 E3 F1  F2 Y  - 6 E2  F1  F2 Y  + 3 E1  E2 F1  F2 Y
            4  2
+ 2 E1 E3 F1  Y
    2   4  2                 2         2         2
+ E2  F1  Y  - 27 E2 E3 F1 F2  Y + 9 E1  E3 F1 F2  Y
         2      2
+ 3 E1 E2  F1 F2  Y
    3         2                3            2      3
- E1  E2 F1 F2  Y + 15 E2 E3 F1  F2 Y - 2 E1  E3 F1  F2 Y
       2   3
- E1 E2  F1  F2 Y
    5          2   3                 3       3      3
- 2 E2 E3 F1  Y - 27 E3  F2  + 18 E1 E2 E3 F2  - 4 E1  E3 F2

       3   3
 - 4 E2  F2
    2   2   3        2   2   2                2   2     3      2   2
+ E1  E2  F2  + 27 E3  F1  F2  - 9 E1 E2 E3 F1  F2  + E1  E3 F1  F2
    3   2   2       2   4                 4        2   6
+ E2  F1  F2  - 9 E3  F1  F2 + E1 E2 E3 F1  F2 + E3  F1

Recherche du polyno^me dont les racines sont les somme a+u ou a est racine de z^2 - e1* z + e2 et u est racine de z^2 - f1* z + f2

DIRECT([z^2  - e1* z + e2,z^2  - f1* z + f2], z,a+u,[[u],[a]]);
 4         3         3         2     2  2            2         2
Y  - 2 F1 Y  - 2 E1 Y  + 2 F2 Y  + F1  Y  + 3 E1 F1 Y  + 2 E2 Y
    2  2
+ E1  Y
                               2                   2
- 2 F1 F2 Y - 2 E1 F2 Y - E1 F1  Y - 2 E2 F1 Y - E1  F1 Y
                2
- 2 E1 E2 Y + F2
                         2           2                2
+ E1 F1 F2 - 2 E2 F2 + E1  F2 + E2 F1  + E1 E2 F1 + E2

DIRECT peut prendre deux drapeaux possibles : ELEMENTAIRES et PUISSANCES (valeur par de’faut) qui permettent de de’composer les polyno^mes syme’triques apparaissant dans ce calcul par les fonctions syme’triques e’le’mentaires ou les fonctions puissances respectivement.

fonctions de SYM utilis’ees dans cette fonction :

MULTI_ORBIT (donc ORBIT), PUI_DIRECT, MULTI_ELEM
(donc ELEM), MULTI_PUI (donc PUI), PUI2ELE, ELE2PUI
(si le drapeau DIRECT est a` PUISSANCES).

Function ELE2COMP (m , l)

passe des fonctions syme’triques e’le’mentaires aux fonctions comple‘tes. Similaire a‘ COMP2ELE et COMP2PUI.

autres fonctions de changements de bases :

COMP2ELE, COMP2PUI, ELE2PUI, ELEM, MON2SCHUR, MULTI_ELEM,
MULTI_PUI, PUI, PUI2COMP, PUI2ELE, PUIREDUC, SCHUR2COMP.

Function ELE2POLYNOME (l,z)

donne le polyno^me en z dont les fonctions syme’triques e’le’mentaires des racines sont dans la liste l. l=[n,e1,...,en] ou‘ n est le degre’ du polyno^me et ei la i-ie‘me fonction syme’trique e’le’mentaire.

      ele2polynome([2,e1,e2],z);
                                       2
                                      Z  - E1 Z + E2
      polynome2ele(x^7-14*x^5  + 56*x^3  - 56*X + 22,x);
                   [7, 0, - 14, 0, 56, 0, - 56, - 22]
      ele2polynome( [7, 0, - 14, 0, 56, 0, - 56, - 22],x);
                               7       5       3
                              X  - 14 X  + 56 X  - 56 X + 22

la re’ciproque : POLYNOME2ELE(p,z) autres fonctions a‘ voir : POLYNOME2ELE, PUI2POLYNOME.

Function ELE2PUI (m, l)

passe des fonctions syme’triques e’le’mentaires aux fonctions comple‘tes. Similaire a‘ COMP2ELE et COMP2PUI. autres fonctions de changements de bases :

     COMP2ELE, COMP2PUI, ELE2COMP, ELEM, MON2SCHUR, MULTI_ELEM,
     MULTI_PUI, PUI, PUI2COMP, PUI2ELE, PUIREDUC, SCHUR2COMP.

Function ELEM (ele,sym,lvar)

de’compose le polyno^me syme’trique sym, en les variables contenues de la liste lvar, par les fonctions syme’triques e’le’mentaires contenues dans la liste ele. Si le premier e’le’ment de ele est donne’ ce sera le cardinal de l’alphabet sinon on prendra le degre’ du polyno^me sym. Si il manque des valeurs a‘ la liste ele des valeurs formelles du type "ei" sont rajoute’es. Le polyno^me sym peut etre donne’ sous 3 formes diffe’rentes : contracte’e (ELEM doit alors valoir 1 sa valeur par de’faut), partitionne’e (ELEM doit alors valoir 3) ou e’tendue (i.e. le polyno^me en entier) (ELEM doit alors valoir 2). L’utilsation de la fonction PUI se re’alise sur le me^me mode‘le. Sur un alphabet de cardinal 3 avec e1, la premie‘re fonction syme’trique e’le’mentaire, valant 7, le polyno^me syme’trique en 3 variables dont la forme contracte’e (ne de’pendant ici que de deux de ses variables) est x^4-2*x*y se de’compose ainsi en les fonctions syme’triques e’le’mentaires :

     ELEM([3,7],x^4-2*x*y,[x,y]);
                                    2
                        28 e3 + 2 e2  - 198 e2 + 2401

autres fonctions de changements de bases : COMP2ELE, COMP2PUI, ELE2COMP, ELE2PUI, MON2SCHUR, MULTI ELEM, MULTI PUI, PUI, PUI2COMP, PUI2ELE, PUIREDUC, SCHUR2COMP.

Function EXPLOSE (pc,lvar)

rend le polyno^me syme’trique associe’ a‘ la forme contracte’e pc. La liste lvar contient les variables.

     EXPLOSE(a*x +1,[x,y,z]);
                          (x + y + z) a + 1

Autres fonctions de changements de repre’sentations : CONTRACT, CONT2PART, PART2CONT, PARTPOL, TCONTRACT, TPART-POL.

Function KOSTKA (part1,part2)

e’crite par P. ESPERET) calcule le nombre de kostka associe’ aux partition part1 et part2

     kostka([3,3,3],[2,2,2,1,1,1]);
                                       6

Function LGTREILLIS (n,m)

rend la liste des partitions de poids n et de longueur m. LGTREILLIS(4,2); [[3, 1], [2, 2]] Voir e’galement : LTREILLIS, TREILLIS et TREINAT.

Function LTREILLIS (n,m)

rend la liste des partitions de poids n et de longueur infe’rieure ou e’gale a‘ m.

      ltreillis(4,2);
                              [[4, 0], [3, 1], [2, 2]]

Voir e’galement : LGTREILLIS, TREILLIS et TREINAT.

Function MON2SCHUR (l)

la liste l repre’sente la fonction de Schur S l : On a l=[i1,i2,...,iq] avec i1 <= i2 <= ... <= iq . La fonction de Schur est S [i1,i2...,iq] est le mineur de la matrice infinie (h {i-j}) i>=1, j>=1 compose’ des q premie‘res lignes et des colonnes i1+1,i2+2,...,iq+q.

On e’crit cette fonction de Schur en fonction des formes monomiales en utilisant les fonctions TREINAT et KOSTKA. La forme rendue est un polyno^me syme’trique dans une de ses repre’sentations contracte’es avec les variables x1, x2, ...

      mon2schur([1,1,1]);
                                    X1 X2 X3
      mon2schur([3]);
                                              2        3
                                 X1 X2 X3 + X1  X2 + X1
      MON2SCHUR([1,2]);
          2
             2 x1 x2 x3 + x1  x2

ce qui veut dire que pour 3 variables cela donne :

        2 x1 x2 x3 + x1^2 x2 + x2^2 x1 + x1^2 x3 + x3^2 x1
         + x2^2 x3 + x3^2 x2

autres fonctions de changements de bases :

     COMP2ELE, COMP2PUI, ELE2COMP, ELE2PUI, ELEM, MULTI_ELEM,
     MULTI_PUI, PUI, PUI2COMP, PUI2ELE, PUIREDUC, SCHUR2COMP.

Function MULTI ELEM (l elem,multi pc,l var)

de’compose un polyno^me multi-syme’trique sous la forme multi-contracte’e multi pc en les groupes de variables contenue dans la liste de listes l var sur les groupes de fonctions syme’triques e’le’mentaires contenues dans l elem.

MULTI ELEM([[2,e1,e2],[2,f1,f2]],a*x+a^2+x^3,[[x,y],[a,b]]);

2 3 - 2 f2 + f1 + e1 f1 - 3 e1 e2 + e1

autres fonctions de changements de bases :

     COMP2ELE, COMP2PUI, ELE2COMP, ELE2PUI, ELEM,
     MON2SCHUR, MULTI_PUI, PUI, PUI2COMP, PUI2ELE,
     PUIREDUC, SCHUR2COMP.

Function MULTI ORBIT (P,[lvar1, lvar2,...,lvarp])

P est un polyno^me en l’ensemble des variables contenues dans les listes lvar1, lvar2 ... lvarp. Cette fonction rame‘ne l’orbite du polyno^me P sous l’action du produit des groupes syme’triques des ensembles de variables repre’sente’s par ces p LISTES.

      MULTI_ORBIT(a*x+b*y,[[x,y],[a,b]]);
               [b y + a x, a y + b x]
      multi_orbit(x+y+2*a,[[x,y],[a,b,c]]);
                   [Y + X + 2 C, Y + X + 2 B, Y + X + 2 A]

voir e’galement : ORBIT pour l’action d’un seul groupe syme’trique

Function MULTI PUI

est a‘ la fonction PUI ce que la fonction MULTI ELEM est a‘ la fonction ELEM.

     MULTI_PUI([[2,p1,p2],[2,t1,t2]],a*x+a^2+x^3,[[x,y],[a,b]]);
                                                   3
                                            3 P1 P2   P1
                               T2 + P1 T1 + ------- - ---
                                               2       2

Function MULTINOMIAL (r,part)

ou‘ r est le poids de la partition part. Cette fonction rame‘ne le coefficient multinomial associe’ : si les parts de la partitions part sont i1, i2, ..., ik, le re’sultat de MULTINOMIAL est r!/(i1!i2!...ik!).

Function MULTSYM (ppart1, ppart2,N)

re’alise le produit de deux polyno^mes syme’triques de N variables en ne travaillant que modulo l’action du groupe syme’trique d’ordre N. Les polyno^mes sont dans leur repre’sentation partitionne’e. Soient les 2 polyno^mes syme’triques en x, y : 3*(x+y) + 2*x*y et 5*(x^2+y^2) dont les formes partitionne’es sont respectivement [[3,1],[2,1,1]] et [[5,2]], alors leur produit sera donne’ par :

      MULTSYM([[3,1],[2,1,1]],[[5,2]],2);
                 [[10, 3, 1], [15, 2, 1], [15, 3, 0]]

soit 10*(x^3*y+y^3*x)+15*(x^2*y +y^2*x) +15(x^3+y^3) Fonctions de changements de repre’sentations d’un polyno^me syme’trique : CONTRACT, CONT2PART, EXPLOSE, PART2CONT, PARTPOL, TCONTRACT, TPARTPOL.

Function ORBIT (P,lvar)

calcul l’orbite du polyno^me P en les variables de la liste lvar sous l’action du groupe syme’trique de l’ensemble des variables contenues dans la liste lvar.

      orbit(a*x+b*y,[x,y]);
                             [A Y + B X, B Y + A X]
      orbit(2*x+x^2,[x,y]);
                                     2         2
                                   [Y  + 2 Y, X  + 2 X]

voir e’galement : MULTI ORBIT pour l’action d’un produit de groupes syme’triques sur un polyno^me.

Function PART2CONT (ppart,lvar)

passe de la forme partitionne’e a‘ la forme contracte’e d’un polyno^me syme’trique. La forme contracte’e est rendue avec les variables contenues dans lvar.

     PART2CONT([[2*a^3*b,4,1]],[x,y]);
                             3    4
                          2 a  b x  y

Autres fonctions de changements de repre’sentations : CONTRACT, CONT2PART, EXPLOSE, PARTPOL, TCONTRACT, TPARTPOL.

Function PARTPOL (psym, lvar)

psym est un polyno^me syme’trique en les variables de lvar. Cette fonction rame‘ne sa repre’sentation partitionne’e.

     PARTPOL(-a*(x+y)+3*x*y,[x,y]);
                        [[3, 1, 1], [- a, 1, 0]]

Autres fonctions de changements de repre’sentations : CONTRACT, CONT2PART, EXPLOSE, PART2CONT, TCONTRACT, TPART-POL.

Function PERMUT (l)

rame‘ne la liste des permutations de la liste l.

Function POLYNOME2ELE (p,x)

donne la liste l=[n,e1,...,en] ou‘ n est le degre’ du polyno^me p en la variable x et ei la i-ieme fonction syme’trique e’le’mentaire des racines de p.

      POLYNOME2ELE(x^7-14*x^5  + 56*x^3  - 56*X + 22,x);
                   [7, 0, - 14, 0, 56, 0, - 56, - 22]
      ELE2POLYNOME( [7, 0, - 14, 0, 56, 0, - 56, - 22],x);
                               7       5       3
                              X  - 14 X  + 56 X  - 56 X + 22

la re’ciproque : ELE2POLYNOME(l,x)

Function PRODRAC (L,K)

L est une liste contenant les fonctions syme’triques e’le’mentaires sur un ensemble A. PRODRAC rend le polyno^me dont les racines sont les produits K a‘ K des e’le’ments de A.

Function PUI (pui,sym,lvar)

de’compose le polyno^me syme’trique sym, en les variables contenues de la liste lvar, par les fonctions puissances contenues dans la liste pui. Si le premier e’le’ment de pui est donne’ ce sera le cardinal de l’alphabet sinon on prendra le degre’ du polyno^me sym. Si il manque des valeurs a‘ la liste pui, des valeurs formelles du type "pi" sont rajoute’es. Le polyno^me sym peut etre donne’ sous 3 formes diffe’rentes : contracte’e (PUI doit alors valoir 1 sa valeur par de’faut), partitionne’e (PUI doit alors valoir 3) ou e’tendue (i.e. le polyno^me en entier) (PUI doit alors valoir 2). La fonction ELEM s’utilise de la me^me manie‘re.

     PUI;
            1
     PUI([3,a,b],u*x*y*z,[x,y,z]);
            3
          (a  - 3 b a + 2 p3) u
                    ---------------------
           6

autres fonctions de changements de bases :

     COMP2ELE, COMP2PUI, ELE2COMP, ELE2PUI, ELEM, MON2SCHUR,
     MULTI_ELEM, MULTI_PUI, PUI2COMP, PUI2ELE, PUIREDUC,
     SCHUR2COMP.

Function PUI2COMP (N,LPUI)

rend la liste des N premie‘res fonctions comple‘tes (avec en te^te le cardinal) en fonction des fonctions puissance donne’es dans la liste LPUI. Si la liste LPUI est vide le cardinal est N sinon c’est son premier e’le’ment similaire a‘ COMP2ELE et COMP2PUI.

 PUI2COMP(2,[]);
 2
       p1  + p2
                [2, p1, --------]
  2
 PUI2COMP(3,[2,a1]);
      2         3
    a1  + p2  a1  + 3 p2 a1 + 2 p3
               [2, a1, --------, --------------------]
       2        6

Autres fonctions de changements de bases :

COMP2ELE, COMP2PUI, ELE2COMP, ELE2PUI, ELEM,
MON2SCHUR, MULTI_ELEM, MULTI_PUI, PUI, PUI2ELE,
PUIREDUC, SCHUR2COMP.

Function PUI2ELE (N,LPUI)

re’alise le passage des fonctions puissances aux fonctions syme’triques e’le’mentaires. Si le drapeau PUI2ELE est GIRARD, on re’cupe‘re la liste des fonctions syme’triques e’le’mentaires de 1 a‘ N, et s’il est e’gal a‘ CLOSE, la Nie‘me fonction syme’trique e’le’mentaire. Autres fonctions de changements de bases :

COMP2ELE, COMP2PUI, ELE2COMP, ELE2PUI, ELEM,
MON2SCHUR, MULTI_ELEM, MULTI_PUI, PUI, PUI2COMP,
PUIREDUC, SCHUR2COMP.

Function PUI2POLYNOME (X,LPUI)

calcul le polyno^me en X dont les fonctions puissances des racines sont donne’es dans la liste LPUI.

(C6) polynome2ele(x^3-4*x^2+5*x-1,x);
(D6)                             [3, 4, 5, 1]
(C7) ele2pui(3,%);
(D7)                             [3, 4, 6, 7]
(C8) pui2polynome(x,%);
                               3      2
(D8)                          X  - 4 X  + 5 X - 1

Autres fonctions a‘ voir : POLYNOME2ELE, ELE2POLYNOME.

Function PUI DIRECT (ORBITE,[lvar1,...,lvarn],[d1,d2,...,dn])

Soit f un polynome en n blocs de variables lvar1,...,lvarn. Soit ci le nombre de variables dans lvari . Et SC le produit des n groupes syme’triques de degre’ c1,...,cn. Ce groupe agit naturellement sur f La liste ORBITE est l’orbite, note’e SC(f), de la fonction f sous l’action de SC. (Cette liste peut e^tre obtenue avec la fonction : MULTI ORBIT). Les di sont des entiers tels que c1<=d1, c2<=d2,...,cn<=dn. Soit SD le produit des groupes syme’triques S d1 x S d2 x...x S dn. la fonction pui direct rame‘ne les N premie‘res fonctions puissances de SD(f) de’duites des fonctions puissances de SC(f) ou‘ N est le cardinal de SD(f). Le re’sultat est rendue sous forme multi-contracte’e par rapport a SD. i.e. on ne conserve qu’un e’le’ment par orbite sous l’action de SD).

     L:[[x,y],[a,b]]$
     PUI_DIRECT(MULTI_ORBIT(a*x+b*y, L), L,[2,2]);
                                         2  2
                      [a x, 4 a b x y + a  x ]
     PUI_DIRECT(MULTI_ORBIT(a*x+b*y, L), L,[3,2]);
                              2  2     2    2        3  3
       [2 A X, 4 A B X Y + 2 A  X , 3 A  B X  Y + 2 A  X ,
         2  2  2  2      3    3        4  4
     12 A  B  X  Y  + 4 A  B X  Y + 2 A  X ,
         3  2  3  2      4    4        5  5
     10 A  B  X  Y  + 5 A  B X  Y + 2 A  X ,
         3  3  3  3       4  2  4  2      5    5        6  6
     40 A  B  X  Y  + 15 A  B  X  Y  + 6 A  B X  Y + 2 A  X ]
      PUI_DIRECT([y+x+2*c, y+x+2*b, y+x+2*a],[[x,y],[a,b,c]],[2,3]);
                                  2              2
           [3 x + 2 a, 6 x y + 3 x  + 4 a x + 4 a ,
                   2                   3        2       2        3
                9 x  y + 12 a x y + 3 x  + 6 a x  + 12 a  x + 8 a ]

     PUI_DIRECT([y+x+2*c, y+x+2*b, y+x+2*a],[[x,y],[a,b,c]],[3,4]);

Function PUIREDUC (N,LPUI)

LPUI est une liste dont le premier e’le’ment est un entier M. PUIREDUC donne les N premie‘res fonctions puissances en fonction des M premie‘res.

   PUIREDUC(3,[2]);
     3
         3 p1 p2 - p1
               [2, p1, p2, -------------]
       2

Function RESOLVANTE (p,x,f,[x1,...,xd])

calcule la re’solvante du polyno^me p de la variable x et de degre’ n >= d par la fonction f exprime’e en les variables x1,...,xd. Il est important pour l’efficacite’ des calculs de ne pas mettre dans la liste [x1,...,xd] les variables n’intervenant pas dans la fonction de transformation f. Afin de rendre plus efficaces les calculs on peut mettre des drapeaux a‘ la variable RESOLVANTE afin que des algorithmes ade’quates soient utilise’s : Si la fonction f est unitaire :

• un polyno^me d’une variable,

• line’aire ,

• alterne’e,

• une somme de variables,

• syme’trique en les variables qui apparaissent dans son expression,

• un produit de variables,

• la fonction de la re’solvante de Cayley (utilisable qu’en degre’ 5)

       (x1*x2+x2*x3+x3*x4+x4*x5+x5*x1 -
            (x1*x3+x3*x5+x5*x2+x2*x4+x4*x1))^2

generale, le drapeau de RESOLVANTE pourra e^tre respectivement :

• unitaire,

• lineaire,

• alternee,

• somme,

• produit,

• cayley,

• generale.

    resolvante:unitaire;
   resolvante(x^7-14*x^5  + 56*x^3  - 56*X + 22,x,x^3-1,[x]);
     7      6        5         4          3           2
   Y  + 7 Y  - 539 Y  - 1841 Y  + 51443 Y  + 315133 Y  + 376999 Y

 + 125253
resolvante : lineaire;
resolvante(x^4-1,x,x1+2*x2+3*x3,[x1,x2,x3]);
 24       20         16            12             8              4
Y   + 80 Y   + 7520 Y   + 1107200 Y   + 49475840 Y  + 344489984 Y
 + 655360000
       Meme solution pour :
resolvante : general;
resolvante(x^4-1,x,x1+2*x2+3*x3,[x1,x2,x3]);
resolvante(x^4-1,x,x1+2*x2+3*x3,[x1,x2,x3,x4])
direct([x^4-1],x,x1+2*x2+3*x3,[[x1,x2,x3]]);
resolvante:lineaire$
resolvante(x^4-1,x,x1+x2+x3,[x1,x2,x3);
       4
      Y  - 1
resolvante:symetrique$
resolvante(x^4-1,x,x1+x2+x3,[x1,x2,x3]);
       4
      Y  - 1
resolvante(x^4+x+1,x,x1-x2,[x1,x2]);
  12      8       6        4        2
 Y   + 8 Y  + 26 Y  - 112 Y  + 216 Y  + 229
resolvante:alternee$
resolvante(x^4+x+1,x,x1-x2,[x1,x2]);
  12      8       6        4        2
 Y   + 8 Y  + 26 Y  - 112 Y  + 216 Y  + 229

resolvante:produit;
resolvante(x^7-7*x+3,x,x1*x2*x3,[x1,x2,x3]);
   35      33         29        28         27        26         24
  Y   - 7 Y   - 1029 Y   + 135 Y   + 7203 Y   - 756 Y   + 1323 Y
          23          22            21           20          19
+ 352947 Y   - 46305 Y   - 2463339 Y   + 324135 Y   - 30618 Y
  18
- 453789 Y

    17              15             14              12
- 40246444 Y   + 282225202 Y   - 44274492 Y   + 155098503 Y
            11
+ 12252303 Y
   10              9            8            7             6
+ 2893401 Y   - 171532242 Y  + 6751269 Y  + 2657205 Y  - 94517766 Y
   5             3
- 3720087 Y  + 26040609 Y  + 14348907
   resolvante:symetrique$
  resolvante(x^7-7*x+3,x,x1*x2*x3,[x1,x2,x3]);
  35      33         29        28         27        26         24
 Y   - 7 Y   - 1029 Y   + 135 Y   + 7203 Y   - 756 Y   + 1323 Y
          23          22            21           20          19
+ 352947 Y   - 46305 Y   - 2463339 Y   + 324135 Y   - 30618 Y
           18
 - 453789 Y
            17              15             14              12
- 40246444 Y   + 282225202 Y   - 44274492 Y   + 155098503 Y
             11
 + 12252303 Y
           10              9            8            7             6
+ 2893401 Y   - 171532242 Y  + 6751269 Y  + 2657205 Y  - 94517766 Y
           5             3
- 3720087 Y  + 26040609 Y  + 14348907
resolvante:cayley$
resolvante(x^5-4*x^2+x+1,x,a,[]);
" resolvante de Cayley "
 6       5         4          3            2
X  - 40 X  + 4080 X  - 92928 X  + 3772160 X  + 37880832 X + 93392896

Pour la re’solvante de Cayley, les 2 derniers arguments sont neutres et le polyno^me donne’ en entre’e doit ne’cessairement e^tre de degre’ 5.

Voir e’galement :

RESOLVANTE_BIPARTITE, RESOLVANTE_PRODUIT_SYM,
RESOLVANTE_UNITAIRE, RESOLVANTE_ALTERNEE1, RESOLVANTE_KLEIN,

RESOLVANTE_KLEIN3, RESOLVANTE_VIERER, RESOLVANTE_DIEDRALE.

Function RESOLVANTE ALTERNEE1 (p,x)

calcule la transformation de p(x) de degre n par la fonction $\prod {1\leq i<j\leq n-1} (x i-x j)$. Voir e’galement :

RESOLVANTE_PRODUIT_SYM, RESOLVANTE_UNITAIRE,
RESOLVANTE , RESOLVANTE_KLEIN, RESOLVANTE_KLEIN3,
RESOLVANTE_VIERER, RESOLVANTE_DIEDRALE, RESOLVANTE_BIPARTITE.

Function RESOLVANTE BIPARTITE (p,x)

calcule la transformation de p(x) de degre n (n pair) par la fonction $x 1x 2\ldots x {n/2}+x {n/2+1}\ldotsx n$ Voir e’galement :

RESOLVANTE_PRODUIT_SYM, RESOLVANTE_UNITAIRE,
RESOLVANTE , RESOLVANTE_KLEIN, RESOLVANTE_KLEIN3,
RESOLVANTE_VIERER, RESOLVANTE_DIEDRALE,RESOLVANTE_ALTERNEE1
 RESOLVANTE_BIPARTITE(x^6+108,x);
                 10        8           6             4
                Y   - 972 Y  + 314928 Y  - 34012224 Y

Voir e’galement :

RESOLVANTE_PRODUIT_SYM, RESOLVANTE_UNITAIRE,
RESOLVANTE, RESOLVANTE_KLEIN, RESOLVANTE_KLEIN3,
RESOLVANTE_VIERER, RESOLVANTE_DIEDRALE,
RESOLVANTE_ALTERNEE1.

Function RESOLVANTE DIEDRALE (p,x)

calcule la transformation de p(x) par la fonction x 1x 2+x 3x 4.

resolvante_diedrale(x^5-3*x^4+1,x);
 15       12       11       10        9         8         7        6
X   - 21 X   - 81 X   - 21 X   + 207 X  + 1134 X  + 2331 X  - 945 X
           5          4          3          2
   - 4970 X  - 18333 X  - 29079 X  - 20745 X  - 25326 X - 697

Voir e’galement :

RESOLVANTE_PRODUIT_SYM, RESOLVANTE_UNITAIRE,
RESOLVANTE_ALTERNEE1, RESOLVANTE_KLEIN, RESOLVANTE_KLEIN3,
RESOLVANTE_VIERER, RESOLVANTE.

Function RESOLVANTE KLEIN (p,x)

calcule la transformation de p(x) par la fonction x 1x 2x 4+x 4. Voir e’galement :

RESOLVANTE_PRODUIT_SYM, RESOLVANTE_UNITAIRE,
RESOLVANTE_ALTERNEE1, RESOLVANTE, RESOLVANTE_KLEIN3,
RESOLVANTE_VIERER, RESOLVANTE_DIEDRALE.

Function RESOLVANTE KLEIN3 (p,x)

calcule la transformation de p(x) par la fonction x 1x 2x 4+x 4. Voir e’galement :

RESOLVANTE_PRODUIT_SYM, RESOLVANTE_UNITAIRE,
RESOLVANTE_ALTERNEE1, RESOLVANTE_KLEIN, RESOLVANTE,
RESOLVANTE_VIERER, RESOLVANTE_DIEDRALE.

Function RESOLVANTE PRODUIT SYM (p,x)

calcule la liste toutes les r\’esolvantes produit du polyn\^ome p(x).

  resolvante_produit_sym(x^5+3*x^4+2*x-1,x);
  5      4             10      8       7       6       5    4
[Y  + 3 Y  + 2 Y - 1, Y   - 2 Y  - 21 Y  - 31 Y  - 14 Y  - Y
      3
+ 14 Y
      2       10      8       7    6       5       4       3      2
+ 3 Y  + 1, Y   + 3 Y  + 14 Y  - Y  - 14 Y  - 31 Y  - 21 Y  - 2 Y
      5      4
+ 1, Y  - 2 Y  - 3 Y - 1, Y - 1]

resolvante:produit$
esolvante(x^5+3*x^4+2*x-1,x,a*b*c,[a,b,c]);
 10      8       7    6       5       4       3      2
Y   + 3 Y  + 14 Y  - Y  - 14 Y  - 31 Y  - 21 Y  - 2 Y  + 1

Voir e’galement :

RESOLVANTE, RESOLVANTE_UNITAIRE,
RESOLVANTE_ALTERNEE1, RESOLVANTE_KLEIN, RESOLVANTE_KLEIN3,

     RESOLVANTE_VIERER, RESOLVANTE_DIEDRALE.

Function RESOLVANTE UNITAIRE (p,q,x)

calcule la r\’esolvante du polyn\^ome p(x) par le polyn\^ome q(x). Voir e’galement :

     RESOLVANTE_PRODUIT_SYM, RESOLVANTE,
     RESOLVANTE_ALTERNEE1, RESOLVANTE_KLEIN, RESOLVANTE_KLEIN3,
     RESOLVANTE_VIERER, RESOLVANTE_DIEDRALE.

Function RESOLVANTE VIERER (p,x)

calcule la transformation de p(x) par la fonction x 1x 2-x 3x 4. Voir e’galement :

     RESOLVANTE_PRODUIT_SYM, RESOLVANTE_UNITAIRE,
     RESOLVANTE_ALTERNEE1, RESOLVANTE_KLEIN, RESOLVANTE_KLEIN3,
     RESOLVANTE, RESOLVANTE_DIEDRALE.

Function SCHUR2COMP (P,l var)

: P est un polyno^mes en les variables contenues dans la liste l var. Chacune des variables de l var repre’sente une fonction syme’trique comple‘te. On repre’sente dans l var la ie‘me fonction syme’trique comple‘te comme la concate’nation de la lettre h avec l’entier i : hi. Cette fonction donne l’expression de P en fonction des fonctions de Schur.

       SCHUR2COMP(h1*h2-h3,[h1,h2,h3]);

           s
          1, 2
      SCHUR2COMP(a*h3,[h3]);
             s  a
            3

Function SOMRAC (liste,K)

la liste contient les fonctions syme’triques e’le’mentaires d’un polyno^me P . On calcul le polyno^mes dont les racines sont les sommes K a‘ K distinctes des racines de P. Voir e’galement PRODRAC.

Function TCONTRACT (pol,lvar)

teste si le polyno^me pol est syme’trique en les variables contenues dans la liste lvar. Si oui il rend une forme contracte’e comme la fonction CONTRACT.

Autres fonctions de changements de repre’sentations : CONTRACT, CONT2PART, EXPLOSE, PART2CONT, PARTPOL, TPARTPOL.

Function TPARTPOL (pol,lvar)

teste si le polyno^me pol est syme’trique en les variables contenues dans la liste lvar. Si oui il rend sa forme partionne’e comme la fonction PARTPOL. Autres fonctions de changements de repre’sentations : CONTRACT, CONT2PART, EXPLOSE, PART2CONT, PARTPOL, TCONTRACT.

Function TREILLIS (n)

rame‘ne toutes les partitions de poids n.

      treillis(4);
                 [[4], [3, 1], [2, 2], [2, 1, 1], [1, 1, 1, 1]]

Voir e’galement : LGTREILLIS, LTREILLIS et TREINAT.

Function TREINAT

TREINAT(part) rame‘ne la liste des partitions infe’rieures a‘ la partition part pour l’ordre naturel.

        treinat([5]);
                                    [[5]]
        treinat([1,1,1,1,1]);
         [[5], [4, 1], [3, 2], [3, 1, 1], [2, 2, 1], [2, 1, 1, 1],
          [1, 1, 1, 1, 1]]
        treinat([3,2]);
                             [[5], [4, 1], [3, 2]]

Voir e’galement : LGTREILLIS, LTREILLIS et TREILLIS.

18 Groups

18.1 Definitions for Groups

Function TODD COXETER (relations,subgroup)

Find the order of G/H where G is the Free Group modulo RELATIONS, and H is the subgroup of G generated by SUBGROUP. SUBGROUP is an optional argument, defaulting to []. In doing this it produces a multiplication table for the right action of G on G/H, where the cosets are enumerated [H,Hg2,Hg3,...] This can be seen internally in the $todd coxeter state. The multiplication tables for the variables are in table:todd coxeter state[2] Then table[i] gives the table for the ith variable. mulcoset(coset,i) := table[varnum][coset]; Example:

(C1) symet(n):=create_list(if (j - i) = 1 then (p(i,j))^^3 else
     if (not i = j) then (p(i,j))^^2 else p(i,i) , j,1,n-1,i,1,j);
      <3>
(D1) SYMET(N) := CREATE_LIST(IF J - I = 1 THEN P(I, J)
       <2>
 ELSE (IF NOT I = J THEN P(I, J)
 ELSE P(I, I)), J, 1, N - 1, I, 1, J)
(C2) p(i,j) :=concat(x,i).concat(x,j);
(D2)                P(I, J) := CONCAT(X, I) . CONCAT(X, J)
(C3) symet(5);
                        <3>                    <2>           <3>
(D3) [X1 . X1, (X1 . X2)   , X2 . X2, (X1 . X3)   , (X2 . X3)   ,
                      <2>                <2>           <3>
X3 . X3,     (X1 . X4)  , (X2 . X4)   , (X3 . X4)   , X4 . X4]
(C4) todd_coxeter(d3);
Rows tried 426
(D4)        120
(C5) todd_coxeter(d3,[x1]);
Rows tried 213
(D5)        60
(C6) todd_coxeter(d3,[x1,x2]);
Rows tried 71
(D6)        20
(C7) table:todd_coxeter_state[2]$
(C8) table:todd_coxeter_state[2]$
(C9) table[1];
(D9) {Array: FIXNUM #(0 2 1 3 7 6 5 4 8 11 17 9 12 14 13 20
           16 10 18 19 15 0 0 0 0 0 0 0 0  0 0 0 0 0 0 0)}

148 Maxima Manual Note only the elements 1 thru 20 of this array d9 are meaningful. table[1][4] = 7 indicates coset4.var1 = coset7

19 Operators

19.1 NARY

- An NARY operator is used to denote a function of any number of arguments, each of which is separated by an occurrence of the operator, e.g. A+B or A+B+C. The NARY("x") function is a syntax extension function to declare x to be an NARY operator. Do DE-SCRIBE(SYNTAX); for more details. Functions may be DECLAREd to be NARY. If DECLARE(J,NARY); is done, this tells the simplifier to simplify, e.g. J(J(A,B),J(C,D)) to J(A, B, C, D).

19.2 NOFIX

- NOFIX operators are used to denote functions of no arguments. The mere presence of such an operator in a command will cause the corresponding function to be evaluated. For example, when one types "exit;" to exit from a MACSYMA break, "exit" is behaving similar to a NOFIX operator. The function NOFIX("x") is a syntax extension function which declares x to be a NOFIX operator. Do DESCRIBE(SYNTAX); for more details.

19.3 OPERATOR

- See OPERATORS

19.4 POSTFIX

- POSTFIX operators like the PREFIX variety denote functions of a single argument, but in this case the argument immediately precedes an occurrence of the operator in the input string, e.g. 3! . The POSTFIX("x") function is a syntax extension function to declare x to be a POSTFIX operator. Do DESCRIBE(SYNTAX); for details.

19.5 PREFIX

- A PREFIX operator is one which signifies a function of one argument, which argument immediately follows an occurrence of the operator. PREFIX("x") is a syntax extension function to declare x to be a PREFIX operator. Do DESCRIBE(SYNTAX); for more details.

19.6 Definitions for Operators

operator "!"

The factorial operator, which is the product of all the integers from 1 up to its argument. Thus 5! = 1*2*3*4*5 = 120. The value of /the option FACTLIM (default: [-1]) gives the highest factorial which is automatically expanded. If it is -1 then all integers are expanded. See also the FACTORIAL, MINFACTORIAL, and FACTCOMB commands.

operator "!!"

Stands for double factorial which is defined as the product of all the consecutive odd (or even) integers from 1 (or 2) to the odd (or even) argument. Thus 8!! is 2*4*6*8 = 384.

operator "#"

The logical operator "Not equals".

operator "."

The dot operator, for matrix (non-commutative) multiplication. When "." is used in this way, spaces should be left on both sides of it, e.g. A . B. This distinguishes it plainly from a decimal point in a floating point number. Do APROPOS(DOT); for a list of the switches which affect the dot operator. DESCRIBE(switch-name); will explain them.

operator ":"

The assignment operator. E.g. A:3 sets the variable A to 3.

operator "::"

Assignment operator. :: assigns the value of the expression on its right to the value of the quantity on its left, which must evaluate to an atomic variable or subscripted variable.

operator "::="

The "::=" is used instead of ":=" to indicate that what follows is a macro definition, rather than an ordinary functional definition. See DESCRIBE(MACROS).

operator ":="

The function definition operator. E.g. F(X):=SIN(X) defines a function F.

operator "="

denotes an equation to MACSYMA. To the pattern matcher in MACSYMA it denotes a total relation that holds between two expressions if and only if the expressions are syntactically identical.

special symbol ADDITIVE

- If DECLARE(F,ADDITIVE) has been executed, then: (1) If F is univariate, whenever the simplifier encounters F applied to a sum, F will be distributed over that sum. I.e. F(Y+X); will simplify to F(Y)+F(X). (2) If F is a function of 2 or more arguments, additivity is defined as additivity in the first argument to F, as in the case of ’SUM or ’INTEGRATE, i.e. F(H(X)+G(X),X); will simplify to F(H(X),X)+F(G(X),X). This simplification does not occur when F is applied to expressions of the form SUM(X[I],I,lower-limit,upper-limit).

keyword ALLBUT

works with the PART commands (i.e. PART, INPART, SUBSTPART, SUBSTIN-PART, DPART, and LPART). For example,

     if EXPR is E+D+C+B+A,
     then PART(EXPR,[2,5]);
     ==> D+A

while

     PART(EXPR,ALLBUT(2,5))==>E+C+B

It also works with the KILL command,

     KILL(ALLBUT(name1,...,namek))

will do a KILL(ALL) except it will not KILL the names specified. Note: namei means a name such as function name such as U, F, FOO, or G, not an infolist such as FUNCTIONS.

declaration ANTISYMMETRIC

- If DECLARE(H,ANTISYMMETRIC); is done, this tells the simplifier that H is antisymmetric. E.g. H(X,Z,Y) will simplify to - H(X, Y, Z). That is, it will give (-1)^n times the result given by SYMMETRIC or COMMUTATIVE, where n is the number of interchanges of two arguments necessary to convert it to that form.

Function CABS (exp)

returns the complex absolute value (the complex modulus) of exp.

declaration COMMUTATIVE

- If DECLARE(H,COMMUTATIVE); is done, this tells the simplifier that H is a commutative function. E.g. H(X,Z,Y) will simplify to H(X, Y, Z). This is the same as SYMMETRIC.

Function ENTIER (X)

largest integer <= X where X is numeric. FIX (as in FIXnum) is a synonym for this, so FIX(X); is precisely the same.

Function EQUAL (expr1,expr2)

used with an "IS", returns TRUE (or FALSE) if and only if expr1 and expr2 are equal (or not equal) for all possible values of their variables (as determined by RAT-SIMP). Thus IS(EQUAL((X+1)**2,X**2+2*X+1)) returns TRUE whereas if X is unbound IS((X+1)**2=X**2+2*X+1) returns FALSE. Note also that IS(RAT(0)=0) gives FALSE but IS(EQUAL(RAT(0),0)) gives TRUE. If a determination can’t be made with EQUAL then a simplified but equivalent form is returned whereas = always causes either TRUE or FALSE to be returned. All variables occurring in exp are presumed to be real valued. EV(exp,PRED) is equivalent to IS(exp).

     (C1) IS(X**2 >= 2*X-1);
     (D1)                               TRUE
     (C2) ASSUME(A>1);
     (D2)                               DONE

     (C3) IS(LOG(LOG(A+1)+1)>0 AND A^2+1>2*A);
     (D3)                               TRUE

Function EVAL

causes an extra post-evaluation of exp to occur.

Function EVENP (exp)

is TRUE if exp is an even integer. FALSE is returned in all other cases.

Function FIX (x)

a synonym for ENTIER(X) - largest integer <= X where X is numeric.

Function FULLMAP (fn, exp1, ...)

is similar to MAP but it will keep mapping down all subexpressions until the main operators are no longer the same. The user should be aware that FULLMAP is used by the MACSYMA simplifier for certain matrix manipulations; thus, the user might see an error message concerning FULLMAP even though FULLMAP was not explicitly called by the user.

     (C1) A+B*C$
     (C2) FULLMAP(G,%);
     (D2)              G(B) G(C) + G(A)
     (C3)  MAP(G,D1);
     (D3)                G(B C) + G(A)

Function FULLMAPL (fn, list1, ...)

is similar to FULLMAP but it only maps onto lists and matrices

     (C1) FULLMAPL("+",[3,[4,5]],[[A,1],[0,-1.5]]);
     (D1)                      [[A + 3, 4], [4, 3.5]]

Function IS (exp)

attempts to determine whether exp (which must evaluate to a predicate) is provable from the facts in the current data base. IS returns TRUE if the predicate is true for all values of its variables consistent with the data base and returns FALSE if it is false for all such values. Otherwise, its action depends on the setting of the switch PREDERROR (default: TRUE). IS errs out if the value of PREDERROR is TRUE and returns UNKNOWN if PREDERROR is FALSE.

Function ISQRT (X)

takes one integer argument and returns the "integer SQRT" of its absolute value.

Function MAX (X1, X2, ...)

yields the maximum of its arguments (or returns a simplified form if some of its arguments are non-numeric).

Function MIN (X1, X2, ...)

yields the minimum of its arguments (or returns a simplified form if some of its arguments are non-numeric).

Function MOD (poly)

converts the polynomial poly to a modular representation with respect to the current modulus which is the value of the variable MODULUS. MOD(poly,m) specifies a MODULUS m to be used for converting poly, if it is desired to override the current global value of MODULUS. See DESCRIBE(MODULUS); .

Function ODDP (exp)

is TRUE if exp is an odd integer. FALSE is returned in all other cases.

operator PRED

(EVFLAG) causes predicates (expressions which evaluate to TRUE or FALSE) to be evaluated.

Function RANDOM (X)

returns a random integer between 0 and X-1. If no argument is given then a random integer between -2^(29) and 2^(29) -1 is returned. If X is FALSE then the random sequence is restarted from the beginning. Note that the range of the returned result when no argument is given differs in NIL MACSYMA from that of PDP-10 and Multics MACSYMA, which is -2^(35) to 2^(35) -1. This range is the range of the FIXNUM datatype of the underlying LISP.

Function SIGN (exp)

attempts to determine the sign of its specified expression on the basis of the facts in the current data base. It returns one of the following answers: POS (positive), NEG (negative), ZERO, PZ (positive or zero), NZ (negative or zero), PN (positive or negative), or PNZ (positive, negative, or zero, i.e. nothing known).

Function SIGNUM (X)

if X<0 then -1 else if X>0 then 1 else 0. If X is not numeric then a simplified but equivalent form is returned. For example, SIGNUM(-X) gives -SIGNUM(X).

Function SORT (list,optional-predicate)

sorts the list using a suitable optional-predicate of two arguments (such as "<" or ORDERLESSP). If the optional-predicate is not given, then MACSYMA’s built-in ordering predicate is used.

Function SQRT (X)

the square root of X. It is represented internally by X^(1/2). Also see ROOTSCON-TRACT. RADEXPAND[TRUE] - if TRUE will cause nth roots of factors of a product which are powers of n to be pulled outside of the radical, e.g. SQRT(16*X^2) will become 4*X only if RADEXPAND is TRUE.

Variable SQRTDISPFLAG

default: [TRUE] - if FALSE causes SQRT to display with exponent 1/2.

Function SUBLIS (list,expr)

allows multiple substitutions into an expression in parallel. Sample syntax:

             SUBLIS([A=B,B=A],SIN(A)+COS(B));
              => SIN(B) + COS(A)

The variable SUBLIS APPLY LAMBDA[TRUE] controls simplification after SUB-LIS. For full documentation, see the file SHARE2;SUBLIS INFO.

Function SUBLIST (L,F)

returns the list of elements of the list L for which the function F returns TRUE. E.g., SUBLIST([1,2,3,4],EVENP); returns [2,4].

Variable SUBLIS APPLY LAMBDA

default:[TRUE] - controls whether LAMBDA’s substituted are applied in simplification after SUBLIS is used or whether you have to do an EV to get things to apply. TRUE means do the application.

Function SUBST (a, b, c)

substitutes a for b in c. b must be an atom, or a complete subexpression of c. For example, X+Y+Z is a complete subexpression of 2*(X+Y+Z)/W while X+Y is not. When b does not have these characteristics, one may sometimes use SUB-STPART or RATSUBST (see below). Alternatively, if b is of the form e/f then one could use SUBST(a*f,e,c) while if b is of the form e**(1/f) then one could use SUBST(a**f,e,c). The SUBST command also discerns the X^Y in X^-Y so that SUBST(A,SQRT(X),1/SQRT(X)) yields 1/A. a and b may also be operators of an expression enclosed in "s or they may be function names. If one wishes to substitute for the independent variable in derivative forms then the AT function (see below) should be used. Note: SUBST is an alias for SUBSTITUTE. SUBST(eq1,exp) or SUBST([eq1,...,eqk],exp) are other permissible forms. The eqi are equations indicating substitutions to be made. For each equation, the right side will be substituted for the left in the expression exp. EXPTSUBST[FALSE] if TRUE permits substitutions like Y for %E**X in %E**(A*X) to take place. OPSUBST[TRUE] if FALSE, SUBST will not attempt to substitute into the operator of an expression. E.g. (OP-SUBST:FALSE, SUBST(X^2,R,R+R[0])); will work.

     (C1) SUBST(A,X+Y,X+(X+Y)**2+Y);
                                      2
     (D1)                    Y + X + A
     (C2) SUBST(-%I,%I,A+B*%I);

(D2)                             A - %I B

(Note that C2 is one way of obtaining the complex conjugate of an analytic expression.) For further examples, do EXAMPLE(SUBST);

Function SUBSTINPART (x, exp, n1, ...)

is like SUBSTPART but works on the internal representation of exp.

(C1) X.'DIFF(F(X),X,2);
                                   2
                                  d
(D1)                         X . (--- F(X))
                                    2
                                  dX
(C2) SUBSTINPART(D**2,%,2);
                                      2
(D2)                             X . D
(C3) SUBSTINPART(F1,F[1](X+1),0);
(D3)                            F1(X + 1)
              Additional Information
If the last argument to a part function is a list of indices then
several subexpressions are picked out, each one corresponding to an
index of the list.  Thus PART(X+Y+Z,[1,3]) is Z+X.
PIECE holds the value of the last expression selected when using the
part functions.  It is set during the execution of the function and
thus may be referred to in the function itself as shown below.
If PARTSWITCH[FALSE] is set to TRUE then END is returned when a
selected part of an expression doesn't exist, otherwise an error
message is given.
(C1)  27*Y**3+54*X*Y**2+36*X**2*Y+Y+8*X**3+X+1;
            3         2       2            3
(D1)    27 Y  + 54 X Y  + 36 X  Y + Y + 8 X  + X + 1
(C2)  PART(D1,2,[1,3]);
                  2
(D2)          54 Y
(C3)  SQRT(PIECE/54);
(D3)            Y
(C4)  SUBSTPART(FACTOR(PIECE),D1,[1,2,3,5]);
                          3
(D4)           (3 Y + 2 X) + Y + X + 1
(C5) 1/X+Y/X-1/Z;
                                 1   Y   1
(D5)                           - - + - + -
                                 Z   X   X
(C6) SUBSTPART(XTHRU(PIECE),%,[2,3]);
                                Y + 1   1
(D6)                            ----- - -
                                  X     Z

Also, setting the option INFLAG to TRUE and calling PART/SUBSTPART is the same as calling INPART/SUBSTINPART.

Function SUBSTPART (x, exp, n1, ..., nk)

substitutes x for the subexpression picked out by the rest of the arguments as in PART. It returns the new value of exp. x may be some operator to be substituted for an operator of exp. In some cases it needs to be enclosed in "s (e.g. SUBST-PART("+",A*B,0); -> B + A ).

     (C1) 1/(X**2+2);
                                         1
     (D1)                              ------
                                        2
                                       X  + 2
     (C2) SUBSTPART(3/2,%,2,1,2);
                                         1
     (D2)                             --------
                                       3/2
                                      X    + 2
     (C3) A*X+F(B,Y);
     (D3)                          A X + F(B, Y)
     (C4) SUBSTPART("+",%,1,0);
     (D4)                         X + F(B, Y) + A

Also, setting the option INFLAG to TRUE and calling PART/SUBSTPART is the same as calling INPART/SUBSTINPART.

Function SUBVARP (exp)

is TRUE if exp is a subscripted variable, for example A[I].

Function SYMBOLP (exp)

returns TRUE if "exp" is a "symbol" or "name", else FALSE. I.e., in effect, SYM-BOLP(X):=ATOM(X) AND NOT NUMBERP(X)$ .

Function UNORDER ()

stops the aliasing created by the last use of the ordering commands ORDERGREAT and ORDERLESS. ORDERGREAT and ORDERLESS may not be used more than one time each without calling UNORDER. Do DESCRIBE(ORDERGREAT); and DESCRIBE(ORDERLESS);, and also do EXAMPLE(UNORDER); for specifics.

Function VECTORPOTENTIAL (givencurl)

Returns the vector potential of a given curl vector, in the current coordinate system. POTENTIALZEROLOC has a similar role as for POTENTIAL, but the order of the left-hand sides of the equations must be a cyclic permutation of the coordinate variables.

Function XTHRU (exp)

combines all terms of exp (which should be a sum) over a common denominator without expanding products and exponentiated sums as RATSIMP does. XTHRU cancels common factors in the numerator and denominator of rational expressions but only if the factors are explicit. Sometimes it is better to use XTHRU before RATSIMPing an expression in order to cause explicit factors of the gcd of the numerator and denominator to be canceled thus simplifying the expression to be RATSIMPed.

     (C1) ((X+2)**20-2*Y)/(X+Y)**20+(X+Y)**-19-X/(X+Y)**20;
                                                     20
                          1           X       (X + 2)   - 2 Y
     (D1)             --------- - --------- + ---------------
                             19          20             20
                      (Y + X)     (Y + X)        (Y + X)
     (C2) XTHRU(%);
                                          20
                                   (X + 2)   - Y
     (D2)                          -------------
                                            20
                                     (Y + X)

Function ZEROEQUIV (exp,var)

tests whether the expression exp in the variable var is equivalent to zero. It returns either TRUE, FALSE, or DONTKNOW. For example ZEROEQUIV(SIN(2*X)

- 2*SIN(X)*COS(X),X) returns TRUE and ZEROEQUIV(%E^X+X,X) returns FALSE. On the other hand ZEROEQUIV(LOG(A*B) - LOG(A) - LOG(B),A) will return DONTKNOW because of the presence of an extra parameter. The restrictions are: (1) Do not use functions that MACSYMA does not know how to differentiate and evaluate. (2) If the expression has poles on the real line, there may be errors in the result (but this is unlikely to occur). (3) If the expression contains functions which are not solutions to first order differential equations (e.g. Bessel functions) there may be incorrect results. (4) The algorithm uses evaluation at randomly chosen points for carefully selected subexpressions. This is always a somewhat hazardous business, although the algorithm tries to minimize the potential for error.

158 Maxima Manual

20 Constants

20.1 Definitions for Constants

Variable E

- The base of natural logarithms, e, is represented in MACSYMA as %E.

Variable FALSE

- the Boolean constant, false. (NIL in LISP)

Variable MINF

- real minus infinity.

Variable PI

- "pi" is represented in MACSYMA as %PI.

Variable TRUE

- the Boolean constant, true. (T in LISP)

160 Maxima Manual

21 Special Functions

21.1 Introduction to Special Functions

%J [index](expr) - Bessel Funct 1st Kind (in SPECINT)

%K [index](expr)

Bessel Funct 2nd Kind (in SPECINT)

Constant, in ODE2

21.2 GAMALG

- A Dirac gamma matrix algebra program which takes traces of and does manipulations on gamma matrices in n dimensions. It may be loaded into MACSYMA by LOAD-FILE("gam"); A preliminary manual is contained in the file SHARE;GAM USAGE and may be printed using PRINTFILE(GAM,USAGE,SHARE);

21.3 SPECINT

- The Hypergeometric Special Functions Package HYPGEO is still under development. At the moment it will find the Laplace Transform or rather, the integral from 0 to INF of some special functions or combinations of them. The factor, EXP(-P*var) must be explicitly stated. The syntax is as follows: SPECINT(EXP(-P*var)*expr,var); where var is the variable of integration and expr may be any expression containing special functions (at your own risk). Special function notation follows:

%J[index](expr)         Bessel Funct 1st Kind
%K[index](expr)           "     "    2nd Kind
%I[     ](    )         Modified Bessel
%HE[     ](  )          Hermite Poly
%P[  ]( )               Legendre Funct
%Q[  ]( )               Legendre of second kind
HSTRUVE[ ]( )           Struve H Function
LSTRUVE[ ]( )             "    L Function
%F[ ]([],[],expr)       Hypergeometric Function
GAMMA()
GAMMAGREEK()
GAMMAINCOMPLETE()
SLOMMEL
%M[]()                  Whittaker Funct 1st Kind
%W[]()                     "       "    2nd  "

For a better feeling for what it can do, do DEMO(HYPGEO,DEMO,SHARE1); .

21.4 Definitions for Special Functions

Function AIRY (X)

returns the Airy function Ai of real argument X. The file SHARE1;AIRY FASL contains routines to evaluate the Airy functions Ai(X), Bi(X), and their derivatives dAi(X), dBi(X). Ai and Bi satisfy the AIRY eqn diff(y(x),x,2)-x*y(x)=0. Read SHARE1;AIRY USAGE for details.

Function ASYMP

- A preliminary version of a program to find the asymptotic behavior of Feynman diagrams has been installed on the SHARE1; directory. For further information, see the file SHARE1;ASYMP USAGE. (For Asymptotic Analysis functions, see ASYMPA.)

Function ASYMPA

- Asymptotic Analysis - The file SHARE1;ASYMPA > contains simplification functions for asymptotic analysis, including the big-O and little-o functions that are widely used in complexity analysis and numerical analysis. Do BATCH("asympa.mc"); . (For asymptotic behavior of Feynman diagrams, see ASYMP.)

Function BESSEL (Z,A)

returns the Bessel function J for complex Z and real A > 0.0 . Also an array BESSE-LARRAY is set up such that BESSELARRAY[I] = J[I+A- ENTIER(A)](Z).

Function BETA (X, Y)

same as GAMMA(X)*GAMMA(Y)/GAMMA(X+Y).

Function GAMMA (X)

the gamma function. GAMMA(I)=(I-1)! for I a positive integer. For the Euler-Mascheroni constant, see %GAMMA. See also the MAKEGAMMA function. The variable GAMMALIM[1000000] (which see) controls simplification of the gamma function.

Variable GAMMALIM

default: [1000000] controls simplification of the gamma function for integral and rational number arguments. If the absolute value of the argument is not greater than GAMMALIM, then simplification will occur. Note that the FACTLIM switch controls simplification of the result of GAMMA of an integer argument as well.

Function INTOPOIS (A)

converts A into a Poisson encoding.

Function MAKEFACT (exp)

transforms occurrences of binomial,gamma, and beta functions in exp to factorials.

Function MAKEGAMMA (exp)

transforms occurrences of binomial,factorial, and beta functions in exp to gamma functions.

Function NUMFACTOR (exp)

gives the numerical factor multiplying the expression exp which should be a single term. If the gcd of all the terms in a sum is desired the CONTENT function may be used.

     (C1) GAMMA(7/2);
     (D1)               15 SQRT(%PI)
                        ------------
                             8
     (C2) NUMFACTOR(%)
                         15
     (D2)                --
                          8

Function OUTOFPOIS (A)

converts A from Poisson encoding to general representation. If A is not in Poisson form, it will make the conversion, i.e. it will look like the result of OUTOF-POIS(INTOPOIS(A)). This function is thus a canonical simplifier for sums of powers of SIN’s and COS’s of a particular type.

Function POISDIFF (A, B)

differentiates A with respect to B. B must occur only in the trig arguments or only in the coefficients.

Function POISEXPT (A, B)

B a positive integer) is functionally identical to INTOPOIS(A**B).

Function POISINT (A, B)

integrates in a similarly restricted sense (to POISDIFF). Non-periodic terms in B are dropped if B is in the trig arguments.

Variable POISLIM

default: [5] - determines the domain of the coefficients in the arguments of the trig functions. The initial value of 5 corresponds to the interval [-2^(5-1)+1,2^(5-1)], or [-15,16], but it can be set to [-2^(n-1)+1, 2^(n-1)].

Function POISMAP (series, sinfn, cosfn)

will map the functions sinfn on the sine terms and cosfn on the cosine terms of the poisson series given. sinfn and cosfn are functions of two arguments which are a coefficient and a trigonometric part of a term in series respectively.

Function POISPLUS (A, B)

is functionally identical to INTOPOIS(A+B).

Function POISSIMP (A)

converts A into a Poisson series for A in general representation.

special symbol POISSON

- The Symbol /P/ follows the line label of Poisson series expressions.

Function POISSUBST (A, B, C)

substitutes A for B in C. C is a Poisson series. (1) Where B is a variable U, V, W, X, Y, or Z then A must be an expression linear in those variables (e.g. 6*U+4*V). (2) Where B is other than those variables, then A must also be free of those variables, and furthermore, free of sines or cosines. POISSUBST(A, B, C, D, N) is a special type of substitution which operates on A and B as in type (1) above, but where D is a Poisson series, expands COS(D) and SIN(D) to order N so as to provide the result of substituting A+D for B in C. The idea is that D is an expansion in terms of a small parameter. For example, POISSUBST(U,V,COS(V),E,3) results in COS(U)*(1-E^2/2) - SIN(U)*(E-E^3/6).

Function POISTIMES (A, B)

is functionally identical to INTOPOIS(A*B).

Function POISTRIM ()

is a reserved function name which (if the user has defined it) gets applied during Poisson multiplication. It is a predicate function of 6 arguments which are the coefficients of the U, V,..., Z in a term. Terms for which POISTRIM is TRUE (for the coefficients of that term) are eliminated during multiplication.

Function PRINTPOIS (A)

prints a Poisson series in a readable format. In common with OUTOFPOIS, it will convert A into a Poisson encoding first, if necessary.

Function PSI (X)

derivative of LOG(GAMMA(X)). At this time, MACSYMA does not have numerical evaluation capabilities for PSI. For information on the PSI[N](X) notation, see POLYGAMMA.

22 Command Line

22.1 Introduction to Command Line

Function %TH (i)

is the ith previous computation. That is, if the next expression to be computed is D(j) this is D(j-i). This is useful in BATCH files or for referring to a group of D expressions. For example, if SUM is initialized to 0 then FOR I:1 THRU 10 DO SUM:SUM+%TH(I) will set SUM to the sum of the last ten D expressions.

operator "’"

- (single quote) has the effect of preventing evaluation. E.g. ’(F(X)) means do not evaluate the expression F(X). ’F(X) means return the noun form of F applied to [X].

operator "”"

- (two single quotes) causes an extra evaluation to occur. E.g. ”c4; will re-execute line C4. ”(F(X)) means evaluate the expression F(X) an extra time. ”F(X) means return the verb form of F applied to [X].

22.2 Definitions for Command Line

Function ALIAS (newname1, oldname1, newname2, oldname2, ...)

provides an alternate name for a (user or system) function, variable, array, etc. Any even number of arguments may be used.

Function DEBUG ()

LISPDEBUGMODE(); DEBUGPRINTMODE(); and DEBUG(); make available to the user debugging features used by systems programmers. These tools are powerful, and although some conventions are different from the usual macsyma level it is felt their use is very intuitive. [Some printout may be verbose for slow terminals, there are switches for controlling this.] These commands were designed for the user who must debug translated macsyma code, as such they are a boon. See MACDOC;TRDEBG USAGE for more information. For more help, consult GJC.

Variable DEBUGMODE

default: [FALSE] - causes MACSYMA to enter a MACSYMA break loop whenever a MACSYMA error occurs if it is TRUE and to terminate that mode if it is FALSE. If it is set to ALL then the user may examine BACKTRACE for the list of functions currently entered.

Function DEBUGPRINTMODE ()

LISPDEBUGMODE(); DEBUGPRINTMODE(); and DEBUG(); make available to the user debugging features used by systems programmers. These tools are powerful, and although some conventions are different from the usual macsyma level it is felt their use is very intuitive. [Some printout may be verbose for slow terminals, there are switches for controlling this.] These commands were designed for the user who must debug translated macsyma code, as such they are a boon. See MACDOC;TRDEBG USAGE for more information. For more help, consult GJC.

Function EV (exp, arg1, ..., argn)

is one of MACSYMA’s most powerful and versatile commands. It evaluates the expression exp in the environment specified by the argi. This is done in steps, as follows:

• (1) First the environment is set up by scanning the argi which may be as follows: SIMP causes exp to be simplified regardless of the setting of the switch SIMP which inhibits simplification if FALSE. NOEVAL supresses the evaluation phase of EV (see step (4) below). This is useful in conjunction with the other switches and in causing exp to be resimplified without being reevaluated. EXPAND causes expansion. EXPAND(m,n) causes expansion, setting the values of MAXPOSEX and MAXNEGEX to m and n respectively. DETOUT causes any matrix inverses computed in exp to have their determinant kept outside of the inverse rather than dividing through each element. DIFF causes all differentiations indicated in exp to be performed. DERIVLIST(var1,...,vark) causes only differentiations with respect to the indicated variables. FLOAT causes non-integral rational numbers to be converted to floating point. NUMER causes some mathematical functions (including exponentiation) with numerical arguments to be evaluated in floating point. It causes variables in exp which have been given numervals to be replaced by their values. It also sets the FLOAT switch on. PRED causes predicates (expressions which evaluate to TRUE or FALSE) to be evaluated. EVAL causes an extra post-evaluation of exp to occur. (See step (5) below.) E where E is an atom declared to be an EVFLAG causes E to be bound to TRUE during the evaluation of exp. V:expression (or alternately V=expression) causes V to be bound to the value of expression during the evaluation of exp. Note that if V is a MACSYMA option, then expression is used for its value during the evaluation of exp. If more than one argument to EV is of this type then the binding is done in parallel. If V is a non-atomic expression then a substitution rather than a binding is performed. E where E, a function name, has been declared to be an EVFUN causes E to be applied to exp. Any other function names (e.g. SUM) cause evaluation of occurrences of those names in exp as though they were verbs. In addition a function occurring in exp (say F(args)) may be defined locally for the purpose of this evaluation of exp by giving F(args):=body as an argument to EV. If an atom not mentioned above or a subscripted variable or subscripted expression was given as an argument, it is evaluated and if the result is an equation or assignment then the indicated binding or substitution is performed. If the result is a list then the members of the list are treated as if they were additional arguments given to EV. This permits a list of equations to be given (e.g. [X=1, Y=A**2] ) or a list of names of equations (e.g. [E1,E2] where E1 and E2 are equations) such as that returned by SOLVE. The argi of EV may be given in any order with the exception of substitution equations which are handled in sequence, left to right, and EVFUNS which are composed, e.g. EV(exp,RATSIMP,REALPART) is handled as REALPART(RATSIMP(exp)). The SIMP, NUMER, FLOAT, and PRED switches may also be set locally in a block, or globally at the "top level" in MACSYMA so that they will remain in effect until being reset. If exp is in CRE form then EV will return a result in CRE form provided the NUMER and FLOAT switches are not both TRUE.

• (2) During step (1), a list is made of the non-subscripted variables appearing on the left side of equations in the argi or in the value of some argi if the value is an equation. The variables (both subscripted variables which do not have associated array functions, and non-subscripted variables) in the expression exp are replaced by their global values, except for those appearing in this list. Usually, exp is just a label or % (as in (C2) below), so this step simply retrieves the expression named by the label, so that EV may work on it.

• (3) If any substitutions are indicated by the argi, they are carried out now.

• (4) The resulting expression is then re-evaluated (unless one of the argi was NOEVAL) and simplified according the the argi. Note that any function calls in exp will be carried out after the variables in it are evaluated and that EV(F(X)) thus may behave like F(EV(X)).

• (5) If one of the argi was EVAL, steps (3) and (4) are repeated.

                     Examples
(C1) SIN(X)+COS(Y)+(W+1)**2+'DIFF(SIN(W),W);
                        d                 2
(D1)  COS(Y) + SIN(X) + -- SIN(W) + (W + 1)
                        dW
(C2) EV(%,SIN,EXPAND,DIFF,X=2,Y=1);
                    2
(D2)      COS(W) + W  + 2 W + COS(1) + 1.90929742

An alternate top level syntax has been provided for EV, whereby one may just type in its arguments, without the EV(). That is, one may write simply

exp, arg1, ...,argn.

This is not permitted as part of another expression, i.e. in functions, blocks, etc.

(C4) X+Y,X:A+Y,Y:2;
(D4)                Y + A + 2
(Notice the parallel binding process)
(C5) 2*X-3*Y=3$
(C6) -3*X+2*Y=-4$
(C7) SOLVE([D5,D6]);
SOLUTION
                                1
(E7)                     Y =  - -
                                5
                          6
(E8)                  X = -
                          5
(D8)               [E7, E8]
(C9) D6,D8;

     (D9)               - 4 =  - 4
     (C10) X+1/X > GAMMA(1/2);
                         1
     (D10)           X + - > SQRT(%PI)
                         X
     (C11) %,NUMER,X=1/2;
     (D11)            2.5 > 1.7724539
     (C12) %,PRED;
     (D12)                  TRUE

Variable EVFLAG

default: [] - the list of things known to the EV function. An item will be bound to TRUE during the execution of EV if it is mentioned in the call to EV, e.g. EV(%,numer);. Initial evflags are

     FLOAT, PRED, SIMP, NUMER, DETOUT, EXPONENTIALIZE, DEMOIVRE,
     KEEPFLOAT, LISTARITH, TRIGEXPAND, SIMPSUM, ALGEBRAIC,
     RATALGDENOM, FACTORFLAG, %EMODE, LOGARC, LOGNUMER,
     RADEXPAND, RATSIMPEXPONS, RATMX, RATFAC, INFEVAL, %ENUMER,
     PROGRAMMODE, LOGNEGINT, LOGABS, LETRAT, HALFANGLES,
     EXPTISOLATE, ISOLATE_WRT_TIMES, SUMEXPAND, CAUCHYSUM,
     NUMER_PBRANCH, M1PBRANCH, DOTSCRULES, and LOGEXPAND.

Variable EVFUN

- the list of functions known to the EV function which will get applied if their name is mentioned. Initial evfuns are FACTOR, TRIGEXPAND, TRIGREDUCE, BFLOAT, RATSIMP, RATEXPAND, RADCAN, LOGCONTRACT, RECTFORM, and PO-LARFORM.

special symbol INFEVAL

leads to an "infinite evaluation" mode. EV repeatedly evaluates an expression until it stops changing. To prevent a variable, say X, from being evaluated away in this mode, simply include X=’X as an argument to EV. Of course expressions such as EV(X,X=X+1,INFEVAL); will generate an infinite loop. CAVEAT EVALUATOR.

Function KILL (arg1, arg2, ...)

eliminates its arguments from the MACSYMA system. If argi is a variable (including a single array element), function, or array, the designated item with all of its properties is removed from core. If argi=LABELS then all input, intermediate, and output lines to date (but not other named items) are eliminated. If argi=CLABELS then only input lines will be eliminated; if argi=ELABELS then only intermediate E-lines will be eliminated; if argi=DLABELS only the output lines will be eliminated. If argi is the name of any of the other information lists (the elements of the MACSYMA variable INFOLISTS), then every item in that class (and its properties) is KILLed and if argi=ALL then every item on every information list previously defined as well as LABELS is KILLed. If argi=a number (say n), then the last n lines (i.e. the lines with the last n line numbers) are deleted. If argi is of the form [m,n] then all lines with numbers between m and n inclusive are killed. Note that KILL(VALUES) or KILL(variable) will not free the storage occupied unless the labels which are pointing to the same expressions are also KILLed. Thus if a large expression was assigned to X on line C7 one should do KILL(D7) as well as KILL(X) to release the storage occupied. KILL(ALLBUT(name1,...,namek) will do a KILL(ALL) except it will not KILL the names specified. (Note: namei means a name such as U, V, F, G, not an infolist such as FUNCTIONS.) KILL removes all properties from the given argument thus KILL(VALUES) will kill all properties associated with every item on the VALUES list whereas the REMOVE set of functions (REMVALUE,REMFUNCTION,REMARRAY,REMRULE) remove a specific property. Also the latter print out a list of names or FALSE if the specific argument doesn’t exist whereas KILL always has value "DONE" even if the named item doesn’t exist. Note that killing expressions will not help the problem which occurs on MC indicated by "NO CORE - FASLOAD" which results when either too many FASL files have been loaded in or when allocation level has gotten too high. In either of these cases, no amount of killing will cause the size of these spaces to decrease. Killing expressions only causes some spaces to get emptied out but not made smaller.

Function LABELS (char)

takes a char C, D, or E as arg and generates a list of all C-labels, D-labels, or Elabels, respectively. If you’ve generated many E- labels via SOLVE, then

     FIRST(REST(LABELS(C)))

reminds you what the last C-label was. LABELS will take as arg any symbolic name, so if you have reset INCHAR, OUTCHAR, or LINECHAR, it will return the list of labels whose first character matches the first character of the arg you give to LABELS. The variable, LABELS, default: [], is a list of C, D, and E lines which are bound.

Variable LASTTIME

- the time to compute the last expression in milliseconds presented as a list of "time" and "gctime".

Variable LINENUM

- the line number of the last expression.

Variable MYOPTIONS

default: [] - all options ever reset by the user (whether or not they get reset to their default value).

Variable NOLABELS

default: [FALSE] - if TRUE then no labels will be bound except for E lines generated by the solve functions. This is most useful in the "BATCH" mode where it eliminates the need to do KILL(LABELS) in order to free up storage.

Variable OPTIONSET

default: [FALSE] - if TRUE, MACSYMA will print out a message whenever a MAC-SYMA option is reset. This is useful if the user is doubtful of the spelling of some option and wants to make sure that the variable he assigned a value to was truly an option variable.

Function PLAYBACK (arg)

"plays back" input and output lines. If arg=n (a number) the last n expressions (Ci, Di, and Ei count as 1 each) are "played-back", while if arg is omitted, all lines are. If arg=INPUT then only input lines are played back. If arg=[m,n] then all lines with numbers from m to n inclusive are played-back. If m=n then [m] is sufficient for arg. Arg=SLOW places PLAYBACK in a slow-mode similar to DEMO’s (as opposed to the "fast" BATCH). This is useful in conjunction with SAVE or STRINGOUT when creating a secondary-storage file in order to pick out useful expressions. If arg=TIME then the computation times are displayed as well as the expressions. If arg=GCTIME or TOTALTIME, then a complete breakdown of computation times are displayed, as with SHOWTIME:ALL;. Arg=STRING strings-out (see STRING function) all input lines when playing back rather than displaying them. If ARG=GRIND "grind" mode can also be turned on (for processing input lines) (see GRIND). One may include any number of options as in PLAYBACK([5,10],20,TIME,SLOW).

Function PRINTPROPS (a, i)

will display the property with the indicator i associated with the atom a. a may also be a list of atoms or the atom ALL in which case all of the atoms with the given property will be used. For example, PRINTPROPS([F,G],ATVALUE). PRINTPROPS is for properties that cannot otherwise be displayed, i.e. for ATVALUE, ATOMGRAD, GRADEF, and MATCHDECLARE.

Variable PROMPT

default: [ ] is the prompt symbol of the DEMO function, PLAYBACK(SLOW) mode, and (MACSYMA-BREAK).

Function QUIT ()

kills the current MACSYMA but doesn’t affect the user’s other jobs; equivalent to exiting to DCL and stopping the MACSYMA process. One may "quit" to MACSYMA top-level by typing Control-C Control-G; Control-C gets NIL’s interrupt prompt, at which one types either Control-G or just G. Typing X at the Interrupt prompt will cause a quit in a computation started within a MACSYMA-BREAK without disrupting the suspended main computation.

Function REMFUNCTION (f1, f2, ...)

removes the user defined functions f1,f2,... from MACSYMA. If there is only one argument of ALL then all functions are removed.

Function RESET ()

causes all MACSYMA options to be set to their default values. (Please note that this does not include features of terminals such as LINEL which can only be changed by assignment as they are not considered to be computational features of MACSYMA.)

Function RESTORE (file-specification)

reinitializes all quantities filed away by a use of the SAVE or STORE functions, in a prior MACSYMA session, from the file given by file-specification without bringing them into core.

Variable SHOWTIME

default: [FALSE] - if TRUE then the computation time will be printed automatically with each output expression. By setting SHOWTIME:ALL, in addition to the cpu time MACSYMA now also prints out (when not zero) the amount of time spent in garbage collection (gc) in the course of a computation. This time is of course included in the time printed out as "time=" . (It should be noted that since the "time=" time only includes computation time and not any intermediate display time or time it takes to load in out-of-core files, and since it is difficult to ascribe "responsibility" for gc’s, the gctime printed will include all gctime incurred in the course of the computation and hence may in rare cases even be larger than "time=").

Function SSTATUS (feature,package)

- meaning SET STATUS. It can be used to SSTATUS(FEATURE,HACK PACKAGE) so that STATUS(FEATURE,HACK PACKAGE) will then return TRUE. This can be useful for package writers, to keep track of what FEATURES they have loaded in.

Function TOBREAK ()

causes the MACSYMA break which was left by typing TOPLEVEL; to be re-entered. If TOBREAK is given any argument whatsoever, then the break will be exited, which is equivalent to typing TOBREAK() immediately followed by EXIT;.

Function TOPLEVEL ()

During a break one may type TOPLEVEL;. This will cause top-level MACSYMA to be entered recursively. Labels will now be bound as usual. Everything will be identical to the previous top-level state except that the computation which was interrupted is saved. The function TOBREAK() will cause the break which was left by typing TOPLEVEL; to be re-entered. If TOBREAK is given any argument whatsoever, then the break will be exited, which is equivalent to typing TOBREAK() immediately followed by EXIT;.

Function TO LISP ()

enters the LISP system under MACSYMA. This is useful on those systems where control-uparrow is not available for this function.

Variable TTYINTFUN

default: [FALSE] - Governs the function which will be run whenever the Userinterrupt-character is typed. To use this feature, one sets TTYINTFUN (default FALSE meaning feature not in use) to a function of no arguments. Then whenever (e.g.) ^U (control-U) is typed, this function is run. E.g. suppose you have a FOR statement loop which increments I, and you want an easy way of checking on the value of I while the FOR statement is running. You can do: TTYINTFUN:PRINTI$ PRINTI():=PRINT(I)$ , then whenever you type (e.g.) ^U you get the check you want.

Variable TTYINTNUM

default: [21] (the ascii value of Control-U (^U), U being the 21st letter of the alphabet). This controls what character becomes the User-interrupt-character. ^U was chosen for it mnemonic value. Most users should not reset TTYINTNUM unless they are already using ^U for something else.

Variable VALUES

default:[] - all bound atoms, i.e. user variables, not MACSYMA Options or Switches, (set up by : , :: , or functional binding).

23 Contexts

23.1 Definitions for Contexts

Function ACTIVATE (cont1, cont2, ...)

causes the specified contexts conti to be activated. The facts in these contexts are used in making deductions and retrieving information. The facts in these contexts are not listed when FACTS(); is done. The variable ACTIVECONTEXTS is the list of contexts which are active by way of the ACTIVATE function.

Variable ACTIVECONTEXTS

default: [] is a list of the contexts which are active by way of the ACTIVATE function, as opposed to being active because they are subcontexts of the current context.

Function ASSUME (pred1, pred2, ...)

First checks the specified predicates for redundancy and consistency with the current data base. If the predicates are consistent and non-redundant, they are added to the data base; if inconsistent or redundant, no action is taken. ASSUME returns a list whose entries are the predicates added to the data base and the atoms REDUNDANT or INCONSISTENT where applicable.

Variable ASSUMESCALAR

default: [TRUE] - helps govern whether expressions exp for which

     NONSCALARP(exp) is FALSE

are assumed to behave like scalars for certain transformations as follows: Let exp represent any non-list/non-matrix, and [1,2,3] any list or matrix.

     exp.[1,2,3]; ==>
      [exp,2*exp,3*exp]

if ASSUMESCALAR is TRUE or SCALARP(exp) is TRUE or CONSTANTP(exp) is TRUE. If ASSUMESCALAR is TRUE, such expressions will behave like scalars only for the commutative operators, but not for ".". If ASSUMESCALAR is FALSE, such expressions will behave like non-scalars. If ASSUMESCALAR is ALL, such expressions will behave like scalars for all the operators listed above.

Variable ASSUME POS

default:[FALSE] - When using INTEGRATE, etc. one often introduces parameters which are real and positive or one’s calculations can often be constructed so that this is true. There is a switch ASSUME POS (default FALSE) such that if set to TRUE, MACSYMA will assume one’s parameters are positive. The intention here is to cut down on the number of questions MACSYMA needs to ask. Obviously, ASSUME information or any contextual information present will take precedence. The user can control what is considered to be a parameter for this purpose. Parameters by default are those which satisfy SYMBOLP(x) OR SUBVARP(x). The user can change this by setting the option ASSUME POS PRED [default FALSE] to the name of a predicate function of one argument. E.g. if you want only symbols to be parameters, you can do ASSUME POS:TRUE$ ASSUME POS PRED:’SYMBOLP$ SIGN(A); -> POS, SIGN(A[1]); -> PNZ.

Variable ASSUME POS PRED

default:[FALSE] - may be set to one argument to control what will be considered a parameter for the "assumptions" that INTEGRATE will make... see ASSUME and ASSUME POS .

Variable CONTEXT

default: INITIAL. Whenever a user assumes a new fact, it is placed in the context named as the current value of the variable CONTEXT. Similarly, FORGET references the current value of CONTEXT. To change contexts, simply bind CONTEXT to the desired context. If the specified context does not exist it will be created by an invisible call to NEWCONTEXT. The context specified by the value of CONTEXT is automatically activated. (Do DESCRIBE(CONTEXTS); for a general description of the CONTEXT mechanism.)

Variable CONTEXTS

default: [INITIAL,GLOBAL] is a list of the contexts which currently exist, including the currently active context. The context mechanism makes it possible for a user to bind together and name a selected portion of his data base, called a context. Once this is done, the user can have MACSYMA assume or forget large numbers of facts merely by activating or deactivating their context. Any symbolic atom can be a context, and the facts contained in that context will be retained in storage until the user destroys them individually by using FORGET or destroys them as a whole by using KILL to destroy the context to which they belong. Contexts exist in a formal hierarchy, with the root always being the context GLOBAL, which contains information about MACSYMA that some functions need. When in a given context, all the facts in that context are "active" (meaning that they are used in deductions and retrievals) as are all the facts in any context which is an inferior of that context. When a fresh MACSYMA is started up, the user is in a context called INITIAL, which has GLOBAL as a subcontext. The functions which deal with contexts are: FACTS, NEWCONTEXT, SUPCONTEXT, KILLCONTEXT, ACTIVATE, DEACTIVATE, ASSUME, and FORGET.

Function DEACTIVATE (cont1, cont2, ...)

causes the specified contexts conti to be deactivated.

Function FACTS (item)

If ’item’ is the name of a context then FACTS returns a list of the facts in the specified context. If no argument is given, it lists the current context. If ’item’ is not the name of a context then it returns a list of the facts known about ’item’ in the current context. Facts that are active, but in a different context, are not listed.

declaration FEATURES

- MACSYMA has built-in properties which are handled by the data base. These are called FEATURES. One can do DECLARE(N,INTEGER), etc. One can also DECLARE one’s own FEATURES by e.g. DECLARE(INCREASING,FEATURE); which will then allow one to say DECLARE(F,INCREASING);. One can then check if F is INCREASING by using the predicate FEATUREP via FEATUREP(F,INCREASING). There is an infolist FEATURES which is a list of known FEATURES. At present known FEATURES are: INTEGER, NONINTEGER, EVEN, ODD, RATIONAL, IRRATIONAL, REAL, IMAGINARY, COMPLEX, ANALYTIC, INCREASING, DECREASING, ODDFUN, EVENFUN, POSFUN, COMMUTATIVE, LASSOCIA-TIVE, RASSOCIATIVE, SYMMETRIC, and ANTISYMMETRIC. [Note: system "features" may be checked with STATUS(FEATURE,...); See DESCRIBE(STATUS); or DESCRIBE(FEATURE); for details.]

Function FORGET (pred1, pred2, ...)

removes relations established by ASSUME. The predicates may be expressions equivalent to (but not necessarily identical to) those previously ASSUMEd. FORGET(list) is also a legal form.

Function KILLCONTEXT (context1,context2,...,contextn)

kills the specified contexts. If one of them is the current context, the new current context will become the first available subcontext of the current context which has not been killed. If the first available unkilled context is GLOBAL then INITIAL is used instead. If the INITIAL context is killed, a new INITIAL is created, which is empty of facts. KILLCONTEXT doesn’t allow the user to kill a context which is currently active, either because it is a subcontext of the current context, or by use of the function ACTIVATE.

Function NEWCONTEXT (name)

creates a new (empty) context, called name, which has GLOBAL as its only subcontext. The new context created will become the currently active context.

Function SUPCONTEXT (name,context)

will create a new context (called name) whose subcontext is context. If context is not specified, the current context will be assumed. If it is specified, context must exist.

176 Maxima Manual

24 Matrices and Linear Algebra

24.1 Introduction to Matrices and Linear Algebra

24.1.1 DOT

- . The dot operator, for matrix (non-commutative) multiplication. When "." is used in this way, spaces should be left on both sides of it, e.g. A . B. This distinguishes it plainly from a decimal point in a floating point number. Do APROPOS(DOT); for a list of the switches which affect the dot operator.

24.1.2 VECTORS

- The file SHARE;VECT > contains a vector analysis package, share/vect.dem contains a corresponding demonstration, and SHARE;VECT ORTH contains definitions of various orthogonal curvilinear coordinate systems. LOAD(VECT); will load this package for you. The vector analysis package can combine and simplify symbolic expressions including dot products and cross products, together with the gradient, divergence, curl, and Laplacian operators. The distribution of these operators over sums or products is under user control, as are various other expansions, including expansion into components in any specific orthogonal coordinate systems. There is also a capability for deriving the scalar or vector potential of a field. The package contains the following commands: VECTORSIMP, SCALEFAC-TORS, EXPRESS, POTENTIAL, and VECTORPOTENTIAL. Do DESCRIBE(cmd) on these command names, or PRINTFILE(VECT,USAGE,SHARE); for details. Warning: The VECT package declares "." to be a commutative operator.

24.2 Definitions for Matrices and Linear Algebra

Function ADDCOL (M,list1,list2,...,listn)

appends the column(s) given by the one or more lists (or matrices) onto the matrix M.

Function ADDROW (M,list1,list2,...,listn)

appends the row(s) given by the one or more lists (or matrices) onto the matrix M.

Function ADJOINT (matrix)

computes the adjoint of a matrix.

Function AUGCOEFMATRIX ([eq1, ...], [var1, ...])

the augmented coefficient matrix for the variables var1,... of the system of linear equations eq1,.... This is the coefficient matrix with a column adjoined for the constant terms in each equation (i.e. those not dependent upon var1,...). Do EXAM-PLE(AUGCOEFMATRIX); for an example.

Function CHARPOLY (M, var)

computes the characteristic polynomial for Matrix M with respect to var. That is, DE-TERMINANT(M - DIAGMATRIX(LENGTH(M),var)). For examples of this command, do EXAMPLE(CHARPOLY); .

Function COEFMATRIX ([eq1, ...], [var1, ...])

the coefficient matrix for the variables var1,... of the system of linear equations eq1,...

Function COL (M,i)

gives a matrix of the ith column of the matrix M.

Function COLUMNVECTOR (X)

a function in the EIGEN package. Do LOAD(EIGEN) to use it. COLUMNVECTOR takes a LIST as its argument and returns a column vector the components of which are the elements of the list. The first element is the first component,...etc...(This is useful if you want to use parts of the outputs of the functions in this package in matrix calculations.)

Function CONJUGATE (X)

a function in the EIGEN package on the SHARE directory. It returns the complex conjugate of its argument. This package may be loaded by LOAD(EIGEN); . For a complete description of this package, do PRINTFILE("eigen.usg"); .

Function COPYMATRIX (M)

creates a copy of the matrix M. This is the only way to make a copy aside from recreating M elementwise. Copying a matrix may be useful when SETELMX is used.

Function DETERMINANT (M)

computes the determinant of M by a method similar to Gaussian elimination. The form of the result depends upon the setting of the switch RATMX. There is a special routine for dealing with sparse determininants which can be used by setting the switches RATMX:TRUE and SPARSE:TRUE.

Variable DETOUT

default: [FALSE] if TRUE will cause the determinant of a matrix whose inverse is computed to be kept outside of the inverse. For this switch to have an effect DOALLMXOPS and DOSCMXOPS should be FALSE (see their descriptions). Alternatively this switch can be given to EV which causes the other two to be set correctly.

Function DIAGMATRIX (n, x)

returns a diagonal matrix of size n by n with the diagonal elements all x. An identity matrix is created by DIAGMATRIX(n,1), or one may use IDENT(n).

Variable DOALLMXOPS

default: [TRUE] if TRUE all operations relating to matrices are carried out. If it is FALSE then the setting of the individual DOT switches govern which operations are performed.

Variable DOMXEXPT

default: [TRUE] if TRUE,

     %E^MATRIX([1,2],[3,4]) ==>
     MATRIX([%E,%E^2],[%E^3,%E^4])

In general, this transformation affects expressions of the form <base>^<power> where <base> is an expression assumed scalar or constant, and <power> is a list or matrix. This transformation is turned off if this switch is set to FALSE.

Variable DOMXMXOPS

default: [TRUE] if TRUE then all matrix-matrix or matrix-list operations are carried out (but not scalar-matrix operations); if this switch is FALSE they are not.

Variable DOMXNCTIMES

default: [FALSE] Causes non-commutative products of matrices to be carried out.

Variable DONTFACTOR

default: [] may be set to a list of variables with respect to which factoring is not to occur. (It is initially empty). Factoring also will not take place with respect to any variables which are less important (using the variable ordering assumed for CRE form) than those on the DONTFACTOR list.

Variable DOSCMXOPS

default: [FALSE] if TRUE then scalar-matrix operations are performed.

Variable DOSCMXPLUS

default: [FALSE] if TRUE will cause SCALAR + MATRIX to give a matrix answer. This switch is not subsumed under DOALLMXOPS.

Variable DOT0NSCSIMP

default: [TRUE] Causes a non-commutative product of zero and a nonscalar term to be simplified to a commutative product.

Variable DOT0SIMP

default: [TRUE] Causes a non-commutative product of zero and a scalar term to be simplified to a commutative product.

Variable DOT1SIMP

default: [TRUE] Causes a non-commutative product of one and another term to be simplified to a commutative product.

Variable DOTASSOC

default: [TRUE] when TRUE causes (A.B).C to simplify to A.(B.C)

Variable DOTCONSTRULES

default: [TRUE] Causes a non-commutative product of a constant and another term to be simplified to a commutative product. Turning on this flag effectively turns on DOT0SIMP, DOT0NSCSIMP, and DOT1SIMP as well.

Variable DOTDISTRIB

default: [FALSE] if TRUE will cause A.(B+C) to simplify to A.B+A.C

Variable DOTEXPTSIMP

default: [TRUE] when TRUE causes A.A to simplify to A^^2

Variable DOTIDENT

default: [1] The value to be returned by X^^0.

Variable DOTSCRULES

default: [FALSE] when TRUE will cause A.SC or SC.A to simplify to SC*A and A.(SC*B) to simplify to SC*(A.B)

Function ECHELON (M)

produces the echelon form of the matrix M. That is, M with elementary row operations performed on it such that the first non-zero element in each row in the resulting matrix is a one and the column elements under the first one in each row are all zero.

                             [2  1 - A  -5 B ]
     (D2)                    [               ]
                             [A    B      C  ]
     (C3) ECHELON(D2);
                      [      A - 1        5 B      ]
                      [1   - -----      - ---      ]
                      [        2           2       ]
     (D3)             [                            ]
                      [                2 C + 5 A B ]
                      [0     1         ------------]
                      [                       2    ]
                      [                2 B + A  - A]

Function EIGENVALUES (mat)

There is a package on the SHARE; directory which contains functions for computing EIGENVALUES and EIGENVECTORS and related matrix computations. For information on it do PRINTFILE(EIGEN,USAGE,SHARE); . EIGENVALUES(mat) takes a MATRIX as its argument and returns a list of lists the first sublist of which is the list of eigenvalues of the matrix and the other sublist of which is the list of the multiplicities of the eigenvalues in the corresponding order. [ The MACSYMA function SOLVE is used here to find the roots of the characteristic polynomial of the matrix. Sometimes SOLVE may not be able to find the roots of the polynomial;in that case nothing in this package except CONJUGATE, INNERPRODUCT, UNITVECTOR, COLUMNVECTOR and GRAMSCHMIDT will work unless you know the eigenvalues. In some cases SOLVE may generate very messy eigenvalues. You may want to simplify the answers yourself before you go on. There are provisions for this and they will be explained below. ( This usually happens when SOLVE returns a notso-obviously real expression for an eigenvalue which is supposed to be real...)] The EIGENVALUES command is available directly from MACSYMA. To use the other functions you must have loaded in the EIGEN package, either by a previous call to EIGENVALUES, or by doing LOADFILE("eigen"); .

Function EIGENVECTORS (MAT)

takes a MATRIX as its argument and returns a list of lists the first sublist of which is the output of the EIGENVALUES command and the other sublists of which are the eigenvectors of the matrix corresponding to those eigenvalues respectively. This function will work directly from MACSYMA, but if you wish to take advantage of the flags for controlling it (see below), you must first load in the EIGEN package from the SHARE; directory. You may do that by LOADFILE("eigen");. The flags that affect this function are: NONDIAGONALIZABLE[FALSE] will be set to TRUE or FALSE depending on whether the matrix is nondiagonalizable or diagonalizable after an EIGENVECTORS command is executed. HERMITIANMATRIX[FALSE] If set to TRUE will cause the degenerate eigenvectors of the hermitian matrix to be orthogonalized using the Gram-Schmidt algorithm. KNOWNEIGVALS[FALSE] If set to TRUE the EIGEN package will assume the eigenvalues of the matrix are known to the user and stored under the global name LISTEIGVALS. LISTEIGVALS should be set to a list similar to the output of the EIGENVALUES command. ( The MACSYMA function ALGSYS is used here to solve for the eigenvectors. Sometimes if the eigenvalues are messy, ALGSYS may not be able to produce a solution. In that case you are advised to try to simplify the eigenvalues by first finding them using EIGENVALUES command and then using whatever marvelous tricks you might have to reduce them to something simpler. You can then use the KNOWNEIGVALS flag to proceed further. )

Function EMATRIX (m, n, x, i, j)

will create an m by n matrix all of whose elements are zero except for the i,j element which is x.

Function ENTERMATRIX (m, n)

allows one to enter a matrix element by element with MACSYMA requesting values for each of the m*n entries.

     (C1) ENTERMATRIX(3,3);
     Is the matrix  1. Diagonal  2. Symmetric  3. Antisymmetric
      4. General
     Answer 1, 2, 3 or 4
     1;
     Row 1 Column 1:  A;
     Row 2 Column 2:  B;
     Row 3 Column 3:  C;
     Matrix entered.
                                      [ A  0  0 ]
                                      [         ]

     (D1)                             [ 0  B  0 ]
                                      [         ]
                                      [ 0  0  C ]

Function GENMATRIX (array, i2, j2, i1, j1)

generates a matrix from the array using array(i1,j1) for the first (upper-left) element and array(i2,j2) for the last (lower-right) element of the matrix. If j1=i1 then j1 may be omitted. If j1=i1=1 then i1 and j1 may both be omitted. If a selected element of the array doesn’t exist a symbolic one will be used.

     (C1) H[I,J]:=1/(I+J-1)$
     (C2) GENMATRIX(H,3,3);
                                [   1  1]
                                [1  -  -]
                                [   2  3]
                                [       ]
                                [1  1  1]
     (D2)                       [-  -  -]
                                [2  3  4]
                                [       ]
                                [1  1  1]
                                [-  -  -]
                                [3  4  5]

Function GRAMSCHMIDT (X)

a function in the EIGEN package. Do LOAD(EIGEN) to use it. GRAMSCHMIDT takes a LIST of lists the sublists of which are of equal length and not necessarily orthogonal (with respect to the innerproduct defined above) as its argument and returns a similar list each sublist of which is orthogonal to all others. (Returned results may contain integers that are factored. This is due to the fact that the MACSYMA function FACTOR is used to simplify each substage of the Gram-Schmidt algorithm. This prevents the expressions from getting very messy and helps to reduce the sizes of the numbers that are produced along the way.)

Function HACH (a,b,m,n,l)

An implementation of Hacijan’s linear programming algorithm is available by doing BATCH("kach.mc"$. Details of use are available by doing BATCH("kach.dem");

Function IDENT (n)

produces an n by n identity matrix.

Function INNERPRODUCT (X,Y)

a function in the EIGEN package. Do LOAD(EIGEN) to use it. INNERPRODUCT takes two LISTS of equal length as its arguments and returns their inner (scalar) product defined by (Complex Conjugate of X).Y (The "dot" operation is the same as the usual one defined for vectors).

Function INVERT (matrix)

finds the inverse of a matrix using the adjoint method. This allows a user to compute the inverse of a matrix with bfloat entries or polynomials with floating pt. coefficients without converting to cre-form. The DETERMINANT command is used to compute cofactors, so if RATMX is FALSE (the default) the inverse is computed without changing the representation of the elements. The current implementation is inefficient for matrices of high order. The DETOUT flag if true keeps the determinant factored out of the inverse. Note: the results are not automatically expanded. If the matrix originally had polynomial entries, better appearing output can be generated by EXPAND(INVERT(mat)),DETOUT. If it is desirable to then divide through by the determinant this can be accomplished by XTHRU(%) or alternatively from scratch by EXPAND(ADJOINT(mat))/EXPAND(DETERMINANT(mat)). IN-VERT(mat):=ADJOINT(mat)/DETERMINANT(mat). See also DESCRIBE("^^"); for another method of inverting a matrix.

Variable LMXCHAR

default: [[] - The character used to display the (left) delimiter of a matrix (see also RMXCHAR).

Function MATRIX (row1, ..., rown)

defines a rectangular matrix with the indicated rows. Each row has the form of a list of expressions, e.g. [A, X**2, Y, 0] is a list of 4 elements. There are a number of MACSYMA commands which deal with matrices, for example: DETER-MINANT, CHARPOLY, GENMATRIX, ADDCOL, ADDROW, COPYMATRIX, TRANSPOSE, ECHELON, and RANK. There is also a package on the SHARE directory for computing EIGENVALUES. Try DESCRIBE on these for more information. Matrix multiplication is effected by using the dot operator, ".", which is also convenient if the user wishes to represent other non-commutative algebraic operations. The exponential of the "." operation is "^^" . Thus, for a matrix A, A.A = A^^2 and, if it exists, A^^-1 is the inverse of A. The operations +,-,*,** are all element-by-element operations; all operations are normally carried out in full, including the . (dot) operation. Many switches exist for controlling simplification rules involving dot and matrix-list operations. Options Relating to Matrices: LMXCHAR, RMXCHAR, RATMX, LISTARITH, DETOUT, DOALLMXOPS, DOMXEXPT DOMXMXOPS, DOSCMXOPS, DOSCMXPLUS, SCALARMATRIX, and SPARSE. Do DESCRIBE(option) for details on them.

Function MATRIXMAP (fn, M)

will map the function fn onto each element of the matrix M.

Function MATRIXP (exp)

is TRUE if exp is a matrix else FALSE.

Variable MATRIX ELEMENT ADD

default: [+] - May be set to "?"; may also be the name of a function, or a LAMBDA expression. In this way, a rich variety of algebraic structures may be simulated. For more details, do DEMO("matrix.dem1"); and DEMO("matrix.dem2");.

Variable MATRIX ELEMENT MULT

default: [*] - May be set to "."; may also be the name of a function, or a LAMBDA expression. In this way, a rich variety of algebraic structures may be simulated. For more details, do DEMO("matrix.dem1"); and DEMO("matrix.dem2");

Variable MATRIX ELEMENT TRANSPOSE

default: [FALSE] - Other useful settings are TRANSPOSE and NONSCALARS; may also be the name of a function, or a LAMBDA expression. In this way, a rich variety of algebraic structures may be simulated. For more details, do DEMO("matrix.dem1"); and DEMO("matrix.dem2");.

Function MATTRACE (M)

computes the trace [sum of the elements on the main diagonal] of the square matrix M. It is used by NCHARPOLY, an alternative to MACSYMA’s CHARPOLY. It is used by doing LOADFILE("nchrpl");

Function MINOR (M, i, j)

computes the i,j minor of the matrix M. That is, M with row i and column j removed.

Function NCEXPT (A,B)

if an (non-commutative) exponential expression is too wide to be displayed as A^^B it will appear as NCEXPT(A,B).

Function NCHARPOLY (M,var)

finds the characteristic polynomial of the matrix M with respect to var. This is an alternative to MACSYMA’s CHARPOLY. NCHARPOLY works by computing traces of powers of the given matrix, which are known to be equal to sums of powers of the roots of the characteristic polynomial. From these quantities the symmetric functions of the roots can be calculated, which are nothing more than the coefficients of the characteristic polynomial. CHARPOLY works by forming the determinant of VAR * IDENT [N] - A. Thus NCHARPOLY wins, for example, in the case of large dense matrices filled with integers, since it avoids polynomial arithmetic altogether. It may be used by doing LOADFILE("nchrpl");

Function NEWDET (M,n)

also computes the determinant of M but uses the Johnson-Gentleman tree minor algorithm. M may be the name of a matrix or array. The argument n is the order; it is optional if M is a matrix.

declaration NONSCALAR

- makes ai behave as does a list or matrix with respect to the dot operator.

Function NONSCALARP (exp)

is TRUE if exp is a non-scalar, i.e. it contains atoms declared as non-scalars, lists, or matrices.

Function PERMANENT (M,n)

computes the permanent of the matrix M. A permanent is like a determinant but with no sign changes.

Function RANK (M)

computes the rank of the matrix M. That is, the order of the largest non-singular subdeterminant of M. Caveat: RANK may return the wrong answer if it cannot determine that a matrix element that is equivalent to zero is indeed so.

Variable RATMX

default: [FALSE] - if FALSE will cause determinant and matrix addition, subtraction, and multiplication to be performed in the representation of the matrix elements and will cause the result of matrix inversion to be left in general representation. If it is TRUE, the 4 operations mentioned above will be performed in CRE form and the result of matrix inverse will be in CRE form. Note that this may cause the elements to be expanded (depending on the setting of RATFAC) which might not always be desired.

Function ROW (M, i)

gives a matrix of the ith row of matrix M.

Variable SCALARMATRIXP

default: [TRUE] - if TRUE, then whenever a 1 x 1 matrix is produced as a result of computing the dot product of matrices it will be converted to a scalar, namely the only element of the matrix. If set to ALL, then this conversion occurs whenever a 1 x 1 matrix is simplified. If set to FALSE, no conversion will be done.

Function SETELMX (x, i, j, M)

changes the i,j element of M to x. The altered matrix is returned as the value. The notation M[i,j]:x may also be used, altering M in a similar manner, but returning x as the value.

Function SIMILARITYTRANSFORM (MAT)

a function in the EIGEN package. Do LOAD(EIGEN) to use it. SIMILARITY-TRANSFORM takes a MATRIX as its argument and returns a list which is the output of the UNITEIGENVECTORS command. In addition if the flag NON-DIAGONALIZABLE is FALSE two global matrices LEFTMATRIX and RIGHT-MATRIX will be generated. These matrices have the property that LEFTMA-TRIX.MAT.RIGHTMATRIX is a diagonal matrix with the eigenvalues of MAT on the diagonal. If NONDIAGONALIZABLE is TRUE these two matrices will not be generated. If the flag HERMITIANMATRIX is TRUE then LEFTMATRIX is the complex conjugate of the transpose of RIGHTMATRIX. Otherwise LEFTMATRIX is the inverse of RIGHTMATRIX. RIGHTMATRIX is the matrix the columns of which are the unit eigenvectors of MAT. The other flags (see DESCRIBE(EIGENVALUES); and DESCRIBE(EIGENVECTORS);) have the same effects since SIMILARITYTRANS-FORM calls the other functions in the package in order to be able to form RIGHT-MATRIX.

Variable SPARSE

default: [FALSE] - if TRUE and if RATMX:TRUE then DETERMINANT will use special routines for computing sparse determinants.

Function SUBMATRIX (m1, ..., M, n1, ...)

creates a new matrix composed of the matrix M with rows mi deleted, and columns ni deleted.

Function TRANSPOSE (M)

produces the transpose of the matrix M.

Function TRIANGULARIZE (M)

produces the upper triangular form of the matrix M which needn’t be square.

Function UNITEIGENVECTORS (MAT)

a function in the EIGEN package. Do LOAD(EIGEN) to use it. UNITEIGENVEC-TORS takes a MATRIX as its argument and returns a list of lists the first sublist of which is the output of the EIGENVALUES command and the other sublists of which are the unit eigenvectors of the matrix corresponding to those eigenvalues respectively. The flags mentioned in the description of the EIGENVECTORS command have the same effects in this one as well. In addition there is a flag which may be useful : KNOWNEIGVECTS[FALSE] - If set to TRUE the EIGEN package will assume that the eigenvectors of the matrix are known to the user and are stored under the global name LISTEIGVECTS. LISTEIGVECTS should be set to a list similar to the output of the EIGENVECTORS command. (If KNOWNEIGVECTS is set to TRUE and the list of eigenvectors is given the setting of the flag NONDIAGONALIZABLE may not be correct. If that is the case please set it to the correct value. The author assumes that the user knows what he is doing and will not try to diagonalize a matrix the eigenvectors of which do not span the vector space of the appropriate dimension...)

Function UNITVECTOR (X)

a function in the EIGEN package. Do LOAD(EIGEN) to use it. UNITVECTOR takes a LIST as its argument and returns a unit list. (i.e. a list with unit magnitude).

Function VECTORSIMP (vectorexpression)

This function employs additional simplifications, together with various optional expansions according to the settings of the following global flags:

     EXPANDALL, EXPANDDOT, EXPANDDOTPLUS, EXPANDCROSS, EXPANDCROSSPLUS,
     EXPANDCROSSCROSS, EXPANDGRAD, EXPANDGRADPLUS, EXPANDGRADPROD,
     EXPANDDIV, EXPANDDIVPLUS, EXPANDDIVPROD, EXPANDCURL, EXPANDCURLPLUS,
     EXPANDCURLCURL, EXPANDLAPLACIAN, EXPANDLAPLACIANPLUS,
     EXPANDLAPLACIANPROD.

All these flags have default value FALSE. The PLUS suffix refers to employing additivity or distributivity. The PROD suffix refers to the expansion for an operand that is any kind of product. EXPANDCROSSCROSS refers to replacing p~(q~r) with (p.r)*q-(p.q)*r, and EXPANDCURLCURL refers to replacing CURL CURL p with GRAD DIV p + DIV GRAD p. EXPANDCROSS:TRUE has the same effect as EXPANDCROSSPLUS:EXPANDCROSSCROSS:TRUE, etc. Two other flags, EXPANDPLUS and EXPANDPROD, have the same effect as setting all similarly suffixed flags true. When TRUE, another flag named EXPANDLAPLACIANTO-DIVGRAD, replaces the LAPLACIAN operator with the composition DIV GRAD. All of these flags are initially FALSE. For convenience, all of these flags have been declared EVFLAG. For orthogonal curvilinear coordinates, the global variables CO-ORDINATES[[X,Y,Z]], DIMENSION[3], SF[[1,1,1]], and SFPROD[1] are set by the function invocation

Variable VECT CROSS

default:[FALSE] - If TRUE allows DIFF(X~Y,T) to work where ~ is defined in SHARE;VECT (where VECT CROSS is set to TRUE, anyway.)

Function ZEROMATRIX (m,n)

takes integers m,n as arguments and returns an m by n matrix of 0’s.

special symbol "["

- [ and ] are the characters which MACSYMA uses to delimit a list.

188 Maxima Manual

25 Runtime Environment

25.1 Introduction for Runtime Environment

- A file which is loaded automatically for you when you start up a MACSYMA, to customize MACSYMA for you. It is possible to have an init file written as a BATCH file of macsyma commands. We hope this makes it easier for users to customize their macsyma environment. Here is an example init file

             /*-*-macsyma-*-*/
             setup_autoload("share\;bessel",j0,j1,jn);
             showtime:all; comgrind:true;

The strange looking comment at the top of the file "/*-*-macsyma-*-*/" tells that it is a macsyma-language file. Also: "SETUP AUTOLOAD" can be used to make functions in BATCH files autoloading, meaning that you can then use (for instance, here) the functions J0, J1 and Jn from the BESSEL package directly because when you use the function the BESSEL package will be loaded in for you automatically. If the second file name in the argument to SETUP AUTOLOAD is not specified (the preferred usage) then the standard search for second file names of "FASL", "TRLISP", and ">" is done.

25.2 INTERRUPTS

- There are several ways the user can interrupt a MACSYMA computation, usually with a control character. Do DESCRIBE(CHARACTERS); for details. MACSYMA will also be interrupted if ^Z (control-Z) is typed, as this will exit back to Unix shell level Usually Control-C interrupts the computation putting you in a break loop. Typing :t should give you top level maxima back again.

25.3 Definitions for Runtime Environment

Function ALARMCLOCK (arg1, arg2, arg3)

will execute the function of no arguments whose name is arg3 when the time specified by arg1 and arg2 elapses. If arg1 is the atom "TIME" then arg3 will be executed after arg2 seconds of real-time has elapsed while if arg1 is the atom "RUNTIME" then arg3 will be executed after arg2 milliseconds of cpu time. If arg2 is negative then the arg1 timer is shut off.

Function ALLOC

takes any number of arguments which are the same as the replies to the "run out of space" question. It increases allocations accordingly. E.g. If the user knows initially that his problem will require much space, he can say ALLOC(4); to allocate the maximum amount initially. See also the DYNAMALLOC switch.

Function BUG ("message")

similar to mail, sends a message to MACSYMA Mail. This may be used for reporting bugs or suspected bugs in MACSYMA. Expressions may be included by referring to them, outside double quotes, e.g. BUG("I am trying to integrate",D3,"but it asks for more list space. What should I do?");

Function CLEARSCREEN ()

Clears the screen. The same as typing control-L.

Function CONTINUE

- Control-^typed while in MACSYMA causes LISP to be entered. The user can now type any LISP S-expression and have it evaluated. Typing (CONTINUE) or ^G (control-G) causes MACSYMA to be re-entered.

Function DDT ()

Exits from MACSYMA to the operating system level. (The same as control-Z on ITS, or control-C on Tops-20.)

Function DELFILE (file-specification)

will delete the file given by the file-specification (i.e. firstname, secondname, device, user) from the given device.

Function DISKFREE ()

With no args or an arg of TRUE, will return the total number of free blocks of disk space in the system. With an arg of 0, 1, or 13, it will return the number of free blocks of diskspace on the respective disk pack. With an arg of SECONDARY or PRIMARY, it will return the total number of free blocks of disk space on the secondary or primary disk pack respectively.

declaration FEATURE

- A nice adjunct to the system. STATUS(FEATURE) gives you a list of system features. At present the list for MC is: MACSYMA, NOLDMSG, MACLISP, PDP10, BIGNUM, FASLOAD, HUNK, FUNARG, ROMAN, NEWIO, SFA, PAGING, MC, and ITS. Any of these "features" may be given as a second argument to STA-TUS(FEATURE,...); If the specified feature exists, TRUE will be returned, else FALSE. Note: these are system features, and not really "user related". See also DESCRIBE(features); for more user-oriented features.

Function FEATUREP (a,f)

attempts to determine whether the object a has the feature f on the basis of the facts in the current data base. If so, it returns TRUE, else FALSE. See DE-SCRIBE(FEATURES); .

     (C1) DECLARE(J,EVEN)$
     (C2) FEATUREP(J,INTEGER);
     (D2)                                TRUE

Function ROOM ()

types out a verbose description of the state of storage and stack management in the Macsyma. This simply utilizes the Lisp ROOM function. ROOM(FALSE) - types out a very terse description, containing most of the same information.

Function STATUS (arg)

will return miscellaneous status information about the user’s MACSYMA depending upon the arg given. Permissible arguments and results are as follows:

• TIME - the time used so far in the computation.

• DAY - the day of the week.

• DATE - a list of the year, month, and day.

• DAYTIME - a list of the hour, minute, and second.

• RUNTIME - accumulated cpu time times the atom "MILLISECONDS" in the current MACSYMA.

• REALTIME - the real time (in sec) elapsed since the user started up his MAC-SYMA.

• GCTIME - the garbage collection time used so far in the current computation.

• TOTALGCTIME - gives the total garbage collection time used in MACSYMA so far.

• FREECORE - the number of blocks of core your MACSYMA can expand before it runs out of address space. (A block is 1024 words.) Subtracting that value from 250*BLOCKS (the maximum you can get on MC) tells you how many blocks of core your MACSYMA is using up. (A MACSYMA with no "fix" file starts at approx. 191 blocks.)

• FEATURE - gives you a list of system features. At present the list for MC is: MACSYMA, NOLDMSG, MACLISP, PDP10, BIGNUM, FASLOAD, HUNK, FUNARG, ROMAN, NEWIO, SFA, PAGING, MC, and ITS. Any of these "features" may be given as a second argument to STATUS(FEATURE,...); If the specified feature exists, TRUE will be returned, else FALSE. Note: these are system features, and not really "user related".

For information about your files, see the FILEDEFAULTS(); command.

Function TIME (Di1, Di2, ...)

gives a list of the times in milliseconds taken to compute the Di. (Note: the Variable SHOWTIME, default: [FALSE], may be set to TRUE to have computation times printed out with each D-line.)

192 Maxima Manual

26 Miscellaneous Options

26.1 Introduction to Miscellaneous Options

In this section various options are discussed which have a global effect on the operation of maxima. Also various lists such as the list of all user defined functions, are discussed.

26.2 SHARE

- The SHARE directory on MC or on a DEC20 version of MACSYMA contains programs, information files, etc. which are considered to be of interest to the MACSYMA community. Most files on SHARE; are not part of the MACSYMA system per se and must be loaded individually by the user, e.g. LOADFILE("array");. Many files on SHARE; were contributed by MACSYMA users. Do PRINTFILE(SHARE,USAGE,SHARE); for more details and the conventions for contributing to the SHARE directory. For an annotated "table of contents" of the directory, do: PRINTFILE(SHARE,>,SHARE);

26.3 Definitions for Miscellaneous Options

Variable ALIASES

default: [] atoms which have a user defined alias (set up by the ALIAS, ORDER-GREAT, ORDERLESS functions or by DECLAREing the atom a NOUN).

Variable ALLSYM

default: [TRUE] - If TRUE then all indexed objects are assumed symmetric in all of their covariant and contravariant indices. If FALSE then no symmetries of any kind are assumed in these indices. Derivative indices are always taken to be symmetric.

declaration ALPHABETIC

Adds to MACSYMA’s alphabet which initially contains the letters A-Z, % and . Thus, DECLARE("~",ALPHABETIC) enables NEW~VALUE to be used as a name.

Function APROPOS (string)

takes a character string as argument and looks at all the MACSYMA names for ones with that string appearing anywhere within them. Thus, APROPOS(EXP); will return a long list of all the flags and functions which have EXP as part of their names, such as EXPAND, EXP, EXPONENTIALIZE. Thus if you can only remember part of the name of something you can use this command to find the rest of the name. Similarily, you could say APROPOS(TR ); to find a list of many of the switches relating to the TRANSLATOR (most of which begin with TR ).

Function ARGS (exp)

returns a list of the args of exp. I.e. it is essentially equivalent to

     SUBSTPART("[",exp,0)

Both ARGS and SUBSTPART depend on the setting of INFLAG.

Function DUMMY (i1,i2,...)

will set each index i1,i2,... to name of the form !n where n is a positive integer. This guarantees that dummy indices which are needed in forming expressions will not conflict with indices already in use. COUNTER[default 1] determines the numerical suffix to be used in generating the next dummy index. The prefix is determined by the option DUMMYX[!].

Variable GENINDEX

default: [I] is the alphabetic prefix used to generate the next variable of summation when necessary.

Variable GENSUMNUM

[0] is the numeric suffix used to generate the next variable of summation. If it is set to FALSE then the index will consist only of GENINDEX with no numeric suffix.

Variable INF

- real positive infinity.

Variable INFINITY

- complex infinity, an infinite magnitude of arbitrary phase angle. (See also INF and MINF.)

Variable INFOLISTS

default: [] a list of the names of all of the information lists in MACSYMA. These are: LABELS - all bound C,D, and E labels. VALUES - all bound atoms, i.e. user variables, not MACSYMA Options or Switches, (set up by : , :: , or functional binding). FUNCTIONS - all user defined functions (set up by f(x):=...). ARRAYS - declared and undeclared arrays (set up by : , :: , or :=...) MACROS - any Macros defined by the user. MYOPTIONS - all options ever reset by the user (whether or not they get reset to their default value). RULES - user defined pattern matching and simplification rules (set up by TELLSIMP, TELLSIMPAFTER, DEFMATCH, or, DEFRULE.) ALIASES - atoms which have a user defined alias (set up by the ALIAS, ORDER-GREAT, ORDERLESS functions or by DECLAREing the atom a NOUN). DEPEN-DENCIES - atoms which have functional dependencies (set up by the DEPENDS or GRADEF functions). GRADEFS - functions which have user defined derivatives (set up by the GRADEF function). PROPS - atoms which have any property other than those mentioned above, such as atvalues, matchdeclares, etc. as well as properties specified in the DECLARE function. LET RULE PACKAGES - a list of all the userdefined let rule packages plus the special package DEFAULT LET RULE PACKAGE. (DEFAULT LET RULE PACKAGE is the name of the rule package used when one is not explicitly set by the user.)

Function INTEGERP (exp)

is TRUE if exp is an integer else FALSE.

Variable M1PBRANCH

default: [FALSE] - "principal branch for -1 to a power". Quantities such as (-1)^(1/3) [i.e. "odd" rational exponent] and (-1)^(1/4) [i.e. "even" rational exponent] are now handled as indicated in the following chart:

                  DOMAIN:REAL(default)
     (-1)^(1/3):      -1
     (-1)^(1/4):   (-1)^(1/4)
                     DOMAIN:COMPLEX
     M1PBRANCH:FALSE(default)   M1PBRANCH:TRUE
     (-1)^(1/3)               1/2+%i*sqrt(3)/2
     (-1)^(1/4)              sqrt(2)/2+%i*sqrt(2)/2

Function NUMBERP (exp)

is TRUE if exp is an integer, a rational number, a floating point number or a bigfloat else FALSE.

Function PROPERTIES (a)

will yield a list showing the names of all the properties associated with the atom a.

special symbol PROPS

- atoms which have any property other than those explicitly mentioned in IN-FOLISTS, such as atvalues, matchdeclares, etc. as well as properties specified in the DECLARE function.

Function PROPVARS (prop)

yields a list of those atoms on the PROPS list which have the property indicated by prop. Thus PROPVARS(ATVALUE) will yield a list of atoms which have atvalues.

Function PUT (a, p, i)

associates with the atom a the property p with the indicator i. This enables the user to give an atom any arbitrary property.

Function QPUT (a, p, i)

is similar to PUT but it doesn’t have its arguments evaluated.

Function REM (a, i)

removes the property indicated by i from the atom a.

Function REMOVE (args)

will remove some or all of the properties associated with variables or functions. RE-MOVE(a1, p1, a2, p2, ...) removes the property pi from the atom ai. Ai and pi may also be lists as with DECLARE. Pi may be any property e.g. FUNCTION, MODE DECLARE, etc. It may also be TRANSFUN implying that the translated LISP version of the function is to be removed. This is useful if one wishes to have the MACSYMA version of the function executed rather than the translated version. Pi may also be OP or OPERATOR to remove a syntax extension given to ai (see Appendix II). If ai is "ALL" then the property indicated by pi is removed from all atoms which have it. Unlike the more specific remove functions (REMVALUE, RE-MARRAY, REMFUNCTION, and REMRULE) REMOVE does not indicate when a given property is non-existent; it always returns "DONE".

Function REMVALUE (name1, name2, ...)

removes the values of user variables (which can be subscripted) from the system. If name is ALL then the values of all user variables are removed. Values are those items given names by the user as opposed to those which are automatically labeled by MACSYMA as Ci, Di, or Ei.

Function RENAME (exp)

returns an expression equivalent to exp but with the dummy indices in each term chosen from the set [!1,!2,...]. Each dummy index in a product will be different; for a sum RENAME will try to make each dummy index in a sum the same. In addition, the indices will be sorted alphanumerically.

Function RNCOMBINE (exp)

transforms exp by combining all terms of exp that have identical denominators or denominators that differ from each other by numerical factors only. This is slightly different from the behavior of COMBINE, which collects terms that have identical denominators. Setting PFEFORMAT:TRUE and using COMBINE will achieve results similar to those that can be obtained with RNCOMBINE, but RNCOMBINE takes the additional step of cross-multiplying numerical denominator factors. This results in neater forms, and the possiblity of recognizing some cancellations. Bugs to ASB.

Function SCALARP (exp)

is TRUE if exp is a number, constant, or variable DECLAREd SCALAR, or composed entirely of numbers, constants, and such variables, but not containing matrices or lists.

Function SCALEFACTORS (coordinatetransform)

Here coordinatetransform evaluates to the form [[expression1, expression2, ...], indeterminate1, indeterminat2, ...], where indeterminate1, indeterminate2, etc. are the curvilinear coordinate variables and where a set of rectangular Cartesian components is given in terms of the curvilinear coordinates by [expression1, expression2, ...]. COORDINATES is set to the vector [indeterminate1, indeterminate2,...], and DI-MENSION is set to the length of this vector. SF[1], SF[2], ..., SF[DIMENSION] are set to the coordinate scale factors, and SFPROD is set to the product of these scale factors. Initially, COORDINATES is [X, Y, Z], DIMENSION is 3, and SF[1]=SF[2]=SF[3]=SFPROD=1, corresponding to 3-dimensional rectangular Cartesian coordinates. To expand an expression into physical components in the current coordinate system, there is a function with usage of the form Function SETUP AUTOLOAD (file,func1,...,funcN)

which takes two or more arguments: a file specification, and one or more function names, "funcI", and which indicates that if a call to "funcI" is made and "funcI" is not defined, that the file specified by "file" is to be automatically loaded in via LOAD, which file should contain a definition for "funcI". (This is the process by which calling e.g. INTEGRATE in a fresh MACSYMA causes various files to be loaded in.) As with the other file-handling commands in MACSYMA, the arguments to SETUP AUTOLOAD are not evaluated. Example: SETUP AUTOLOAD("bessel")$ J1(0.0); . Note: SETUP AUTOLOAD does not work for array functions.

198 Maxima Manual

27 Rules and Patterns

27.1 Introduction to Rules and Patterns

This section discusses user defined pattern matching and simplification rules (set up by TELLSIMP, TELLSIMPAFTER, DEFMATCH, or, DEFRULE.) You may affect the main simplification procedures, or else have your rules applied explicityly using APPLY1 and APPLY2. There are additional mechanisms for polynomials rules under TELLRAT, and for commutative and non commutative algebra in chapter on AFFINE.

27.2 Definitions for Rules and Patterns

Function APPLY1 (exp, rule1, ..., rulen)

repeatedly applies the first rule to exp until it fails, then repeatedly applies the same rule to all subexpressions of exp, left-to-right, until the first rule has failed on all subexpressions. Call the result of transforming exp in this manner exp’. Then the second rule is applied in the same fashion starting at the top of exp’. When the final rule fails on the final subexpression, the application is finished.

Function APPLY2 (exp, rule1, ..., rulen)

differs from APPLY1 in that if the first rule fails on a given subexpression, then the second rule is repeatedly applied, etc. Only if they all fail on a given subexpression is the whole set of rules repeatedly applied to the next subexpression. If one of the rules succeeds, then the same subexpression is reprocessed, starting with the first rule. MAXAPPLYDEPTH[10000] is the maximum depth to which APPLY1 and APPLY2 will delve.

Function APPLYB1 (exp, rule1, ..., rulen)

is similar to APPLY1 but works from the "bottom up" instead of from the "top down". That is, it processes the smallest subexpression of exp, then the next smallest, etc. MAXAPPLYHEIGHT[10000] - is the maximum height to which APPLYB1 will reach before giving up.

Variable CURRENT LET RULE PACKAGE

default:[DEFAULT LET RULE PACKAGE] - the name of the rule package that is presently being used. The user may reset this variable to the name of any rule package previously defined via the LET command. Whenever any of the functions comprising the let package are called with no package name the value of

     CURRENT_LET_RULE_PACKAGE

is used. If a call such as LETSIMP(expr,rule pkg name); is made, the rule package rule pkg name is used for that LETSIMP command only, i.e. the value of CUR-RENT LET RULE PACKAGE is not changed.

Variable DEFAULT LET RULE PACKAGE

- the name of the rule package used when one is not explicitly set by the user with LET or by changing the value of CURRENT LET RULE PACKAGE.

Function DEFMATCH (progname, pattern, parm1, ..., parmn)

creates a function of n+1 arguments with the name progname which tests an expression to see if it can match a particular pattern. The pattern is some expression containing pattern variables and parameters. The parms are given explicitly as arguments to DEFMATCH while the pattern variables (if supplied) were given implicitly in a previous MATCHDECLARE function. The first argument to the created function progname, is an expression to be matched against the "pattern" and the other n arguments are the actual variables occurring in the expression which are to take the place of dummy variables occurring in the "pattern". Thus the parms in the DEFMATCH are like the dummy arguments to the SUBROUTINE statement in FORTRAN. When the function is "called" the actual arguments are substituted. For example:

     (C1)  NONZEROANDFREEOF(X,E):=  IF E#0 AND FREEOF(X,E)
                 THEN TRUE ELSE FALSE$
     (IS(E#0 AND FREEOF(X,E)) is an  equivalent function
     definition)
     (C2)  MATCHDECLARE(A,NONZEROANDFREEOF(X),B,FREEOF(X))$
     (C3)  DEFMATCH(LINEAR,A*X+B,X)$
         This has caused the function LINEAR(exp,var1) to be defined.  It

tests exp to see if it is of the form A*var1+B where A and B do not contain var1 and A is not zero. DEFMATCHed functions return (if the match is successful) a list of equations whose left sides are the pattern variables and parms and whose right sides are the expressions which the pattern variables and parameters matched. The pattern variables, but not the parameters, are set to the matched expressions. If the match fails, the function returns FALSE. Thus LINEAR(3*Z+(Y+1)*Z+Y**2,Z) would return [B=Y**2, A=Y+4, X=Z]. Any variables not declared as pattern variables in MATCHDECLARE or as parameters in DEFMATCH which occur in pattern will match only themselves so that if the third argument to the DEFMATCH in (C4) had been omitted, then LINEAR would only match expressions linear in X, not in any other variable. A pattern which contains no parameters or pattern variables returns TRUE if the match succeeds. Do EXAMPLE(DEFMATCH); for more examples.

Function DEFRULE (rulename, pattern, replacement)

defines and names a replacement rule for the given pattern. If the rule named rulename is applied to an expression (by one of the APPLY functions below), every subexpression matching the pattern will be replaced by the replacement. All variables in the replacement which have been assigned values by the pattern match are assigned those values in the replacement which is then simplified. The rules themselves can be treated as functions which will transform an expression by one operation of the pattern match and replacement. If the pattern fails, the original expression is returned.

Function DISPRULE (rulename1, rulename2, ...)

will display rules with the names rulename1, rulename2, as were given by DEFRULE, TELLSIMP, or TELLSIMPAFTER or a pattern defined by DEFMATCH. For example, the first rule modifying SIN will be called SINRULE1. DISPRULE(ALL); will display all rules.

Function LET (prod, repl, predname, arg1, arg2, ..., argn)

defines a substitution rule for LETSIMP such that prod gets replaced by repl. prod is a product of positive or negative powers of the following types of terms:

• (1) Atoms which LETSIMP will search for literally unless previous to calling LETSIMP the MATCHDECLARE function is used to associate a predicate with the atom. In this case LETSIMP will match the atom to any term of a product satisfying the predicate.

• (2) Kernels such as SIN(X), N!, F(X,Y), etc. As with atoms above LETSIMP will look for a literal match unless MATCHDECLARE is used to associate a predicate with the argument of the kernel. A term to a positive power will only match a term having at least that power in the expression being LETSIMPed. A term to a negative power on the other hand will only match a term with a power at least as negative. In the case of negative powers in "product" the switch LETRAT must be set to TRUE (see below). If a predicate is included in the LET function followed by a list of arguments, a tentative match (i.e. one that would be accepted if the predicate were omitted) will be accepted only if predname(arg1’,...,argn’) evaluates to TRUE where argi’ is the value matched to argi. The argi may be the name of any atom or the argument of any kernel appearing in prod. repl may be any rational expression. If any of the atoms or arguments from prod appear in repl the appropriate substitutions will be made.

LETRAT[FALSE] when FALSE, LETSIMP will simplify the numerator and denominator of expr independently and return the result. Substitutions such as N!/N goes to (N-1)! will fail. To handle such situations LETRAT should be set to TRUE, then the numerator, denominator, and their quotient will be simplified in that order. These substitution functions allow you to work with several rulepackages at once. Each rulepackage can contain any number of LETed rules and is referred to by a user supplied name. To insert a rule into the rulepackage name, do LET([prod,repl,pred,arg1,...],name). To apply the rules in rulepackage name, do LETSIMP(expr, name). The function LETSIMP(expr,name1,name2,...) is equivalent to doing LETSIMP(expr,name1) followed by LETSIMP(%,name2) etc. CUR-RENT LET RULE PACKAGE is the name of the rule package that is presently being used. The user may reset this variable to the name of any rule package previously defined via the LET command. Whenever any of the functions comprising the let package are called with no package name the value of CURRENT LET RULE PACKAGE is used. If a call such as LETSIMP(expr,rule pkg name); is made, the rule package rule pkg name is used for that LETSIMP command only, i.e. the value of CUR-RENT LET RULE PACKAGE is not changed. There is a DEFAULT LET RULE PACKAGE which is assumed when no other name is supplied to any of the functions. Whenever a LET includes a rulepackage name that is used as the CURRENT LET RULE PACKAGE.

Variable LETRAT

default: [FALSE] - when FALSE, LETSIMP will simplify the numerator and denominator of expr independently and return the result. Substitutions such as N!/N goes to (N-1)! will fail. To handle such situations LETRAT should be set to TRUE, then the numerator, denominator, and their quotient will be simplified in that order.

Function LETRULES ()

displays the rules in the current rulepackage. LETRULES(name) displays the rules in the named rulepackage. The current rulepackage is the value of

     CURRENT_LET_RULE_PACKAGE

The initial value of the rules is

     DEFAULT_LET_RULE_PACKAGE

Function LETSIMP (exp)

will continually apply the substitution rules previously defined by the function LET until no further change is made to exp. LETSIMP(expr,rule pkg name); will cause the rule package rule pkg name to be used for that LETSIMP command only, i.e. the value of CURRENT LET RULE PACKAGE is not changed.

Variable LET RULE PACKAGES

default:[DEFAULT LET RULE PACKAGE] - The value of LET RULE PACKAGES is a list of all the user-defined let rule packages plus the special package

     DEFAULT_LET_RULE_PACKAGE

This is the name of the rule package used when one is not explicitly set by the user.

Function MATCHDECLARE (patternvar, predicate, ...)

associates a predicate with a pattern variable so that the variable will only match expressions for which the predicate is not FALSE. (The matching is accomplished by one of the functions described below). For example after

     MATCHDECLARE(Q,FREEOF(X,%E))

is executed, Q will match any expression not containing X or %E. If the match succeeds then the variable is set to the matched expression. The predicate (in this case FREEOF) is written without the last argument which should be the one against which the pattern variable is to be tested. Note that the patternvar and the arguments to the predicate are evaluated at the time the match is performed. The odd numbered argument may also be a list of pattern variables all of which are to have the associated predicate. Any even number of arguments may be given. For pattern matching, predicates refer to functions which are either FALSE or not FALSE (any non FALSE value acts like TRUE). MATCHDECLARE(var,TRUE) will permit var to match any expression.

Function MATCHFIX

- MATCHFIX operators are used to denote functions of any number of arguments which are passed to the function as a list. The arguments occur between the main operator and its "matching" delimiter. The MATCHFIX("x",...) function is a syntax extension function which declares x to be a MATCHFIX operator. Do DE-SCRIBE(SYNTAX); for more details.

Function REMLET (prod, name)

deletes the substitution rule, prod –> repl, most recently defined by the LET function. If name is supplied the rule is deleted from the rule package name. REMLET() and REMLET(ALL) delete all substitution rules from the current rulepackage. If the name of a rulepackage is supplied, e.g. REMLET(ALL,name), the rulepackage, name, is also deleted. If a substitution is to be changed using the same product, REMLET need not be called, just redefine the substitution using the same product (literally) with the LET function and the new replacement and/or predicate name. Should REMLET(product) now be called the original substitution rule will be revived.

Function REMRULE (function, rulename)

will remove a rule with the name rulename from the function which was placed there by DEFRULE, DEFMATCH, TELLSIMP, or TELLSIMPAFTER. If rule-name is ALL, then all rules will be removed.

Function TELLSIMP (pattern, replacement)

is similar to TELLSIMPAFTER but places new information before old so that it is applied before the built-in simplification rules. TELLSIMP is used when it is important to modify the expression before the simplifier works on it, for instance if the simplifier "knows" something about the expression, but what it returns is not to your liking. If the simplifier "knows" something about the main operator of the expression, but is simply not doing enough for you, you probably want to use TELLSIMPAFTER. The pattern may not be a sum, product, single variable, or number. RULES is a list of names having simplification rules added to them by DEFRULE, DEFMATCH, TELLSIMP, or TELLSIMPAFTER. Do EXAMPLE(TELLSIMP); for examples.

Function TELLSIMPAFTER (pattern, replacement)

defines a replacement for pattern which the MACSYMA simplifier uses after it applies the built-in simplification rules. The pattern may be anything but a single variable or a number.

204 Maxima Manual

28 Number Theory

28.1 Definitions for Number Theory

Function BERN (x)

gives the Xth Bernoulli number for integer X. ZEROBERN[TRUE] if set to FALSE excludes the zero BERNOULLI numbers. (See also BURN).

Function BERNPOLY (v, n)

generates the nth Bernoulli polynomial in the variable v.

Function BFZETA (exp,n)

BFLOAT version of the Riemann Zeta function. The 2nd argument is how many digits to retain and return, it’s a good idea to request a couple of extra. This function is available by doing LOAD(BFFAC); .

Function BGZETA (S, FPPREC)

BGZETA is like BZETA, but avoids arithmetic overflow errors on large arguments, is faster on medium size arguments (say S=55, FPPREC=69), and is slightly slower on small arguments. It may eventually replace BZETA. BGZETA is available by doing LOAD(BFAC);.

Function BHZETA (S,H,FPPREC)

gives FPPREC digits of

     SUM((K+H)^-S,K,0,INF)

This is available by doing LOAD(BFFAC);.

Function BINOMIAL (X, Y)

the binomial coefficient X*(X-1)*...*(X-Y+1)/Y!. If X and Y are integers, then the numerical value of the binomial coefficient is computed. If Y, or the value X-Y, is an integer, the binomial coefficient is expressed as a polynomial.

Function BURN (N)

is like BERN(N), but without computing all of the uncomputed Bernoullis of smaller index. So BURN works efficiently for large, isolated N. (BERN(402) takes about 645 secs vs 13.5 secs for BURN(402). BERN’s time growth seems to be exponential, while BURN’s is about cubic. But if next you do BERN(404), it only takes 12 secs, since BERN remembers all in an array, whereas BURN(404) will take maybe 14 secs or maybe 25, depending on whether MACSYMA needs to BFLOAT a better value of %PI.) BURN is available by doing LOAD(BFFAC);. BURN uses an observation of WGD that (rational) Bernoulli numbers can be approximated by (transcendental) zetas with tolerable efficiency.

Function BZETA

- This function is obsolete, see BFZETA.

Function CF (exp)

converts exp into a continued fraction. exp is an expression composed of arithmetic operators and lists which represent continued fractions. A continued fraction a+1/(b+1/(c+...)) is represented by the list [a,b,c,...]. a,b,c,.. must be integers. Exp may also involve SQRT(n) where n is an integer. In this case CF will give as many terms of the continued fraction as the value of the variable CFLENGTH[1] times the period. Thus the default is to give one period. (CF binds LISTARITH to FALSE so that it may carry out its function.)

Function CFDISREP (list)

converts the continued fraction represented by list into general representation.

     (C1) CF([1,2,-3]+[1,-2,1]);
     (D1)                 [1, 1, 1, 2]
     (C2) CFDISREP(%);
                                 1
     (D2)                1 + ---------
                                   1
                             1 + -----
                                     1
                                 1 + -
                                     2

Function CFEXPAND (x)

gives a matrix of the numerators and denominators of the next-to-last and last convergents of the continued fraction x.

     (C1) CF(SQRT(3));
     (D1)                 [1, 1, 2, 1, 2, 1, 2, 1]
     (C2) CFEXPAND(%);
                                  [71  97]
     (D2)                         [      ]
                                  [41  56]
     (C3) D2[1,2]/D2[2,2],NUMER;
     (D3)                        1.7321429

Variable CFLENGTH

default: [1] controls the number of terms of the continued fraction the function CF will give, as the value CFLENGTH[1] times the period. Thus the default is to give one period.

Function CGAMMA

- The Gamma function in the complex plane. Do LOAD(CGAMMA) to use these functions. Functions Cgamma, Cgamma2, and LogCgamma2. These functions evaluate the Gamma function over the complex plane using the algorithm of Kuki, CACM algorithm 421. Calculations are performed in single precision and the relative error is typically around 1.0E-7; evaluation at one point costs less than 1 msec. The algorithm provides for an error estimate, but the Macsyma implementation currently does not use it. Cgamma is the general function and may be called with a symbolic or numeric argument. With symbolic arguments, it returns as is; with real floating or rational arguments, it uses the Macsyma Gamma function; and for complex numeric arguments, it uses the Kuki algorithm. Cgamma2 of two arguments, real and imaginary, is for numeric arguments only; LogCgamma2 is the same, but the log-gamma function is calculated. These two functions are somewhat more efficient.

Function CGAMMA2

- See CGAMMA.

Function DIVSUM (n,k)

adds up all the factors of n raised to the kth power. If only one argument is given then k is assumed to be 1.

Function EULER (X)

gives the Xth Euler number for integer X. For the Euler-Mascheroni constant, see %GAMMA.

Function FACTORIAL (X)

The factorial function. FACTORIAL(X) = X! . See also MINFACTORIAL and FACTCOMB. The factorial operator is !, and the double factorial operator is !!.

Function FIB (X)

the Xth Fibonacci number with FIB(0)=0, FIB(1)=1, and FIB(-N)=(-1)^(N+1) *FIB(N). PREVFIB is FIB(X-1), the Fibonacci number preceding the last one computed.

Function FIBTOPHI (exp)

converts FIB(n) to its closed form definition. This involves the constant %PHI (= (SQRT(5)+1)/2 = 1.618033989). If you want the Rational Function Package to know About %PHI do TELLRAT(%PHI^2-%PHI-1)$ ALGEBRAIC:TRUE$ .

Function INRT (X,n)

takes two integer arguments, X and n, and returns the integer nth root of the absolute value of X.

Function JACOBI (p,q)

is the Jacobi symbol of p and q.

Function LCM (exp1,exp2,...)

returns the Least Common Multiple of its arguments. Do LOAD(FUNCTS); to access this function.

Variable MAXPRIME

default: [489318] - the largest number which may be given to the PRIME(n) command, which returns the nth prime.

Function MINFACTORIAL (exp)

examines exp for occurrences of two factorials which differ by an integer. It then turns one into a polynomial times the other. If exp involves binomial coefficients then they will be converted into ratios of factorials.

     (C1) N!/(N+1)!;
                                         N!
     (D1)                             --------
                                      (N + 1)!
     (C2) MINFACTORIAL(%);
                                          1
     (D2)                               -----
                                        N + 1

Function PARTFRAC (exp, var)

expands the expression exp in partial fractions with respect to the main variable, var. PARTFRAC does a complete partial fraction decomposition. The algorithm employed is based on the fact that the denominators of the partial fraction expansion (the factors of the original denominator) are relatively prime. The numerators can be written as linear combinations of denominators, and the expansion falls out. See EXAMPLE(PARTFRAC); for examples.

Function PRIME (n)

gives the nth prime. MAXPRIME[489318] is the largest number accepted as argument. Note: The PRIME command does not work in maxima, since it required a large file of primes, which most users do not want. PRIMEP does work however.

Function PRIMEP (n)

returns TRUE if n is a prime, FALSE if not.

Function QUNIT (n)

gives the principal unit of the real quadratic number field SQRT(n) where n is an integer, i.e. the element whose norm is unity. This amounts to solving Pell’s equation A**2- n*B**2=1.

     (C1) QUNIT(17);
     (D1)              SQRT(17)+4
     (C2)  EXPAND(%*(SQRT(17)-4));
     (D2)               1

Function TOTIENT (n)

is the number of integers less than or equal to n which are relatively prime to n.

Variable ZEROBERN

default: [TRUE] - if set to FALSE excludes the zero BERNOULLI numbers. (See the BERN function.)

Function ZETA (X)

gives the Riemann zeta function for certain integer values of X.

Variable ZETA%PI

default: [TRUE] - if FALSE, suppresses ZETA(n) giving coeff*%PI^n for n even.

210 Maxima Manual

29 Help

29.1 Introduction to Help

The most useful online help command is DESCRIBE which obtains help on all commands containing a particular string. Here by command we mean a built in operator such as INTEGRATE or FACTOR etc.

Online documentation is also available using INFO and printed documentation is available by using tex on the file info/maxima.texi in the maxima distribution.

29.2 LISP-AND-MACSYMA

All of Maxima is of course written in lisp. There is a naming convention for functions and variables: All symbols which begin with a "$" sign at lisp level, are read with the "$" sign stripped off at Macsyma level. For example, there are two lisp functions TRANSLATE and $TRANSLATE. If at macsyma level you enter TRANSLATE(FOO); the function which is called is the $translate function. To access the other function you must prefix with a "?":

     (C1) ?TRANSLATE(FOO);

Of course, this may well not do what you wanted it to do since it is a completely different function.

To enter lisp level on the lisp machine, you should type Suspend or Control-Suspend. This will cause a lisp break loop to be entered. You could now evaluate $d2 (or macsyma:$d2 if you have not done (pkg-goto ’macsyma)) and view the value of the line label D2, in its internal lisp format. Resume or abort will return to Macsyma top level.

Thus if you intend to write lisp functions to be called at macsyma level you should name them by names beginning with a "$".

If you are at lisp level and wish to read a macsyma expression you may enter

       (setq $foo #$[x,y]$)

This will have the same effect as entering

     (C1)FOO:[X,Y];

except that foo will not appear in the VALUES list. In order to view foo in macsyma printed format you may type

     (displa $foo)

In this documentation when we wish to refer to a macsyma symbol we shall generally omit the $ just as you would when typing at macsyma level. This will cause confusion when we also wish to refer to a lisp symbol. In this case we shall usually try to use lower case for the lisp symbol and upper case for the macsyma symbol. For example LIST for $list and list for the lisp symbol whose printname is "list".

To enter more complicated lisp expressions at macsyma level you may need the quoting character \ eg. for the lisp function make-array

     AA:?MAKE\-ARRAY(3,?:FILL\-POINTER,0);

would be have the same result as the lisp form

     (setq $aa (make-array 3  :fill-pointer 0)).

Since functions defined using the MAXIMA language are not ordinary lisp functions, you must use mfuncall to call them. For example:

     (D2)      FOO(X, Y) := X + Y + 3

then at lisp level

     CL-MAXIMA>>(mfuncall '$foo 4 5)
     12

A number of lisp functions are shadowed in the maxima package. This is because their use within maxima is not compatible with the definition as a system function. For example typep behaves differently common lisp than it did in Maclisp. If you want to refer to the zeta lisp typep while in the maxima package you should use global:typep (or cl:typep for common lisp). Thus

       (macsyma:typep '(1 2)) ==> 'list
       (lisp:typep '(1 2))==> error (lisp:type-of '(1 2))==> 'cons

To see which symbols are shadowed look in "Maxima-source:maxima;sysdef.lisp" or do a describe of the package.

29.3 GARBAGE COLLECTION

Symbolic computation tends to create a good deal of garbage, and effective handling of this can be crucial to successful completion of some programs.

Under GCL, on UNIX systems where the mprotect system call is available (including SUN OS 4.0 and some variants of BSD) a stratified garbage collection is available. This limits the collection to pages which have been recently written to. See the GCL documentation under ALLOCATE and GBC. At the lisp level doing (setq si::*notify-gbc* t) will help you determine which areas might need more space.

29.4 DOCUMENTATION

The function DESCRIBE(item) will offer a list of topics which include the substring item. The documentation will then be printed on the screen. It may also be put into a buffer by using the command MACSYMA DESCRIBE in the zmacs editor.

The variable ALL MACSYMA DOCUMENTATION is a list of file names which contain documentation in a format suitable to be read by DESCRIBE. You may of course add to any of these files, or add another file to the list. Each separate item should begin with "&" on a new line separated by a blank line. The key which will be used for searching, is the string up to the first blank space.

The function set-up-index will then be called to make an index and add it to *DESCRIBE-INDICES*, so that future calls to DESCRIBE will also see the additional documentation.

(set-up-index "MAXIMA-source:maxima;foo.bar" :make-new-one t)

would cause a new index to be made even if foo-index.bin already existed containing the appropriate index.

29.5 Definitions for Help

Function DEMO (file)

this is the same as BATCH but pauses after each command line and continues when a space is typed. In NIL macsyma, DEMO does file searching like BATCH does, and in addition searches for files on the [DEMO] directory in the Macsyma directory hierarchy, and also for files with extensions of .DEM, .DM1, and .DM2 in addition to .MAC. This is to make it easier to run demo files which come supplied with NIL macsyma, which are named things like MAX$DISK:[DEMO]MAT.DEM.

Function DESCRIBE (cmd)

prints out information about "cmd", which may be any MACSYMA command, switch or variable. Certain key words have also been included, where they seem appropriate, thus DESCRIBE(TRIG); will print out a list of the trig functions implemented in MACSYMA. Some function names or operators may require quotation marks around them, e.g. DESCRIBE("DO"); or DESCRIBE(".");. See also APROPOS(string) which allows you to locate command names even if you aren’t sure of the full name. Do DESCRIBE(APROPOS); for details.

Function EXAMPLE (command)

will start up a demonstration of how command works on some expressions. After each command line it will pause and wait for a space to be typed, as in the DEMO command.

214 Maxima Manual

30 Differential Equations

30.1 Definitions for Differential Equations

Function DESOLVE ([eq1,...,eqn],[var1,...,varn])

where the eq’s are differential equations in the dependent variables var1,...,varn. The functional relationships must be explicitly indicated in both the equations and the variables. For example

(C1) 'DIFF(F,X,2)=SIN(X)+'DIFF(G,X);
(C2) 'DIFF(F,X)+X^2-F=2*'DIFF(G,X,2);
is NOT the proper format.  The correct way is:
(C3) 'DIFF(F(X),X,2)=SIN(X)+'DIFF(G(X),X);
(C4) 'DIFF(F(X),X)+X^2-F(X)=2*'DIFF(G(X),X,2);
The call is then DESOLVE([D3,D4],[F(X),G(X)]);
If initial conditions at 0 are known, they should be supplied before
calling DESOLVE by using ATVALUE.
(C11) 'DIFF(F(X),X)='DIFF(G(X),X)+SIN(X);
                        d         d
(D11)                   -- F(X) = -- G(X) + SIN(X)
                        dX        dX
(C12) 'DIFF(G(X),X,2)='DIFF(F(X),X)-COS(X);
                         2
                        d          d
(D12)                   --- G(X) = -- F(X) - COS(X)
                          2        dX
                        dX
(C13) ATVALUE('DIFF(G(X),X),X=0,A);
(D13)                                A
(C14) ATVALUE(F(X),X=0,1);
(D14)                                1
(C15) DESOLVE([D11,D12],[F(X),G(X)]);
                X                            X
(D16) [F(X)=A %E  - A+1, G(X) = COS(X) + A %E  - A + G(0) - 1]
/* VERIFICATION */
(C17) [D11,D12],D16,DIFF;
                  X       X      X                X
(D17)        [A %E  = A %E , A %E  - COS(X) = A %E  - COS(X)]

If DESOLVE cannot obtain a solution, it returns "FALSE".

Function IC1 (exp,var,var)

In order to solve initial value problems (IVPs) and boundary value problems (BVPs), the routine IC1 is available in the ODE2 package for first order equations, and IC2 and BC2 for second order IVPs and BVPs, respectively. Do LOAD(ODE2) to access these. They are used as in the following examples:

     (C3) IC1(D2,X=%PI,Y=0);
                              COS(X) + 1
     (D3)               Y = - ----------
                                   3
                                  X
     (C4) 'DIFF(Y,X,2) + Y*'DIFF(Y,X)^3 = 0;
                            2
                           d Y      dY 3
     (D4)                  --- + Y (--)  = 0
                             2      dX
                           dX
     (C5) ODE2(%,Y,X);
                      3
                     Y  - 6 %K1 Y - 6 X
     (D7)            ------------------ = %K2
                             3
     (C8) RATSIMP(IC2(D7,X=0,Y=0,'DIFF(Y,X)=2));
                          3
                       2 Y  - 3 Y + 6 X
     (D9)            - ---------------- = 0
                              3
     (C10) BC2(D7,X=0,Y=1,X=1,Y=3);
                      3
                     Y  - 10 Y - 6 X
     (D11)           --------------- = - 3
                            3

Function ODE (equation,y,x)

a pot-pourri of Ordinary Differential solvers combined in such a way as to attempt more and more difficult methods as each fails. For example, the first attempt is with ODE2, so therefore, a user using ODE can assume he has all the capabilities of ODE2 at the very beginning and if he has been using ODE2 in programs they will still run if he substitutes ODE (the returned values, and calling sequence are identical). In addition, ODE has a number of user features which can assist an experienced ODE solver if the basic system cannot handle the equation. The equation is of the same form as required for ODE2 (which see) and the y and x are dependent and independent variables, as with ODE2. For more details, do PRINTFILE(ODE,USAGE,SHARE); .

Function ODE2 (exp,dvar,ivar)

takes three arguments: an ODE of first or second order (only the left hand side need be given if the right hand side is 0), the dependent variable, and the independent variable. When successful, it returns either an explicit or implicit solution for the dependent variable. %C is used to represent the constant in the case of first order equations, and %K1 and %K2 the constants for second order equations. If ODE2 cannot obtain a solution for whatever reason, it returns FALSE, after perhaps printing out an error message. The methods implemented for first order equations in the order in which they are tested are: linear, separable, exact - perhaps requiring an integrating factor, homogeneous, Bernoulli’s equation, and a generalized homogeneous method. For second order: constant coefficient, exact, linear homogeneous with non-constant coefficients which can be transformed to constant coefficient, the Euler or equidimensional equation, the method of variation of parameters, and equations which are free of either the independent or of the dependent variable so that they can be reduced to two first order linear equations to be solved sequentially. In the course of solving ODEs, several variables are set purely for informational purposes: METHOD denotes the method of solution used e.g. LINEAR, INTFACTOR denotes any integrating factor used, ODEINDEX denotes the index for Bernoulli’s method or for the generalized homogeneous method, and YP denotes the particular solution for the variation of parameters technique.

218 Maxima Manual

31 Statistics

31.1 Definitions for Statistics

Function GAUSS (mean,sd)

returns a random floating point number from a normal distribution with mean MEAN and standard deviation SD. This is part of the BESSEL function package, do LOAD(BESSEL); to use it.

220 Maxima Manual

32 Limits

32.1 Definitions for Limits

Variable LHOSPITALLIM

default: [4] - the maximum number of times L’Hospital’s rule is used in LIMIT. This prevents infinite looping in cases like LIMIT(COT(X)/CSC(X),X,0).

Function LIMIT (exp, var, val, dir)

finds the limit of exp as the real variable var approaches the value val from the direction dir. Dir may have the value PLUS for a limit from above, MINUS for a limit from below, or may be omitted (implying a two-sided limit is to be computed). For the method see Wang, P., "Evaluation of Definite Integrals by Symbolic Manipulation" - Ph.D. Thesis - MAC TR-92 October 1971. LIMIT uses the following special symbols: INF (positive infinity) and MINF (negative infinity). On output it may also use UND (undefined), IND (indefinite but bounded) and INFINITY (complex infinity). LHOSPITALLIM[4] is the maximum number of times L’Hospital’s rule is used in LIMIT. This prevents infinite looping in cases like LIMIT(COT(X)/CSC(X),X,0). TLIMSWITCH[FALSE] when true will cause the limit package to use Taylor series when possible. LIMSUBST[FALSE] prevents LIMIT from attempting substitutions on unknown forms. This is to avoid bugs like LIMIT(F(N)/F(N+1),N,INF); giving 1. Setting LIMSUBST to TRUE will allow such substitutions. Since LIMIT is often called upon to simplify constant expressions, for example, INF-1, LIMIT may be used in such cases with only one argument, e.g. LIMIT(INF-1); Do EXAMPLE(LIMIT); for examples.

Function TLIMIT (exp,var,val,dir)

is just the function LIMIT with TLIMSWITCH set to TRUE.

Variable TLIMSWITCH

default: [FALSE] - if true will cause the limit package to use Taylor series when possible.

222 Maxima Manual

33 Logarithms

33.1 Definitions for Logarithms

Function LOG (X)

the natural logarithm. LOGEXPAND[TRUE] - causes LOG(A^B) to become B*LOG(A). If it is set to ALL, LOG(A*B) will also simplify to LOG(A)+LOG(B). If it is set to SUPER, then LOG(A/B) will also simplify to LOG(A)-LOG(B) for rational numbers a/b, a#1. (LOG(1/B), for B integer, always simplifies.) If it is set to FALSE, all of these simplifications will be turned off. LOGSIMP[TRUE] - if FALSE then no simplification of %E to a power containing LOG’s is done. LOGNUMER[FALSE] - if TRUE then negative floating point arguments to LOG will always be converted to their absolute value before the log is taken. If NUMER is also TRUE, then negative integer arguments to LOG will also be converted to their absolute value. LOGNEGINT[FALSE] - if TRUE implements the rule LOG(-n) -> LOG(n)+%i*%pi for n a positive integer. %E TO NUMLOG[FALSE] - when TRUE, "r" some rational number, and "x" some expression, %E^(r*LOG(x)) will be simplified into x^r . It should be noted that the RADCAN command also does this transformation, and more complicated transformations of this ilk as well. The LOGCONTRACT command "contracts" expressions containing LOG.

Variable LOGABS

default: [FALSE] - when doing indefinite integration where logs are generated, e.g. INTEGRATE(1/X,X), the answer is given in terms of LOG(ABS(...)) if LOGABS is TRUE, but in terms of LOG(...) if LOGABS is FALSE. For definite integration, the LOGABS:TRUE setting is used, because here "evaluation" of the indefinite integral at the endpoints is often needed.

Variable LOGARC

default: [FALSE] - if TRUE will cause the inverse circular and hyperbolic functions to be converted into logarithmic form. LOGARC(exp) will cause this conversion for a particular expression without setting the switch or having to re-evaluate the expression with EV.

Variable LOGCONCOEFFP

default:[FALSE] - controls which coefficients are contracted when using LOGCON-TRACT. It may be set to the name of a predicate function of one argument. E.g. if you like to generate SQRTs, you can do LOGCONCOEFFP:’LOGCONFUN$ LOG-CONFUN(M):=FEATUREP(M,INTEGER) OR RATNUMP(M)$ . Then LOG-CONTRACT(1/2*LOG(X)); will give LOG(SQRT(X)).

Function LOGCONTRACT (exp)

recursively scans an exp, transforming subexpressions of the form a1*LOG(b1) + a2*LOG(b2) + c into LOG(RATSIMP(b1^a1 * b2^a2)) + c

     (C1) 2*(A*LOG(X) + 2*A*LOG(Y))$
     (C2) LOGCONTRACT(%);
                                             2  4
     (D3)                             A LOG(X  Y )

If you do DECLARE(N,INTEGER); then LOGCONTRACT(2*A*N*LOG(X)); gives A*LOG(X^(2*N)). The coefficients that "contract" in this manner are those such as the 2 and the N here which satisfy FEATUREP(coeff,INTEGER). The user can control which coefficients are contracted by setting the option LOGCONCO-EFFP[FALSE] to the name of a predicate function of one argument. E.g. if you like to generate SQRTs, you can do LOGCONCOEFFP:’LOGCONFUN$ LOGCON-FUN(M):=FEATUREP(M,INTEGER) OR RATNUMP(M)$ . Then LOGCON-TRACT(1/2*LOG(X)); will give LOG(SQRT(X)).

Variable LOGEXPAND

default: [TRUE] - causes LOG(A^B) to become B*LOG(A). If it is set to ALL, LOG(A*B) will also simplify to LOG(A)+LOG(B). If it is set to SUPER, then LOG(A/B) will also simplify to LOG(A)-LOG(B) for rational numbers a/b, a#1. (LOG(1/B), for B integer, always simplifies.) If it is set to FALSE, all of these simplifications will be turned off.

Variable LOGNEGINT

default: [FALSE] - if TRUE implements the rule LOG(-n) -> LOG(n)+%i*%pi for n a positive integer.

Variable LOGNUMER

default: [FALSE] - if TRUE then negative floating point arguments to LOG will always be converted to their absolute value before the log is taken. If NUMER is also TRUE, then negative integer arguments to LOG will also be converted to their absolute value.

Variable LOGSIMP

default: [TRUE] - if FALSE then no simplification of %E to a power containing LOG’s is done.

Function PLOG (X)

the principal branch of the complex-valued natural logarithm with -%PI < CARG(X) <= +%PI .

Function POLARFORM (exp)

returns R*%E^(%I*THETA) where R and THETA are purely real.

34 Debugging

34.1 Definitions for Debugging

Variable REFCHECK

default: [FALSE] - if TRUE causes a message to be printed each time a bound variable is used for the first time in a computation.

Function REMTRACE ()

This function is no longer used with the new TRACE package.

Variable SETCHECK

default: [FALSE] - if set to a list of variables (which can be subscripted) will cause a printout whenever the variables, or subscripted occurrences of them, are bound (with : or :: or function argument binding). The printout consists of the variable and the value it is bound to. SETCHECK may be set to ALL or TRUE thereby including all variables. Note: No printout is generated when a SETCHECKed variable is set to itself, e.g. X:’X.

Variable SETCHECKBREAK

default: [FALSE] - if set to TRUE will cause a (MACSYMA-BREAK) to occur whenever the variables on the SETCHECK list are bound. The break occurs before the binding is done. At this point, SETVAL holds the value to which the variable is about to be set. Hence, one may change this value by resetting SETVAL.

Variable SETVAL

- holds the value to which a variable is about to be set when a SETCHECKBREAK occurs. Hence, one may change this value by resetting SETVAL. (See SETCHECK-BREAK).

Function TIMER (F)

will put a timer-wrapper on the function F, within the TRACE package, i.e. it will print out the time spent in computing F.

Variable TIMER DEVALUE

default: [FALSE] - when set to TRUE then the time charged against a function is the time spent dynamically inside the function devalued by the time spent inside other TIMED functions.

Function TIMER INFO (F)

will print the information on timing which is stored also as GET(’F,’CALLS); GET(’F,’RUNTIME); and GET(’F,’GCTIME); . This is a TRACE package function.

Function TRACE (name1, name2, ...)

gives a trace printout whenever the functions mentioned are called. TRACE() prints a list of the functions currently under TRACE. On MC see MACDOC;TRACE US-AGE for more information. Also, DEMO("trace.dem"); . To remove tracing, see UNTRACE.

Function TRACE OPTIONS (F,option1,option2,...)

gives the function F the options indicated. An option is either a keyword or an expression. The possible Keywords are: Keyword Meaning of return value ——— ——————————- NOPRINT If TRUE do no printing. BREAK If TRUE give a breakpoint. LISP PRINT If TRUE use lisp printing. INFO Extra info to print. ERRORCATCH If TRUE errors are caught. A keyword means that the option is in effect. Using a keyword as an expression, e.g. NOPRINT(predicate function) means to apply the predicate function (which is user-defined) to some arguments to determine if the option is in effect. The argument list to this predicate function is always [LEVEL, DIRECTION, FUNCTION, ITEM] where LEVEL is the recursion level for the function. DIRECTION is either ENTER or EXIT. FUNCTION is the name of the function. ITEM is either the argument list or the return value. On MC see DEMO("trace.dem"); for more details.

Function UNTRACE (name1, ...)

removes tracing invoked by the TRACE function. UNTRACE() removes tracing from all functions.

35 Affine

35.1 Definitions for Affine

Function FAST LINSOLVE (eqns,variables)

Solves the linear system of equations EQNS for the variables VARIABLES and returns a result suitable to SUBLIS. The function is faster than linsolve for system of equations which are sparse.

Function GROBNER BASIS (eqns)

Takes as argument a macsyma list of equations and returns a grobner basis for them. The function POLYSIMP may now be used to simplify other functions relative to the equations. GROBNER BASIS([3*X^2+1,Y*X])$ POLYSIMP(Y^2*X+X^3*9+2)==> -3*x+2 Polysimp(f)==> 0 if and only if f is in the ideal generated by the EQNS ie. if and only if f is a polynomial combination of the elements of EQNS.

Function SET UP DOT SIMPLIFICATIONS (eqns,[check-thru-degree])

The eqns are polynomial equations in non commutative variables. The value of CUR-RENT VARIABLES is the list of variables used for computing degrees. The equations must be homogeneous, in order for the procedure to terminate. If you have checked overlapping simplifications in DOT SIMPLIFICATIONS above the degree of f, then the following is true: DOTSIMP(f)==> 0 if and only if f is in the ideal generated by the EQNS ie. if and only if f is a polynomial combination of the elements of EQNS. The degree is that returned by NC DEGREE. This in turn is influenced by the weights of individual variables.

Function DECLARE WEIGHT (var1,wt1,var2,wt2,...)

Assigns VAR1 weight WT1, VAR2 weight wt2.. These are the weights used in computing NC DEGREE.

Function NC DEGREE (poly)

Degree of a non commutative polynomial. See DECLARE WEIGHTS.

Function DOTSIMP (f)

==> 0 if and only if f is in the ideal generated by the EQNS ie. if and only if f is a polynomial combination of the elements of EQNS.

Function FAST CENTRAL ELEMENTS (variables,degree)

if SET UP DOT SIMPLIFICATIONS has been previously done, finds the central polynomials in the variables in the given degree, For example:

     set_up_dot_simplifications([y.x+x.y],3);
     fast_central_elements([x,y],2);
     [y.y,x.x];

Function CHECK OVERLAPS (degree,add-to-simps)

checks the overlaps thru degree, making sure that you have sufficient simplification rules in each degree, for dotsimp to work correctly. This process can be speeded up if you know before hand what the dimension of the space of monomials is. If it is of finite global dimension, then HILBERT should be used. If you don’t know the monomial dimensions, do not specify a RANK FUNCTIION. An optional third argument RESET, false says don’t bother to query about resetting things.

Function MONO (vari,n)

VARI is a list of variables. Returns the list of independent monomials relative to the current dot simplifications, in degree N

Function MONOMIAL DIMENSIONS (n)

Compute the hilbert series through degreen n for the current algebra.

Function EXTRACT LINEAR EQUATIONS (List nc polys,monoms)

Makes a list of the coefficients of the polynomials in list nc polys of the monoms. MONOMS is a list of noncommutative monomials. The coefficients should be scalars. Use LIST NC MONOMIALS to build the list of monoms.

Function LIST NC MONOMIALS (polys or list)

returns a list of the non commutative monomials occurring in a polynomial or a collection of polynomials.

Function PCOEFF (poly monom [variables-to-exclude-from-cof (list-variables

monom)]) This function is called from lisp level, and uses internal poly format.

     CL-MAXIMA>>(setq me (st-rat #$x^2*u+y+1$))
     (#:Y 1 1 0 (#:X 2 (#:U 1 1) 0 1))
     CL-MAXIMA>>(pcoeff me (st-rat #$x^2$))
     (#:U 1 1)

Rule: if a variable appears in monom it must be to the exact power, and if it is in variables to exclude it may not appear unless it was in monom to the exact power. (pcoeff pol 1 ..) will exclude variables like substituting them to be zero.

Function NEW-DISREP (poly)

From lisp this returns the general maxima format for an arg which is in st-rat form:

     (displa(new-disrep (setq me (st-rat #$x^2*u+y+1$))))
            2
     Y + U X  + 1

Function CREATE LIST (form,var1,list1,var2,list2,...)

Create a list by evaluating FORM with VAR1 bound to each element of LIST1, and for each such binding bind VAR2 to each element of LIST2,... The number of elements in the result will be length(list1)*length(list2)*... Each VARn must actually be a symbol–it will not be evaluated. The LISTn args will be evaluated once at the beginning of the iteration.

     (C82) create_list1(x^i,i,[1,3,7]);
     (D82) [X,X^3,X^7]

With a double iteration:

     (C79) create_list([i,j],i,[a,b],j,[e,f,h]);
     (D79) [[A,E],[A,F],[A,H],[B,E],[B,F],[B,H]]

Instead of LISTn two args maybe supplied each of which should evaluate to a number. These will be the inclusive lower and upper bounds for the iteration.

     (C81) create_list([i,j],i,[1,2,3],j,1,i);
     (D81) [[1,1],[2,1],[2,2],[3,1],[3,2],[3,3]]

Note that the limits or list for the j variable can depend on the current value of i.

Variable ALL DOTSIMP DENOMS

if its value is FALSE the denominators encountered in getting dotsimps will not be collected. To collect the denoms

     ALL_DOTSIMP_DENOMS:[];

and they will be nconc’d onto the end of the list.

230 Maxima Manual

36 Tensor

36.1 Introduction to Tensor

- Indicial Tensor Manipulation package. It may be loaded by LOADFILE("itensr"); A manual for the Tensor packages is available in share/tensor.descr. A demo is available by DEMO("itenso.dem1"); (and additional demos are in ("itenso.dem2"), ("itenso.dem3") and following).

- There are two tensor packages in MACSYMA, CTENSR and ITENSR. CTENSR is Component Tensor Manipulation, and may be accessed with LOAD(CTENSR); . ITENSR is Indicial Tensor Manipulation, and is loaded by doing LOAD(ITENSR); A manual for CTENSR AND ITENSR is available from the LCS Publications Office. Request MIT/LCS/TM-167. In addition, demos exist on the TENSOR; directory under the filenames CTENSO DEMO1, DEMO2, etc. and ITENSO DEMO1, DEMO2, etc. Do DEMO("ctenso.dem1"); or DEMO("itenso.dem2"); Send bugs or comments to RP or TENSOR.

36.2 Definitions for Tensor

Function CANFORM (exp)

[Tensor Package] Simplifies exp by renaming dummy indices and reordering all indices as dictated by symmetry conditions imposed on them. If ALLSYM is TRUE then all indices are assumed symmetric, otherwise symmetry information provided by DEC-SYM declarations will be used. The dummy indices are renamed in the same manner as in the RENAME function. When CANFORM is applied to a large expression the calculation may take a considerable amount of time. This time can be shortened by calling RENAME on the expression first. Also see the example under DECSYM. Note: CANFORM may not be able to reduce an expression completely to its simplest form although it will always return a mathematically correct result.

Function CANTEN (exp)

[Tensor Package] Simplifies exp by renaming (see RENAME) and permuting dummy indices. CANTEN is restricted to sums of tensor products in which no derivatives are present. As such it is limited and should only be used if CANFORM is not capable of carrying out the required simplification.

Function CARG (exp)

returns the argument (phase angle) of exp. Due to the conventions and restrictions, principal value cannot be guaranteed unless exp is numeric.

Variable COUNTER

default: [1] determines the numerical suffix to be used in generating the next dummy index in the tensor package. The prefix is determined by the option DUMMYX[#].

Function DEFCON (tensor1,<tensor2,tensor3>)

gives tensor1 the property that the contraction of a product of tensor1 and tensor2 results in tensor3 with the appropriate indices. If only one argument, tensor1, is given, then the contraction of the product of tensor1 with any indexed object having the appropriate indices (say tensor) will yield an indexed object with that name, i.e.tensor, and with a new set of indices reflecting the contractions performed. For example, if METRIC: G, then DEFCON(G) will implement the raising and lowering of indices through contraction with the metric tensor. More than one DEFCON can be given for the same indexed object; the latest one given which applies in a particular contraction will be used. CONTRACTIONS is a list of those indexed objects which have been given contraction properties with DEFCON.

Function FLUSH (exp,tensor1,tensor2,...)

Tensor Package - will set to zero, in exp, all occurrences of the tensori that have no derivative indices.

Function FLUSHD (exp,tensor1,tensor2,...)

Tensor Package - will set to zero, in exp, all occurrences of the tensori that have derivative indices.

Function FLUSHND (exp,tensor,n)

Tensor Package - will set to zero, in exp, all occurrences of the differentiated object tensor that have n or more derivative indices as the following example demonstrates.

     (C1) SHOW(A([I],[J,R],K,R)+A([I],[J,R,S],K,R,S));
                                    J R S      J R
     (D1)                          A        + A
                                    I,K R S    I,K R
     (C2) SHOW(FLUSHND(D1,A,3));
                                          J R
     (D2)                                A
                                          I,K R

Function KDELTA (L1,L2)

is the generalized Kronecker delta function defined in the Tensor package with L1 the list of covariant indices and L2 the list of contravariant indices. KDELTA([i],[j]) returns the ordinary Kronecker delta. The command EV(EXP,KDELTA) causes the evaluation of an expression containing KDELTA([],[]) to the dimension of the manifold.

Function LC (L)

is the permutation (or Levi-Civita) tensor which yields 1 if the list L consists of an even permutation of integers, -1 if it consists of an odd permutation, and 0 if some indices in L are repeated.

Function LORENTZ (exp)

imposes the Lorentz condition by substituting 0 for all indexed objects in exp that have a derivative index identical to a contravariant index.

Function MAKEBOX (exp)

will display exp in the same manner as SHOW; however, any tensor d’Alembertian occurring in exp will be indicated using the symbol []. For example, []P([M],[N]) represents G([],[I,J])*P([M],[N],I,J).

Function METRIC (G)

specifies the metric by assigning the variable METRIC:G; in addition, the contraction properties of the metric G are set up by executing the commands DEFCON(G), DEFCON(G,G,KDELTA). The variable METRIC, default: [], is bound to the metric, assigned by the METRIC(g) command.

Function NTERMSG ()

gives the user a quick picture of the "size" of the Einstein tensor. It returns a list of pairs whose second elements give the number of terms in the components specified by the first elements.

Function NTERMSRCI ()

returns a list of pairs, whose second elements give the number of terms in the RICCI component specified by the first elements. In this way, it is possible to quickly find the non-zero expressions and attempt simplification.

Function NZETA (Z)

returns the complex value of the Plasma Dispersion Function for complex Z.

     NZETAR(Z) ==> REALPART(NZETA(Z))

NZETAI(Z) returns IMAGPART(NZETA(Z)). This function is related to the complex error function by

     NZETA(Z) = %I*SQRT(%PI)*EXP(-Z^2)*(1-ERF(-%I*Z)).

Function RAISERIEMANN (dis)

returns the contravariant components of the Riemann curvature tensor as array elements UR[I,J,K,L]. These are displayed if dis is TRUE.

Variable RATEINSTEIN

default: [] - if TRUE rational simplification will be performed on the non-zero components of Einstein tensors; if FACRAT:TRUE then the components will also be factored.

Variable RATRIEMAN

- This switch has been renamed RATRIEMANN.

Variable RATRIEMANN

default: [] - one of the switches which controls simplification of Riemann tensors; if TRUE, then rational simplification will be done; if FACRAT:TRUE then each of the components will also be factored.

Function REMCON (tensor1,tensor2,...)

removes all the contraction properties from the tensori. REMCON(ALL) removes all contraction properties from all indexed objects.

Function RICCICOM (dis)

Tensor package) This function first computes the covariant components LR[i,j] of the Ricci tensor (LR is a mnemonic for "lower Ricci"). Then the mixed Ricci tensor is computed using the contravariant metric tensor. If the value of the argument to RICCICOM is TRUE, then these mixed components, RICCI[i,j] (the index i is covariant and the index j is contravariant), will be displayed directly. Otherwise, RICCICOM(FALSE) will simply compute the entries of the array RICCI[i,j] without displaying the results.

Function RINVARIANT ()

Tensor package) forms the invariant obtained by contracting the tensors

     R[i,j,k,l]*UR[i,j,k,l].

This object is not automatically simplified since it can be very large.

Function SCURVATURE ()

returns the scalar curvature (obtained by contracting the Ricci tensor) of the Riemannian manifold with the given metric.

Function SETUP ()

this has been renamed to TSETUP(); Sets up a metric for Tensor calculations.

Function WEYL (dis)

computes the Weyl conformal tensor. If the argument dis is TRUE, the non-zero components W[I,J,K,L] will be displayed to the user. Otherwise, these components will simply be computed and stored. If the switch RATWEYL is set to TRUE, then the components will be rationally simplified; if FACRAT is TRUE then the results will be factored as well.

37 Ctensor

37.1 Introduction to Ctensor

- Component Tensor Manipulation Package. To use the CTENSR package, type TSETUP(); which automatically loads it from within MACSYMA (if it is not already loaded) and then prompts the user to input his coordinate system. The user is first asked to specify the dimension of the manifold. If the dimension is 2, 3 or 4 then the list of coordinates defaults to [X,Y], [X,Y,Z] or [X,Y,Z,T] respectively. These names may be changed by assigning a new list of coordinates to the variable OMEGA (described below) and the user is queried about this. ** Care must be taken to avoid the coordinate names conflicting with other object definitions **. Next, the user enters the metric either directly or from a file by specifying its ordinal position. As an example of a file of common metrics, see TENSOR;METRIC FILE. The metric is stored in the matrix LG. Finally, the metric inverse is computed and stored in the matrix UG. One has the option of carrying out all calculations in a power series. A sample protocol is begun below for the static, spherically symmetric metric (standard coordinates) which will be applied to the problem of deriving Einstein’s vacuum equations (which lead to the Schwarzschild solution) as an example. Many of the functions in CTENSR will be displayed for the standard metric as examples.

(C2) TSETUP();
Enter the dimension of the coordinate system:
4;
Do you wish to change the coordinate names?
N;
Do you want to
1. Enter a new metric?
2. Enter a metric from a file?
3. Approximate a metric with a Taylor series?
Enter 1, 2 or 3
1;
Is the matrix  1. Diagonal  2. Symmetric  3. Antisymmetric  4. General
Answer 1, 2, 3 or 4
1;
Row 1 Column 1:  A;
Row 2 Column 2:  X^2;
Row 3 Column 3:  X^2*SIN(Y)^2;
Row 4 Column 4:  -D;
Matrix entered.
Enter functional dependencies with the DEPENDS function or 'N' if none
DEPENDS([A,D],X);
Do you wish to see the metric?
Y;
                          [ A  0       0        0  ]
                          [                        ]
                          [     2                  ]
                          [ 0  X       0        0  ]

                          [                        ]
                          [         2    2         ]
                          [ 0  0   X  SIN (Y)   0  ]
                          [                        ]
                          [ 0  0       0       - D ]
Do you wish to see the metric inverse?
N;

37.2 Definitions for Ctensor

Function CHR1 ([i,j,k])

yields the Christoffel symbol of the first kind via the definition

            (g      + g      - g     )/2 .
              ik,j     jk,i     ij,k

To evaluate the Christoffel symbols for a particular metric, the variable METRIC must be assigned a name as in the example under CHR2.

Function CHR2 ([i,j],[k])

yields the Christoffel symbol of the second kind defined by the relation

                            ks
         CHR2([i,j],[k]) = g    (g      + g      - g     )/2
                                  is,j     js,i     ij,s

Function CHRISTOF (arg)

A function in the CTENSR (Component Tensor Manipulation) package. It computes the Christoffel symbols of both kinds. The arg determines which results are to be immediately displayed. The Christoffel symbols of the first and second kinds are stored in the arrays LCS[i,j,k] and MCS[i,j,k] respectively and defined to be symmetric in the first two indices. If the argument to CHRISTOF is LCS or MCS then the unique non-zero values of LCS[i,j,k] or MCS[i,j,k], respectively, will be displayed. If the argument is ALL then the unique non-zero values of LCS[i,j,k] and MCS[i,j,k] will be displayed. If the argument is FALSE then the display of the elements will not occur. The array elements MCS[i,j,k] are defined in such a manner that the final index is contravariant.

Function COVDIFF (exp,v1,v2,...)

yields the covariant derivative of exp with respect to the variables vi in terms of the Christoffel symbols of the second kind (CHR2). In order to evaluate these, one should use EV(exp,CHR2).

Function CURVATURE ([i,j,k],[h])

Indicial Tensor Package) yields the Riemann curvature tensor in terms of the Christoffel symbols of the second kind (CHR2). The following notation is used:

                    h            h           h        %1        h
           CURVATURE     = - CHR2      - CHR2     CHR2    + CHR2
                    i j k        i k,j       %1 j     i k       i j,k
                                   h         %1
                             + CHR2      CHR2
                                   %1 k      i j

Variable DIAGMETRIC

default:[] - An option in the CTENSR (Component Tensor Manipulation) package. If DIAGMETRIC is TRUE special routines compute all geometrical objects (which contain the metric tensor explicitly) by taking into consideration the diagonality of the metric. Reduced run times will, of course, result. Note: this option is set automatically by TSETUP if a diagonal metric is specified.

Variable DIM

default:[4] - An option in the CTENSR (Component Tensor Manipulation) package. DIM is the dimension of the manifold with the default 4. The command DIM:N; will reset the dimension to any other integral value.

Function EINSTEIN (dis)

A function in the CTENSR (Component Tensor Manipulation) package. EINSTEIN computes the mixed Einstein tensor after the Christoffel symbols and Ricci tensor have been obtained (with the functions CHRISTOF and RICCICOM). If the argument dis is TRUE, then the non-zero values of the mixed Einstein tensor G[i,j] will be displayed where j is the contravariant index. RATEINSTEIN[TRUE] if TRUE will cause the rational simplification on these components. If RATFAC[FALSE] is TRUE then the components will also be factored.

Function LRICCICOM (dis)

A function in the CTENSR (Component Tensor Manipulation) package. LRICCI-COM computes the covariant (symmetric) components LR[i,j] of the Ricci tensor. If the argument dis is TRUE, then the non-zero components are displayed.

Function MOTION (dis)

A function in the CTENSR (Component Tensor Manipulation) package. MOTION computes the geodesic equations of motion for a given metric. They are stored in the array EM[i]. If the argument dis is TRUE then these equations are displayed.

Variable OMEGA

default:[] - An option in the CTENSR (Component Tensor Manipulation) package. OMEGA assigns a list of coordinates to the variable. While normally defined when the function TSETUP is called, one may redefine the coordinates with the assignment OMEGA:[j1,j2,...jn] where the j’s are the new coordinate names. A call to OMEGA will return the coordinate name list. Also see DESCRIBE(TSETUP); .

Function RIEMANN (dis)

A function in the CTENSR (Component Tensor Manipulation) Package. RIEMANN computes the Riemann curvature tensor from the given metric and the corresponding Christoffel symbols. If dis is TRUE, the non-zero components R[i,j,k,l] will be displayed. All the indicated indices are covariant. As with the Einstein tensor, various switches set by the user control the simplification of the components of the Riemann tensor. If RATRIEMAN[TRUE] is TRUE then rational simplification will be done. If RATFAC[FALSE] is TRUE then each of the components will also be factored.

Function TRANSFORM

- The TRANSFORM command in the CTENSR package has been renamed to TTRANSFORM.

Function TSETUP ()

A function in the CTENSR (Component Tensor Manipulation) package which automatically loads the CTENSR package from within MACSYMA (if it is not already loaded) and then prompts the user to make use of it. Do DESCRIBE(CTENSR); for more details.

Function TTRANSFORM (matrix)

A function in the CTENSR (Component Tensor Manipulation) package which will perform a coordinate transformation upon an arbitrary square symmetric matrix. The user must input the functions which define the transformation. (Formerly called TRANSFORM.)

Appendix A: Function and Variable Index 239

Appendix A Function and Variable Index

%

% . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 %% . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 %EDISPFLAG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 %RNUM_LIST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 %TH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67, 165

?

?ROUND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 ?TRUNCATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

[

[index](expr) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161

"

"!!" . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 "!" . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 "#" . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 "''" . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 "'" . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 "." . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 "::=" . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 "::" . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 ":=" . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 ":" . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 "=" . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 "?" . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 "[" . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187

A

240 Maxima Manual

ABSBOXCHAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 ACOS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 ACOSH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 ACOT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 ACOTH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 ACSC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 ACSCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 ACTIVATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 ACTIVECONTEXTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 ADDCOL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 ADDITIVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 ADDROW. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 ADJOINT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 AIRY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 ALARMCLOCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 ALGEBRAIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 ALGEPSILON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 ALGEXACT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 ALGSYS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 ALIAS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 ALIASES. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 ALL_DOTSIMP_DENOMS. . . . . . . . . . . . . . . . . . . . . . . . 229 ALLBUT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 ALLOC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 ALLROOTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 ALLSYM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 ALPHABETIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 ANTID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ANTIDIFF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ANTISYMMETRIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 APPEND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 APPENDFILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 APPLY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 APPLY_NOUNS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 APPLY1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 APPLY2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 APPLYB1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 APROPOS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 ARGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 ARRAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 ARRAYAPPLY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 ARRAYINFO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 ARRAYMAKE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 ARRAYS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 ASEC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 ASECH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 ASIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 ASINH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 ASKEXP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 ASKINTEGER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 ASKSIGN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 ASSUME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 ASSUME_POS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 ASSUME_POS_PRED . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 ASSUMESCALAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 ASYMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 ASYMPA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 AT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 ATAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 ATAN2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 ATANH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 ATOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 ATOMGRAD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ATRIG1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 ATVALUE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 AUGCOEFMATRIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177

B

BACKSUBST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 BACKTRACE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 BACKUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 BASHINDICES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 BATCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 BATCHKILL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 BATCHLOAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 BATCON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 BATCOUNT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 BERLEFACT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 BERN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 BERNPOLY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 BESSEL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 BETA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 BEZOUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 BFFAC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 BFLOAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 BFLOATP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 BFPSI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 BFTORAT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 BFTRUNC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 BFZETA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 BGZETA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 BHZETA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 BINDTEST. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 BINOMIAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 BLOCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 BOTHCASES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 BOTHCOEF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 BOX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 BOXCHAR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 BREAK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 BREAKUP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 BUG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 BUILDQ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 BURN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 BZETA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205

C

CABS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 CANFORM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 CANTEN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 CARG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 CARTAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 CATCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 CAUCHYSUM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 CBFAC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 CF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 CFDISREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 CFEXPAND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 CFLENGTH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 CGAMMA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 CGAMMA2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 CHANGE_FILEDEFAULTS. . . . . . . . . . . . . . . . . . . . . . . . 69 CHANGEVAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 CHARPOLY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 CHECK_OVERLAPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 CHR1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 CHR2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 CHRISTOF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 CLEARSCREEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 CLOSEFILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 CLOSEPS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 COEFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 COEFMATRIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 COL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 COLLAPSE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 COLUMNVECTOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 COMBINE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 COMMUTATIVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 COMP2PUI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 COMPFILE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 COMPGRIND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 COMPILE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 COMPILE_FILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 COMPILE_LISP_FILE . . . . . . . . . . . . . . . . . . . . . . . . . . 59 CONCAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 CONJUGATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 CONS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 CONSTANT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 CONSTANTP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 CONT2PART . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 CONTENT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 CONTEXT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 CONTEXTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 CONTINUE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 CONTRACT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46, 129 COPYLIST. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 COPYMATRIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 COS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 COSH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 COT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 COTH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 COUNTER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 COVDIFF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 CREATE_LIST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 CSC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 CSCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

CURRENT_LET_RULE_PACKAGE . . . . . . . . . . . . . . . . . 199

CURSORDISP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 CURVATURE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236

D

DBLINT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 DDT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 DEACTIVATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 DEBUG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 DEBUGMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 DEBUGPRINTMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 DECLARE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 DECLARE_TRANSLATED . . . . . . . . . . . . . . . . . . . . . . . . . 66 DECLARE_WEIGHT . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227

DEFAULT_LET_RULE_PACKAGE . . . . . . . . . . . . . . . . . 200

DEFCON. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 DEFINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 DEFINE_VARIABLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 DEFINT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 DEFMATCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 DEFRULE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 DEFTAYLOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 DELETE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 DELFILE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 DELTA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 DEMO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 DEMOIVRE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 DENOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 DEPENDENCIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 DEPENDS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 DERIVABBREV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 DERIVDEGREE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 DERIVLIST. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 DERIVSUBST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 DESCRIBE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 DESOLVE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 DETERMINANT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 DETOUT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 DIAGMATRIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 DIAGMETRIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 DIFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 DIM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 DIMENSION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 DIREC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 DIRECT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 DISKFREE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 DISOLATE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 DISP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 DISPCON. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 DISPFLAG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 DISPFORM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 DISPFUN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 DISPLAY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

DISPLAY_FORMAT_INTERNAL . . . . . . . . . . . . . . . . . . . 70

DISPLAY2D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 DISPRULE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 DISPTERMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 DISTRIB. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 DIVIDE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 DIVSUM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 DO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 DOALLMXOPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 DOMAIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 DOMXEXPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 DOMXMXOPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 DOMXNCTIMES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 DONTFACTOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 DOSCMXOPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 DOSCMXPLUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 DOT0NSCSIMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 DOT0SIMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 DOT1SIMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 DOTASSOC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 DOTCONSTRULES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 DOTDISTRIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 DOTEXPTSIMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 DOTIDENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 DOTSCRULES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 DOTSIMP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 DPART . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 DSCALAR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 DSKALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 DUMMY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194

E

E . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 ECHELON. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 EIGENVALUES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 EIGENVECTORS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 EINSTEIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 ELE2COMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 ELE2POLYNOME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 ELE2PUI. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 ELEM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 ELIMINATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 EMATRIX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 ENDCONS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ENTERMATRIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 ENTIER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 EQUAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 ERF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ERFFLAG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ERRCATCH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 ERREXP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 ERRINTSCE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ERROR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 ERROR_SIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 ERROR_SYMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 ERRORFUN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 ERRORMSG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 EULER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 EV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 EVAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 EVENP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 EVFLAG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 EVFUN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 EXAMPLE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 EXP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 EXPAND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 EXPANDWRT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 EXPANDWRT_DENOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 EXPANDWRT_FACTORED . . . . . . . . . . . . . . . . . . . . . . . . . 38 EXPLOSE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 EXPON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 EXPONENTIALIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 EXPOP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 EXPRESS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 EXPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 EXPTDISPFLAG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 EXPTISOLATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 EXPTSUBST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

                                                EXTRACT_LINEAR_EQUATIONS . . . . . . . . . . . . . . . . . 228

EZGCD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

F

FACEXPAND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 FACTCOMB. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 FACTLIM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 FACTOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 FACTORFLAG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 FACTORIAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 FACTOROUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 FACTORSUM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 FACTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 FALSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 FASSAVE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 FAST_CENTRAL_ELEMENTS . . . . . . . . . . . . . . . . . . . . 227 FAST_LINSOLVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 FASTTIMES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 FEATURE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 FEATUREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 FEATURES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 FFT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 FIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 FIBTOPHI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 FILE_SEARCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 FILE_STRING_PRINT . . . . . . . . . . . . . . . . . . . . . . . . . . 73 FILE_TYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 FILEDEFAULTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 FILENAME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 FILENAME_MERGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 FILENUM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 FILLARRAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 FIRST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 FIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 FLOAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 FLOAT2BF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 FLOATDEFUNK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 FLOATNUMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 FLUSH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 FLUSHD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 FLUSHND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 FOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 FORGET. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 FORTINDENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 FORTMX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 FORTRAN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 FORTSPACES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 FPPREC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 FPPRINTPREC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 FREEOF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 FULLMAP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 FULLMAPL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 FULLRATSIMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 FULLRATSUBST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 FUNCSOLVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 FUNCTIONS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 FUNDEF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 FUNMAKE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

G

GAMMA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 GAMMALIM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 GAUSS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 GCD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 GCFACTOR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 GENDIFF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 GENFACT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 GENINDEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 GENMATRIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 GENSUMNUM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 GET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 GETCHAR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 GFACTOR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 GFACTORSUM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 GLOBALSOLVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 GO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 GRADEF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 GRADEFS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 GRAMSCHMIDT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 GRIND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 GROBNER_BASIS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227

H

HACH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 HALFANGLES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 HIPOW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 HORNER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

I

IBASE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 IC1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 IDENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 IEQN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 IEQNPRINT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 IF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 IFT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 ILT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 IMAGPART. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 INCHAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 INDICES. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 INF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 INFEVAL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 INFINITY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 INFIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 INFLAG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 INFOLISTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 INNERPRODUCT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 INPART . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 INRT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 INTEGERP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 INTEGRATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

INTEGRATION_CONSTANT_COUNTER . . . . . . . . . . . . . . 17

INTERPOLATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 INTFACLIM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 INTOPOIS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 INTOSUM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 INTPOLABS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 INTPOLERROR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 INTPOLREL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 INTSCE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 INVERT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 IS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 ISOLATE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 ISOLATE_WRT_TIMES . . . . . . . . . . . . . . . . . . . . . . . . . . 50 ISQRT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152

J

JACOBI. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207

K

KDELTA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 KEEPFLOAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 KILL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 KILLCONTEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 KOSTKA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133

L

LABELS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 LAPLACE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 LASSOCIATIVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 LAST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LASTTIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 LC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 LCM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 LDEFINT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 LDISP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 LDISPLAY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 LENGTH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 LET_RULE_PACKAGES . . . . . . . . . . . . . . . . . . . . . . . . . 202 LETRAT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 LETRULES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 LETSIMP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 LGTREILLIS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 LHOSPITALLIM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 LHS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 LIMIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 LINEAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 LINECHAR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 LINEDISP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 LINEL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 LINENUM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 LINSOLVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 LINSOLVE_PARAMS . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 LINSOLVEWARN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 LISPDEBUGMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 LIST_NC_MONOMIALS . . . . . . . . . . . . . . . . . . . . . . . . . 228 LISTARITH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LISTARRAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 LISTCONSTVARS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 LISTDUMMYVARS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 LISTOFVARS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 LISTP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LMXCHAR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 LOAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 LOADFILE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 LOADPRINT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 LOCAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 LOG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 LOGABS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 LOGARC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 LOGCONCOEFFP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 LOGCONTRACT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 LOGEXPAND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 LOGNEGINT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 LOGNUMER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 LOGSIMP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 LOPOW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 LORENTZ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 LPART . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 LRATSUBST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 LRICCICOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 LTREILLIS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133

M

M1PBRANCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 MACROEXPANSION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 MAINVAR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 MAKE_ARRAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 MAKEBOX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 MAKEFACT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 MAKEGAMMA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 MAKELIST. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 MAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 MAPATOM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 MAPERROR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 MAPLIST. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 MATCHDECLARE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 MATCHFIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 MATRIX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 MATRIX_ELEMENT_ADD. . . . . . . . . . . . . . . . . . . . . . . . 183 MATRIX_ELEMENT_MULT. . . . . . . . . . . . . . . . . . . . . . . 184

                                                MATRIX_ELEMENT_TRANSPOSE . . . . . . . . . . . . . . . . . 184

MATRIXMAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 MATRIXP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 MATTRACE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 MAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 MAXAPPLYDEPTH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 MAXAPPLYHEIGHT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 MAXNEGEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 MAXPOSEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 MAXPRIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 MAXTAYORDER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 MEMBER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 METRIC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 MIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 MINF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 MINFACTORIAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 MINOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 MOD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 MODE_CHECK_ERRORP . . . . . . . . . . . . . . . . . . . . . . . . . . 61 MODE_CHECK_WARNP . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 MODE_CHECKP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 MODE_DECLARE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 MODE_IDENTITY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 MODULUS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 MON2SCHUR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 MONO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 MONOMIAL_DIMENSIONS. . . . . . . . . . . . . . . . . . . . . . . 228 MOTION. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 MULTI_ELEM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 MULTI_ORBIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 MULTI_PUI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 MULTINOMIAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 MULTIPLICATIVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 MULTIPLICITIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 MULTSYM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 MULTTHRU. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 MYOPTIONS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169

N

NC_DEGREE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 NCEXPT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 NCHARPOLY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 NEGDISTRIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 NEGSUMDISPFLAG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 NEW-DISREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 NEWCONTEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 NEWDET. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 NEWFAC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 NEWTON. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 NICEINDICES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 NICEINDICESPREF . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 NOEVAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 NOLABELS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 NONSCALAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 NONSCALARP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 NOSTRING. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 NOUN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 NOUNDISP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 NOUNIFY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 NOUNS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 NROOTS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 NTERMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 NTERMSG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 NTERMSRCI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 NTHROOT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 NUM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 NUMBERP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 NUMER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 NUMERVAL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 NUMFACTOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 NUSUM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 NZETA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233

O

OBASE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 ODDP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 ODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 ODE2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 OMEGA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 OPPROPERTIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 OPSUBST. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 OPTIMIZE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 OPTIMPREFIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 OPTIONSET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 ORBIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 ORDERGREAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 ORDERGREATP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 ORDERLESS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 ORDERLESSP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 OUTATIVE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 OUTCHAR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 OUTOFPOIS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163

P

PACKAGEFILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 PADE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 PARSEWINDOW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 PART . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 PART2CONT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 PARTFRAC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 PARTITION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 PARTPOL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 PARTSWITCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 PCOEFF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 PERMANENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 PERMUT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 PFEFORMAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 PI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 PICKAPART . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 PIECE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 PLAYBACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 PLOG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 PLOT_OPTIONS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 PLOT2D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 PLOT2D_PS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 PLOT3D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 POISDIFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 POISEXPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 POISINT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 POISLIM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 POISMAP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 POISPLUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 POISSIMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 POISSON. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 POISSUBST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 POISTIMES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 POISTRIM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 POLARFORM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 POLARTORECT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 POLYNOME2ELE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 POSFUN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 POTENTIAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 POWERDISP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 POWERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 POWERSERIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 PRED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 PREDERROR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 PRIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 PRIMEP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 PRINT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 PRINTPOIS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 PRINTPROPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 PRODHACK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 PRODRAC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 PRODUCT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 PROGRAMMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 PROMPT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 PROPERTIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 PROPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 PROPVARS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 PSCOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 PSDRAW_CURVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 PSEXPAND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 PSI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 PUI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 PUI_DIRECT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 PUI2COMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 PUI2ELE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 PUI2POLYNOME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 PUIREDUC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 PUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195

Q

QPUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 QQ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 QUANC8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 QUIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 QUNIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 QUOTIENT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

R

RADCAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 RADEXPAND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 RADPRODEXPAND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 RADSUBSTFLAG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 RAISERIEMANN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 RANDOM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 RANK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 RASSOCIATIVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 RAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 RATALGDENOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 RATCOEF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 RATDENOM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 RATDENOMDIVIDE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 RATDIFF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 RATDISREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 RATEINSTEIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 RATEPSILON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 RATEXPAND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 RATFAC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 RATMX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 RATNUMER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 RATNUMP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 RATP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 RATPRINT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 RATRIEMAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 RATRIEMANN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 RATSIMP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 RATSIMPEXPONS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 RATSUBST. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 RATVARS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 RATWEIGHT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 RATWEIGHTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 RATWEYL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 RATWTLVL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 READ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 READONLY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 REALONLY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 REALPART. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 REALROOTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 REARRAY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 RECTFORM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 RECTTOPOLAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 REFCHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 REM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 REMAINDER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 REMARRAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 REMBOX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 REMCON. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 REMFUNCTION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 REMLET. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 REMOVE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 REMRULE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 REMTRACE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 REMVALUE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 RENAME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 RESET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 RESIDUE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 RESOLVANTE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 RESOLVANTE_ALTERNEE1. . . . . . . . . . . . . . . . . . . . . . 143 RESOLVANTE_BIPARTITE. . . . . . . . . . . . . . . . . . . . . . 143 RESOLVANTE_DIEDRALE. . . . . . . . . . . . . . . . . . . . . . . 143 RESOLVANTE_KLEIN . . . . . . . . . . . . . . . . . . . . . . . . . . 144 RESOLVANTE_KLEIN3 . . . . . . . . . . . . . . . . . . . . . . . . . 144

RESOLVANTE_PRODUIT_SYM . . . . . . . . . . . . . . . . . . . 144

RESOLVANTE_UNITAIRE. . . . . . . . . . . . . . . . . . . . . . . 145 RESOLVANTE_VIERER . . . . . . . . . . . . . . . . . . . . . . . . . 145 REST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 RESTORE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 RESULTANT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 RETURN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 REVEAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 REVERSE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 REVERT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 RHS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 RICCICOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 RIEMANN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 RINVARIANT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 RISCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 RMXCHAR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 RNCOMBINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 ROMBERG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ROMBERGABS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ROMBERGIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ROMBERGMIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ROMBERGTOL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ROOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 ROOTSCONMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 ROOTSCONTRACT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 ROOTSEPSILON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 ROW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185

S

SAVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 SAVEDEF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 SAVEFACTORS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 SCALARMATRIXP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 SCALARP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 SCALEFACTORS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 SCANMAP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 SCHUR2COMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 SCONCAT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 SCSIMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 SCURVATURE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 SEC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 SECH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 SET_PLOT_OPTION . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

                                                SET_UP_DOT_SIMPLIFICATIONS . . . . . . . . . . . . . . . 227

SETCHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 SETCHECKBREAK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 SETELMX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 SETUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 SETUP_AUTOLOAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 SETVAL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 SHOW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 SHOWRATVARS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 SHOWTIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 SIGN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 SIGNUM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 SIMILARITYTRANSFORM. . . . . . . . . . . . . . . . . . . . . . . 185 SIMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 SIMPSUM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 SIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 SINH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 SOLVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116

                                                SOLVE_INCONSISTENT_ERROR . . . . . . . . . . . . . . . . . 117

SOLVEDECOMPOSES . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 SOLVEEXPLICIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 SOLVEFACTORS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 SOLVENULLWARN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 SOLVERADCAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 SOLVETRIGWARN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 SOMRAC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 SORT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 SPARSE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 SQFR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 SQRT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 SQRTDISPFLAG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 SRRAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 SSTATUS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 STARDISP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 STATUS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 STRING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 STRINGOUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 SUBLIS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 SUBLIS_APPLY_LAMBDA. . . . . . . . . . . . . . . . . . . . . . . 154 SUBLIST. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 SUBMATRIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 SUBST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 SUBSTINPART . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 SUBSTPART . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 SUBVARP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 SUM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 SUMCONTRACT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 SUMEXPAND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 SUMHACK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 SUMSPLITFACT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 SUPCONTEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 SYMBOLP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 SYMMETRIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 SYSTEM(command) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

T

TAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 TANH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 TAYLOR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 TAYLOR_LOGEXPAND . . . . . . . . . . . . . . . . . . . . . . . . . . 127

TAYLOR_ORDER_COEFFICIENTS . . . . . . . . . . . . . . . . 127

TAYLOR_SIMPLIFIER . . . . . . . . . . . . . . . . . . . . . . . . . 127

TAYLOR_TRUNCATE_POLYNOMIALS . . . . . . . . . . . . . . 127

TAYLORDEPTH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 TAYLORINFO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 TAYLORP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 TAYTORAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 TCONTRACT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 TELLRAT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 TELLSIMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 TELLSIMPAFTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 THROW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 TIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 TIMER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 TIMER_DEVALUE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 TIMER_INFO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 TLDEFINT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 TLIMIT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 TLIMSWITCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 TO_LISP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 TOBREAK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 TODD_COXETER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 TOPLEVEL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 TOTALDISREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 TOTIENT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 TPARTPOL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 TR_ARRAY_AS_REF . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

TR_BOUND_FUNCTION_APPLYP . . . . . . . . . . . . . . . . . . 64

TR_FILE_TTY_MESSAGESP . . . . . . . . . . . . . . . . . . . . . 64

TR_FLOAT_CAN_BRANCH_COMPLEX . . . . . . . . . . . . . . . 64
TR_FUNCTION_CALL_DEFAULT . . . . . . . . . . . . . . . . . . 64

TR_GEN_TAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 TR_NUMER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 TR_OPTIMIZE_MAX_LOOP. . . . . . . . . . . . . . . . . . . . . . . 65

TR_OUTPUT_FILE_DEFAULT . . . . . . . . . . . . . . . . . . . . 65
TR_PREDICATE_BRAIN_DAMAGE . . . . . . . . . . . . . . . . . 65

TR_SEMICOMPILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 TR_STATE_VARS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

TR_TRUE_NAME_OF_FILE_BEING_TRANSLATED . . . . 65

TR_VERSION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

TR_WARN_BAD_FUNCTION_CALLS . . . . . . . . . . . . . . . . 65

TR_WARN_FEXPR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 TR_WARN_MEVAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 TR_WARN_MODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 TR_WARN_UNDECLARED . . . . . . . . . . . . . . . . . . . . . . . . . 66

TR_WARN_UNDEFINED_VARIABLE . . . . . . . . . . . . . . . . 66

TR_WARNINGS_GET . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 TR_WINDY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 TRACE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 TRACE_OPTIONS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 TRANSBIND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 TRANSCOMPILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 TRANSFORM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 TRANSLATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 TRANSLATE_FILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 TRANSPOSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 TRANSRUN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 TREILLIS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 TREINAT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 TRIANGULARIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 TRIGEXPAND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 TRIGEXPANDPLUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 TRIGEXPANDTIMES . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 TRIGINVERSES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 TRIGRAT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 TRIGREDUCE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 TRIGSIGN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 TRIGSIMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 TRUE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 TRUNC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 TSETUP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 TTRANSFORM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 TTYINTFUN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 TTYINTNUM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 TTYOFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

U

UNDECLAREDWARN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 UNDIFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 UNITEIGENVECTORS . . . . . . . . . . . . . . . . . . . . . . . . . . 186 UNITVECTOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 UNKNOWN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 UNORDER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 UNSUM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 UNTELLRAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 UNTRACE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 USE_FAST_ARRAYS . . . . . . . . . . . . . . . . . . . . . . . . . . . 120

V

VALUES. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 VECT_CROSS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 VECTORPOTENTIAL . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 VECTORSIMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 VERB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 VERBIFY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 VERBOSE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

W

WEYL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234

                                                WITH_STDOUT(file,stmt1,stmt2,...). . . . . . . . . 78

WRITEFILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

X

xgraph_curves(list). . . . . . . . . . . . . . . . . . . . . . . . 91 XTHRU . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157

Z

ZEROBERN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 ZEROEQUIV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 ZEROMATRIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 ZETA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 ZETA%PI. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 ZRPOLY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 ZSOLVE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 ZUNDERFLOW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

250 Maxima Manual

Short Contents

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1 Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2 Differentiation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3 Integration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 4 Polynomials . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 5 Simplification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 6 Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 7 Function Definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 8 Input and Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 9 Program Flow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 10 Plotting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 11 Floating Point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 12 Numerical . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 13 Trigonometric . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 14 Equations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 15 Arrays and Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 16 Series . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 17 Symmetries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 18 Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 19 Operators. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 20 Constants. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 21 Special Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 22 Command Line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 23 Contexts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 24 Matrices and Linear Algebra . . . . . . . . . . . . . . . . . . . . . . 177 25 Runtime Environment . . . . . . . . . . . . . . . . . . . . . . . . . . 189 26 Miscellaneous Options . . . . . . . . . . . . . . . . . . . . . . . . . . 193 27 Rules and Patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 28 Number Theory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 29 Help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 30 Differential Equations . . . . . . . . . . . . . . . . . . . . . . . . . . 215 31 Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 32 Limits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 33 Logarithms. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 34 Debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 ii Maxima Manual

35 Affine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 36 Tensor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 37 Ctensor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 Appendix A Function and Variable Index . . . . . . . . . . . . . . . 239

Table of Contents

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1 Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1.1 Introduction to Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 Definitions for Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2 Differentiation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.1 Definitions for Differentiation . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

3 Integration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

3.1 Introduction to Integration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.2 Definitions for Integration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

4 Polynomials . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

4.1 Introduction to Polynomials. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.2 Definitions for Polynomials. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

5 Simplification. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

5.1 Definitions for Simplification . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

6 Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

6.1 Introduction to Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 6.2 ASSIGNMENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 6.3 COMPLEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 6.4 INEQUALITY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 6.5 SYNTAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 6.6 Definitions for Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

7 Function Definition . . . . . . . . . . . . . . . . . . . . . . . 57

7.1 Introduction to Function Definition . . . . . . . . . . . . . . . . . . . . . . 57 7.2 FUNCTION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 7.3 MACROS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 7.4 OPTIMIZATION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 7.5 Definitions for Function Definition . . . . . . . . . . . . . . . . . . . . . . . 57

8 Input and Output . . . . . . . . . . . . . . . . . . . . . . . . 67

8.1 Introduction to Input and Output . . . . . . . . . . . . . . . . . . . . . . . 67 8.2 FILES. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 8.3 PLAYBACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 8.4 Definitions for Input and Output . . . . . . . . . . . . . . . . . . . . . . . . 67

9 Program Flow . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

9.1 Introduction to Program Flow. . . . . . . . . . . . . . . . . . . . . . . . . . . 81 9.2 Definitions for Program Flow. . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

10 Plotting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

10.1 Definitions for Plotting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

11 Floating Point . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

11.1 Definitions for Floating Point . . . . . . . . . . . . . . . . . . . . . . . . . . 93

12 Numerical. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

12.1 Introduction to Numerical . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 12.2 DCADRE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 12.3 ELLIPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 12.4 FOURIER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 12.5 NDIFFQ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 12.6 Definitions for Numerical . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

13 Trigonometric . . . . . . . . . . . . . . . . . . . . . . . . . . 105

13.1 Introduction to Trigonometric. . . . . . . . . . . . . . . . . . . . . . . . . 105 13.2 Definitions for Trigonometric. . . . . . . . . . . . . . . . . . . . . . . . . . 105

14 Equations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

14.1 Definitions for Equations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

15 Arrays and Tables . . . . . . . . . . . . . . . . . . . . . . 119

15.1 Definitions for Arrays and Tables . . . . . . . . . . . . . . . . . . . . . . 119

16 Series . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

16.1 Introduction to Series . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 16.2 Definitions for Series . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

17 Symmetries . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

17.1 Definitions for Symmetries . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

18 Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147

18.1 Definitions for Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147

19 Operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149

19.1 NARY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 19.2 NOFIX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 19.3 OPERATOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 19.4 POSTFIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 19.5 PREFIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 19.6 Definitions for Operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149

20 Constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159

20.1 Definitions for Constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159

21 Special Functions . . . . . . . . . . . . . . . . . . . . . . . 161

21.1 Introduction to Special Functions . . . . . . . . . . . . . . . . . . . . . 161 21.2 GAMALG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 21.3 SPECINT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 21.4 Definitions for Special Functions . . . . . . . . . . . . . . . . . . . . . . 162

22 Command Line . . . . . . . . . . . . . . . . . . . . . . . . . 165

22.1 Introduction to Command Line . . . . . . . . . . . . . . . . . . . . . . . 165 22.2 Definitions for Command Line . . . . . . . . . . . . . . . . . . . . . . . . 165

23 Contexts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

23.1 Definitions for Contexts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

24 Matrices and Linear Algebra . . . . . . . . . . . . 177

24.1 Introduction to Matrices and Linear Algebra . . . . . . . . . . . 177 24.1.1 DOT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 24.1.2 VECTORS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177

24.2 Definitions for Matrices and Linear Algebra . . . . . . . . . . . . 177

25 Runtime Environment . . . . . . . . . . . . . . . . . . 189

25.1 Introduction for Runtime Environment . . . . . . . . . . . . . . . . 189 25.2 INTERRUPTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 25.3 Definitions for Runtime Environment . . . . . . . . . . . . . . . . . . 189

26 Miscellaneous Options . . . . . . . . . . . . . . . . . . 193

26.1 Introduction to Miscellaneous Options . . . . . . . . . . . . . . . . . 193 26.2 SHARE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 26.3 Definitions for Miscellaneous Options . . . . . . . . . . . . . . . . . . 193

27 Rules and Patterns . . . . . . . . . . . . . . . . . . . . . 199

27.1 Introduction to Rules and Patterns . . . . . . . . . . . . . . . . . . . . 199 27.2 Definitions for Rules and Patterns . . . . . . . . . . . . . . . . . . . . . 199

28 Number Theory . . . . . . . . . . . . . . . . . . . . . . . . 205

28.1 Definitions for Number Theory . . . . . . . . . . . . . . . . . . . . . . . . 205

29 Help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211

29.1 Introduction to Help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 29.2 LISP-AND-MACSYMA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 29.3 GARBAGE COLLECTION. . . . . . . . . . . . . . . . . . . . . . . . . . . 212 29.4 DOCUMENTATION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 29.5 Definitions for Help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213

30 Differential Equations. . . . . . . . . . . . . . . . . . . 215

30.1 Definitions for Differential Equations . . . . . . . . . . . . . . . . . . 215

31 Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219

31.1 Definitions for Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219

32 Limits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221

32.1 Definitions for Limits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221

33 Logarithms. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223

33.1 Definitions for Logarithms . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223

34 Debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225

34.1 Definitions for Debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225

35 Affine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227

35.1 Definitions for Affine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227

36 Tensor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231

36.1 Introduction to Tensor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 36.2 Definitions for Tensor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231

37 Ctensor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235

37.1 Introduction to Ctensor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 37.2 Definitions for Ctensor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236

Appendix A Function and Variable Index … 239